### v1

# The Analysis of Willmore Surfaces and its Generalizations in Higher Dimensions

Abstract
We review recent progress concerning the analysis of Lagrangians on immersions into $\mathbb{R}^{d}$ depending on the first and second fundamental forms and their covariant derivatives.
In honor of Professor Gang Tian on the occasion of his 65th birthday
I Introduction
The study of topological and geometrical properties of manifolds through the search of special metrics “equipping” these manifolds and solving some special Partial Differential Equation has a long and rich history which takes its origin maybe in the construction of special curves (like Brachistochrone curves in the XVIIth century, or Euler’s elastica in the XVIIIth century) and constant Gauss curvature metrics on closed surfaces in relation with the uniformization theorem for Riemann Surfaces in the XVIIIth century. The development of what could be called “intrinsic geometric analysis” has been the source of spectacular results in differential topology, differential geometry and complex geometry with the search of constant scalar curvature metric, Einstein metrics, Kähler Einstein metrics, solutions to the Ricci flow… Another branch of geometric analysis is dealing with the study of “special submanifolds” within a given Riemannian manifold and its interaction with the geometry of the manifold itself. This branch of geometric analysis is maybe rooted originally both in the calculus of variations with the variational constructions of closed geodesics, the resolution of the Lagrange-Plateau problem in Euclidean space as well in the “explicit differential geometry” of special submanifolds such as plane algebraic curves, algebraic surfaces, etc. The central objects of what can be described as “extrinsic geometric analysis” are minimal surfaces and their generalizations (constant and prescribed mean curvature surfaces). The underlying PDEs are related to the area variations under various constraints. In the present work, we consider a branch of extrinsic variational geometric analysis that has seen substantial development in recent decades, and concerns immersions that arise as critical points of Lagrangians depending on the first and second fundamental form, and their covariant derivatives. The simplest and maybe most studied model for such a Lagrangian is the Willmore functional of immersed surfaces into ${\mathbb{R}}^{m}$. Let $\Sigma$ be closed oriented two dimensional manifold and let $\vec{\Phi}$ be a $C^{2}$ immersion of this surface into the three dimensional Euclidean space ${\mathbb{R}}^{m}$. Denoting by $g_{\vec{\Phi}}$ the first fundamental form of this immersion and by $\vec{\mathrm{I\!I}}_{\vec{\Phi}}$ its second fundamental form, the mean curvature of the immersion is given by
$\vec{H}_{\vec{\Phi}}\coloneqq\frac{1}{2}\mbox{tr}_{g_{\vec{\Phi}}}\vec{\mathrm{I\!I}}_{\vec{\Phi}}.$
The Willmore energy is
$W(\vec{\Phi})\coloneqq\int_{\Sigma}|\vec{H}_{\vec{\Phi}}|^{2}\ d\textup{vol}_{g_{\vec{\Phi}}}.$
This energy has been initially introduced by Sophie Germain and Simeon Poisson in an attempt to generalize to two dimensional elastic membranes the famous Euler Elastica modelizing the free energy of a beam [64, 160]. The very first derivation of the Euler–Lagrange equation of $W$ was made by Poisson himself around 1814 in the case of a graph (see page 60 of [160] and chapter 10 of [137]) in $\mathbb{R}^{3}$. Using concepts which were not completely clarified at the time (such as the Gauss curvature and the Laplace–Beltrami operator), we can rewrite the Poisson’s Euler–Lagrange equation for the Lagrangian $W$ in the following form
$\Delta_{g}H_{\vec{\Phi}}+2H_{\vec{\Phi}}\,(H^{2}_{\vec{\Phi}}-K_{\vec{\Phi}})=0,$ (I.1)
where $H_{\vec{\Phi}}=\vec{n}_{\vec{\Phi}}\cdot\vec{H}_{\vec{\Phi}}$ is the mean curvature and $\vec{n}_{\vec{\Phi}}$ the unit Gauss map to the immersion, $\Delta_{g}$ is the negative Laplace–Beltrami operator which reads in local coordinates $(x^{1},x^{2})$ $\Delta_{g}\cdot\coloneqq(\det(g))^{-1/2}\partial_{x_{i}}(g^{ij}\,(\det(g))^{1/2}\partial_{x_{j}}\cdot)$ and $K_{\vec{\Phi}}\coloneqq\mbox{det}_{g_{\vec{\Phi}}}(\vec{n}_{\vec{\Phi}}\cdot\vec{\mathrm{I\!I}}_{\vec{\Phi}})$ is the Gauss curvature. Throughout this paper, we shall often write $g=g_{\vec{\Phi}}$ when there is no ambiguity. With these notations we have also
$\vec{H}_{\vec{\Phi}}=H_{\vec{\Phi}}\ \vec{n}_{\vec{\Phi}}=\frac{1}{2}\,\left(\vec{n}_{\vec{\Phi}}\cdot\Delta_{g}\vec{\Phi}\right)\ \vec{n}_{\vec{\Phi}}=\frac{1}{2}\,\Delta_{g}\vec{\Phi}.$ (I.2)
Comparing (I.2) with (I.1), it appears that Poisson’s Euler–Lagrange equation is a degenerate nonlinear fourth-order equation combining analytical difficulties which were way too advanced for the early XIXth century. Hence in a natural way, in the following 100 years, the attention has been exclusively devoted to special solutions solving a simpler second order version of (I.1):
$H_{\vec{\Phi}}=0.$
This is the minimal surface equation corresponding to immersions that are critical points of the area.
Poisson equation (I.1) has been rediscovered by Wilhelm Balschke [21] and his student Gerhard Thomsen [183] a century after (around 1924) in the framework of conformal geometry. Hence, the study of the Lagrangian $W$ was motivated by the merging of two theories: conformal geometry and minimal surfaces theory. On the one hand they proved that $W$ is invariant under the action of generic [^1] Precisely, for any conformal transformation $\Psi$ from ${\mathbb{R}}^{3}\cup\{\infty\}$ into itself and for any immersion $\vec{\Phi}$ of an oriented closed two dimensional manifold $\Sigma$ into ${\mathbb{R}}^{3}$ such that $\Psi^{-1}(\{\infty\})\cap\vec{\Phi}(\Sigma)=\emptyset$ there holds $W(\Psi\circ\vec{\Phi})=W(\vec{\Phi})\ .$ conformal transformations on the other hand, as it has been observed above, minimal surfaces are stable critical points of $W$. Probably because of the absence of known examples that were not just composition of minimal immersions with conformal transformations, the interest in functional $W$ disappeared again from the mathematical literature for a few decades until the famous paper [188] by Thomas Willmore in 1965. In this paper Willmore proved that for any immersion $\vec{\Phi}$ of a closed surface $\Sigma$ the following lower bound holds
$W(\vec{\Phi})\geq 4\pi.$
Moreover, there is equality if and only if $\vec{\Phi}(\Sigma)$ is a round sphere. This lower bound could be interpreted as a Fenchel type theorem for surfaces or as a variant of the famous Chern–Lashof inequality (see [37]) stating that
$\int_{\Sigma}|K_{\vec{\Phi}}|\ d\textup{vol}_{g_{\vec{\Phi}}}\geq 4\pi.$
There is equality if and only if $\vec{\Phi}(\Sigma)$ is a convex surface. Willmore then conjectured that if $\Sigma$ is not diffeomorphic to $\mathbb{S}^{2}$ then the lower bound is increased and should be equal to $2\pi^{2}$, with equality if and only if $\vec{\Phi}(\Sigma)$ is a compact conformal transformation of the 2-torus obtained by rotating the vertical circle included in the xzx-z plane of radius 1 and centered at the point $(\sqrt{2},0,0)$. Since this paper the Sophie Germain Functional was called Willmore Functional.
The first analytical work on the Willmore functional $W$ is the paper by Leon Simon [180] in which he shows that the infimum of $W$ is achieved among all possible $C^{2}$ immersions of the torus by a smooth immersion. Relying for a large part on the analysis of [180], Matthias Bauer and Ernst Kuwert in [12], by proving some clever strict inequalities on connected sums of immersions of surfaces, have been able to extend Simon’s result to any genus, that is, for any genus $g$ the infimum of $W$ among all immersions of the surface of genus $g$ is achieved. What should be noticed regarding the analysis in [180] is that first the minimization is considered in measure theoretical sense using the theory of varifolds, then the regularity is relying a lot on comparison arguments with local competitors which are graphs satisfying the biharmonic equation (i.e. the linearised version of (I.1)). In this sense, the analysis in [180] is very much restricted to minimization operations.
The third author of the present work has been considering in the early 2000 the possibility to use more classical functional analysis to perform the minimisation of $W$ with the perspective of considering more general variational arguments such as min-max operations, or the study of the associated gradient flow. One of the first difficulties was to obtain an Euler–Lagrange equation compatible with the functional space in which the variations of $W$ are studied. In order to illustrate this difficulty, it could be interesting to look at a simpler framework and to go one dimension lower. We consider the variations of the Euler Elastica for immersions $\vec{\Phi}$ of the segment $[0,1]$ into the plane $\mathbb{R}^{2}$ given by
$E(\vec{\Phi})=\int_{[0,1]}\kappa^{2}_{\vec{\Phi}}\ dl_{\vec{\Phi}},$
where $\kappa_{\vec{\Phi}}$ is the curvature of the immersion and $dl_{\vec{\Phi}}$ is the length element of the immersion. The critical points to $E$ are known to satisfy the Euler–Lagrange equation
$2\,\ddot{\kappa}_{\vec{\Phi}}+\kappa^{3}_{\vec{\Phi}}=0.$ (I.3)
This can be seen as the one-dimensional counterpart of the Willmore equation (I.1). Obviously, this equation does not present any difficulty by itself and can be solved almost explicitly by multiplying by $\dot{\kappa}_{\vec{\Phi}}$, integrating and solving the following first-order ODE using elliptic integrals
$\dot{\kappa}^{2}_{\vec{\Phi}}+\frac{\kappa^{4}_{\vec{\Phi}}}{4}=c^{2}_{0}\qquad\Longleftrightarrow\qquad\frac{\dot{\kappa}_{\vec{\Phi}}}{\sqrt{{c^{2}_{0}}-\frac{\kappa^{4}_{\vec{\Phi}}}{4}}}=\pm 1.$
But this is not the solvability of (I.3) which is addressed at this stage, but rather the compatibility between the Euler–Lagrange equation and the Lagrangian from a purely variational perspective. In unit speed parameterization $\vec{\Phi}(t)$ for $s\in[0,L]$ and $L=\int dl_{\vec{\Phi}}$ we have $\kappa_{\vec{\Phi}}\,\vec{n}_{\vec{\Phi}}=\frac{d^{2}\vec{\Phi}}{ds^{2}}$ and thus it holds
$E(\vec{\Phi})=\int_{0}^{L}\left|\frac{d^{2}\vec{\Phi}}{ds^{2}}\right|^{2}ds.$
Hence, obviously, the natural space in which one should consider the variations of $E$ is the Sobolev space $W^{2,2}([0,L],{\mathbb{R}}^{2})$. The problem is that the nonlinearity in the Euler–Lagrange equation involves the cube of the second derivative of $\vec{\Phi}$ and a priori
$\kappa^{3}_{\vec{\Phi}}=\left(\vec{n}_{\vec{\Phi}}\cdot\frac{d^{2}\vec{\Phi}}{dt^{2}}\right)^{3}\qquad\notin L^{1}_{\textup{loc}}([0,L]).$
This does not define a distribution for an arbitrary $\vec{\Phi}$ in $W^{2,2}([0,L],{\mathbb{R}}^{2})$. As paradoxal as it appears, the Euler–Lagrange equation (I.3) is not compatible with the Lagrangian itself!
The same paradox is present one dimension higher for the Willmore functional. Using (I.2) we have
$W(\vec{\Phi})=\frac{1}{4}\int_{\Sigma}|\Delta_{g}\vec{\Phi}|^{2}\ d\textup{vol}_{g}.$
Naturally the function space to be considered is the space $W^{2,2}(\Sigma,{\mathbb{R}}^{m})$ and, even assuming that the metric $g_{\vec{\Phi}}$ defined on $T\Sigma$ and its inverse $g^{-1}_{\vec{\Phi}}$ would be bounded in $L^{\infty}(\Sigma)$, the non-linearity in the Euler–Lagrange (I.1) is containing terms such as
$2\,H_{\vec{\Phi}}\,|H_{\vec{\Phi}}|^{3}\qquad\notin L^{1}_{\textup{loc}}(\Sigma).$
As well, this does not define a distribution a priori. Hence, the question of constructing an ad-hoc framework to study the variations of $W$ was open. To that aim, the third author introduced in 2010 (paper published in 2014 [166]) the notion of $W^{2,2}$ weak immersions:
A map $\vec{\Phi}\in W^{1,\infty}\cap W^{2,2}(\Sigma,{\mathbb{R}}^{m})$ is said to be a $W^{2,2}$ weak immersion [^2] In the original work as well as in subsequent works such as [168] the author introduced the space of weak immersions ${\mathcal{E}}_{\Sigma}$ which only requires $\vec{\Phi}\in W^{1,\infty}$ as well as (I.4) and the fact that the Gauss map is in $W^{1,2}(\Sigma,G_{2}({\mathbb{R}}^{m}))$. This later hypothesis is implied by the assumption $\vec{\Phi}\in W^{1,\infty}\cap W^{2,2}(\Sigma,{\mathbb{R}}^{m})$ together with (I.4). The reverse is not true as observed in [104], that is weak immersions are not necessarily $W^{2,2}$ in a new bi-Lipschitz chart and hence the space of $W^{2,2}$ weak immersions we are defining here is smaller than the original space defined in [166]. It enjoys nevertheless the desired almost weak closure property (see Theorem I.2). if for a given smooth reference metric $g_{0}$ on the closed surface $\Sigma$ there exists $C_{\vec{\Phi}}>1$ such that
$C_{\vec{\Phi}}^{-1}\,|X|_{g_{0}}^{2}\leq|X|_{g_{\vec{\Phi}}}^{2}=|\vec{\Phi}_{\ast}X|^{2}_{{\mathbb{R}}^{m}}\leq C_{\vec{\Phi}}\,|X|_{g_{0}}^{2},\qquad\text{for a.e. }p\in\Sigma\,\text{ and all }X\in T_{p}\Sigma.$ (I.4)
The interest of considering $W^{2,2}$ weak immersions is due to an “almost weak closure” property which is going to be a consequence of the existence of an underlying smooth conformal structure for any $W^{2,2}$ weak immersion. Precisely we have the following result.
Theorem I.1.
Let $\vec{\Phi}$ be a $W^{2,2}$ weak immersion defined on a smooth closed orientable surface $\Sigma$ then there exists a constant Gauss curvature metric $h$ on $\Sigma$ and a function $\alpha\in C^{0}\cap W^{1,2}(\Sigma,{\mathbb{R}})$ such that
$g_{\vec{\Phi}}=e^{2\alpha}\,h.$ (I.5)
Moreover $\alpha$ satisfies the Liouville equation
$\Delta_{h}\alpha=e^{2\alpha}\,K_{g_{\vec{\Phi}}}-K_{h}.$ (I.6)
The Gauss–Bonnet theorem holds, denoting $\gamma(\Sigma)$ is the genus of $\Sigma$,
$\int_{\Sigma}K_{g_{\vec{\Phi}}}\,d\textup{vol}_{g_{\vec{\Phi}}}=4\pi\,(1-\gamma(\Sigma)).$ (I.7)
A proof of the first part of this result was proposed in [168] but a very last ingredient which requires $\vec{\Phi}\in W^{2,2}(\Sigma,{\mathbb{R}}^{m})$ and not only $\vec{n}_{\vec{\Phi}}\in W^{1,2}(\Sigma,G_{2}({\mathbb{R}}^{m}))$ was missing to complete the argument and to ensure that $\phi$ given by (1.72) in the middle of page 323 is a bi-Lipschitz homeomorphism. The complete argument for proving Theorem I.1 is given in Section IV of the present paper. Observe that the existence of bi-Lipschitz isothermal charts does not hold if one only assumes that the Gauss map is in $W^{1,2}$ without assuming that the weak immersion is in $W^{2,2}$: in the local sense, a counterexample $z\rightarrow(z^{2}/|z|,0)\in\mathbb{R}^{3}$ is provided by the first author in [104], and when $\Sigma=\mathbb{S}^{2}$, Plotnikov [159] construct a global counterexample by composing the map $z\rightarrow z^{2}/|z|$ with the stereographic projection.
A consequence of Theorem I.1, that is, the existence of an underlying smooth conformal structure, is the following almost sequential weak closure property for $W^{2,2}$ weak immersions proved first in [166] when the underlying conformal class is precompact and in the general case in [107]. It relies on Deligne–Mumford compactification of the moduli space of conformal structures (see for instance [88] for the definition of nodal surfaces).
Theorem I.2.
Let $\Sigma$ be a closed surface and let $\vec{\Phi}_{k}$ be a sequence of $W^{2,2}$ weak immersion into ${\mathbb{R}}^{m}$ with uniformly bounded $L^{2}$ norm of the second fundamental form. Then, up to a subsequence, for any connected component of the limiting punctured nodal surface $\tilde{\Sigma}$ there exist a Möbius transformation $\Xi_{k}$ of $\mathbb{R}^{m}$, a sequence of $W^{2,2}$ bi-Lipschitz homeomorphisms $\psi_{k}$ and at most finitely many points $a_{1},\dots,a_{L}$} of $\tilde{\Sigma}$ such that
$\vec{\xi}_{k}\coloneqq\Xi_{k}\circ\vec{\Phi}_{k}\circ\psi_{k}\ \rightharpoonup\ \vec{\xi}_{\infty}\qquad\mbox{ weakly in }\quad W^{2,2}_{\textup{loc}}(\tilde{\Sigma}\setminus\{a_{1},\dots,a_{L}\}),$
moreover $\vec{\xi}_{k}$ extends as a possibly branched $W^{2,2}$ weak immersion into ${\mathbb{R}}^{m}$.
Some weeks after the proof of Theorem I.2 has been posted on arxiv, a proof of this result has appeared on arxiv as well but assuming the sequence $\vec{\Phi}_{k}$ to be conformal with respect to some sequence of Riemann surface (see [100]). It is important to insist at this stage that the restriction to conformal weak immersions is missing the goal posed in [166] of developing a variational theory for Willmore. It does not fulfill the need of exploring the neighborhood of a weak immersion in order to deduce an Euler–Lagrange equation for instance and Theorem I.1 is an essential tool in that respect.
The notion of branched $W^{2,2}$ weak immersion into ${\mathbb{R}}^{m}$ is an extension of $W^{2,2}$ weak immersion allowing for isolated branched points $\{a_{1},\dots,a_{L}\}$ at which the gradient of the weak immersion is degenerating. Branched $W^{2,2}$ weak immersion have an underlying conformal structure that is there exists a constant Gauss curvature metric $h$ on $\Sigma$ such that
$g_{\vec{\Phi}}=e^{2\alpha}\,h.$
But the conformal factor $\alpha$ is not bounded in $L^{\infty}$ anymore at each branched point but has a relatively well understood asymptotic expansion: in a local conformal chart $x=(x^{1},x^{2})\in D^{2}$ for $h$ containing exactly one of the branched point, say $a_{i}$, and for which $x(a_{i})=0$, there exists $d_{i}\in{\mathbb{N}}$ such that (see Lemma A.5 of [166])
$\|\alpha(x)-d_{i}\log|x|\|_{L^{\infty}(D^{2})}<+\infty,$
The maximal number of branched points for a given closed two-dimensional manifold $\Sigma$ is bounded by the $L^{2}$ norm of the second fundamental form. Indeed, it satisfies the generalized Liouville equation
$\Delta_{h}\alpha=e^{2\alpha}\,K_{g_{\vec{\Phi}}}-K_{h}-\sum_{i=1}^{L}\ d_{i}\,\delta_{a_{i}}.$ (I.8)
Once a weak closure result is known, only part of the task of developing a variational theory for Willmore is fulfilled. Indeed, the paradox of having an Euler–Lagrange equation not compatible with the Lagrangian and the assumption remains. This paradox has been solved by the third author some years before Theorem I.2 was known. He proved in [164] that the Willmore equation can be rewritten in conservative form: In 3 dimension for instance (I.1) is equivalent to (omitting the subscript $\vec{\Phi}$)
$d\ast_{g}\left[d\vec{H}-\,2\,H\,d\vec{n}-H^{2}\,d\vec{\Phi}\right]=0.$ (I.9)
In local conformal coordinates this equation is equivalent to
$\mbox{div}\left[\nabla\vec{H}-\,2\,H\,\nabla\vec{n}-H^{2}\,\nabla\vec{\Phi}\right]=0.$ (I.10)
One can observe at this stage that for a weak immersion there holds
$\vec{H}\in L^{2},\quad\nabla\vec{n}\in L^{2},\quad\mbox{ and }\quad\nabla\vec{\Phi}\in L^{\infty}.$
Therefore, the following quantity is a well-defined distribution
$\nabla\vec{H}-\,2\,H\,\nabla\vec{n}-H^{2}\,\nabla\vec{\Phi}\in W^{-1,2}+L^{1}.$
Using Poincaré lemma together with a result by Jean Bourgain and Haïm Brezis [24] we deduce that a weak immersion is a critical point to $W$ if and only if there exists $\vec{L}\in L^{2}$ such that
$\nabla^{\perp}\vec{L}=2\,H\,\nabla\vec{n}+H^{2}\,\nabla\vec{\Phi}-\nabla\vec{H},$
where $\nabla^{\perp}\coloneqq(-\partial_{2}\,,\partial_{1})$. At this stage the equation, though meaningful for weak immersions, was presenting the difficulty of having a non linearity $(2\,H\,\nabla\vec{n}+H^{2}\,\nabla\vec{\Phi})$ a-priori only in $L^{1}$, whose Hodge decomposition (omitting to precise boundary condition) is given by
$2\,H\,\nabla\vec{n}+H^{2}\,\nabla\vec{\Phi}=\nabla\vec{H}+\nabla^{\perp}\vec{L}.$
This difficulty has been overcome by identifying further conserved quantities which have been later on interpreted as direct consequences of Noether theorem by Yann Bernard [14]. The Euler–Lagrange equation in divergence form (I.10) corresponds to the invariance of the Lagrangian by translation. The invariance by dilation following Noether theorem reads
$\mbox{div}\left[\vec{L}\cdot\nabla^{\perp}\vec{\Phi}\right]=0.$ (I.11)
The invariance by rotation is giving
$\mbox{div}\left[\vec{L}\times\nabla^{\perp}\vec{\Phi}+H\ \nabla^{\perp}\vec{\Phi}\right]=0.$ (I.12)
The next step in [164] consists in introducing the “primitives” of the two last conservation laws: there exist $S\in W^{1,2}$ and $\vec{R}\in W^{1,2}$ such that
$\nabla^{\perp}S\coloneqq\vec{L}\cdot\nabla^{\perp}\vec{\Phi}\qquad\mbox{ and }\qquad\nabla^{\perp}\vec{R}\coloneqq\vec{L}\times\nabla^{\perp}\vec{\Phi}+H\ \nabla^{\perp}\vec{\Phi}.$
Finally another result in [164] is the discovery of an equation which relates $S$ and $\vec{R}$ called “the $\vec{R}-S$ system”
$\nabla\vec{R}=\vec{n}\times\nabla^{\perp}\vec{R}+\nabla^{\perp}S\ \vec{n}.$ (I.13)
It is proved in [164] that the solutions to the $\vec{R}-S$ system combined with (I.14) are smooth by applying integrability by compensation to this first order elliptic system combined with the structural equation
$2\vec{H}=\nabla S\cdot\nabla^{\perp}\vec{\Phi}+\nabla\vec{R}\,\dot{\times}\,\nabla^{\perp}\vec{\Phi}.$ (I.14)
While the use of local conformal coordinates was centrally used in [164], Section V of the present paper is presenting a proof of the regularity without using conformal coordinates. The motivation being to “prepare the ground” for considering regularity questions for solution to the Euler–Lagrange equation to higher dimensional counterpart to the Willmore Lagrangian which is the subject of the second part of the paper. Before coming to this second part we prove the following approximation theorem (in Section IV.4.3), which is one of the main new result of the present paper.
Theorem I.3.
Let $(\Sigma,h)$ be a closed Riemann surface, where $h$ is a Riemannian metric on $\Sigma$ of constant Gaussian curvature $1$, 1-1, or 0. Let $\vec{\Phi}\in W^{2,2}(\Sigma,\mathbb{R}^{m})$ be a conformal weak immersion and $\alpha\in W^{1,2}(\Sigma)$ be such that
$g_{\vec{\Phi}}=e^{2\alpha}\mkern 1.0muh.$
Then there exist sequences of smooth immersions $(\vec{\Phi}_{k})_{k\in\mathbb{N}}\subset C^{\infty}(\Sigma,\mathbb{R}^{m})$, functions $(\alpha_{k})_{k\in\mathbb{N}}\subset C^{\infty}(\Sigma)$, and constant Gaussian curvature metrics $(h_{k})_{k\in\mathbb{N}}$ on $\Sigma$ such that
$\begin{dcases}(i)\ g_{\vec{\Phi}_{k}}=e^{2\alpha_{k}}\mkern 1.0muh_{k},\\[4.30554pt] (ii)\ \vec{\Phi}_{k}\rightarrow\vec{\Phi}&\text{in }\,W^{2,2}(\Sigma,\mathbb{R}^{m}),\\[4.30554pt] (iii)\>\alpha_{k}\rightarrow\alpha&\text{in }\,C^{0}(\Sigma),\\[4.30554pt] (iv)\ h_{k}\rightarrow h&\text{in }\,C^{\infty}(\Sigma,T^{*}\Sigma\otimes T^{*}\Sigma).\end{dcases}$
This approximability property shows for instance that in order to prove an analytical property for weak immersions, it is sufficient to prove this property for $C^{\infty}$ immersions and to show that one can pass to the limit. This allows for instance, to define the Willmore flow for arbitrary initial data in the space of $W^{2,2}$ weak immersions. We prove that the density of $C^{\infty}$ immersions into weak immersions with VMO derivatives remains valid in higher dimensions, see Theorem IV.23. This result is fundamental in the recent work [129], where the second and third authors study the closure of the set of weak immersions under a certain Sobolev-type bound on the second fundamental form whose motivation arises from the following context.
In 2018, Andrea Mondino and Huy Nguyen [146] proved that the Willmore functional is the only (up to a topological constant) conformally invariant functional among all the functionals defined on hypersurfaces of 3-dimensional manifolds and depending only on the first and second fundamental form. However, this is not the case in higher dimensions and codimensions. Easy examples can be found in dimension 4, with for instance, functionals defined on immersed submanifolds $\Sigma^{4}\subset\mathbb{R}^{5}$ such as
$\int_{\Sigma}\big|\mathring{\mathrm{I\!I}}\big|^{4}_{g}\,d\textup{vol}_{g},\quad\int_{\Sigma}\mathrm{tr}_{g}\left(\mathring{\mathrm{I\!I}}^{4}\right)\,d\textup{vol}_{g},\quad\int_{\Sigma}\det_{g}\big(\mathring{\mathrm{I\!I}}\big)\,d\textup{vol}_{g}.$
Such conformally invariant functionals on submanifolds naturally arise in the context of the AdS/CFT correspondence in physics introduced by Juan Maldacena [119]. Broadly speaking, this correspondence states the existence of a duality between some gravitational theories on an asymptotically Anti De Sitter space-time $(X^{d+2},g)$ and some conformal field theories on the conformal boundary $\partial_{\infty}X$, see for instance [194] for an introduction. In this context, many questions require a priori to compute the volume of a noncompact minimal hypersurface such as the computation of the entanglement entropy or the expectation value of Wilson lines, see for instance [73] for various applications. In order to give meaning to this volume, one proceeds to a volume renormalization introduced by Henningson–Skenderis [83] to study the Weyl anomaly. This method can be roughly summarized as follows: first compute the volume in a ball of size $R$, then compute an asymptotic expansion as $R\to+\infty$. Graham–Witten [75] proved in 1999 that one of the term in this expansion provides a conformally invariant functional on submanifolds of $\partial_{\infty}X$. For minimal hypersurfaces of $X^{d+2}=\mathbb{H}^{4}$, they obtained the Willmore energy for surfaces in $\mathbb{R}^{3}=\partial_{\infty}\mathbb{H}^{4}$. In light of [146], this result is not surprising, since Graham–Witten proved that one has to find a conformally invariant functional on 2-dimensional surfaces by this procedure, and the Willmore energy is the only possibility. This renormalization procedure has been fully generalized by Gover–Waldron [68] in 2017.
The volume renormalization procedure for minimal hypersurfaces has the advantage of producing conformally invariant functionals having minimal hypersurfaces as critical points. However, as discussed in Section II.2.1, some of these functionals are not bounded from below or from above. Returning to the original idea of Germain and Poisson [64, 160], one would like to study functionals that measure in some sense the umbilicity of a hypersurface. Such a functional should be able to provide ”best representation” of submanifolds and in particular, this requires to have a lower bound. In [70], Gover–Waldron proved that conformally invariant functionals can also be stated in terms of singular Yamabe-type problem, a problem raised by Loewner–Nirenberg [113] in 1974. In [67, 70], the authors prove that given any separating submanifold $\Sigma^{d-1}$ of some Riemannian manifold $(M^{d},g)$, if one look for the ”best” asymptotic expansion of a defining function [^3] A defining function $s\in C^{\infty}(M)$ is defining for $\Sigma$ if $s=0$ and $|ds|_{g}=1$ on $\Sigma$. The question raised in [70] is to find the highest number $\ell\in\mathbb{N}$ and a defining function $s$ such that we have an expansion of the form $|ds|_{g}=1+O(s^{\ell})$ near $\Sigma$. $s$ for $\Sigma$, then there is an obstruction, called the obstruction density. This obstruction density for d1=2d-1=2 (i.e. when $\Sigma$ is a surface) and $M^{d}=\mathbb{R}^{3}$ turns out to be exactly the left-hand side of (I.1). As proved by [72, 68], this obstruction can be understood in general as the Euler–Lagrange equation of a coefficient in some renormalized volume expansion reminiscent to [75, 74]. These alternatives will be discussed in Section II.2.2. One common point of these functionals is that if $\Sigma^{n}\subset\mathbb{R}^{d}$ has even dimension $n\geq 4$, then the Euler–Lagrange equation of any of this generalized Willmore energy has a leading-order term of the form $\Delta_{g_{\Sigma}}^{\frac{n}{2}}\vec{H}_{\Sigma}$. This term is also the leading-order term of the functional
$\vec{\Phi}\in\mathrm{Imm}(\Sigma^{n};\mathbb{R}^{d})\mapsto\int_{\Sigma^{n}}\big|\nabla^{\frac{n-2}{2}}\vec{\mathrm{I\!I}}_{\vec{\Phi}}\big|^{2}_{g_{\vec{\Phi}}}\ d\textup{vol}_{g_{\vec{\Phi}}}.$
The study of compactness and regularity questions for such functionals has been the subject of recent works that we will describe in Section II.2.
Structure of the article.
In Section II, we review some recent progress surrounding the analysis of Willmore surfaces and their generalizations in higher dimensions. In Section III, we set some notations and record some estimates that will be used later. In Section IV, we introduce the setting of weak immersions. We prove that $W^{2,2}$ weak immersions induce a controlled complex structure and obtain Theorem I.3. We also prove that weak immersions have integer densities. In Section V, we provide a new proof of the regularity of Willmore surfaces without the use of conformal coordinates. In Section VI, we consider the case of scale-invariant Lagrangians on immersed 4-dimensional submanifolds in $\mathbb{R}^{5}$ with leading order term of the form $\int|dH|^{2}$. We apply the Noether theorem and compute the associated conservation laws.
Acknowledgments.
This project is financed by Swiss National Science Foundation, project SNF 200020_219429.
II Open questions
II.1 In dimension 2
In this section, we review some recent advances in the theory of Willmore surfaces (concerning mainly closed Willmore surfaces in codimension 1, but one can ask similar questions in higher codimension) and state some open questions. The field is subject to a fast development, and thus we might not record all the recent literature.
II.1.1 Construction of Willmore surfaces
Since the Willmore energy is conformally invariant, minimal surfaces of $\mathbb{R}^{3}$, $\mathbb{S}^{3}$ and $\mathbb{H}^{3}$ and their conformal transformations are Willmore. One of the major problems is the understanding of closed Willmore surfaces, which are not conformal transformations of minimal surfaces. In 1985, Pinkall [158] proved that such surfaces exist. Despite the work of Hélein [79], where he constructed the equivalent of Weierstrass–Enneper parametrization in the theory of minimal surfaces, only few methods for constructing explicit examples of Willmore surfaces are known in codimension 1, and most of them are not conformally invariant. For instance, Babich–Bobenko [10] constructed in 1993 examples of Willmore surfaces with a line of umbilic points by gluing minimal tori of $\mathbb{H}^{3}$ along their common boundary in $\partial_{\infty}\mathbb{H}^{3}=\mathbb{R}^{2}$. In 2024, Dall’Acqua–Schätzle [51] proved that any rotationally symmetric Willmore surface with a line of umbilic points is actually a Willmore surface obtained by such construction. To do so, they prove that in this case, the umbilic line should be contained in a plane intersecting the surface orthogonally. By an application of Cauchy–Kovalevskya theorem, they prove that the mean curvature of the parts of the surface above and below this plane, seen as submanifolds of $\mathbb{H}^{3}$, completely vanishes. The $\mathbb{S}^{1}$-equivariant Willmore tori have been classified by Ferus–Pedit [60] (and actually contain the construction of [158]) by analyzing the geodesic flow on these surfaces. The understanding of Willmore graphs has also been subject of numerous works. For example, it is known that entire Willmore graphs with finite energy and entire Willmore graphs of radial functions are flat planes [35, 36, 114]. Simply connected, orientable, complete Willmore surfaces with vanishing Gaussian curvature have been fully classified in [190]. The loop group methods introduced by Hélein [79], has later been developed by Ma [116] and Xia–Shen [191] (see also [54] for a recent survey) and provide a way of constructing Willmore surfaces. However, this is a very abstract method and it would be interesting to know whether one can find more explicit constructions.
Open Question II.1.
Is it possible to glue a given closed Willmore surface into another given closed Willmore surface?
A first work in this direction has been done by Marque [122], where he exhibited the first example of bubbling for Willmore surfaces. Using a parametrization of Willmore spheres introduced by Bryant [28] (discussed in Section II.1.2 below), Marque constructed a sequence of Willmore spheres converging to the gluing of the inversions of a López surface and an Enneper surface. Li [110] has also proved that embedded Willmore surfaces verify some rigidity properties. For instance, an embedded Willmore sphere must be a round sphere. Hence, the setting of immersions (opposed to embeddings) is natural and cannot be avoided.
An alternative way for such constructions is provided by variational methods, such as minimization or min-max procedures in a given class of surfaces. As an example of application, the Willmore conjecture for genus 1 in codimension 1 has been solved by Marques–Neves [124] in 2013 using a min-max scheme for minimal surfaces combined with Urbano theorem, but remains open for higher genus and general codimension. This conjecture states that up to conformal transformations and in all codimensions, the only minimizer of the Willmore functional in genus 1 is the Clifford torus, whereas in genus larger than 1, the likely candidates are the Lawson surfaces [99, Conjecture 8.4]. Even if there is some numerical evidence [86], its complete resolution remains largely open.
Conjecture II.2 (Willmore conjecture in codimension 1 [99]).
Given an integer $g>1$, the only minimizers of the Willmore energy among all the smooth surfaces of $\mathbb{R}^{3}$ are the conformal transformations of the Lawson surfaces of genus $g$.
On the other hand, min-max procedures have recently become an active topic. One motivation for the study of the Willmore energy, is to apply Morse theory to understand the topology of the space $\mathrm{Imm}(\Sigma,\mathbb{R}^{m})$ of immersions of $\Sigma$ into $\mathbb{R}^{m}$. For instance, a basic problem would be to understand min-max procedures for $W$: if $\gamma\in\pi_{k}(\mathrm{Imm}(\Sigma,\mathbb{R}^{m}))$ is a generator of regular homotopy of immersions, we consider the quantity
$\beta_{\gamma}\coloneqq\inf\left\{\sup_{t\in\mathbb{S}^{k}}W(\vec{\Phi}_{t}):(\vec{\Phi}_{t})_{t\in\mathbb{S}^{k}}\in\gamma\right\}.$
One can ask a few natural questions: can we bound these numbers? Does there exist any immersion realizing these optimization problems? If so, how many are they? Can we classify them? A first issue is that the map $W\colon\mathrm{Imm}(\Sigma,\mathbb{R}^{m})\to\mathbb{R}$ cannot be a Morse function, partly due to the conformal invariance. To overcome this, the third author adopted a viscosity approach in [170], previously introduced for the construction of geodesics in [140]. Instead of considering $W$ alone, he adds a ”smoother” times a small parameter $\sigma>0$ and then let $\sigma\to 0$. He applied successfully this method for a well chosen smoother, in the case $\Sigma=\mathbb{S}^{2}$, and obtained that the values of min-max procedures for $W$ on $\mathbb{S}^{2}$ is the energy of a bubble tree. We refer to [169, 170, 138] for more information about the viscosity method for geodesics or minimal surfaces. A particular case of [170] is the case $m=3$, i.e. the codimension 1 case, which can be applied to the sphere eversion and constitutes a new step towards the $16\pi$-conjecture.
[…]
Let $\Omega$ be the set of path in $\mathrm{Imm}(\mathbb{S}^{2},\mathbb{R}^{3})$ joining the standard $\mathbb{S}^{2}$ to the sphere $\mathbb{S}^{2}$ with opposite orientation. Consider the following number
$\beta_{0}\coloneqq\inf_{\omega\in\Omega}\sup_{\vec{\Phi}\in\omega}W(\vec{\Phi}).$
Then $\beta_{0}$ is reached by a path $\omega\in C^{0}([0,1],\mathrm{Imm}(\mathbb{S}^{2},\mathbb{R}^{3}))$ such that $\omega_{|[0,1/2]}$ and $\omega_{|[1/2,1]}$ are Willmore flows, and $\omega(1/2)$ is the inversion of the Morin surface, a minimal surface with 4 planar ends.
Max–Banchoff [131] proved that any path $\omega\in\Omega$ must contain an immersion $\Phi\in\omega$ having a quadruple point (another proof is given by Hughes [87]). Combining this with Li–Yau inequality shows that $\beta_{0}\geq 16\pi$. As a consequence of [170], we know that the number $\beta_{0}$ is equal to the sum of the energies of finitely many branched Willmore spheres. Hence, the need for a classification of branched Willmore spheres became strongly motivated. As we will explain in the following Section II.1.2, Bryant [28] proved that any smooth Willmore sphere is a conformal transformation of a minimal surface in $\mathbb{R}^{3}$. A first generalization to the branched Willmore spheres has been obtained by Lamm–Nguyen [103] when the total number of branched points counted with multiplicites is at most 3. Michelat and the third author [143] successfully classified all ”variational branched Willmore spheres” which include in particular weak limits of immersed Willmore spheres. The second author in [126] proved that every branched Willmore sphere in ${\mathbb{R}}^{3}$ is the inversion of a minimal surface. This implies in particular that $\beta_{0}$ is a multiple of $4\pi$.
In order to pursue the understanding of such min-max schemes, one can study the Morse index of closed Willmore surfaces. The Morse index of a Willmore immersion $\vec{\Phi}$ is, by definition, the number of negative eigenvalues of the quadratic form $\delta^{2}W(\vec{\Phi})$, this is a measure of how far is $\vec{\Phi}$ from being a minimizer. The study of the Morse index for Willmore spheres has been initiated by Alexis Michelat in his PhD thesis [137] by proving that the Morse index of a Willmore immersion $\vec{\Phi}\coloneqq\frac{\vec{\Psi}}{|\vec{\Psi}|^{2}}$, for some minimal immersion $\vec{\Psi}\colon\Sigma\setminus\{p_{1},\ldots,p_{n}\}\to\mathbb{R}^{3}$ with $n$ ends, is bounded from above by $n$ in full generality, and by n1n-1 when $\vec{\Phi}$ has some branch points of order at most 1 and some branch points of order 2 without any flux. However, Michelat conjectured in [138] that whether the Willmore surface $\vec{\Phi}$ is branched or not, the Morse index $\mathrm{Ind}_{W}(\vec{\Phi})$ of $\vec{\Phi}$ should be linear in $n$. He also conjectured that if $\vec{\Phi}$ has no branch point, then $\mathrm{Ind}_{W}(\vec{\Phi})\leq n-3$. To do so, he proved that the computation of the index of $\vec{\Phi}$ can be reduced to the computation of the index of a matrix of size $n\times n$ whose entries are defined by fluxes of $\vec{\Psi}$ at each of its ends, together with some coefficients in the asymptotic expansion of some Jacobi fields. This result can be summarized as follows. We denote
$V\coloneqq\left\{u\in C^{\infty}_{loc}\left(\Sigma\setminus\{p_{1},\ldots,p_{n}\}\right):\frac{u}{|\vec{\Phi}|^{2}}\in W^{2,2}(\Sigma)\right\}.$
Since $W^{2,2}(\Sigma)\hookrightarrow C^{0}(\Sigma)$, we can define the following quantity for any $u\in V$,
$\mathrm{Eval}(u,p_{i})\coloneqq\left(\frac{u}{|\vec{\Phi}|^{2}}\right)(p_{i})\in\mathbb{R}.$
We denote $\mathcal{L}\coloneqq\Delta_{g_{\vec{\Psi}}}-2K_{g_{\vec{\Psi}}}$ the Jacobi operator of the minimal surface $\vec{\Psi}$. If $u\in V$ verifies $\mathcal{L}^{2}u=0$, then $u$ verifies the following expansion in a conformal charts near $p_{i}$ where $|\vec{\Psi}|^{-2}=|z|^{2m}(1+O(|z|))$, for some numbers $\mathrm{Log}(u,p_{i})\in\mathbb{R}$:
$u(z)=\frac{\mathrm{Eval}(u,p_{i})}{|z|^{2m}}+\sum_{1\geq k+l\geq 1-2m}\Re\left(\alpha_{k,l}\,z^{k}\,\bar{z}^{l}\right)+\mathrm{Log}(u,p_{i})\,\log|z|+O(1).$
The main result of [138] can be stated as (see for instance [138, Remark 6.5])
$\mathrm{Ind}_{W}(\vec{\Phi})=\dim\left\{u\in V:\mathcal{L}^{2}u=0,\ \sum_{i=1}^{n}\mathrm{Eval}(u,p_{i})\,\mathrm{Log}(u,p_{i})<0\right\}.$
In the case where $\Sigma=\mathbb{S}^{2}$ and the ends of $\vec{\Psi}$ are planar (so that $\vec{\Phi}$ as no branch point), Hirsch, Kusner and Mäder-Baumdicker [85, 84], revisiting Michelat’s computations, proved that the Willmore index of $\vec{\Phi}$ is exactly n3n-3, and thus verified Michelat’s conjecture in the case of unbranched Willmore spheres. To do so, they restricted the study of the admissible variations to the following set
[…]
They show (still when $\vec{\Phi}$ is not branched) in [85] that
$\mathrm{Ind}_{W}(\vec{\Phi})=\dim\left\{u\in V^{\prime}:\exists i\in\{1,\ldots,n\},\ \gamma_{i}\neq 0\right\}\in\{n-3,n-2\}.$
In [84], they obtain $\mathrm{Ind}_{W}(\vec{\Phi})=n-3$ by combining a perturbative argument together with the lower semi-continuity of the Morse index for inversions of minimal surfaces.
Concerning branched Willmore spheres, the situation is more complex and the bound n3n-3 is no longer valid. In fact, one can show that the Morse index of the inversion of the López surface (one planar end and one end of multiplicity 3) is exactly 1. It would be interesting to know whether the number n1n-1 is sharp or not for the Morse index of branched Willmore spheres.
Open Question II.4.
Consider $\vec{\Phi}:\mathbb{S}^{2}\setminus\{p_{1},\ldots,p_{n}\}\to\mathbb{R}^{3}$ a complete minimal surface with $n$ arbitrary ends such that $0\notin\vec{\Phi}(\mathbb{S}^{2}\setminus\{p_{1},\ldots,p_{n}\})$ and $\frac{\vec{\Phi}}{|\vec{\Phi}|^{2}}$ is a branched Willmore surface. Is it true that the Willmore Morse index of $\frac{\vec{\Phi}}{|\vec{\Phi}|^{2}}$ is linear in $n$? What if we replace $\mathbb{S}^{2}$ by any closed Riemann surface of higher genus?
In the context of min-max procedures, Michelat [136] proved that the Morse index of Willmore spheres obtained by realizing some width of Willmore sweepout with the viscosity method is bounded by the number of parameters of this sweepout. The ”smoother” used in for this result is the Onofri energy whose second variation is given by a non-local operator. Michelat–Rivière [144] then proved the upper semicontinuity of the nullity plus the Morse index. Michelat extended this result recently to the case of branched Willmore spheres in [139] by developing a sharp analysis of weighted Rellich-type inequalities for fourth-order elliptic operators on degenerating annuli. In order to develop min-max procedures when the underlying surface is not a sphere, it would be interesting to have a generic way of estimating the index of a given Willmore surface which is not necessarily a conformal transformation of a minimal surface in $\mathbb{R}^{3}$. Thanks to the works of Chodosh–Maximo [39, 40], it is known that the Morse index (for the area functional) is bounded from above and below by the topology of the surface. Since this also holds for smooth Willmore spheres (by the results of Michelat and Hirsch–Kusner–Mäder-Baumdicker), one can wonder whether such a property also holds or not for Willmore surfaces.
Open Question II.5.
Can we estimate the Morse index of a given smooth Willmore surface of genus $g\geq 1$ by its topology and its Willmore energy?
Open Question II.6.
Can we relate the Willmore index of conformal transformations of minimal surfaces in $\mathbb{R}^{3}$, $\mathbb{H}^{3}$ or $\mathbb{S}^{3}$ with their area index?
Now that the Willmore spheres have been classified, one can start working on a possible use of the Willmore functional as a quasi-Morse functional to describe the set of immersions of $\mathbb{S}^{2}$ into $\mathbb{R}^{3}$, in the same manner as geodesics and minimal surfaces are used to describe the fundamental groups of manifolds, see for instance [97, 149].
Open Question II.7.
Can we describe the topology of the set $\mathrm{Imm}(\mathbb{S}^{2},\mathbb{R}^{3})$ using the Willmore functional as a pseudo-Morse function?
II.1.2 Conformal Gauss map approach
In 1984, Bryant [28] (see also [29]) proved that to each smooth immersion $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{3}$, one can associate a map $Y\colon\Sigma\to\mathbb{S}^{3,1}$, where $\mathbb{S}^{3,1}$ is the De Sitter space
$\mathbb{S}^{3,1}\coloneq\left\{p\in\mathbb{R}^{5}:|p|^{2}_{\eta}=1\right\},\qquad\eta\coloneq(dx^{1})^{2}+\cdots+(dx^{4})^{2}-(dx^{5})^{2}.$
Moreover, the map $Y\colon(\Sigma,g_{\vec{\Phi}})\to(\mathbb{S}^{3,1},\eta)$ is harmonic if and only if $\vec{\Phi}$ is Willmore, and is a space-like immersion away from the umbilic points of $\vec{\Phi}$. We refer to [120] for an introduction and to [17] for $\varepsilon$-regularity properties in this context. Thanks to this duality, Bryant exhibited the quartic $\mathcal{Q}\coloneq\eta(\partial^{2}_{zz}Y,\partial^{2}_{zz}Y)\,(dz)^{4}$ which is holomorphic when $\vec{\Phi}$ is Willmore. This quartic can be seen as a measure of the distance of $\vec{\Phi}$ to the set of conformal transformations of minimal surfaces in $\mathbb{R}^{3}$. Indeed, if $\omega\coloneq\mathring{\mathrm{I\!I}}(\partial_{z},\partial_{z})(dz)^{2}$ is the complex tracefree curvature of $\vec{\Phi}$, $K_{Y}$ is the Gaussian curvature of the metric $g_{Y}\coloneq Y^{*}\eta$ on $\Sigma$ and $K_{Y}^{\perp}$ is the curvature of the normal bundle of $Y$ in $\mathbb{S}^{3,1}$ (this is a 2-dimensional vector bundle), then Palmer [155] proved in 1994 that
$\frac{\mathcal{Q}}{\omega^{2}}=(1-K_{Y}+i\,K_{Y}^{\perp})\,\frac{|\mathring{\mathrm{I\!I}}|^{2}_{g_{\vec{\Phi}}}}{2}.$
The above formula is well-defined outside of the umbilic set of $\vec{\Phi}$, that is to say when $\omega\neq 0$. By the work of Bryant [28], we have that if $\mathcal{Q}=0$ (i.e. $K_{Y}^{\perp}=0$ and $K_{Y}=1$), then $\vec{\Phi}$ is a conformal transformation of a minimal surface in $\mathbb{R}^{3}$. If we relax this information, then Marque [121] proved that $K_{Y}^{\perp}=0$ if and only if $\vec{\Phi}$ is the conformal transformation of a minimal surface in $\mathbb{R}^{3}$, $\mathbb{S}^{3}$ or $\mathbb{H}^{3}$. This approach seems very promising, but first requires a good understanding of the umbilic set and only a few advances have been made in understanding the conformal Gauss map of Willmore surfaces in codimension 1.
One of the major achievement in [28] is the classification of smooth Willmore spheres. Indeed, since there is no holomorphic quartic on $\mathbb{S}^{2}$, every smooth Willmore sphere must satisfy $\mathcal{Q}=0$ and thus, be a conformal transformation of a minimal surface in $\mathbb{R}^{3}$. If $\vec{\Phi}\colon\mathbb{S}^{2}\to\mathbb{R}^{3}$ is a branched Willmore immersion (for instance obtained in a bubble tree), then the equation $\bar{\partial}\mathcal{Q}=0$ is a-priori not verified at the branch points of $\vec{\Phi}$, hence $\mathcal{Q}$ is a-priori merely meromorphic and it is not clear that Bryant’s classification extends to branched Willmore immersions. Lamm–Nguyen [103] computed an asymptotic expansion of $\mathcal{Q}$ across branch points and proved that $\mathcal{Q}=0$ if the number of branch points of $\vec{\Phi}$ is at most 3, when counted with multiplicities. This idea can also be seen as an application of Liouville theorem for holomorphic functions, as explained in [120, Theorem 2.6.3]. Then Michelat–Rivière [143] proved that a branched Willmore sphere obtained as a weak limit of smooth Willmore surfaces or obtained in a bubbling process of a sequence of smooth Willmore surfaces (called a ”variational branched Willmore sphere”) also verify $\mathcal{Q}=0$. To do so, they compute a refined asymptotic expansion of $\mathcal{Q}$ near a branch point and prove that $\mathcal{Q}$ is actually bounded near branch points where one of the residue of $\vec{\Phi}$ verifies a smallness assumption. They proved in an earlier work [141, 142] that this condition is full-filled for variational branched Willmore spheres, see also [122] for a similar property. The full classification has recently been obtained in [126], by adapting the analysis of the conformal Gauss map developed in [17] in order to obtain $\varepsilon$-regularity on the conformal Gauss map and [127] in order to prove an energy quantization result for Willmore surfaces with bounded Morse index.
The next step, which would be an interesting starting point for the Willmore conjecture, is to understand the geometry of smooth Willmore surfaces $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{3}$ when $\Sigma$ has positive genus. On the one hand, Palmer [155] proved that the Bryant’s quartic could be interpreted as the curvature of the conformal Gauss map. Moreover, he proved that if $\mathcal{Q}$ verifies the pointwise condition $|\mathcal{Q}|_{g_{Y}}>1$, then $\Sigma$ is a torus and $\vec{\Phi}$ is unstable. On the other hand, Fisher-Colbrie [61] proved that stable minimal surfaces in positively curved manifold have finite total Gauss curvature. Now that some regularity properties have been understood in [17, 123, 127], it would be intersting to study stability questions from the view-point of the conformal Gauss map.
Open Question II.8.
Can we prove that the Gauss curvature of the conformal Gauss map of a stable Willmore surface in $\mathbb{R}^{3}$ is integrable?
In higher codimension, the conformal Gauss map has already been studied [148, 56, 117, 115] and there are Willmore spheres which are not conformal transformation of minimal surfaces. For instance, Montiel [148] classified all the Willmore spheres in $\mathbb{R}^{4}$ and proved that in addition to conformal transformations of minimal surfaces, Willmore spheres can also be obtained as a rational curve in $\mathbb{CP}^{3}$ via the Penrose twistor fibration $\mathbb{CP}^{3}\to\mathbb{S}^{4}$. This classification has also been extended in [143] to branched Willmore spheres satisfying some smallness assumption on a residue.
Open Question II.9.
Is it true that any branched Willmore spheres in $\mathbb{R}^{4}$ belongs to one of these two classes: either a conformal transformations of a branched minimal surface of $\mathbb{R}^{4}$, or via the Penrose twistor fibration $\mathbb{CP}^{3}\to\mathbb{S}^{4}$ a branched rational curve in $\mathbb{CP}^{3}$?
One of the difficulty to such an approach comes from the size of the umbilic set of Willmore surfaces. Contrarily to minimal surfaces, Babich–Bobenko [10] proved that Willmore surfaces can have curves of umbilic points. At these points, the geometry of the conformal Gauss map degenerate. Examples of Willmore surfaces having such curves have been constructed by Babich–Bobenko [10] and by Dall’Acqua–Schätzle [51]. However, the structure of such sets is still not fully understood. Schätzle [177] proved that the umbilic set of a smooth Willmore immersion is composed of isolated points and one dimensional real-analytic manifold without boundary. In the recent work [123], Marque and the second author proved that the structure of the umbilic set provides strong information on the geometry of Willmore immersions and on the value of its energy. For instance if a Willmore surface has a geodesic umbilic line, then it is of Babich–Bobenko type.
Open Question II.10.
What properties does the umbilic lines of closed Willmore surfaces satisfy? Do they satisfy any particular equation? Can we describe their location on the surface? Can we estimate their size in terms of the Willmore energy or any other geometric quantities?
One can also wonder about the geometric interpretation of the Bryant’s quartic. A natural question would be to understand the set of Willmore surfaces having the same given Bryant’s quartic, or even the set of quartics on a given surface that can be achieved as the Bryant’s quartic of some Willmore immersion.
Open Question II.11.
Given two Willmore surfaces $\vec{\Phi}_{1},\vec{\Phi}_{2}\colon\Sigma\to\mathbb{R}^{3}$ such that their Bryant’s quartic satisfies $\mathcal{Q}_{1}=\mathcal{Q}_{2}\neq 0$. Is it possible to relate $\vec{\Phi}_{1}$ to a conformal transformation of $\vec{\Phi}_{2}$? Since the Bryant’s quartic roughly describes the second fundamental form of their conformal Gauss map, the map $\vec{\Phi}_{1}$ should be a conformal transformation of $\vec{\Phi}_{2}$ in the case where the induced metrics of their conformal Gauss map coincide. But it seems wrong if this is not the case. It would be interesting to study the possible differences.
Open Question II.12.
Given a complex structure on a given closed surface $\Sigma$, can we describe the set of holomorphic quartics which are the Bryant’s quartic of some Willmore immersion?
The Poisson problem has been introduced in [47] (analog of the Plateau problem for Willmore surfaces) where the authors solve the Plateau problem for datas given for the immersion and the Gauss map, see also [161]. Due to the conformal invariance of the Willmore energy, it would be natural to solve the Plateau problem for conformally invariant data on the boundary.
Open Question II.13.
Solve the Poisson problem for boundary data given by a boundary curve $\Gamma\subset\mathbb{R}^{3}$ and the image of its conformal Gauss map described as a closed curve in $\mathbb{S}^{3,1}$.
Open Question II.14.
What is the optimal regularity assumption on the given boundary datas in order to solve the Poisson problem? How many solutions are there?
Up to now, the conformal Gauss map approach has been mainly developed for surfaces of a Euclidean space. It would be interesting to study its generalization in manifolds. A possibility is provided by the normal tractor, see for instance [45].
II.1.3 Willmore flow
Geometric flows are designed to produce explicit paths leading to critical points of a given functional. They are useful both from a theoretic and from a practical view-point. For instance, Conjecture II.3 is an example of application of the Willmore flow, which roughly says that there is no smooth Willmore sphere of energy strictly between the energy of $4\pi$ and $16\pi$. One could also think about the Willmore flow as a first step before studying the Canham–Helfrich flow modelling the evolution of membranes of the cells, see for instance [25, 173] and the references therein.
The main analytical difficulty of the Willmore flow (despite the fact that it is a geometric flow, and thus has to be degenerate) is the fact that it is of order 4, meaning that there exists no comparison principle in the spirit of the mean curvature flow. For instance, if an initial data is included in the interior of a given sphere $S$, one can a-priori not say that the flow remains inside the interior of $S$. Geometric flows of order 4 have original been explored in the context of biharmonic surfaces [57]. The Willmore flow has first been introduced in 2001 independentely by Kuwert–Schätzle [101] and Simonett [181]. They proved the existence of a solution in short time, and proved that if the initial data has small energy, then the flow smoothly converges to the round sphere. In 2003, Mayer–Simonett [132] proved that the Willmore flow does not preserve embeddedness of the initial data. Blatt [22] constructed in 2009 an example of initial data of spherical type for which a singularity occurs in the Willmore flow. More recently, Dall’Acqua–Müller–Schätzle–Spener [50] produced examples of flows made of tori of revolutions, and also produced degenerating examples. This flow is still widely unexplored. The parametric approach has been introduced by Palmurella–Rivière [156, 157], where the authors show that there are only two possible kind of singularities in finite time: either a concentration of energy with a possible branch point, or a degeneration of the conformal class.
Open Question II.15.
Is it possible for the conformal class to degenerate in finite time along the Willmore flow?
In order to study the possible degeneracies, one can rely on the existing literature on the bubbling analysis of sequences $\vec{\Phi}_{k}\colon\Sigma\to\mathbb{R}^{3}$ of Willmore surfaces. When the underlying complex structure of $(\Sigma,g_{\vec{\Phi}_{k}})$ remains bounded in the moduli space, Bernard–Rivière [19] proved in 2014 that some bubbles can appear, but the energy is always preserved. Then Laurain–Rivière [106] proved in 2018 that when the underlying complex structure degenerates, then only one residue is responsible for a possible loss of energy. This has recently been extended in [112], where Li–Yin–Zhou prove that this residue precisely describes the loss of energy. In 2021, Marque [122] proved that the set of Willmore immersions with energy less than or equal to $12\pi$ is compact. In 2025, the second author [127] proved that the energy is preserved if $\vec{\Phi}_{k}$ has bounded Morse index, independently of the underlying complex structure. Recently, Michelat [139] proved the lower semi-continuity of the Morse index under some smallness assumption of the residue (which corresponds to the assumption in [106]). Since the analysis of Willmore immersions is now understood, it should be possible to study the case of the flow.
Open Question II.16.
Can we prove an energy quantization property for the Willmore flow?
Once a singularity has occurred, one would like to continue the flow. However, the initial data is not smooth and the existence of a solution to the Willmore flow is not clear. In particular, the setting of smooth surfaces cannot be used and one has to find the proper notion of solution.
Open Question II.17.
Is it possible to define a solution of the Willmore flow starting from a branched surface? From a nodal surface?
As an application, one can think about the $16\pi$-conjecture. In a recent work [118], Mäder-Baumdicker–Seidel studied the possible singularities that occur for an initial data $\vec{\Phi}_{0}\colon\mathbb{S}^{2}\to\mathbb{R}^{3}$ with Willmore energy strictly less than $16\pi$. They proved that if $W(\vec{\Phi}_{0})\leq 12\pi$, then either the flow starting from $\vec{\Phi}_{0}$ converges to a round sphere or the flow converges to two round spheres glued by a catenoid neck. If $W(\vec{\Phi}_{0})<16\pi$, then other singularities might arise.
Open Question II.18.
Given a immersion $\vec{\Phi}_{0}\colon\mathbb{S}^{2}\to\mathbb{R}^{m}$ satisfying $W(\vec{\Phi}_{0})<16\pi$, under which conditions the Willmore flow $(\vec{\Phi}_{t})_{t>0}$ starting from $\vec{\Phi}_{0}$ does not develop any singularity as long as $12\pi<W(\vec{\Phi}_{t})<16\pi$?
II.1.4 Critical points of the Willmore energy under various constraints
Instead of looking at critical points of the Willmore functional among the set of all the immersions, one can restrict the set of admissible variations (for example, one need to fix some geometric constraints for the study of the Canham–Helfrich functional in cell biology). Three main possibilities are curently studied.
One possibility, is to study the minimizers (and more generally the critical points) of the Willmore functional under the constraints of fixing both the area of the immersed surface and the volume enclosed it encloses. Since the Willmore functional is scale-invariant, this question is equivalent to prescribing the isoperimetric ratio of the immersion $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{3}$ given by
$\mathcal{V}(\vec{\Phi})\coloneqq-\frac{1}{3}\int_{\Sigma}\vec{\Phi}\cdot\vec{n}_{\vec{\Phi}}\,d\textup{vol}_{g_{\vec{\Phi}}},$
$\mathcal{I}(\vec{\Phi})\coloneqq\frac{\textup{vol}_{g_{\vec{\Phi}}}(\Sigma)}{\mathcal{V}(\vec{\Phi})^{\frac{2}{3}}}.$
By the isoperimetric inequality, we have
$\mathcal{I}(\vec{\Phi})\geq\mathcal{I}(\mathbb{S}^{2})=(36\pi)^{\frac{1}{3}}.$
We can therefore study the following minimization problem for any $R\geq(36\pi)^{\frac{1}{3}}$:
$\beta^{\mathcal{I}}(R)\coloneqq\inf\left\{W(\vec{\Phi}):\vec{\Phi}\in\mathrm{Imm}(\Sigma;\mathbb{R}^{3}),\ \mathcal{I}(\vec{\Phi})=R\right\}.$
Keller–Mondino–Rivière [96] proved in 2014 that a minimizer of $\beta^{\mathcal{I}}(R)$ exists provided that this number verifies some strict inequality. This inequality has been proved for all $R\in\left((36\pi)^{\frac{1}{3}},+\infty\right)$ when $\Sigma=\mathbb{S}^{2}$ by Schygulla [179] in 2012, when $\Sigma=\mathbb{T}^{2}$ by Scharrer [174] in 2022 and by Ndiaye–Schätzle [153] in all the remaining cases in 2015.
Open Question II.19.
Describe the solutions of $\beta^{\mathcal{I}}(R)$ for $R\in\left((36\pi)^{\frac{1}{3}},+\infty\right)$. Are they unique? Do they have some symmetries?
Another possibility, is to study the minimization of the Willmore energy under a constraint on the total mean curvature. Given an immersion $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{3}$, we denote
$\mathcal{T}(\vec{\Phi})\coloneqq\frac{1}{vol_{g_{\vec{\Phi}}}(\Sigma)^{\frac{1}{2}}}\int_{\Sigma}H_{\vec{\Phi}}\,d\textup{vol}_{g_{\vec{\Phi}}}.$
By Hölder inequality, it holds
$\mathcal{T}(\vec{\Phi})^{2}\leq W(\vec{\Phi}).$
We then study the following problem for any $R\in\mathbb{R}$
$\beta^{\mathcal{T}}(R)\coloneqq\inf\left\{W(\vec{\Phi}):\vec{\Phi}\in\mathrm{Imm}(\Sigma;\mathbb{R}^{3}),\ \mathcal{T}(\vec{\Phi})=R\right\}.$
Scharrer–West [176] proved the existence of a minimizer for $R\in\left(0,\sqrt{2}\,\mathcal{T}(\mathbb{S}^{2})\right)\setminus\{\mathcal{T}(\mathbb{S}^{2})\}$ and studied the properties of the map $(R\in\mathbb{R}\mapsto\beta^{\mathcal{T}}(R))$. For the case $R=\mathcal{T}(\mathbb{S}^{2})$, they argue that the arguments developed in [96] do not apply.
Open Question II.20.
Is $\beta^{\mathcal{T}}(\mathcal{T}(\mathbb{S}^{2}))$ reached? Is there a minimizer for $\beta^{\mathcal{T}}(R)$ with $R\notin\left(0,\sqrt{2}\,\mathcal{T}(\mathbb{S}^{2})\right)\setminus\{\mathcal{T}(\mathbb{S}^{2})\}$?
Open Question II.21.
Can we study the evolution of minimizers of $\beta^{\mathcal{T}}(R)$ when $R$ varies?
The study of $\beta^{\mathcal{T}}$ is related to the study of the following functional, for a given parameter $\bar{H}\in\mathbb{R}$ called the spontaneous curvature
$\forall\vec{\Phi}\in\mathrm{Imm}(\Sigma;\mathbb{R}^{3}),\qquad W_{\bar{H}}(\vec{\Phi})\coloneqq\int_{\Sigma}(H_{\vec{\Phi}}-\bar{H})^{2}\,d\textup{vol}_{g_{\vec{\Phi}}}.$
The Helfrich flow has been introduced in 2006 [98], where Kohsaka–Nagasawa prove the short term existence of solution to the gradient flow of $W_{\bar{H}}$ with the enclosed volume and the area prescribed by some given constants $V_{0}$ and $A_{0}$. The cases when the initial data has small energy or when it is radially symmetric have been studied by Rupp–Scharrer–Schlierf [173]. This work extends earlier results of Rupp [171, 172] concerning the Willmore flow (i.e. the case $\bar{H}=0$) under a prescription of the isoperimetric ratio instead of a prescription of both the enclosed volume and the area. In order to study the flow for more general initial data, where singularities might appear, Scharrer–West [175] developed a bubbling analysis for constrained Willmore surfaces in the spirit of [19, 106].
Open Question II.22.
Can we prove an energy quantization property for the Helfrich flow?
Another important Willmore constrained variational problem consists in studying the variation of Willmore energy under constraints on the Riemann Surface the immersion is inducing.
Let $\vec{\Phi}$ be a weak immersion of a surface $\Sigma$, as mentioned above it defines a smooth conformal class — i.e. a Riemann Surface — which is the one induced by the pull back metric metric $g_{\vec{\Phi}}=\vec{\Phi}^{\,\ast}g_{{\mathbb{R}}^{3}}$. The conformal class defined by this immersion is said to be a degenerate point of the map $\mathcal{C}$ assigning the conformal class of $(\Sigma,g_{\vec{\Phi}})$ in the Teichmüller space if and only if there exists a non zero holomorphic quadratic form [^4] An holomorphic quadratic form is an holomorphic section of the bundle given by $\wedge^{1,0}T^{\ast}\Sigma\otimes\wedge^{1,0}T^{\ast}\Sigma$ where $\Sigma$ is equipped with the conformal structure induced by $g_{\vec{\Phi}}=\vec{\Phi}^{\,\ast}g_{{\mathbb{R}}^{m}}$ $q$ such that
$q,\vec{h}_{0}>_{WP}=0.$ (II.1)
In the above equation, we denoted $\vec{h}_{0}$ the Weingarten operator given by
$\vec{h}_{0}:=-\,\partial\pi_{\vec{n}}\,{\otimes}\,\partial\vec{\Phi},$
where $\pi_{\vec{n}}$ is the orthogonal projection linear map from ${\mathbb{R}}^{m}$ onto the normal space to the immersion $N_{x}\Sigma\coloneq(\vec{\Phi}_{\ast}T_{x}\Sigma)^{\perp}$ and $\cdot,\cdot>_{WP}$ is denoting the Weil–Peterson scalar product given in local conformal coordinates for which
$g_{\vec{\Phi}}=e^{2\lambda}\ [dx_{1}^{2}+dx_{2}^{2}],$
by (see [167])
$\left<\phi\ dz\otimes dz,\psi\ dz\otimes dz\right>:=e^{-4\lambda}\ \Re\left[\overline{\phi}\,{\psi}\right]\ .$
Immersions in ${\mathbb{R}}^{3}$ which are degenerate points of the map $\mathcal{C}$} have been first studied in the late XIXth century under the name of isothermic surfaces (see [165]). There are the surfaces, such that, away from umbilic points there exists locally conformal coordinates which give the principal directions.
Assuming now that $(\Sigma,g_{\vec{\Phi}})$ is a non degenerate point under some constraint on the conformal class of the induced metric then there exists an holomorphic quadratic form $q$ such that
$d\ast_{g}\left[d\vec{H}-\,3\,\pi_{\vec{n}}(d\vec{H})+\star(\ast_{g}\ d\vec{n}\wedge\vec{H})\right]=<q,\vec{h}_{0}>_{WP}\ .$ (II.2)
where $\star$ is the standard Hodge operator on multivectors in ${\mathbb{R}}^{m}$ (see [167]). This equation is known as the conformal Willmore equation. In codimension one (i.e. $m=3$) this can be rewritten as follows
$d\ast_{g}\left[d\vec{H}-\,2\,H\,d\vec{n}-H^{2}\,d\vec{\Phi}\right]=<q,\vec{h}_{0}>_{WP}\ .$ (II.3)
The holomorphic quadratic form is a Lagrange multiplier [^5] The holomorphic quadratic form identifies to the tangent space to the moduli space (theorem 4.2.2 [93]) associated to the conformal class constraint. Weak immersions which are solving (II.2) are proved to be smooth ([18]).
Remark II.23.
In [18] it is proved that the Conformal Willmore equation (II.3) is equivalent to the existence of $\vec{L}$ satisfying the set of the two conservation laws (I.11) and (I.12)
$\left\{\begin{array}[]{l}\displaystyle d\left(\vec{L}\cdot d\vec{\Phi}\right)=0\\[14.22636pt] \displaystyle d\left(\vec{L}\times d\vec{\Phi}+H\,d\vec{\Phi}\right)=0\end{array}\right$. (II.4)
coming respectively from Noether theorem for dilations and rotations which is itself equivalent to the so-called $(\vec{R},S)$ system (I.13) combined with the structure equation (I.14). For the higher dimensional counterparts to Willmore functional and in particular for the Dirichlet energy of the mean curvature of a weak immersion of 4-dimensional manifolds into ${\mathbb{R}}^{5}$, we are asking in Open Question VI.4 whether the conservation laws issued from Noether for dilations and rotations do correspond to the variation of this energy under some constraint.
An interesting question is to know whether or not a weak immersion which is both a critical point of the Willmore functional under some constraint on the conformal class and isothermic (i.e. a degenerate point for the constraint) is solving or not equation (II.2) and is then smooth. The answer to this question is proved to be positive in [167] in full generality assuming that the surface has genus less than 3 or for arbitrary larger genus if the underlying conformal class is avoiding the “tiny” subset of the Teichmüller Space of $\Sigma$ made of hyper-elliptic points (a similar result has been obtained in [102] using different methods but with restrictions both on the energy and the co-dimension). One application of this result is the existence of smooth minimizers of the Willmore energy inside a prescribed conformal classes (Theorem 1.5 [167]). Once the existence of a smooth minimizer of the Willmore Energy in a fixed conformal class is known some open questions arise naturally.
Open Question II.24.
Identify minimizers of the Willmore energy in any given conformal class. In particular identify the minimizers of the Willmore energy in the class of any flat torus.
In a first approach one would aim at giving a lower bound of the Willmore energy in a fixed conformal class. This approach has been initiated in the pioneer paper by P.Li and S.T. Yau [109]. The paper of Li and Yau has been then followed by several important contributions such as the one of S.Montiel and A.Ros [147] or more recently by R.Bryant [30] (see also recent results pointing in the direction of the open question II.24 in [153] and [81]).
II.2 In higher dimensions
Even though, geometric properties of generalized Willmore functionals have already been the subject to lots of works, the study of the analytical properties is relatively new. We list below the few known results and propose different directions of research.
II.2.1 Graham–Reichert functional
One advantage of the renormalized volume approach is to provide conformally invariant functionals having minimal surfaces as critical points, which seems to be an important property for physical applications [7]. In dimension 4, this procedure has been applied independently by Zhang [192] and Graham–Reichert [74]. They both end up with the following energy, already found by different methods by Guven [76] in 2005: for an immersion $\vec{\Phi}\colon\Sigma^{4}\to\mathbb{R}^{5}$ the functional is given by
$\mathcal{E}_{GR}(\vec{\Phi})$ $\coloneqq\int_{\Sigma}|dH_{\vec{\Phi}}|^{2}_{g_{\vec{\Phi}}}-H_{\vec{\Phi}}^{2}\,|\mathrm{I\!I}_{\vec{\Phi}}|^{2}_{g_{\vec{\Phi}}}+7H^{4}_{\vec{\Phi}}\,d\textup{vol}_{g_{\vec{\Phi}}}.$
where $g_{\vec{\Phi}}\coloneqq\vec{\Phi}^{*}\xi$, $\xi$ is the flat metric on $\mathbb{R}^{5}$, $\mathrm{I\!I}_{\vec{\Phi}}$ is the second fundamental form of $\vec{\Phi}(\Sigma)$ and $H_{\vec{\Phi}}=\frac{1}{4}\mathrm{tr}_{g_{\vec{\Phi}}}\mathrm{I\!I}_{\vec{\Phi}}$ is its mean curvature. It turns out that the minus sign in the definition is sufficient to make the above functional unbounded from below, as proved in [74, 125]. The example constructed in [125] is based on the observation that the integrand of $\mathcal{E}_{GR}$ is a negative constant for $\mathbb{S}^{2}\times\mathbb{R}^{2}$, however this is not a closed submanifold. As mentioned by Graham in a private communication, a simple way to obtain a degenerating example of immersions from a closed manifold is to consider an ellipsoid with two semi-axis of length 1 and two others equal to some parameter $a\in(0,+\infty)$ that we can either send to 0 or to $\infty$ (since $\mathcal{E}_{GR}$ is scale invariant, both possibilities are equivalent and we recover $\mathbb{S}^{2}\times\mathbb{R}^{2}$ by sending $a\to+\infty$). At first glance, it seems that non-compact CMC immersions can provide an interesting source of examples for possible degeneracy of $\mathcal{E}_{GR}$ towards -$\infty$.
Open Question II.25.
Fix a closed submanifold $\Sigma^{4}\subset\mathbb{R}^{5}$. Is it possible to classify all the behaviour of immersions $\vec{\Phi}_{k}\colon\Sigma^{4}\to\mathbb{R}^{5}$ such that $\mathcal{E}_{GR}(\vec{\Phi}_{k})\xrightarrow[k\to+\infty$$\infty$?
An interesting starting point would be to find bounds in terms of the geometry of the immersion. For example, one can ask the following question.
Open Question II.26.
Let $\vec{\Phi}\colon\Sigma^{4}\to\mathbb{R}^{5}$ be a smooth immersion and $g$ be its first fundamental form. Given bounds on intrinsic conformally invariant quantities of $g$, such as the total $Q$-curvature, the $L^{2}$-norm of its Weyl tensor, or the $L^{1}$-norm of its Bach tensor, can we deduce bounds on $\mathcal{E}_{GR}(\vec{\Phi})$?
Another important question is the construction of critical points. We know that conformal transformations of minimal hypersurfaces are critical points of $\mathcal{E}_{GR}$, but to the best of the authors’ knowledge, no other critical point is known.
Open Question II.27.
Does there exist critical points of $\mathcal{E}_{GR}$ which are not conformal transformations of minimal hypersurfaces?
We now list a few natural questions regarding the two-dimensional case. First, it would be important to know if a generalization of Bryant’s quartic exists or not.
Open Question II.28.
Assume that $\vec{\Phi}\colon\Sigma^{4}\to\mathbb{R}^{5}$ is a smooth critical point of $\mathcal{E}_{GR}$ such that $(\Sigma,g_{\vec{\Phi}})$ carries a complex structure. Can we exhibit a holomorphic differential form?
Another natural question is to know how many critical points $\mathcal{E}_{GR}$ has, or if there is some topological obstruction to the existence of critical points.
Open Question II.29.
Does $\mathcal{E}_{GR}$ have critical points for every topological hypersurface $\Sigma^{4}\subset\mathbb{R}^{5}$? What about the critical points which are not conformal transformations of minimal surfaces?
In order to develop the analysis of $\mathcal{E}_{GR}$, it is important to first understand the regularity of its critical points. First computations in this direction have been performed by Bernard [15]. In these recent works, Bernard rewrote the Euler–Lagrange equation of any conformally invariant functional of the form $\int_{\Sigma^{4}}|dH_{\vec{\Phi}}|^{2}_{g_{\vec{\Phi}}}+l.o.t.$, using the Noether theorem to obtain a system in divergence form. We describe these computations in a simpler setting in Section VI. The full regularity is the subject of the recent work [16]. It would also be interesting to have explicit formulas for higher-dimensional renormalized volumes.
Open Question II.30.
In [74], the authors provide an abstract formula for the generalized Willmore functional in any dimension. Compute it explicitly for hypersurfaces of $\mathbb{R}^{7}$.
II.2.2 Alternatives
In [7], the authors argue that the functional $\mathcal{E}_{GR}$ is relevant for physical purposes. However, it would be interesting to look for other possible functionals that could be of interest for geometric purposes. In [23, 69], the authors define a generalized Willmore energy for hypersurfaces $\Sigma^{4}\subset\mathbb{R}^{5}$ as a conformally invariant functional whose Euler–Lagrange equation has a leading term of the form $\Delta^{2}H$. They use tractor calculus to generate many different functionals arising from different geometric contexts.
The key idea of [23, 69], which has recently been developed in [6], is the link between $Q$-curvature and generalized Willmore energies. They show in [69], that one can define a notion of extrinsic GJMS operators on submanifolds. From these operators, one deduce a notion of extrinsic $Q$-curvature. In [23], the authors prove that for a 4-dimensional submanifold $\Sigma^{4}\subset M^{n}$ with $n\geq 5$, the integral of the difference between the extrinsic $Q$-curvature and the intrinsic $Q$-curvature of $\Sigma$ is a generalized Willmore energy. Hence, one could hope to find similarities between the study of the standard $Q$-curvature and generalized Willmore energies.
In order to study the space of immersions of a given hypersurface $\Sigma^{4}\subset\mathbb{R}^{5}$ into $\mathbb{R}^{5}$ and to provide a notion of ”best representation”, one would need a functional which is bounded from below and, as for the 2-dimensional case, would be a sort of conformally invariant distance of hypersurfaces to totally umbilic sets. One possibility has been introduced in [125] motivated by the generalization of the Dirichlet energy for the conformal Gauss map in dimension 2, see also [6] where Allen–Gover propose an alternative in any codimension using the normal tractor: if $\vec{\Phi}\colon\Sigma^{4}\to\mathbb{R}^{5}$ is a smooth immersion, then
$\mathcal{E}(\vec{\Phi})=\int_{\Sigma}4|dH_{\vec{\Phi}}|^{2}_{g_{\vec{\Phi}}}+\frac{1}{3}|\mathring{\mathrm{I\!I}}_{\vec{\Phi}}|_{g_{\vec{\Phi}}}^{4}+2H^{2}_{\vec{\Phi}}|\mathrm{I\!I}_{\vec{\Phi}}|_{g_{\vec{\Phi}}}^{2}-4H_{\vec{\Phi}}\,\text{tr}_{g_{\vec{\Phi}}}(\mathrm{I\!I}_{\vec{\Phi}}^{3})+2\,\text{tr}_{g_{\vec{\Phi}}}(\mathrm{I\!I}_{\vec{\Phi}}^{4})\ d\textup{vol}_{g_{\vec{\Phi}}}.$
A priori, minimal hypersurfaces of $\mathbb{R}^{5}$ are not critical points of $\mathcal{E}$ anymore. However, it is proved in [125], that $\mathcal{E}(\vec{\Phi})\geq 0$ with equality if and only if $\vec{\Phi}(\Sigma)$ is a round sphere. Furthermore, $\mathcal{E}$ is coercive, in the sense that there exists a universal constant $C>0$ such that for any immersion $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{5}$, it holds
$\int_{\Sigma}|d\mathrm{I\!I}_{\vec{\Phi}}|^{2}_{g_{\vec{\Phi}}}+|\mathrm{I\!I}_{\vec{\Phi}}|^{4}_{g_{\vec{\Phi}}}\,d\textup{vol}_{g_{\vec{\Phi}}}\leq C\,\mathcal{E}(\vec{\Phi}).$ (II.5)
We refer to the appendix of [125] for various relations between the $L^{p}$-norms of $H$, $\mathrm{I\!I}$, $\mathring{\mathrm{I\!I}}$ and their derivatives. Under this coercivity property, one could hope to generalize the known results of Willmore surfaces to the higher dimensional case.
Open Question II.31.
Given a 4-dimensional manifold $\Sigma^{4}$, can we explicitly compute or at least estimate the following number:
$\inf\left\{\mathcal{E}(\vec{\Phi})\colon\vec{\Phi}\in\mathrm{Imm}(\Sigma^{4},\mathbb{R}^{5})\right\}.$
Open Question II.32.
Can we minimize $\mathcal{E}$ among all the immersions of a given closed 4-dimensional manifold $\Sigma$ into $\mathbb{R}^{5}$? If yes, what are the minimizers?
Open Question II.33.
Can we exhibit a generic class of critical points of $\mathcal{E}$? For example, one could consider conformal transformations of hypersurfaces satisfying an equation of the form $P(H)=0$ for some polynomial $P$.
It is possible that $\mathcal{E}$ has a specific structure for which, there is no minimizers. In this case, it would be interesting to find another functional, coercive as well and which does possess minimizers. For instance, one can prove the following lower bound for $\mathcal{E}_{GR}$:
$\mathcal{E}_{GR}(\vec{\Phi})\geq\int_{\Sigma}|dH_{\vec{\Phi}}|^{2}_{g_{\vec{\Phi}}}-\frac{1}{12}|\mathring{\mathrm{I\!I}}_{\vec{\Phi}}|_{g_{\vec{\Phi}}}^{4}\,d\textup{vol}_{g_{\vec{\Phi}}}.$
Thus, there also exists a universal constant $C>0$ for which it holds
$\int_{\Sigma}|d\mathrm{I\!I}_{\vec{\Phi}}|_{g_{\vec{\Phi}}}^{2}+|\mathrm{I\!I}_{\vec{\Phi}}|_{g_{\vec{\Phi}}}^{4}\,d\textup{vol}_{g_{\vec{\Phi}}}\leq C\left(\mathcal{E}_{GR}(\vec{\Phi})+\int_{\Sigma}|\mathring{\mathrm{I\!I}}_{\vec{\Phi}}|^{4}\,d\textup{vol}_{g_{\vec{\Phi}}}\right).$
The energy on the right-hand side is still a generalized Willmore energy. All these generalized Willmore energies are of the form
$\vec{\Phi}\in\mathrm{Imm}(\Sigma^{4},\mathbb{R}^{5})\mapsto\int_{\Sigma}|d\mathrm{I\!I}_{\vec{\Phi}}|_{g_{\vec{\Phi}}}^{2}\,d\textup{vol}_{g}+O\left(\int_{\Sigma}|\mathrm{I\!I}|^{4}_{g_{\vec{\Phi}}}\,d\textup{vol}_{g_{\vec{\Phi}}}\right).$ (II.6)
Open Question II.34.
Can we prove that any generalized Willmore energy satisfying the coercivity condition (II.5) possesses minimizers?
Open Question II.35.
Given a conformally invariant functional of the form (II.6), can we produce a generic class of critical points?
In order to study Open Questions II.32 and II.34, one could follow the strategy introduced by Simon [180] and Bauer–Kuwert [12] in order to show the existence of a minimizer of the Willmore functional among the immersed surfaces of any given genus. In first step in [12] is to prove a key inequality on connected sums of non-umbilic immersed submanifolds. A generalization of this inequality has recently been proved in [189].
II.2.3 Analysis in high dimension
In order to answer the previous questions, one would need to develop a refined analysis for submanifolds in the spirit of the approach presented here and most likely combined with the existing theory of varifolds. An important starting point already raised in [80, Remark 5.4.6], would be to generalize the study of Coulomb frames in higher dimensions but also to generalize the work of Müller–Šverák [151] on the construction of isothermal coordinates for immersed surfaces in a Euclidean space. The Coulomb frame are of major importance due to their relation to isothermal coordinates, see for instance Theorem IV.4 below or [80, Section 5.4]. In higher dimensions, isothermal coordinates do not exist in general, the Cotton and Weyl tensors are the obstruction. However, harmonic coordinates always exist on any manifold of any dimension $n\geq 2$ (isothermal coordinates are harmonic). In recent work [128, 129], the second and third authors show that for an immersion $\vec{\Phi}\colon\Sigma^{4}\to\mathbb{R}^{5}$, an estimate of the following form provides an atlas of harmonic coordinates on $\Sigma$ describing a $C^{1}$-differential structure with quantitative estimates on all the quantities involved
$\int_{\Sigma}|d\mathrm{I\!I}_{\vec{\Phi}}|^{2}_{g_{\vec{\Phi}}}+|\mathrm{I\!I}_{\vec{\Phi}}|^{4}_{g_{\vec{\Phi}}}\ d\textup{vol}_{g_{\vec{\Phi}}}\leq E.$ (II.7)
In particular, the work [128] provides a possible generalization to the work of Müller–Šverák [151] together with an answer to Hélein’s question. The construction of harmonic coordinates is done in two steps. First, we proceed by a continuity argument reminiscent of the one of Uhlenbeck [186] for the construction of a Coulomb gauge for Yang–Mills connections. In this first step, we construct both a controlled system of coordinates and a controlled Coulomb frame. We can then perturb these coordinates to obtain controlled harmonic coordinates. The crucial analytic ingredient that replaces the $\mathrm{div}$-$\operatorname{curl}$ structure in dimension 2 is the use of Lorentz spaces. An estimate of the form (II.7) combined with the Gauss–Codazzi equations implies that the Riemann tensor $\mathrm{Riem}^{g_{\vec{\Phi}}}$ lies in $L^{(2,1)}$. This Lorentz exponent 1 is just enough to be able to close the continuity argument and obtain the existence of harmonic coordinates if $\mathrm{Riem}^{g_{\vec{\Phi}}}$ is small in $L^{(2,1)}$. This construction extends to all dimensions $n\geq 3$ as long as $\mathrm{Riem}^{g_{\vec{\Phi}}}$ lies in $L^{\left(\frac{n}{2},1\right)}$.
Open Question II.36.
Given a closed orientable $n$-dimensional manifold $\Sigma$ and an immersion $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{d}$ with $\mathrm{I\!I}_{\vec{\Phi}}\in L^{n}(\Sigma^{n})$, is it possible to construct counterexamples to the existence of controlled harmonic coordinates? Or can we construct controlled harmonic coordinates on a ”large” subset of $\Sigma$?
The next question is to know whether this atlas of harmonic coordinates can be ”uniformized” to obtain a global metric on $\Sigma$ conformal to $g_{\vec{\Phi}}$ together with some control following from an estimate of the form (II.7). A first attempt would be to find a metric having constant scalar curvature. This question is at the origin of the Yamabe problem and is not solvable in full generality, the topology of $\Sigma$ plays a role. An alternative has been proposed recently in [105], where the authors are able to solve locally $\Delta_{\bar{g}}\mathrm{Scal}_{\bar{g}}=0$ for some metric $\bar{g}$ conformal to $g$ in dimension 4. This equation is always solvable as long as the Sobolev inequalities are verified for the metric $g$ and the following regularity estimate holds
$\forall u\in C^{\infty}_{c}(\Sigma),\qquad\|\nabla u\|_{L^{4}(\Sigma,g_{\vec{\Phi}})}\leq C\,\|\Delta_{g_{\vec{\Phi}}}u\|_{L^{2}(\Sigma,g_{\vec{\Phi}})}.$ (II.8)
Sobolev inequalities are a consequence of the isoperimetric inequality, which always hold for submanifolds having small $L^{n}$-norm of its mean curvature, thanks to Michael–Simon inequality [5, 135]. A global version of the equation $\Delta_{\bar{g}}\mathrm{Scal}_{\bar{g}}=0$ would be that the $Q$-curvature of $\bar{g}$ is pointwise controlled by $|\mathrm{Riem}^{\bar{g}}|^{2}_{\bar{g}}$.
Open Question II.37.
Given an immersion $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{m}$ from a closed $4$-dimensional oriented manifold $\Sigma$ verifying (II.7), can we find a metric $\bar{g}$ conformal to $g_{\vec{\Phi}}$ such that the $Q$-curvature of $\bar{g}$ is constant? Is pointwise controlled by $|\vec{\mathrm{I\!I}}_{\vec{\Phi}}|^{4}_{g_{\vec{\Phi}}}$?
In order to know whether an equation on a manifold is regularizing or not, one needs to develop a refined PDE theory. In particular, one needs to control the constants involved in inequalities such as Sobolev or Poincaré inequalities and elliptic regularity. Such inequalities are directly related to isoperimetric problems, see for instance [33, Chapter IV]. Gallot [62] proved that the isoperimetric inequalities hold on a complete manifold $(\Sigma^{n},g)$ as soon as $\mathrm{Ric}_{g}\in L^{p}(\Sigma,g)$ for some $p>\frac{n}{2}$. Moreover Micheal–Simon inequality (see Brendle [26] for the optimal constants) implies that an isoperimetric inequality always holds for immersed $n$-dimensional submanifolds of arbitrary codimension in a Euclidean space whose mean curvature has small $L^{n}$-norm.
Open Question II.38.
Given a complete manifold $(\Sigma^{n},g)$, does there exists a number $s>0$ such that if $\mathrm{Riem}_{g}\in W^{s,\frac{n}{s+2}}(\Sigma,g)$, then the isoperimetric inequality holds?
In order to develop a variational theory for submanifolds of the Euclidean space, one needs to understand the ”closure” of this space under the norm we are interested in, in the same spirit that the notions of currents or varifolds arise as limits of smooth submanifolds having bounded area. In other words, it would be interesting to generalize Theorem I.2.
Open Question II.39.
Fix a number $s\in[0,n]$. Let $(\vec{\Phi}_{k})_{k\in\mathbb{N}}$ be a sequence of smooth immersions from a given orientable $n$-dimensional submanifold $\Sigma^{n}\subset\mathbb{R}^{n+m}$ into $\mathbb{R}^{n+m}$. Assume that
$\sup_{k\in\mathbb{N}}\left\|\vec{\mathrm{I\!I}}_{\vec{\Phi}_{k}}\right\|_{W^{s,\frac{n}{s+1}}(\Sigma,g_{\vec{\Phi}})}\leq C.$
What is the limit ${\lim_{k\to+\infty}}\vec{\Phi}_{k}$? Can we develop a bubbling analysis adapted to the above energy?
The Sobolev space $W^{s,\frac{n}{s+1}}(\Sigma^{n})$ always embeds in $L^{n}(\Sigma^{n})$. Hence, one should not expect a strong limit in full generality, but rather a weak limit on $\Sigma$ minus a set having small measure. The case $s=\frac{n-2}{2}$ and $n$ even has been treated in [129], based on the approximation result obtained in Theorem IV.23.
Theorem II.40.
Let $(\vec{\Phi}_{k})_{k\in\mathbb{N}}$ be a sequence of smooth immersions from a given orientable $n$-dimensional submanifold $\Sigma^{n}\subset\mathbb{R}^{d}$ into $\mathbb{R}^{d}$. Assume that
$\sup_{k\in\mathbb{N}}\int_{\Sigma}\big|\nabla^{\frac{n}{2}-1}\vec{\mathrm{I\!I}}_{\vec{\Phi}_{k}}\big|^{2}_{g_{\vec{\Phi}_{k}}}+\cdots+\big|\vec{\mathrm{I\!I}}_{\vec{\Phi}_{k}}\big|^{n}_{g_{\vec{\Phi}_{k}}}\,d\textup{vol}_{g_{\vec{\Phi}_{k}}}\leq C.$
Up to rescaling, we can assume that each $\vec{\Phi}_{k}(\Sigma)$ is contained in the unit ball $\mathbb{B}^{d}(0,1)$ and has diameter 1. Then, there exist a closed $C^{1}$ manifold $\Sigma_{\infty}$ (possibly non-connected) and a weak immersion $\vec{\Phi}_{\infty}\colon\Sigma_{\infty}\to\mathbb{R}^{d}$ such that the following holds. There exists a finite number of points $\vec{q}_{1},\ldots,\vec{q}_{K}\in\mathbb{R}^{d}$ such that for each $r>0$, we have the following convergence in the pointed Gromov–Hausdorff topology
$\left(\Sigma\setminus\vec{\Phi}_{k}^{-1}\left(\mathbb{B}^{d}(\vec{q}_{1},r)\cup\cdots\cup\mathbb{B}^{d}(\vec{q}_{1},r)\right),g_{\vec{\Phi}_{k}}\right)\to\left(\Sigma_{\infty}\setminus\vec{\Phi}_{\infty}^{-1}\left(\mathbb{B}^{d}(\vec{q}_{1},r)\cup\cdots\cup\mathbb{B}^{d}(\vec{q}_{1},r)\right),g_{\vec{\Phi}_{\infty}}\right).$
Moreover, $\vec{\Phi}_{k}$ converges to $\vec{\Phi}_{\infty}$ in $C^{0}\left(\Sigma_{\infty}\setminus\vec{\Phi}_{\infty}^{-1}\left(\mathbb{B}^{d}(\vec{q}_{1},r)\cup\cdots\cup\mathbb{B}^{d}(\vec{q}_{1},r)\right),g_{\vec{\Phi}_{\infty}}\right)$ for any $r>0$.
It would be interesting to study the formation of singularities. It would also be useful to study the cases where $s$ is not an integer. Once we have a critical point, it is important to know whether the Euler–Lagrange equation is regularizing or not. In [16], we obtain the following result in dimension $4$.
Theorem II.41.
Let $\vec{\Phi}\colon\mathbb{B}^{4}(0,1)\to\mathbb{R}^{5}$ be a weak immersion critical point of the functional
$E(\vec{\Phi})=\int_{\mathbb{B}^{4}(0,1)}\big|\nabla H_{\vec{\Phi}}\big|^{2}_{g_{\vec{\Phi}}}\ d\textup{vol}_{g_{\vec{\Phi}}}.$
Then $\vec{\Phi}$ is real-analytic in harmonic coordinates.
The generalizations to the cases with higher codimension or with lower order terms as in the Graham–Reichert functional will be the subject of a forthcoming work.
III Preliminaries
III.1 Notation
  1. (i)
    In the Euclidean space $\mathbb{R}^{n}$, we define $B_{r}(x)\coloneqq\{y\in\mathbb{R}^{n}\colon|y-x|<r\}$, and write $B^{n}=B_{1}(0)\subset\mathbb{R}^{n}$ for the unit ball. In the case $n=2$, write $D^{2}=B^{2}$, $D_{r}(x)=B_{r}(x)\subset\mathbb{R}^{2}$. The phrase “for every ball $B_{r}\subset U$” refers to every ball of radius $r$ contained in $U$. Denote $\mathbb{S}^{n-1}=\partial B^{n}$.
  2. (ii)
    We denote by $\mathbb{R}^{m\times n}$ the space of $m\times n$ real matrices, and by $\mathbb{R}^{n\times n}_{\text{sym}}$ the space of $n\times n$ symmetric real matrices.
  3. (iii)
    For a real-valued function $f$, we denote $f^{+}\coloneqq\max(f,0)$, $f^{-}\coloneqq\max(-f,0)$.
  4. (iv)
    Let $\Sigma$ be a closed smooth manifold, $E$ be a vector bundle over $\Sigma$, we denote by $L^{p}(\Sigma,E)$ the space of $L^{p}$ sections of $E$. This convention also applies to the other function/distribution spaces.
  5. (v)
    For an open set $U\subset\mathbb{R}^{n}$, we denote by $\mathcal{D}^{\prime}(U,\mathbb{R}^{m})$ the space of $\mathbb{R}^{m}$-valued distributions on $U$ with the weak topology, and we write $\mathcal{D}^{\prime}(U)=\mathcal{D}^{\prime}(U,\mathbb{R})$.
  6. (vi)
    We denote by $\mathcal{L}^{n}$ the Lebesgue measure on $\mathbb{R}^{n}$. For $U\subset\mathbb{R}^{n}$ with $\mathcal{L}^{n}(U)<\infty$, $f\in L^{1}(U)$, we write $\fint_{U}f\coloneqq(\mathcal{L}^{n}(U))^{-1}\int_{U}f$.
  7. (vii)
    For open sets $U$ and $V$, we write $V\Subset U$ if $\bar{V}$ is compact and $\bar{V}\subset U$.
  8. (viii)
    Let $U$ be a smooth manifold, $f\colon U\rightarrow\mathbb{R}^{m}$. We denote by $df$ the differential of $f$.
    When $U$ is an open subset of $\mathbb{R}^{n}$, we denote $\partial_{i}=\frac{\partial}{\partial x^{i}}$ when there is no ambiguity, and denote the Jacobian matrix of $f$ by $\nabla f$. If $n=2$, we write $\nabla^{\perp}f\coloneqq(-\partial_{2}f,\partial_{1}f)$. If in addition that $f=(\vec{f}_{1},\vec{f}_{2})$ for $\vec{f}_{1},\vec{f}_{2}\colon U\to\mathbb{R}^{m}$, we write $\operatorname{curl}f\coloneqq\partial_{1}\vec{f}_{2}-\partial_{2}\vec{f}_{1}$.
  9. (ix)
    Let $(V,g)$ be an $n$-dimensional inner product space with a positively oriented orthonormal basis $(e_{1},\dots,e_{n})$. The Hodge star operator $_{g}$ is defined as the unique linear operator from $\bigwedge^{k}V$ to $\bigwedge^{n-k}V$ satisfying α∧*_g  β=⟨α,β⟩_g  e_1∧⋯∧e_n,   for all   α,β∈⋀^kV. See for instance [94, Section 3.3].
    When $V=\mathbb{R}^{n}$ and $g=g_{\text{std}}$ is the standard inner product on $\mathbb{R}^{n}$, we write $\star$ instead of $_{g}$.
  10. (x)
    Let $U\subset\mathbb{R}^{n}$ be an open set and let $g$ be a Riemannian metric on $U$. On the space of differential $\ell$-forms on $U$, we define the codifferential (see [94, Definition 3.3.1]) by
    $d^{*_{g}}\coloneqq(-1)^{\ell}*_{g}^{-1}d\,*_{g}=(-1)^{n(\ell+1)+1}*_{g}d\,*_{g}.$ (III.1)
    In particular, when $n=4$, we have $d^{*_{g}}=-*_{g}d\,*_{g}$. If $\alpha\in C^{\infty}(U,\bigwedge^{\ell}T^{*}U)$ and $\beta\in C^{\infty}(U,\bigwedge^{\ell+1}T^{*}U)$ with at least one of $\alpha,\beta$ compactly supported in $U$, then we have
    $\int_{U}\langle d\alpha,\beta\rangle_{g}\,d\textup{vol}_{g}=\int_{U}\langle\alpha,d^{*_{g}}\beta\rangle_{g}\,d\textup{vol}_{g}.$ (III.2)
  11. (xi)
    Levi-Civita symbol:
[…]
  12. (xii)
    Let $U\subset\mathbb{R}^{n}$ be open and bounded. We call $U$ a Lipschitz domain if for each $y\in\partial U$, there exist $r>0$ and a Lipschitz function $f\colon\mathbb{R}^{n-1}\rightarrow\mathbb{R}$ such that, upon rotating and relabeling the coordinate axes if necessary and writing $x=(x^{\prime},x_{n})$, we have
    $U\cap B_{r}(y)=\{x\in B_{r}(y):x_{n}>f(x^{\prime})\}.$
    Equivalently, we say $\partial U$ is Lipschitz or in $C^{0,1}$.
  13. (xiii)
    We define a bilinear map $\mathbin{\dot{\wedge}}\colon\big(\bigwedge^{k_{1}}T^{*}_{p}\Sigma\big)^{m}\times\big(\bigwedge^{k_{2}}T^{*}_{p}\Sigma\big)^{m}\rightarrow\bigwedge^{k_{1}+k_{2}}T^{*}_{p}\Sigma$ by
    $(\vec{u}_{1}\mkern 2.0mudx_{I})\mathbin{\dot{\wedge}}(\vec{u}_{2}\mkern 2.0mudx_{J})=(\vec{u}_{1}\cdot\vec{u}_{2})\mkern 2.0mudx_{I}\wedge dx_{J},$ (III.3)
[…]
    $\vec{v}_{1}\times(\vec{v}_{2}\mkern 2.0mudx_{J})=(\vec{v}_{1}\times\vec{v}_{2})\mkern 2.0mudx_{J},\quad(\vec{v}_{1}\mkern 2.0mudx_{I})\overset{\times}{\wedge}(\vec{v}_{2}\mkern 2.0mudx_{J})=(\vec{v}_{1}\times\vec{v}_{2})\mkern 2.0mudx_{I}\wedge dx_{J},$
    $\vec{w}_{1}\wedge(\vec{w}_{2}\mkern 2.0mudx_{J})=(\vec{w}_{1}\wedge\vec{w}_{2})\mkern 2.0mudx_{J},\quad(\vec{w}_{1}\mkern 2.0mudx_{I})\overset{\text{$\wedge$}}{\wedge}(\vec{w}_{2}\mkern 2.0mudx_{J})=(\vec{w}_{1}\wedge\vec{w}_{2})\mkern 2.0mudx_{I}\wedge dx_{J},$
    where $\vec{v}_{1},\vec{v}_{2}\in\mathbb{R}^{3}$, $\vec{w}_{1},\vec{w}_{2}\in\bigwedge\mathbb{R}^{m}$, $I,J$ are multi-indices.
  14. (xiv)
    For $\vec{u}=(\vec{u}_{1},\vec{u}_{2})$, $\vec{v}=(\vec{v}_{1},\vec{v}_{2})$, where $\vec{u}_{i},\vec{v}_{i}\in\mathbb{R}^{3}$, denote →u⋅→v≔∑_i=1^2 →u_i ⋅→v_i,  →u×→v≔∑_i=1^2 →u_i ×→v_i.
  15. (xv)
    We adopt Einstein summation convention for convenience of computation, and set $\delta_{i}^{j}=\delta_{ij}=\mathbf{1}_{i=j}$.
  16. (xvi)
    We use $C$ to denote positive constants, and $C(\alpha,\beta,\dots)$ to denote a positive constant depending only on $\alpha,\,\beta,\dots$. Similarly, when a term $A$ depends only on $\alpha,\,\beta,\dots$, we write $A=A(\alpha,\beta,\dots)$.
  17. (xvii)
    Let $U\subset\mathbb{R}^{n}$ be open and $\vec{\Phi}\colon U\to\mathbb{R}^{m}$ be an immersion. Set $g=g_{\vec{\Phi}}=\vec{\Phi}^{*}g_{\text{std}}$, where $g_{\textup{std}}$ is the Euclidean metric on $\mathbb{R}^{m}$. Denote g_ij=∂_i→Φ⋅∂_j→Φ, (g^ij)=(g_ij)^-1, detg=det(g_ij), dvol_g=(detg)^12dx^1∧⋯∧dx^n. The pullback metric induces a pairing on $T^{*}U$ with $\langle dx^{i},dx^{j}\rangle_{g}=g^{ij}$; we denote by $|\cdot|_{g}$ the associated pointwise norm. We also write $|\cdot|_{\mathbb{R}^{m}}$ for the Euclidean norm in $\mathbb{R}^{m}$ (abbreviated to $|\cdot|$ when unambiguous).
    For $f\colon U\to\mathbb{R}$, define the Laplacian
    $\Delta_{g}f\coloneqq*_{g}d*_{g}df=(\det g)^{-\frac{1}{2}}\partial_{i}\big(g^{ij}(\det g)^{\frac{1}{2}}\partial_{j}f\big).$
    Let $\pi_{\vec{n}}$ denote the orthogonal projection onto the normal bundle of $\vec{\Phi}(U)\subset\mathbb{R}^{m}$. We define the Gauss map and second fundamental form
    $\vec{n}\coloneqq\star\,\frac{\partial_{1}\vec{\Phi}\wedge\cdots\wedge\partial_{n}\vec{\Phi}}{|\partial_{1}\vec{\Phi}\wedge\cdots\wedge\partial_{n}\vec{\Phi}|},$
    $\vec{\mathrm{I\!I}}(X,Y)\coloneqq\pi_{\vec{n}}X(d\vec{\Phi}(Y)),\qquad\text{for all }X,Y\in T_{p}U\text{ and all }p\in U.$
    The mean curvature vector is defined by
    $\vec{H}\coloneqq n^{-1}\,\mbox{tr}_{g}\,\vec{\mathrm{I\!I}}=\frac{1}{n}\,g^{ij}\,\vec{\mathrm{I\!I}}_{ij}=\frac{1}{n}g^{ij}\,\vec{\mathrm{I\!I}}\left(\frac{\partial}{\partial x^{i}},\frac{\partial}{\partial x^{j}}\right).$
    When $m=n+1$, we write
    $\mathrm{I\!I}_{ij}\coloneqq\vec{\mathrm{I\!I}}_{ij}\cdot\vec{n},\quad H\coloneqq\vec{H}\cdot\vec{n}.$
  18. (xviii)
[…]
    A consequence of (III.5) is that, for $\alpha\in\bigwedge^{p}V$ and $\beta\in\bigwedge^{q}V$,
[…]
  19. (xix)
    Throughout this paper, unless explicitly stated otherwise, we write $\nabla$, $|\cdot|$ for the flat Euclidean derivative and pointwise tensor norms. We reserve $\nabla^{g}$, $\langle\cdot,\cdot\rangle_{g}$, and $|\cdot|_{g}$ for the metric $g$. For an open set $U\subset\mathbb{R}^{n}$, we write $\alpha\in W^{k,p}(U,\bigwedge^{\ell}\mathbb{R}^{n})$ if $\alpha=\sum_{|I|=\ell}\alpha_{I}\mkern 1.0mudx^{I}$ with each $\alpha_{I}\in W^{k,p}(U)$, where $I$ ranges over all strictly increasing multi-indices of length $\ell$. We set
    $|\nabla\alpha|^{2}\coloneqq\sum_{i=1}^{n}\sum_{I}|\partial_{i}\mkern 1.0mu\alpha_{I}|^{2},\qquad\|\alpha\|_{W^{k,p}(U)}\coloneqq\sum_{I}\|\alpha_{I}\|_{W^{k,p}(U)}.$ (III.8)
    Similarly, we write $X\in W^{k,p}(U,TU)$ if $X=X_{j}\mkern 1.0mu\frac{\partial}{\partial x^{j}}$ with each $X_{j}\in W^{k,p}(U)$, and set
    $|\nabla X|^{2}\coloneqq\sum_{i=1}^{n}\sum_{j=1}^{n}|\partial_{i}\mkern 1.0muX_{j}|^{2},\qquad\|X\|_{W^{k,p}(U)}\coloneqq\sum_{j=1}^{n}\|X_{j}\|_{W^{k,p}(U)}.$ (III.9)
    The same convention applies to any Banach function/distribution space on $U$ (e.g. $L^{p}$, $L^{p,q}$, $W^{k,(p,q)}$).
III.2 A useful formula
We provide the proof of an identity here without using conformal coordinates.
Lemma III.1.
[168, Lemma 1.8] Let $\Sigma$ be a $2$-dimensional oriented smooth manifold, and $\vec{\Phi}\colon\Sigma\rightarrow\mathbb{R}^{3}$ be a smooth immersion. Let $_{g}$ be the Hodge star operator with respect to $g=g_{\vec{\Phi}}$. Then we have
$-2H\,d\vec{\Phi}=d\vec{n}+\vec{n}\times*_{g}\,d\vec{n}.$ (III.10)
Proof.
Let $p\in\Sigma$, $(x^{1},x^{2})$ be positive local coordinates around $p$, $\alpha\coloneqq dx^{1}\wedge dx^{2}$, $\mathrm{I\!I}$ denotes the second fundamental form of $\vec{\Phi}$, and define
$g_{ij}\coloneqq\partial_{i}\vec{\Phi}\cdot\partial_{j}\vec{\Phi},\quad(g^{ij})=(g_{ij})^{-1},\quad\det g=\det(g_{ij}).$
For any $k\in\{1,2\}$, note
$\partial_{k}(\vec{n}\cdot\vec{n})=2\,\partial_{k}\vec{n}\cdot\vec{n}=0,$
hence
$\partial_{k}\vec{n}=g^{ij}(\partial_{k}\vec{n}\cdot\partial_{j}\vec{\Phi})\,\partial_{i}\vec{\Phi}.$ (III.11)
Fix $s\in\{1,2\}$, then we have
$dx_{s}\wedge(\vec{n}\times*_{g}\,d\vec{n})$ $=\vec{n}\times\langle dx_{s},d\vec{n}\rangle_{g}\,\sqrt{\det g}\ dx^{1}\wedge dx^{2}$
$=g^{sj}(\partial_{1}\vec{\Phi}\times\partial_{2}\vec{\Phi})\times\partial_{j}\vec{n}\ \alpha$
$=g^{sj}\,\varepsilon^{ki}(\partial_{k}\vec{\Phi}\cdot\partial_{j}\vec{n})\,\partial_{i}\vec{\Phi}\,\alpha$
$=-g^{sj}\,\varepsilon^{ki}\,\mathrm{I\!I}_{kj}\,\partial_{i}\vec{\Phi}\,\alpha,$
where the third equality follows from the identity
$(\vec{u}\times\vec{v})\times\vec{w}=(\vec{u}\cdot\vec{w})\vec{v}-(\vec{v}\cdot\vec{w})\vec{u}$
for all $\vec{u},\vec{v},\vec{w}\in\mathbb{R}^{3}$. By (III.11), we also get
$dx_{s}\wedge d\vec{n}$ $=\varepsilon^{sk}\,\partial_{k}n\,\alpha=-g^{ij}\epsilon^{sk}\,\mathrm{I\!I}_{kj}\,\partial_{i}\vec{\Phi}\,\alpha,$
Therefore,
$\begin{aligned} dx_{s}\wedge(d\vec{n}+\vec{n}\times*_{g}\,d\vec{n})&=-(g^{sj}\varepsilon^{ki}+g^{ij}\varepsilon^{sk})\,\mathrm{I\!I}_{kj}\,\partial_{i}\vec{\Phi}\,\alpha\\[2.84526pt] &=-\varepsilon^{si}g^{kj}\,\mathrm{I\!I}_{kj}\,\partial_{i}\vec{\Phi}\,\alpha\\[2.84526pt] &=-2H\,dx_{s}\wedge d\vec{\Phi},\end{aligned}$ (III.12)
where the second equality follows from the identity
$f(s)\varepsilon^{ki}+f(k)\varepsilon^{is}+f(i)\varepsilon^{sk}=0$
which holds for any $i,k,s\in\{1,2\}$ and real-valued function $f$ defined on $\{1,2\}$. Since (III.12) holds for any $s\in\{1,2\}$, (III.10) follows immediately. ∎
III.3 Some estimates for elliptic equations
We first introduce a standard Morrey estimate for elliptic equations of divergence form with no lower-order terms.
Lemma III.2 ([77, Lemma 4.12]).
Let $n\in\mathbb{N}^{+}$. Suppose $\{a^{ij}\}_{i,j=1}^{n}\subset L^{\infty}(B^{n})$ satisfies
$\Lambda^{-1}|\xi|^{2}\leq a^{ij}(x)\,\xi_{i}\,\xi_{j}\leq\Lambda|\xi|^{2},\qquad\text{for a.e. }x\in B^{n}\text{ and all }\,\xi\in\mathbb{R}^{n},$ (III.13)
where $\Lambda>0$ is a constant. Suppose $u\in W^{1,2}(B^{n})$ is a weak solution of the equation $\partial_{j}(a^{ij}\,\partial_{i}u)=0$ in $B^{n}$, i.e.
$\int_{B^{n}}a^{ij}\,\partial_{i}u\,\partial_{j}\varphi=0,\qquad\text{for all }\,\varphi\in C^{\infty}_{c}(B^{n}).$
Then there exist $\alpha=\alpha(n,\Lambda)\in(0,1)$ and $C=C(n,\Lambda)>0$ such that for any $0<r<1$ there holds
$\int_{B_{r}(0)}|\nabla u|^{2}\leq C\,r^{n-2+2\alpha}\int_{B^{n}}|\nabla u|^{2}.$
Now we assume $n=2$. In 1992, S. Chanillo and Y.Y. Li obtained the following result, which generalized a famous estimate on integrability by compensation proved by H.C. Wente [187], in which it is assumed $a^{ij}=\delta_{ij}$.
Lemma III.3 ([32, Theorem 0.2]).
Suppose $\{a^{ij}\}_{i,j=1}^{2}$ satisfies (III.13) for some $\Lambda\in(0,\infty)$, and $a^{ij}\equiv a^{ji}$. For $u,v\in W^{1,2}(D^{2})$, there exists a unique solution $\phi\in W^{1,2}_{0}(D^{2})$ to the problem
$\begin{cases}\partial_{i}(a^{ij}\,\partial_{j}\phi)=\partial_{1}u\,\partial_{2}v-\partial_{2}u\,\partial_{1}v&\mbox{in }D^{2},\\[5.69054pt] \phi=0&\mbox{on }\partial D^{2},\end{cases}$
and we also have $\phi\in C(\overline{D^{2}})$ with the estimate
$\|\phi\|_{L^{\infty}(D^{2})}+\|\nabla\phi\|_{L^{2}(D^{2})}\leq C(\Lambda)\,\|\nabla u\|_{L^{2}(D^{2})}\,\|\nabla v\|_{L^{2}(D^{2})}.$
We have a stronger version of this result, see Proposition IV.8.
III.4 Bourgain–Brezis inequality
In 2002, J. Bourgain and H. Brezis proved a striking inequality [24]:
$\left\|u-\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{\mathbb{T}^{n}}u\right\|_{L^{\frac{n}{n-1}}(\mathbb{T}^{n})}\leq C\left\|\nabla u\right\|_{L^{1}+W^{-1,\frac{n}{n-1}}(\mathbb{T}^{n})},$ (III.14)
where $C=C(n)$, $\mathbb{T}^{n}$ is the $n$-dimensional torus. In the case $n=2$, F. Da Lio and T. Rivière [48] provided a proof for a stronger version of (III.14) in 2020. Here we denote $\dot{W}^{-1,2}(\mathbb{R}^{2})$ as the dual space of the homogeneous Sobolev space $\dot{W}^{1,2}(\mathbb{R}^{2})$.
Lemma III.4 ([48, Lemma II.3]).
Assume $g\in L^{1}(\mathbb{R}^{2},\mathbb{R}^{2})$, $f\in\dot{W}^{-1,2}(\mathbb{R}^{2},\mathbb{R}^{2})$, $u$ be a tempered distribution on $\mathbb{R}^{2}$ such that $\nabla u=f+g$. Then there exists $c\in\mathbb{R}$, and a universal constant $C$ such that
$\|u-c\|_{L^{2}(\mathbb{R}^{2})}\leq C(\|f\|_{\dot{W}^{-1,2}(\mathbb{R}^{2})}+\|g\|_{L^{1}(\mathbb{R}^{2})}).$
Corollary III.5.
Assume $u\in\mathcal{D}^{\prime}(D^{2})$ such that $\nabla u=f+g$ with $g\in L^{1}(D^{2},\mathbb{R}^{2})$, $f\in W^{-1,2}(D^{2},\mathbb{R}^{2})$. Then for any $r\in(0,1)$, there holds
$\left\|u-\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.98003pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.26338pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.6363pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.45924pt}}\!\int_{D_{r}(0)}u\right\|_{L^{2}(D_{r}(0))}\leq C(r)\big(\|f\|_{W^{-1,2}(D^{2})}+\|g\|_{L^{1}(D^{2})}\big).$
III.5 Quasiconformal maps and Beltrami equations
Definition III.6.
Let $U\subset\mathbb{R}^{2}$ be an open subset. A quasiconformal map $f\colon U\to f(U)\subset\mathbb{C}$ is an orientation-preserving homeomorphism (i.e. $\det(\nabla f)>0$ a.e.) satisfying $f\in W^{1,2}_{\textup{loc}}(U)$ and there exists a constant $k\in[0,1)$ such that
$|\partial_{\bar{z}}f|\leq k|\partial_{z}f|\quad\mbox{a.e. on }U,$
where $\partial_{z}\coloneqq(\partial_{x}-i\partial_{y})/2$, $\partial_{\bar{z}}\coloneqq(\partial_{x}+i\partial_{y})/2$. We call the function $\mu_{f}\coloneqq\partial_{\bar{z}}f/\partial_{z}f$ the complex dilatation of $f$.
Lemma III.7 ([90, Theorem 4.30]).
Let $\mu\in L^{\infty}(\mathbb{C},\mathbb{C})$ satisfies $\|\mu\|_{L^{\infty}(\mathbb{C})}<1$. Then there exists a homeomorphism $f\colon\bar{\mathbb{C}}\rightarrow\bar{\mathbb{C}}$ which is a quasiconformal map of $\mathbb{C}$ and satisfies the Beltrami equation $\partial_{\bar{z}}f=\mu\,\partial_{z}f$. Such $f$ is uniquely determined by the conditions $f(0)=0$, $f(1)=1$, $f(\infty)=f(\infty)$.
Lemma III.8 ([42, Proposition 4]).
Suppose $\mu\in W^{1,2}(\mathbb{C})$ has compact support and $\|\mu\|_{\infty}<1$. Let $f\colon\mathbb{C}\rightarrow\mathbb{C}$ be a quasiconformal map solving the Beltrami equation $\partial_{\bar{z}}f=\mu\,\partial_{z}f$. Then it holds $f\in W_{\textup{loc}}^{2,q}(\mathbb{C})$ for any $q<2$.
III.6 Hardy space $\mathcal{H}^{1}$
For the definition of the Hardy space $\mathcal{H}^{1}(\mathbb{R}^{n})$, see [71, 182].
Lemma III.9 ([43]).
Let $n\geq 2$. Suppose $u\in W^{1,n}(\mathbb{R}^{n},\mathbb{R}^{n})$. Then $\det(\nabla u)\in\mathcal{H}^{1}(\mathbb{R}^{n})$ with
$\|\det(\nabla u)\|_{\mathcal{H}^{1}(\mathbb{R}^{n})}\leq C(n)\|u\|_{W^{1,n}(\mathbb{R}^{n})}.$
Standard singular integral theory (see for instance [182, Theorem III.4]) implies the following results.
Lemma III.10.
Let $n\geq 2$, $f\in\mathcal{H}^{1}(\mathbb{R}^{n})$. Suppose $u\in L^{1}(B^{n})$ satisfies $\Delta u=f$ in $\mathcal{D}^{\prime}(B^{n})$. Then it holds $u\in W^{2,1}_{\textup{loc}}(B^{n})$ and for any $r\in(0,1)$, we have
$\|u\|_{W^{2,1}(B_{r}(0))}\leq C(n,r)\big(\|f\|_{\mathcal{H}^{1}(\mathbb{R}^{n})}+\|u\|_{L^{1}(B^{n})}\big).$ (III.15)
Corollary III.11.
Let $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(D^{2},\mathbb{R}^{m})$ be weakly conformal (see Definitions IV.1 and IV.2) with $g_{\vec{\Phi}}\coloneqq\vec{\Phi}^{*}g_{\text{std}}=e^{2\lambda}(dx_{1}^{2}+dx_{2}^{2})$ for $\lambda\in L^{\infty}(D^{2})$. Then we have $\lambda\in W^{2,1}_{\textup{loc}}(D^{2})$.
Proof.
Let $(\vec{e}_{1},\vec{e}_{2})\coloneqq e^{-\lambda}(\partial_{1}\vec{\Phi},\partial_{2}\vec{\Phi})\in L^{\infty}\cap W^{1,2}(D^{2})$, then we have
$\begin{cases}\vec{e}_{1}\cdot\partial_{1}\vec{e}_{2}=e^{-2\lambda}\partial_{1}\vec{\Phi}\cdot\partial_{1}\partial_{2}\vec{\Phi}=\frac{1}{2}e^{-2\lambda}\partial_{2}|\partial_{1}\vec{\Phi}|^{2}=\partial_{2}\lambda,\\[8.53581pt] \vec{e}_{1}\cdot\partial_{2}\vec{e}_{2}=e^{-2\lambda}\partial_{1}\vec{\Phi}\cdot\partial_{2}\partial_{2}\vec{\Phi}=-e^{-2\lambda}\partial_{2}\partial_{1}\vec{\Phi}\cdot\partial_{2}\vec{\Phi}=-\partial_{1}\lambda.\end{cases}$
In other words, it holds $\nabla^{\perp}\lambda=\vec{e}_{1}\cdot\nabla\vec{e}_{2}$. Thus we have, denoting $\vec{e}_{j}=(e_{j,1},\dots,e_{j,m})$ for $j\in\{1,2\}$,
$\Delta\lambda=\operatorname{curl}(\nabla^{\perp}\lambda)=-\nabla^{\perp}\vec{e}_{1}\cdot\nabla\vec{e}_{2}=-\sum_{i=1}^{m}\nabla^{\perp}e_{1,i}\,\nabla e_{2,i}.$ (III.16)
For $1\leq i\leq m$, define $e^{i}\in W^{1,2}(D^{2},\mathbb{R}^{2})$ by $e^{i}=(e_{1,i},e_{2,i})$. Then we have
$\det(\nabla e^{i})=\nabla^{\perp}e_{1,i}\cdot\nabla e_{2,i}.$
For any fixed $r\in(0,1)$, we choose $\eta\in C_{c}^{\infty}(D^{2})$ such that $\eta\equiv 1$ on $D_{r}(0)$. Then it holds
$\Delta\lambda=-\sum_{i=1}^{m}\det(\nabla(\eta\mkern 1.0mue^{i}))\qquad\text{in }\mathcal{D}^{\prime}(D_{r}(0)).$
Since $\eta\mkern 1.0mue^{i}\in W^{1,2}(\mathbb{R}^{2})$ for each $i$, we have $\det(\nabla(\eta e^{i}))\in\mathcal{H}^{1}(\mathbb{R}^{2})$ by Lemma III.9. This finally implies $\lambda\in W^{2,1}_{\textup{loc}}(D^{2})$ by Lemma III.10. ∎
III.7 Functions of VMO
Definition III.12.
  1. (i)
    Let $n\geq 1$, $U\subset\mathbb{R}^{n}$ be open, $f\in L^{1}_{\textup{loc}}(U)$. For $r>0$, we define
    $\beta_{r}(f)=\beta_{r,U}(f)\coloneqq\sup_{\begin{subarray}{c}\rho\leq r\\ B_{\rho}\subset U\end{subarray}}\fint_{B_{\rho}}\left|f-\fint_{B_{\rho}}f\right|,$ (III.17)
    where $B_{\rho}$ ranges over balls of radius $\rho$ contained in $U$. We say $f$ is of vanishing mean oscillation and write $f\in VMO(U)$ if lim_r0β_r(f)=0.
  2. (ii)
    Let $N$ be an $n$-dimensional smooth manifold without boundary, $U\subset N$ be open and precompact. For $f\in L^{1}_{\textup{loc}}(U)$, we say $f\in VMO(U)$ if for any coordinate system $(V,\phi)$ (an open set $V\subset N$ with a smooth diffeomorphism $\phi\colon V\rightarrow\phi(V)\subset\mathbb{R}^{n}$), any $\xi\in C^{\infty}_{c}(V)$ there holds $(\xi f)\circ\phi^{-1}\in VMO(\phi(V\cap U))$. Similarly, for $f\in W^{1,1}_{\textup{loc}}(U)$ we say $df\in VMO(U)$ if for any coordinate system $(V,\phi)$, $\xi\in C_{c}^{\infty}(V)$ there holds $\nabla((\xi f)\circ\phi^{-1})\in VMO(\phi(V\cap U),\mathbb{R}^{n})$.
Lemma III.13 ([92, Lemma 2.3]).
Let $B_{R}\subset\mathbb{R}^{n}$ be a ball of radius $R$. Suppose there exist constants $\beta,K>0$ such that for any $r>0$ and ball $B_{r}\subset B_{R}$ of radius $r$ satisfying $\operatorname{dist}(B_{r},\partial B_{R})\geq Kr$, there holds
$\int_{B_{r}}\left|f-\fint_{B_{r}}f\right|\leq\beta\,r^{n}.$ (III.18)
Then there exists a positive constant $C=C(n,K)$ such that
$\int_{B_{R}}\left|f-\fint_{B_{R}}f\right|\leq C\beta\,R^{n}.$ (III.19)
Proof.
We can without loss of generality assume $B_{R}=B_{1}(0)$. Let $\sigma\coloneqq 1-(6K+6)^{-1}>\frac{5}{6}$. We construct collections $\{\mathcal{G}^{\prime}_{i}\}_{i=0}^{\infty}$, $\{\mathcal{G}_{i}\}_{i=0}^{\infty}$ of open balls in the following way.
Let $\tilde{B}\coloneqq B_{1-\sigma}(0)$, $\mathcal{G}^{\prime}_{0}\coloneqq\{\tilde{B}\}$. For each $i\geq 1$, there exists a cover $\mathcal{G}_{i}^{\prime}$ of $\{x\in\mathbb{R}^{n}:|x|=1-\sigma^{i}\}$ such that any element $B\in\mathcal{G}^{\prime}_{i}$ is a ball of radius $\sigma^{i}(1-\sigma)$ with center lying on $\{x\in\mathbb{R}^{n}:|x|=1-\sigma^{i}\}$, and $\operatorname{Card}(\mathcal{G}_{i})\leq C(n,K)\,\sigma^{-(n-1)i}$. Let $\mathcal{G}_{i}\coloneqq\{2B:B\in\mathcal{G}^{\prime}_{i}\}$ for $i\geq 0$, where for any constant $c>0$ and any ball $B=B_{r}(x)$ we have denoted $cB=B_{cr}(x)$.
For any $i\geq 0$ and $x\in B_{1}(0)$ with $1-\sigma^{i}\leq|x|\leq 1-\sigma^{i+1}$, there exists $|x^{\prime}|=1-\sigma^{i}$ such that $|x^{\prime}-x|\leq\sigma^{i}(1-\sigma)$. Hence there exists a ball $B\in\mathcal{G}^{\prime}_{i}$ centered at $x^{\prime\prime}\in\{y\in\mathbb{R}^{n}:|y|=1-\sigma^{i}\}$ such that $x^{\prime}\in B$. Thus it holds $|x-x^{\prime\prime}|\leq|x-x^{\prime}|+|x^{\prime}-x^{\prime\prime}|\leq 2\sigma^{i}(1-\sigma)$. It follows that $B_{1}(0)\subset\bigcup_{i=0}^{\infty}\bigcup_{B\in\mathcal{G}_{i}}\bar{B}$.
Now let $i\geq 1$ and $B\in\mathcal{G}_{i}$ centered at $x\in\{y\in\mathbb{R}^{n}:|y|=1-\sigma^{i}\}$. Then we have $B\cap\{y\in\mathbb{R}^{n}:|y|=1-\sigma^{i-1}\}\neq\emptyset$ since $\operatorname{dist}\big(x,\{y\in\mathbb{R}^{n}:|y|=1-\sigma^{i-1}\}\big)=\sigma^{i-1}(1-\sigma)<2\sigma^{i}(1-\sigma)$. Hence there exists $B^{\prime}\in\mathcal{G}_{i-1}$ such that $B^{\prime}\cap B\neq\emptyset$. If we define $B^{\prime\prime}\coloneqq 3B^{\prime}$, then it holds $B\cup B^{\prime}\subset B^{\prime\prime}$ since the radius of $B$ is smaller than that of $B^{\prime}$. Moreover, we have $\operatorname{dist}(B^{\prime\prime},\partial B_{1}(0))=\sigma^{i-1}-6\sigma^{i-1}(1-\sigma)=6K\sigma^{i-1}(1-\sigma)$. Therefore, it holds
$\left|\fint_{B}f-\fint_{B^{\prime\prime}}f\right|\leq\fint_{B}\left|f-\fint_{B^{\prime\prime}}f\right|\leq\left(\frac{3}{\sigma}\right)^{n}\fint_{B^{\prime\prime}}\left|f-\fint_{B^{\prime\prime}}f\right|\leq C(n,K)\,\beta.$
As well, we have
$\left|\fint_{B^{\prime}}f-\fint_{B^{\prime\prime}}f\right|\leq\fint_{B^{\prime}}\left|f-\fint_{B^{\prime\prime}}f\right|\leq 3^{n}\fint_{B^{\prime\prime}}\left|f-\fint_{B^{\prime\prime}}f\right|\leq C(n)\,\beta.$
Consequently, we obtain
$\left|\fint_{B}f-\fint_{B^{\prime}}f\right|\leq C(n,K)\,\beta.$
Then by induction we have
$\left|\fint_{B}f-\fint_{2\tilde{B}}f\right|\leq C(n,K)\,i\beta,\qquad\text{for all }\,i\geq 0\,\text{ and all }B\in\mathcal{G}_{i}.$
Finally, since $\operatorname{dist}(B,\partial B_{1}(0))=\sigma^{i}-2\sigma^{i}(1-\sigma)>2K\sigma^{i}(1-\sigma)$ for any $i\geq 0$ and $B\in\mathcal{G}_{i}$, we have
$\int_{B_{1}(0)}\left|f-\fint_{2\tilde{B}}f\right|$ $\leq\sum_{i=0}^{\infty}\sum_{B\in\mathcal{G}_{i}}\int_{B}\left|f-\fint_{2\tilde{B}}f\right|$
$\leq\sum_{i=0}^{\infty}\sum_{B\in\mathcal{G}_{i}}\left(\int_{B}\left|f-\fint_{B}f\right|+|B|\,\left|\fint_{B}f-\fint_{2\tilde{B}}f\right|\right)$
$\leq C(n,K)\sum_{i=0}^{\infty}\sum_{B\in\mathcal{G}_{i}}|B|\,(i+1)\beta$
$\leq C(n,K)\sum_{i=0}^{\infty}\sigma^{i}(i+1)\beta\leq C(n,K)\beta.$
Lemma III.14 ([162, Theorem 2], [92, Section 3]).
Let $U$ be an open subset of $\mathbb{R}^{n}$, $f\in VMO(U)$, $\varphi\in W^{1,n}_{\textup{loc}}(U,\mathbb{R}^{n})$ be a homeomorphism onto its image. Assume there exists a positive constant $K$ such that
$\det(\nabla\varphi)>K|\nabla\varphi|^{n}\qquad\text{a.e. in }U.$
Then we have $f\circ\varphi^{-1}\in VMO(\varphi(U))$ with
$\beta_{r,\varphi(U)}(f\circ\varphi^{-1})\leq C(K)\,\beta_{r,U}(f)\qquad\text{for all }\,r>0.$
Corollary III.15 ([92]).
Let $U\subset\mathbb{R}^{n}$ be a bounded Lipschitz domain and $f\in VMO(U)\cap L^{\infty}(U)$. Then there exists $\tilde{f}\in VMO(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$ such that $\tilde{f}=f$ on $U$ satisfying the inequality
$\beta_{r,\mathbb{R}^{n}}(\tilde{f})\leq C(U)\big(\beta_{r,U}(f)+r\|f\|_{L^{\infty}(U)}\big),\quad\text{for all }\,r>0.$
III.8 Sobolev–Lorentz spaces
Definition III.16.
Assume $X_{1},X_{2}$ are Banach spaces that continuously embed in the same Hausdorff topological vector space $Z$. Let
$X_{1}+X_{2}\coloneqq\{x_{1}+x_{2}:x_{1}\in X_{1},x_{2}\in X_{2}\}.$
We define norms on $X_{1}+X_{2}$ and $X_{1}\cap X_{2}$ by
$\begin{dcases}\|x\|_{X_{1}\cap X_{2}}\coloneqq\|x\|_{X_{1}}+\|x\|_{X_{2}},\\[8.53581pt] \|x\|_{X_{1}+X_{2}}\coloneqq\inf\{\|x_{1}\|_{X_{1}}+\|x_{2}\|_{X_{2}}:x=x_{1}+x_{2},x_{1}\in X_{1},x_{2}\in X_{2}\}.\end{dcases}$
Equipped with these norms, the sets $X_{1}+X_{2}$ and $X_{1}\cap X_{2}$ are Banach spaces. For brevity we write $L^{p}+L^{q}(\Omega)\coloneqq L^{p}(\Omega)+L^{q}(\Omega)$. The same convention applies to Sobolev–Lorentz spaces.
Definition III.17 (Lorentz spaces).
Let $U\subset\mathbb{R}^{n}$ be a measurable set. Given a measurable function $f\colon U\to\mathbb{R}$, we define the distribution function and the decreasing rearrangement of $f$ as
$d_{f}(\lambda)\coloneqq\mathcal{L}^{n}\bigl\{x\in U:|f(x)|>\lambda\bigr\},\qquad f^{*}(t)\coloneqq\inf\bigl\{\lambda\geq 0:d_{f}(\lambda)\leq t\bigr\}.$
For $1\leq p<\infty$ and $1\leq q\leq\infty$, we define the Lorentz quasinorm
$|f|_{L^{p,q}(U)}\coloneqq\bigl\|t^{\frac{1}{p}}f^{*}(t)\bigr\|_{L^{q}(\mathbb{R}_{+},\,dt/t)}=p^{\frac{1}{q}}\bigl\|\lambda\,d_{f}(\lambda)^{\frac{1}{p}}\bigr\|_{L^{q}(\mathbb{R}_{+},\,d\lambda/\lambda)}.$
The Lorentz space $L^{p,q}(U)$ consists of all measurable $f$ with $|f|_{L^{p,q}(U)}<\infty$. We have (see for instance [13, Chapter 4, Proposition 4.2])
  1. (i)
    $L^{p,p}(U)=L^{p}(U)$,
  2. (ii)
    $L^{p,q}(U)\hookrightarrow L^{p,r}(U)$ if $q<r$,
  3. (iii)
    $L^{p,q}(U)\hookrightarrow L^{s,r}(U)$ if $p>s$ and $\mathcal{L}^{n}(U)<\infty$.
When $p>1$, the Lorentz quasinorm is equivalent to a norm, which we denote by $\|\cdot\|_{L^{p,q}(U)}$.
We record the following form of Hölder’s inequality for Lorentz spaces.
Lemma III.18 ([89, Thm. 4.5]).
Assume $f_{1}\in L^{p_{1},q_{1}}(U)$ and $f_{2}\in L^{p_{2},q_{2}}(U)$, with $p,p_{1},p_{2}\in[1,\infty)$, $q,q_{1},q_{2}\in[1,\infty]$, and $1/p=1/p_{1}+1/p_{2}$, $1/q\leq 1/q_{1}+1/q_{2}$. Then $f_{1}f_{2}\in L^{p,q}(U)$, with
$|f_{1}f_{2}|_{L^{p,q}(U)}\leq C(p_{1},p_{2},q_{1},q_{2})|f_{1}|_{L^{p_{1},q_{1}}(U)}|f_{2}|_{L^{p_{2},q_{2}}(U)}.$
Definition III.19.
Let $k\in\mathbb{N}^{+}$ and $U\subset\mathbb{R}^{n}$ be an open set. For $1<p<\infty$, $1\leq q\leq\infty$, we set
$W^{k,(p,q)}(U)\coloneqq\Big\{f\in L^{p,q}(U)\colon\partial^{\alpha}f\in L^{p,q}(U)\text{ for each }0\leq|\alpha|\leq k\Big\}.$
We also define the negative-order Sobolev and Sobolev–Lorentz spaces by
$W^{-k,p}(U)$ $\coloneqq\bigg\{f\in\mathcal{D}^{\prime}(U)\colon f=\sum_{|\alpha|\leq k}\partial^{\alpha}f_{\alpha}\mbox{ for some }\{f_{\alpha}\}\subset L^{p}(U)\bigg\},$
$W^{-k,(p,q)}(U)$ $\coloneqq\bigg\{f\in\mathcal{D}^{\prime}(U)\colon f=\sum_{|\alpha|\leq k}\partial^{\alpha}f_{\alpha}\mbox{ for some }\{f_{\alpha}\}\subset L^{p,q}(U)\bigg\}.$
The corresponding norms are
$\|f\|_{W^{-k,p}(U)}$ $\coloneqq\inf\bigg\{\sum_{|\alpha|\leq k}\|f_{\alpha}\|_{L^{p}(U)}\colon f=\sum_{|\alpha|\leq k}\partial^{\alpha}f_{\alpha}\bigg\},$
$\|f\|_{W^{-k,(p,q)}(U)}$ $\coloneqq\inf\bigg\{\sum_{|\alpha|\leq k}\|f_{\alpha}\|_{L^{p,q}(U)}\colon f=\sum_{|\alpha|\leq k}\partial^{\alpha}f_{\alpha}\bigg\}.$
When $U$ is bounded, we denote by $W_{0}^{k,p}(U)$ the closure of $C_{c}^{\infty}(U)$ in $W^{k,p}(U)$, equipped with the norm
$\|f\|_{W^{k,p}_{0}(U)}\coloneqq\|\nabla^{k}f\|_{L^{p}(U)}.$
If $1<p<\infty$, $1<q\leq\infty$, and $1/p+1/p^{\prime}=1$, $1/q+1/q^{\prime}=1$, then the same argument as in [133, Section 1.1.15] implies
$W^{-k,(p,q)}(U)=\bigl(W^{k,(p^{\prime},q^{\prime})}_{0}(U)\bigr)^{\!*}.$ (III.20)
Applying estimates for Riesz potentials, we obtain the following embedding results for Sobolev-Lorentz spaces, see for instance [145, Eqs. (1.3)–(1.5)], [13, Ch. 4, Thm. 4.18], and [3, Eq. (1.2.4) and Thm. 3.1.4].
Lemma III.20.
Let $U\subset\mathbb{R}^{n}$ be an open set, $1<p<n$, and $1\leq q\leq\infty$. Suppose $f\colon U\to\mathbb{R}$ is measurable.
  1. (i)
    If $f\in L^{1}(U)$, then $f\in W^{-1,(\frac{n}{n-1},\infty)}(U)$, with
    $\|f\|_{W^{-1,(\frac{n}{n-1},\infty)}(U)}\leq C(n)\|f\|_{L^{1}(U)}.$
  2. (ii)
    If $f\in L^{p,q}(U)$, then $f\in W^{-1,(\frac{np}{n-p},q)}(U)$, with
    $\|f\|_{W^{-1,(\frac{np}{n-p},q)}(U)}\leq C(n,p)\|f\|_{L^{p,q}(U)}.$
IV Weak immersions
We follow the definition as in (I.4).
Definition IV.1.
Let $\Sigma$ be an $n$-dimensional closed smooth manifold with a fixed smooth reference metric $g_{0}$, $m,k\in\mathbb{N}$, $1\leq p\leq\infty$. We say that $\vec{\Phi}$ is a $W^{k,p}$ weak immersion and write $\vec{\Phi}\in W^{k,p}_{\text{imm}}(\Sigma,\mathbb{R}^{m})$ if $\vec{\Phi}\in W^{k,p}\cap W^{1,\infty}(\Sigma,\mathbb{R}^{m})$ and there exists a constant $\Lambda>0$ such that the following holds for a.e. $q\in\Sigma$ and any $X\in T_{q}\Sigma$:
$\Lambda^{-1}\,g_{0}(X,X)\leq|d\vec{\Phi}_{q}(X)|_{\mathbb{R}^{m}}^{2}\leq\Lambda\,g_{0}(X,X).$ (IV.1)
We define $W^{k,p}_{\text{imm}}(B^{n},\mathbb{R}^{m})$ similarly.
In this section, we provide some additional material on the notion of weak immersions developed in the lecture notes [168]. First, we prove the existence of conformal coordinates for a weak immersion in Section IV.1. Then, we prove that weak immersions with VMO derivatives have integral densities and are locally injective in Section IV.3. Finally, we prove that a weak immersion can always be approximated by smooth immersions and that the underlying conformal structure (in the case of surfaces) converges as well in Section IV.4. In all this paper, we will denote $D^{2}$ the unit disk of $\mathbb{R}^{2}$.
IV.1 Existence of conformal coordinates
We start with the definition of conformal coordinates, or isothermal coordinates.
Definition IV.2.
Let $\Sigma$ be a $2$-dimensional closed smooth manifold, $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$, $U\subset\Sigma$ be open. Let $\varphi\colon U\subset\Sigma\rightarrow\varphi(U)\subset\mathbb{R}^{2}$ be a $W^{1,2}_{\textup{loc}}$ bi-Sobolev homeomorphism (i.e. a homeomorphism in $W^{1,2}_{\textup{loc}}$ with a $W^{1,2}_{\textup{loc}}$ inverse). We say $\varphi$ is a (weak) isothermal chart of $\vec{\Phi}$ and $\vec{\Phi}\circ\varphi^{-1}$ is (weakly) conformal if there exists a (Lebesgue) measurable function $\lambda\colon\varphi(U)\rightarrow\mathbb{R}$ such that for $i,j\in\{1,2\}$ there holds
$\partial_{i}(\vec{\Phi}\circ\varphi^{-1})\cdot\partial_{j}(\vec{\Phi}\circ\varphi^{-1})=e^{2\lambda}\delta_{ij}\qquad\text{a.e. in }\varphi(U).$
We define isothermal charts for $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(D^{2},\mathbb{R}^{m})$ similarly.
Throughout this paper, we frequently use the notion of non-smooth metrics and their conformal class.
Definition IV.3.
Let $\Sigma$ be an $n$-dimensional smooth manifold. Let $g$ be a measurable section on the bundle $T^{*}\Sigma\otimes T^{*}\Sigma$, we say $g$ is a metric if $g_{p}$ is symmetric and positive definite on $(T_{p}\Sigma)^{2}$ for a.e. $p\in\Sigma$. Let $g_{1},g_{2}$ be two metrics, we say $g_{1}$ is (weakly) conformal to $g_{2}$ if there exists a measurable function $\lambda\colon\Sigma\rightarrow\mathbb{R}$ such that $g_{1}=e^{\lambda}g_{2}$.
Our main goal in this subsection is to prove the following result:
Theorem IV.4.
Let $\vec{\Phi}\in W_{\textup{imm}}^{2,2}(D^{2},\mathbb{R}^{m})$, then there exists an open neighborhood $U$ of $0\in D^{2}$ with an orientation-preserving weak isothermal chart $\varphi\in W^{2,2}\cap W^{1,\infty}(U,\mathbb{R}^{2})$ of $\vec{\Phi}$. Moreover, we have $\varphi^{-1}\in W^{2,2}\cap W^{1,\infty}(\varphi(U))$.
As a corollary, we obtain the following global result.
Corollary IV.5.
Let $\Sigma$ be a $2$-dimensional closed manifold with a smooth structure $\mathcal{U}$, and $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$. Then we have:
  1. (i)
    there exists a smooth atlas $\mathcal{V}$ on $\Sigma$ consisting of all the weak isothermal charts of $\vec{\Phi}$. If $\Sigma$ is oriented, $\mathcal{V}$ defines a complex structure on $\Sigma$.
  2. (ii)
    For any choice of two charts respectively in $\mathcal{U}$ and $\mathcal{V}$, the transition map between them is a $W^{1,\infty}_{\textup{loc}}\cap W^{2,2}_{\textup{loc}}$ bi-Sobolev homeomorphism.
  3. (iii)
    There exists a Riemannian metric $h$ conformal and smooth on $(\Sigma,\mathcal{V})$ such that $h$ has constant Gaussian curvature $1$, 1-1 or 0, and $g_{\vec{\Phi}}\coloneqq\vec{\Phi}^{*}g_{\text{std}}=e^{2\alpha}h$ for some $\alpha\in W^{2,1}(\Sigma)\subset C^{0}(\Sigma)$.
Proof.
  1. (i)
    We firstly assume $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(D^{2},\mathbb{R}^{m})$ and $\varphi_{1},\varphi_{2}\in W^{1,2}(D^{2},\mathbb{R}^{2})$ be isothermal charts as in Definition IV.2. For $k\in\{1,2\}$, since $\varphi_{k}$ is a $W^{1,2}_{\textup{loc}}$ bi-Sobolev homeomorphism, by standard degree theory we have $\varphi_{k}$ is differentiable a.e. and either $\det(\nabla\varphi_{k})>0$ a.e. or $\det(\nabla\varphi_{k})<0$ a.e., see [82, Theorem 1.7 & Lemma A.28]. Without loss of generality we assume $\det(\nabla\varphi_{k})>0$ a.e. Let $g_{ij}\coloneqq\partial_{i}\vec{\Phi}\cdot\partial_{j}\vec{\Phi}$, then following the computation in [90, Section 1.5.1], since $\varphi_{k}$ is an isothermal chart, we have
    $\frac{\partial_{\bar{z}}\varphi_{k}}{\partial_{z}\varphi_{k}}=\frac{g_{11}-g_{22}+2ig_{12}}{g_{11}+g_{22}+2\sqrt{g_{11}g_{22}-g_{12}^{2}}}\qquad\text{a.e.}$ (IV.2)
    The condition $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(D^{2},\mathbb{R}^{m})$ implies the right hand side has an $L^{\infty}$ norm strictly less than $1$. In particular, $\varphi_{k}$ is quasiconformal, and $\varphi_{1}\circ(\varphi_{2})^{-1}$ is holomorphic on $\varphi_{2}(D^{2})$, see Definition III.6 and [152, Corollary 1.2.8]. Now return to the case when $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$. Let $\mathcal{V}$ be the set of all the weak isothermal charts. Then by the above argument each transition map between elements of $\mathcal{V}$ is either holomorphic or anti-holomorphic, and $\mathcal{V}$ is an atlas by Theorem IV.4.
  2. (ii)
    This follows from $(i)$ and Theorem IV.4.
  3. (iii)
    When $\Sigma$ is orientable, there exists a metric $h$ of constant Gaussian curvature $1$, 1-1 or 0 and conformal under $\mathcal{V}$ by the uniformization theorem for compact Riemann surfaces (see e.g. [93, Section 4.4]). When $\Sigma$ is non-orientable, $h$ is constructed by using a $2$-cover from a closed orientable smooth surface to $\Sigma$.
    The fact $\alpha\in W^{2,1}(\Sigma)$ follows from Corollary III.11. As for the embedding $W^{2,1}(\Sigma)\hookrightarrow C^{0}(\Sigma)$, see for instance [80, Theorem 3.3.4 & 3.3.10].
Remark IV.6.
If $\Sigma$ is oriented and equipped with a metric $g$ such that in any local coordinates the right-hand side in (IV.2) has an $L^{\infty}$ norm less than $1$, then by Lemma III.7 we can still define a $g$-associated conformal structure via the Beltrami equation (IV.2). One major difference is about the regularity of the transition map as in Corollary IV.5 (ii): it may not be Lipschitz, but will be in $W^{1,p}$ for some $p>2$ instead.
Using the conformal structure associated to $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$, the relation (I.6) follows from standard computations, see for instance [168, Theorem 1.6]. By integrating (I.6) with respect to $d\textup{vol}_{h}$ (when $\Sigma$ is orientable), we have (I.7) by the Gauss–Bonnet theorem for the smooth metric $h$. Thus Theorem I.1 is a consequence of Corollary IV.5.
The proof of Theorem IV.4 is similar to the case when $\vec{\Phi}$ is a smooth immersion, while the major difficulty lies in the following analytical lemma, which is a weak inverse function theorem, and its higher-dimensional generalization is proved in [78]. It does not hold if we replace $W^{2,2}$ by some lower regularity from the perspective of Sobolev spaces. One counter example is the map $z\mapsto z^{2}/|z|$. Here we provide a different proof for the $2$-dimensional case. Instead of using properties of BLD maps [130, Lemma 4.6 & Theorem 4.7], we use regularity results for Beltrami equations [42]. Thanks to the embedding $W^{1,2}(D^{2})\hookrightarrow VMO(D^{2})$, we obtain another proof in Corollary IV.13.
Lemma IV.7 ([78, Theorem 1.1]).
Let $\phi\in W^{1,\infty}\cap W^{2,2}(D^{2},\mathbb{R}^{2})$. Assume there exists a constant $c>0$ such that $\det(\nabla\phi)>c$ a.e. in $D$. Then there exists an open neighborhood $U\subset D^{2}$ of 0 such that $\phi$ is open and injective on $U$, and its inverse $\phi^{-1}\in W^{1,\infty}\cap W^{2,2}(\phi(U),\mathbb{R}^{2})$.
Proof.
Without loss of generality assume $\phi(0)=0$ by translation. We denote $h_{ij}\coloneqq\langle\partial_{i}\phi,\partial_{j}\phi\rangle$. Choose $\varphi\in C_{c}^{\infty}(D^{2})$ such that $\varphi(\mathbb{R}^{2})\subset[0,1]$ and $\varphi\equiv 1$ on $D_{1/2}(0)$. Then we regard $\phi$ as a function from $D^{2}\subset\mathbb{C}$ to $\mathbb{C}$ and define
$\mu\coloneqq\varphi\frac{\partial_{\bar{z}}\phi}{\partial_{z}\phi}.$
By direct computation we have
$|\mu|^{2}\leq\frac{|\partial_{\bar{z}}\phi|^{2}}{|\partial_{z}\phi|^{2}}=\frac{h_{11}+h_{22}-2\det(\nabla\phi)}{h_{11}+h_{22}+2\det(\nabla\phi)}\leq\frac{\|\nabla\phi\|^{2}_{L^{\infty}(D^{2})}-c}{\|\nabla\phi\|^{2}_{L^{\infty}(D^{2})}+c}<1\quad\text{ a.e. on $D^{2}$.}$
Hence, it holds $\|\mu\|_{L^{\infty}(\mathbb{C})}<1$. Then we apply Lemma III.7 and obtain a quasiconformal map $f\colon\mathbb{C}\rightarrow\mathbb{C}$ such that $\partial_{\bar{z}}f=\mu\,\partial_{z}f$ almost everywhere.
Let $h\coloneqq\phi\circ f^{-1}$. By the same computation as in [152, Corollary 1.2.8], the map $h$ is holomorphic on $f(D_{1/2}(0))$. Let $k\geq 1$ be the order of zero of $h$ at 0. Then there exists a constant $A\neq 0\in\mathbb{C}$, $0\in U_{1}\subset f(D_{1/2}(0))$, a map $g\in C^{\infty}(\overline{U_{1}},\mathbb{C})$ conformal from the bounded open set $U_{1}$ to $g(U_{1})\subset\mathbb{C}$ with $g(0)=0$ and $g^{\prime}(0)=1$ such that $h=A\,g^{k}$ on $U_{1}$. Let $f_{0}\coloneqq g\circ f$, then since $h=\phi\circ f^{-1}$, we have $\phi=h\circ f=A(g\circ f)^{k}=A\,f_{0}^{k}$ on $f^{-1}(U_{1})$. Since $f$ is quasiconformal and satisfies the equation $\partial_{\bar{z}}f=\mu\,\partial_{z}f$ with $\mu\in W^{1,2}(\mathbb{C})$ of compact support, by Lemma III.8 we have $f\in W^{2,q}_{\textup{loc}}(\mathbb{C})$ for any $q<2$. Then by Sobolev embeddings we have $f\in C^{0,\alpha}_{\textup{loc}}(\mathbb{C})$ for any $\alpha<1$. In particular, $f_{0}=g\circ f\in C^{0,\alpha}(f^{-1}(\overline{U_{1}}))$, and it holds that
$|f_{0}(x)|\leq C(\alpha)|x|^{\alpha},\qquad\text{for all }x\in f^{-1}(\overline{U_{1}})\,\text{ and all }\alpha<1.$ (IV.3)
Now we prove $k=1$. Suppose $k\geq 2$, then there exists a positive constant $C_{1}$ such that
$\sqrt{c}\leq\sqrt{\det(D\phi)}\leq|\partial_{z}\phi|=|kAf_{0}^{k-1}|\,|\partial_{z}f_{0}|\leq C_{1}|f_{0}|\,|\partial_{z}f_{0}|\quad\text{ a.e. on }f^{-1}(\overline{U_{1}}).$ (IV.4)
Fix $\alpha>\frac{1}{2}$. By (IV.3) and (IV.4), there exists a positive constant $C_{2}$ such that for any $x\in f^{-1}(\overline{U_{1}})$, it holds
$|\partial_{z}f_{0}(x)|\geq C_{2}|x|^{-\alpha}.$ (IV.5)
Since $4\alpha>2$, (IV.5) implies
$\int_{f^{-1}(U_{1})}|\partial_{z}f_{0}|^{4}=\infty.$
But by Sobolev embedding again, we have $f\in W^{1,p}_{\textup{loc}}(\mathbb{C})$ for any $p<\infty$, hence we obtain $f_{0}=g\circ f\in W^{1,4}(f^{-1}(U_{1}))$, a contradiction arises.
Let $0\in U\Subset f^{-1}(U_{1})$. Since $k=1$, we have $\phi=A(g\circ f)$ is quasiconformal on $U$, and $\phi^{-1}\in W^{1,\infty}\cap W^{2,2}(\phi(U),\mathbb{R}^{2})$ since $D\phi^{-1}=(D\phi)^{-1}\circ\phi^{-1}$ a.e. on $\phi(U)$. ∎
We are now ready to prove Theorem IV.4
Proof of Theorem IV.4.
We apply the Gram–Schmidt process to the frame $(\partial_{1}\vec{\Phi},\partial_{2}\vec{\Phi})$ and get
$\vec{f}_{1}\coloneqq\frac{\partial_{1}\vec{\Phi}}{|\partial_{1}\vec{\Phi}|},\qquad\text{ and }\qquad\vec{f}_{2}\coloneqq\frac{\partial_{2}\vec{\Phi}-(\partial_{2}\vec{\Phi}\cdot\vec{f}_{1})\vec{f}_{1}}{|\partial_{2}\vec{\Phi}-(\partial_{2}\vec{\Phi}\cdot\vec{f}_{1})\vec{f}_{1}|}.$ (IV.6)
By the definition of weak immersions, there exists a constant $\Lambda>0$ such that for a.e. $p\in D^{2}$ and any $X\in T_{p}(D^{2})$,
$\Lambda^{-1}|X|^{2}\leq|d\vec{\Phi}_{p}(X)|^{2}\leq\Lambda|X|^{2}.$ (IV.7)
In particular, we have $|\partial_{1}\vec{\Phi}|,|\partial_{2}\vec{\Phi}|\in[\Lambda^{-1},\Lambda]$ and $\partial_{2}\vec{\Phi}-(\partial_{2}\vec{\Phi}\cdot\vec{f}_{1})\vec{f}_{1}\in[\Lambda^{-1},\Lambda(1+\Lambda^{2})]$. Hence $(\vec{f}_{1},\vec{f}_{2})\in L^{\infty}\cap W^{1,2}(D^{2},V_{2}({\mathbb{R}}^{m}))$, where $V_{2}({\mathbb{R}}^{m})$ denotes the space of orthonormal 2-frames in ${\mathbb{R}}^{m}$. For $\theta\in W^{1,2}(D^{2},{\mathbb{R}})$, we denote $(\vec{f}^{\,\theta}_{1},\vec{f}^{\,\theta}_{2})$ the rotation by the angle $\theta$ of the original frame $(\vec{f}_{1},\vec{f}_{2})$:
$\vec{f}^{\,\theta}_{1}+i\vec{f}^{\,\theta}_{2}=e^{i\theta}(\vec{f}_{1}+i\vec{f}_{2}),$
then we still have $(\vec{f}^{\,\theta}_{1},\vec{f}^{\,\theta}_{2})\in L^{\infty}\cap W^{1,2}(D^{2},V_{2}(\mathbb{R}^{m}))$. We denote $g=g_{\vec{\Phi}}$. We look now for a rotation of this orthonormal frame realizing the following absolute minimum (for notation, see xvii)
$\begin{array}[]{l}\displaystyle\inf_{\theta\in W^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}|\langle\vec{f}^{\,\theta}_{1},d\vec{f}^{\,\theta}_{2}\rangle|^{2}_{g}\ d\textup{vol}_{g}\\[14.22636pt] \displaystyle=\inf_{\theta\in W^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}\sum_{i,j=1}^{2}g^{ij}\,\langle\vec{f}^{\,\theta}_{1},\partial_{i}\vec{f}^{\,\theta}_{2}\rangle\,\langle\vec{f}^{\,\theta}_{1},\partial_{j}\vec{f}^{\,\theta}_{2}\rangle\,\sqrt{\det g}\ dx^{1}\wedge dx^{2}.\end{array}$ (IV.8)
For $k\in\{1,2\}$, we have
$\langle\vec{f}^{\,\theta}_{1},\partial_{k}\vec{f}^{\,\theta}_{2}\rangle=\partial_{k}\theta+\langle\vec{f}_{1},\partial_{k}\vec{f}_{2}\rangle.$ (IV.9)
Hence the following energy is strictly convex in $W^{1,2}(D^{2},{\mathbb{R}})$:
$E(\theta)\coloneqq\int_{D^{2}}|\langle\vec{f}^{\,\theta}_{1},d\vec{f}^{\,\theta}_{2}\rangle|^{2}_{g}\,d\textup{vol}_{g}=\int_{D^{2}}|d\theta+\langle\vec{f}_{1},d\vec{f}_{2}\rangle|^{2}_{g}\ d\textup{vol}_{g}.$
By a standard application of Mazur’s lemma (see for instance [27, Corollary 3.8]), we obtain a unique $\theta\in W^{1,2}(D^{2},\mathbb{R})$ that achieves the minimum of $E$. It satisfies the Euler–Lagrange equation (see Notation ix for the Hodge star operator $_{g}$):
$\int_{D^{2}}d\phi\wedge*_{g}\left(d\theta+\langle\vec{f}_{1},d\vec{f}_{2}\rangle\right)=0,\qquad\text{for all }\,\phi\in W^{1,2}(D^{2},{\mathbb{R}}).$ (IV.10)
Denote $(\vec{e}_{1},\vec{e}_{2})\coloneqq(\vec{f}^{\,\theta}_{1},\vec{f}^{\,\theta}_{2})$ for this value of $\theta$. From (IV.10) we obtain that
$d\Big(*_{g}\langle\vec{e}_{1},d\vec{e}_{2}\rangle\Big)=0\quad\text{in }\mathcal{D}^{\prime}(D^{2}).$
By the weak Poincaré lemma, there exists $\lambda\in W^{1,2}(D^{2})$ such that
$d\lambda=*_{g}\langle\vec{e}_{1},d\vec{e}_{2}\rangle.$ (IV.11)
Substituting this into (IV.10), we have
$\int_{D^{2}}d\phi\wedge d\lambda=0,\qquad\text{for all }\,\phi\in W^{1,2}(D^{2},{\mathbb{R}}).$
Hence for any $\phi\in C^{\infty}(\overline{D^{2}})$, we have by Stokes theorem
$\int_{\partial D^{2}}\lambda\,d\phi=\int_{D^{2}}d(\lambda\,d\phi)=0.$
It follows that $\lambda\big|_{\mathbb{S}^{1}}$ is constant by fundamental lemma of the calculus of variations [95, Lemma 1.1.1]. Up to adding a constant we can assume $\lambda\in W^{1,2}_{0}(D^{2})$. By applying the operator $_{g}\,d\,*_{g}$ to the equation (IV.11), we obtain
$\Delta_{g}\lambda=*_{g}\,d*_{g}d\lambda=-*_{g}d\langle\vec{e}_{1},d\vec{e}_{2}\rangle\quad\text{in }D^{2}.$
This is written in coordinates as:
$\begin{cases}\displaystyle\,\partial_{i}\left(\sqrt{\det g}\ g^{ij}\,\partial_{j}\lambda\right)=\partial_{2}\vec{e}_{1}\cdot\partial_{1}\vec{e}_{2}-\partial_{1}\vec{e}_{1}\cdot\partial_{2}\vec{e}_{2}\quad&\mbox{ in }D^{2},\\[8.53581pt] \displaystyle\lambda=0\quad&\mbox{ on }\partial D^{2}.\end{cases}$
We deduce from Lemma III.3 that $\lambda\in C^{0}(\overline{D^{2}})\cap W^{1,2}_{0}(D^{2},{\mathbb{R}})$. For $i\in\{1,2\}$, we define pullback of $\vec{e}_{i}$ by $\vec{\Phi}$ and the dual $1$-forms
$\begin{dcases}e_{i}\coloneqq g^{jk}\,\langle\vec{e}_{i},\partial_{j}\vec{\Phi}\rangle\,\dfrac{\partial}{\partial x^{k}}\in L^{\infty}\cap W^{1,2}(D^{2},TD^{2}),\\[5.69054pt] e_{i}^{\ast}\coloneqq\langle\vec{e}_{i},\partial_{j}\vec{\Phi}\rangle\,dx^{j}=\langle\vec{e}_{i},d\vec{\Phi}\rangle\in L^{\infty}\cap W^{1,2}(D^{2},T^{*}D^{2}).\end{dcases}$ (IV.12)
It holds that $e_{i}^{*}(e_{j})=\delta_{ij}$ and $d\vec{\Phi}(e_{i})=\vec{e}_{i}$ for any $i,j\in\{1,2\}$. Since both $e_{i}^{*}$ and $e_{i}$ are in $L^{\infty}\cap W^{1,2}(D^{2})$, by a weak version of Cartan’s magic formula, we have as in the smooth case [168, Equation (1.70)]:
$de_{i}^{*}=-e_{i}^{*}([e_{1},e_{2}])\,e_{1}^{*}\wedge e_{2}^{*}=d\lambda\wedge e_{i}^{\ast}.$ (IV.13)
By direct computation we also have
$d\vec{\Phi}([e_{1},e_{2}])=\pi_{T}(e_{1}(\vec{e}_{2})-e_{2}(\vec{e}_{1}))\in L^{2}(D^{2},\mathbb{R}^{m}),$
where $\pi_{T}$ is the orthogonal projection onto $d\vec{\Phi}(TD^{2})$. Since $\lambda\in L^{\infty}(D^{2})\cap W^{1,2}(D^{2},{\mathbb{R}})$, we deduce from the chain rule that $e^{-\lambda}\in W^{1,2}(D^{2})$ and $d(e^{-\lambda})=-e^{-\lambda}\,d\lambda\in L^{2}(D^{2},T^{*}D^{2})$. Hence the product rule and (IV.13) lead to
$d(e^{-\lambda}\,e_{i}^{*})=0.$
By weak Poincaré lemma, we obtain the existence of a map $\varphi^{i}\in W^{1,\infty}\cap W^{2,2}(D^{2},{\mathbb{R}})$ satisfying
$d\varphi^{i}=e^{-\lambda}\,e_{i}^{*}.$ (IV.14)
Let $\varphi\coloneqq(\varphi^{1},\varphi^{2})\in W^{1,\infty}\cap W^{2,2}(D^{2},{\mathbb{R}^{2}})$. By (IV.12) and (IV.14), we have
$\det(\nabla\varphi)=e^{-2\lambda}\det\big(\langle\vec{e}_{i},\partial_{j}\vec{\Phi}\rangle\big)=e^{-2\lambda}\sqrt{\det(\partial_{i}\vec{\Phi}\cdot\partial_{j}\vec{\Phi})}.$
Since $(e_{1},e_{2})$ is positive and using (IV.7), we obtain that all the eigenvalues of $(\partial_{i}\vec{\Phi}\cdot\partial_{j}\vec{\Phi})_{1\leq i,j\leq 2}$ lie in $\Lambda^{-1},\Lambda]$ and thus, it holds that
$\det(\nabla\varphi)\geq\Lambda^{-1}\,e^{-2\|\lambda\|_{L^{\infty}}}>0.$
It follows that there exists an open neighborhood $U\subset D^{2}$ of 0 such that $\varphi$ is open and injective on $U$, and its inverse $\varphi^{-1}\in W^{1,\infty}\cap W^{2,2}(\varphi(U),\mathbb{R}^{2})$ by Lemma IV.7. We define
$\frac{\partial}{\partial\varphi^{i}}\coloneqq\partial_{i}(\varphi^{-1})^{j}\,\frac{\partial}{\partial x^{j}}\in L^{\infty}\cap W^{1,2}(\varphi(U),\varphi_{*}(TD^{2})),$
where $(\varphi^{-1})^{j}$ is the $j$-th component of $\varphi^{-1}$. Since $e^{\lambda}(d\varphi^{1},d\varphi^{2})$ is orthonormal with respect to $g$, its dual $e^{-\lambda}(\frac{\partial}{\partial\varphi^{1}},\frac{\partial}{\partial\varphi^{2}})$ is also orthonormal with respect to $g$. Therefore, by chain rule, we finally obtain
$\partial_{i}(\vec{\Phi}\circ\varphi^{-1})\cdot\partial_{j}(\vec{\Phi}\circ\varphi^{-1})=d\vec{\Phi}\Big(\frac{\partial}{\partial\varphi^{i}}\Big)\cdot d\vec{\Phi}\Big(\frac{\partial}{\partial\varphi^{j}}\Big)=g\Big(\frac{\partial}{\partial\varphi^{i}},\frac{\partial}{\partial\varphi^{j}}\Big)=e^{2\lambda}\delta_{ij}.$
IV.2 Constant of Wente inequality for a metric and consequence
In 1971, Wente [187] discovered that solutions to elliptic equations of the following form are more regular than expected:
$\begin{cases}\Delta u=\nabla a\cdot\nabla^{\perp}b&\text{in }D^{2},\\ u=0&\text{on }\partial D^{2}.\end{cases}$
Indeed, it holds
$\|u\|_{L^{\infty}(D^{2})}+\|\nabla u\|_{L^{2}(D^{2})}\leq C\|\nabla a\|_{L^{2}(D^{2})}\|\nabla b\|_{L^{2}(D^{2})}.$
Since then, the study of this structure has been intensively studied. Bethuel–Ghidaglia [20] proved that the constant $C$ is independent of the domain. Ge [63], Braraket [11] and Topping [184] studied such equations on surfaces and proved that the constant are independent of the underlying surface. A weighted Wente inequality has also been studied by Da Lio, Gianocca and Rivière in [46, 65]. In this section, we provide a new proof of the independence of the constant with respect to the underlying metric and domain in Proposition IV.8. As an application, we obtain the existence of a Coulomb frame under the smallness of the Gauss curvature in Proposition IV.10.
IV.2.1 Wente inequality
The strategy follows the one introduced in [20, Theorem 1.3] and [49, Lemma A.1], relying on a direct study of the Green kernel.
Proposition IV.8.
Let $\Omega\subset\mathbb{R}^{2}$ be a smooth simply connected open set. Let $g$ be a smooth metric on $\Omega$. Let $a,b\in C^{\infty}(\Omega)$ and $u\in C^{\infty}(\Omega)$ be the solution to
$\left\{\begin{aligned} &-\Delta_{g}u=*_{g}(da\wedge db)\quad&&\text{in }\Omega,\\ &u=0\quad&&\text{on }\partial\Omega.\end{aligned}\right$. (IV.15)
Then, we have the estimate
$\begin{cases}\|u\|_{L^{\infty}(\Omega)}\leq 18\,\|da\|_{L^{2}(\Omega,g)}\|db\|_{L^{2}(\Omega,g)},\\[5.69054pt] \|du\|_{L^{2}(\Omega,g)}\leq 3\sqrt{2}\,\|da\|_{L^{2}(\Omega,g)}\|db\|_{L^{2}(\Omega,g)}.\end{cases}$
Remark IV.9.
  • The first equation of (IV.15) can also be written as
    $-\partial_{i}(g^{ij}\sqrt{\det g}\,\partial_{j}u)=\nabla a\cdot\nabla^{\perp}b.$
  • We can reduce the regularity on $g$, that is to say we can assume $g$ to be a metric of class $L^{\infty}(\Omega)$ such that there exists $C>0$ satisfying $C^{-1}g_{\text{std}}\leq g\leq Cg_{\text{std}}$. By using a mollifier argument, there exists a sequence $(g_{k})_{k\in\mathbb{N}}$ of smooth metrics converging to $g$ in $L^{\infty}$ and almost everywhere such that $C^{-1}g_{\text{std}}\leq g_{k}\leq Cg_{\text{std}}$. Hence, we can pass to the limit in Proposition IV.8.
  • By density, we can assume $a,b\in W^{1,2}$.
Proof.
Up to dividing $a$ and $b$ by $\|da\|_{L^{2}(\Omega,g)}$ and $\|db\|_{L^{2}(\Omega,g)}$ respectively, we can assume that $\|da\|_{L^{2}(\Omega,g)}=\|db\|_{L^{2}(\Omega,g)}=1$. Given $p\in\Omega$, let $G_{p}$ be the solution to
$\begin{cases}-\Delta_{g}G_{p}=\delta_{p}&\text{in }\Omega,\\ G_{p}=0&\text{on }\partial\Omega.\end{cases}$
By the maximum principle, it holds $G_{p}>0$ on $\Omega$, see for instance [9, Chapter 4, Section 2, Theorem 4.17]. Given $0\leq\alpha<\beta$, we define
$\omega_{p}(\alpha,\beta)\coloneqq\{x\in\Omega:\alpha\leq G_{p}(x)\leq\beta\},$ $\Omega_{p}(\alpha)\coloneqq\{x\in\Omega:\alpha\leq G_{p}(x)\}.$
By elliptic regularity, it holds $G_{p}\in C^{\infty}(\Omega\setminus\{p\})$. Thanks to Sard’s theorem (see for instance [108, Chapter 6, Theorem 6.10]), for almost every $\alpha>0$, the set $G_{p}^{-1}(\{\alpha\})$ is a regular curve in $\Omega$. We orient this curve with the vector field
$\nu^{i}=-\frac{g^{ij}\partial_{j}G_{p}}{|dG_{p}|_{g}}.$
Hence we obtain by integration by parts
$\int_{G_{p}^{-1}(\{\alpha\})}|dG_{p}|_{g}\,d\textup{vol}_{g}=-\int_{G_{p}^{-1}(\{\alpha\})}\partial_{\nu}G_{p}\,d\textup{vol}_{g}=-\int_{\Omega_{p}(\alpha)}\Delta_{g}G_{p}\,d\textup{vol}_{g}=1.$
Using the coarea formula, see for instance [33, Chapter IV, Theorem 1], we obtain for almost every $0\leq\alpha<\beta$:
$\int_{\omega_{p}(\alpha,\beta)}|dG_{p}|^{2}_{g}\,d\textup{vol}_{g}=\int_{\alpha}^{\beta}\int_{G_{p}^{-1}(\{s\})}|dG_{p}|_{g}\,d\textup{vol}_{g}\,ds=\beta-\alpha.$ (IV.16)
Now, the solution $u$ to (IV.15) is given by
$u(p)=\int_{\Omega}G_{p}(x)\,da\wedge db.$ (IV.17)
For each $n\in\mathbb{N}$, there exists $\alpha_{n}\in[n,n+1]$ such that
$\int_{G_{p}^{-1}(\{\alpha_{n}\})}|da|_{g}+|db|_{g}\,d\textup{vol}_{g}$ $\leq 2\int_{n}^{n+1}\int_{G_{p}^{-1}(s)}(|da|_{g}+|db|_{g})\,|dG_{p}|_{g}\,d\textup{vol}_{g}$
$\leq 2\int_{\omega_{p}(n,n+1)}(|da|_{g}+|db|_{g})\,|dG_{p}|_{g}\,d\textup{vol}_{g}.$ (IV.18)
From (IV.17), it holds
$u(p)$ $=\sum_{n\in\mathbb{N}}\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}G_{p}(x)\,*_{g}(da\wedge db)\,d\textup{vol}_{g}$
$=\sum_{n\in\mathbb{N}}\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}[G_{p}(x)-n]\,*_{g}(da\wedge db)\,d\textup{vol}_{g}$ (IV.19)
$\qquad+\sum_{n\in\mathbb{N}}n\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}*_{g}(da\wedge db)\,d\textup{vol}_{g}.$ (IV.20)
We estimate the term (IV.19) by Cauchy–Schwarz:
$\left|\sum_{n\in\mathbb{N}}\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}[G_{p}(x)-n]\,*_{g}(da\wedge db)\,d\textup{vol}_{g}\right|$ $\leq 2\sum_{n\in\mathbb{N}}\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}|da|_{g}\,|db|_{g}\,d\textup{vol}_{g}$
$\leq 2\int_{\Omega}|da|_{g}\,|db|_{g}\,d\textup{vol}_{g}.$ (IV.21)
We estimate the term (IV.20) by integration by parts:
$\sum_{n\in\mathbb{N}}n\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}*_{g}(da\wedge db)\,d\textup{vol}_{g}$
$=$ $\ \sum_{n\in\mathbb{N}}n\left(\int_{G_{p}^{-1}(\alpha_{n+1})}(\partial_{\tau}b)a\,d\textup{vol}_{g}-\int_{G_{p}^{-1}(\alpha_{n})}(\partial_{\tau}b)a\,d\textup{vol}_{g}\right)$
$=$ $-\sum_{n\in\mathbb{N}^{*}}\int_{G_{p}^{-1}(\alpha_{n})}(\partial_{\tau}b)a\,d\textup{vol}_{g}.$ (IV.22)
For each $n\in\mathbb{N}^{*}$, we decompose $G_{p}^{-1}(\{\alpha_{n}\})=\bigcup_{i\in I_{n}}\gamma^{i}_{n}$, where every $\gamma^{i}_{n}$ is a closed embedded curve in $\Omega$. Then, it holds
$\left|\int_{G_{p}^{-1}(\alpha_{n})}(\partial_{\tau}b)a\,d\textup{vol}_{g}\right|$ $=\left|\sum_{i\in I_{n}}\int_{\gamma^{i}_{n}}(\partial_{\tau}b)a\,d\textup{vol}_{g}\right|$
$=\left|\sum_{i\in I_{n}}\int_{\gamma^{i}_{n}}(\partial_{\tau}b)\left(a-\fint_{\gamma^{i}_{n}}a\,d\textup{vol}_{g}\right)\,d\textup{vol}_{g}\right|$
$\leq\sum_{i\in I_{n}}\left(\int_{\gamma^{i}_{n}}|\partial_{\tau}b|\,d\textup{vol}_{g}\right)\left(\int_{\gamma^{i}_{n}}|da|_{g}\,d\textup{vol}_{g}\right)$
$\leq\left(\int_{G_{p}^{-1}(\{\alpha_{n}\})}|db|_{g}\,d\textup{vol}_{g}\right)\left(\int_{G_{p}^{-1}(\{\alpha_{n}\})}|da|_{g}\,d\textup{vol}_{g}\right).$
Coming back to (IV.22), we obtain
$\left|\sum_{n\in\mathbb{N}}n\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}*_{g}(da\wedge db)\,d\textup{vol}_{g}\right|\leq\sum_{n\in\mathbb{N}^{*}}\left(\int_{G_{p}^{-1}(\{\alpha_{n}\})}|db|_{g}\,d\textup{vol}_{g}\right)\left(\int_{G_{p}^{-1}(\{\alpha_{n}\})}|da|_{g}\,d\textup{vol}_{g}\right).$
Using the choice of the $\alpha_{n}$ in (IV.18), we obtain
$\left|\sum_{n\in\mathbb{N}}n\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}*_{g}(da\wedge db)\,d\textup{vol}_{g}\right|$
$\leq$ $\ 4\sum_{n\in\mathbb{N}^{*}}\left(\int_{\omega_{p}(n,n+1)}(|da|_{g}+|db|_{g})\,|dG_{p}|_{g}\,d\textup{vol}_{g}\right)^{2}$
$\leq$ $\ 4\sum_{n\in\mathbb{N}^{*}}\|dG_{p}\|_{L^{2}(\omega_{p}(n,n+1),g)}^{2}\,\left(\|da\|_{L^{2}(\omega_{p}(n,n+1),g)}+\|db\|_{L^{2}(\omega_{p}(n,n+1),g)}\right)^{2}.$
Thanks to (IV.16), we obtain
$\left|\sum_{n\in\mathbb{N}}n\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}*_{g}(da\wedge db)\,d\textup{vol}_{g}\right|$
$\leq 4\sum_{n\in\mathbb{N}^{*}}\left(\|da\|_{L^{2}(\omega_{p}(n,n+1),g)}+\|db\|_{L^{2}(\omega_{p}(n,n+1),g)}\right)^{2}$
$\leq 8\,\|da\|_{L^{2}(\Omega,g)}^{2}+8\|db\|_{L^{2}(\Omega,g)}^{2}=16.$ (IV.23)
We used Cauchy–Schwarz and the equalities $\|da\|_{L^{2}(\Omega,g)}=\|db\|_{L^{2}(\Omega,g)}=1$ for the last estimate. Coming back to (IV.17), together with (IV.21) and (IV.23), we deduce that
$\|u\|_{L^{\infty}(\Omega)}\leq 18.$
We multiply (IV.15) by $u$ and integrate by parts:
$\int_{\Omega}|du|^{2}_{g}\,d\textup{vol}_{g}$ $=-\int_{\Omega}u\,\Delta_{g}u\,d\textup{vol}_{g}$
$\leq\|u\|_{L^{\infty}(\Omega)}\int_{\Omega}|da|_{g}\,|db|_{g}\,d\textup{vol}_{g}\leq 18.$
IV.2.2 Estimate of the moving frame
It has been proved in [111] that a Wente-type inequality leads to an estimate on Coulomb frames. Using IV.8, we obtain the following result.
Proposition IV.10.
Let $\Omega\subset\mathbb{R}^{2}$ be a smooth simply connected open set. Consider a weak immersion $\vec{\Phi}:\Omega\to\mathbb{R}^{m}$ such that, with $g=g_{\vec{\Phi}}$,
$\int_{\Omega}|K_{g}|\,d\textup{vol}_{g}\leq\frac{1}{36}.$
Then, there exists an orthonormal frame $(\vec{e}_{1},\vec{e}_{2})\in W^{1,2}(\Omega,g)$ such that
$\int_{\Omega}|d\vec{e}_{1}|^{2}_{g}+|d\vec{e}_{2}|^{2}_{g}\,d\textup{vol}_{g}\leq\frac{3}{2}\int_{\Omega}|d\vec{n}|^{2}_{g}\,d\textup{vol}_{g}.$
Proof.
Thanks to [80, Lemma 4.1.3], there exists a frame $(\vec{e}_{1},\vec{e}_{2})\in W^{1,2}(\Omega,g)$ such that
$\begin{cases}d^{*_{g}}\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}=0&\text{in }\Omega,\\ \left\langle\partial_{\nu}\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}=0&\text{on }\partial\Omega,\end{cases}$ (IV.24)
where $\nu$ is the unit outward-pointing normal for $g$. Moreover, by decomposition each $d\vec{e}_{i}$ in the sum of its tangential part along $\vec{\Phi}$ and its normal part, we obtain the following estimate
$\int_{\Omega}|d\vec{e}_{1}|^{2}_{g,\mathbb{R}^{m}}+|d\vec{e}_{2}|^{2}_{g,\mathbb{R}^{m}}\,d\textup{vol}_{g}\leq 2\int_{\Omega}|\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}|^{2}_{g}\,d\textup{vol}_{g}+\int_{\Omega}|d\vec{n}|^{2}_{g}\,d\textup{vol}_{g}.$ (IV.25)
By Poincaré Lemma and (IV.24), there exists $\lambda\in W^{1,2}(\Omega,g)$ such that
$\begin{cases}d\lambda=*_{g}\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}&\text{in }\Omega,\\ \lambda=0&\text{on }\partial\Omega.\end{cases}$ (IV.26)
Indeed, the boundary condition in (IV.24) implies that $\lambda$ is constant on $\partial\Omega$. Since the first relation of (IV.26) depends only on $d\lambda$ and not on $\lambda$ itself, we can add a constant in order to get the second relation of (IV.26). We apply $d^{*_{g}}$ to (IV.26) to obtain
$\left\{\begin{aligned} &-\Delta_{g}\lambda=-*_{g}d\left(\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}\right)\quad\text{in }\Omega,\\ &\lambda=0\quad\text{on }\partial\Omega.\end{aligned}\right$. (IV.27)
Thanks to Proposition IV.8, it holds
$\begin{cases}\|\lambda\|_{L^{\infty}(\Omega)}\leq 18\,\|d\vec{e}_{1}\|_{L^{2}(\Omega,g)}\,\|d\vec{e}_{2}\|_{L^{2}(\Omega,g)},\\[5.69054pt] \|d\lambda\|_{L^{2}(\Omega,g)}\leq 3\sqrt{2}\,\|d\vec{e}_{1}\|_{L^{2}(\Omega,g)}\,\|d\vec{e}_{2}\|_{L^{2}(\Omega,g)}.\end{cases}$
Denote $\varepsilon_{0}\coloneqq\|K_{g}\|_{L^{1}(\Omega,g)}$. We multiply the first relation of (IV.27) by $\lambda$ and integrate by parts:
$\int_{\Omega}|d\lambda|^{2}_{g}\,d\textup{vol}_{g}$ $\leq\|\lambda\|_{L^{\infty}(\Omega)}\int_{\Omega}\left|d\left(\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}\right)\right|_{g}\,d\textup{vol}_{g}$
$\leq\|\lambda\|_{L^{\infty}(\Omega)}\int_{\Omega}|K_{g}|\,d\textup{vol}_{g}$
$\leq 18\,\varepsilon_{0}\,\|d\vec{e}_{1}\|_{L^{2}(\Omega,g)}\,\|d\vec{e}_{2}\|_{L^{2}(\Omega,g)}.$
Thanks to (IV.26), we obtain
$\int_{\Omega}|\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}|^{2}_{g}\,d\textup{vol}_{g}$ $\leq 18\,\varepsilon_{0}\,\|d\vec{e}_{1}\|_{L^{2}(\Omega,g)}\|d\vec{e}_{2}\|_{L^{2}(\Omega,g)}\leq 9\,\varepsilon_{0}\left(\|d\vec{e}_{1}\|_{L^{2}(\Omega,g)}^{2}+\|d\vec{e}_{2}\|_{L^{2}(\Omega,g)}^{2}\right).$
Using (IV.25), we obtain
$\int_{\Omega}|\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}|^{2}_{g}\,d\textup{vol}_{g}\leq 9\,\varepsilon_{0}\left(2\int_{\Omega}|\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}|^{2}_{g}\,d\textup{vol}_{g}+\int_{\Omega}|d\vec{n}|^{2}_{g}\,d\textup{vol}_{g}\right).$
If $\varepsilon_{0}\leq\frac{1}{36}$, we obtain
$\frac{1}{2}\int_{\Omega}|\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}|^{2}_{g}\,d\textup{vol}_{g}\leq\frac{1}{4}\int_{\Omega}|d\vec{n}|^{2}_{g}\,d\textup{vol}_{g}.$
We conclude thanks to (IV.25). ∎
IV.3 Local injectivity and integral densities
Definition IV.11.
Let $\Sigma$ be an orientable $n$-dimensional closed smooth manifold and consider $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$. For any measurable set $E\subset\mathbb{R}^{m}$, we define the volume of $E\cap\vec{\Phi}(\Sigma)$ as
$\textup{Vol}(E)\coloneqq\int_{\Sigma}\mathbf{1}_{\vec{\Phi}^{-1}(E)}\,d\textup{vol}_{g_{\vec{\Phi}}}.$
For $x\in\mathbb{R}^{m}$, we define the density $\theta_{x}\in\mathbb{N}_{0}$ at $x$ by the following limit (whenever it exists)
$\theta_{x}\coloneqq\lim_{r\rightarrow 0^{+}}\frac{\textup{Vol}(\vec{\Phi}(\Sigma)\cap B_{r}(x))}{r^{n}|B^{n}|}.$
In this section, we prove that weak immersions with VMO derivatives are locally injective and have a well-defined density, which is an integer. For notation on VMO functions, see Section III.7.
Theorem IV.12.
Let $f\in W^{1,\infty}(B^{n},\mathbb{R}^{m})$. Suppose that there exist a function $\omega\colon(0,\infty)\rightarrow[0,\infty]$ with $\lim_{r\rightarrow 0}\omega(r)=0$ and a constant $\Lambda>0$ such that
$\begin{dcases}\beta_{r,B^{n}}(\nabla f)\leq\omega(r),\quad&\text{for all }\,r>0,\\[8.53581pt] \Lambda^{-1}|v|\leq|df_{p}(v)|\leq\Lambda|v|,\quad&\text{for a.e. }p\in B^{n}\text{ and all }\,v\in T_{p}(B^{n}).\end{dcases}$
In particular, $\nabla f\in VMO(B^{n})$. Then there exist a constant $C>0$ and an open neighborhood $U$ of the origin, both depending only on $\Lambda,\,\omega,\,n,\,m$ such that
$|f(x)-f(y)|\geq C|x-y|,\qquad\text{for all }\,x,y\in U.$
Proof.
Let $V$ be the set of full-rank $m\times n$ matrices, then $V$ is open in the space of $m\times n$ matrices. By the definition of weak immersion, there exists a compact subset $K$ of $V$ such that $\nabla f\in K$ a.e. in $B^{n}$. Hence there exists $\varepsilon>0$ such that the set $K^{\prime}\coloneqq\{M\in V\colon\operatorname{dist}(M,K)\leq\varepsilon\}$ is compact and contained in $V$. It follows that there exists a constant $C_{0}>0$ such that for any $v\in\mathbb{R}^{n}$, $M\in K^{\prime}$, it holds
$|Mv|\geq C_{0}|v|.$ (IV.28)
Moreover, $C_{0},\,\varepsilon$, and $K$ can be chosen to depend on $\Lambda,\,\omega,\,n,\,m$ only. We choose $r_{0},\varepsilon_{0}\in(0,1/8)$ small enough (to be determined later) depending on $\Lambda,\,\omega,\,n,\,m$ only such that for every ball $B$ of radius $\leq 4r_{0}$ we have $\fint_{B}\left|\nabla f-\fint_{B}\nabla f\right|<\min(\varepsilon,\varepsilon_{0})$. Now fix $x_{0},y_{0}\in B_{r_{0}}(0)$, we denote $r\coloneqq|x_{0}-y_{0}|$, $B^{\prime}\coloneqq B_{2r}(x_{0})$, and
$M_{0}\coloneqq\fint_{B^{\prime}}\nabla f.$
Since $2r<4r_{0}$ and $\nabla f\in K$ a.e., we have
$\fint_{B^{\prime}}|\nabla f-M_{0}|<\min(\varepsilon,\varepsilon_{0}),$
and hence $M_{0}\in K^{\prime}$. Now let $f_{0}(x)\coloneqq f(x)-M_{0}x$, then apply Morrey’s inequality [58, Theorem 5.6.4] in the case $p=2n$ and John–Nirenberg inequality [77, Theorem 3.5], for $\varepsilon_{0}$ small enough we have
$\begin{aligned} |f_{0}(x_{0})-f_{0}(y_{0})|&\leq C_{1}\mkern 2.0mur^{1/2}\left(\int_{B^{\prime}}|\nabla f_{0}|^{2n}\right)^{\frac{1}{2n}}\\ &=C_{1}\mkern 2.0mur^{1/2}\left(\int_{B^{\prime}}|\nabla f-M_{0}|^{2n}\right)^{\frac{1}{2n}}\\[5.69054pt] &\leq C_{2}\mkern 2.0mur\mkern 2.0mu\varepsilon_{0}=2^{-1}C_{0}\mkern 2.0mur\\[8.53581pt] &=2^{-1}C_{0}|x_{0}-y_{0}|,\end{aligned}$ (IV.29)
where $C_{1},\,C_{2}$ are positive constants depending only on $n$, and we choose $\varepsilon_{0}=(2C_{2})^{-1}C_{0}$. Therefore, by (IV.28) and (IV.29) we have
$|f(x_{0})-f(y_{0})|\geq|M_{0}(x_{0}-y_{0})|-|f_{0}(x_{0})-f_{0}(y_{0})|\geq\frac{C_{0}}{2}|x_{0}-y_{0}|.$
By using the embedding $W^{1,(n,q)}(B^{n})\hookrightarrow VMO(B^{n})$ for $1\leq q<\infty$ (see Definition III.19 and [41, Theorem 6.1(iii)]), we obtain the following corollary.
Corollary IV.13.
Let $m,n\in\mathbb{N}$ with $2\leq n\leq m$, and let $1\leq q<\infty$. Suppose $f\in W^{1,\infty}_{\textup{imm}}(B^{n},\mathbb{R}^{m})$ and $\nabla^{2}f\in L^{n,q}(B^{n})$, then there exist a constant $C>0$ and an open neighborhood $U$ of the origin such that
$|f(x)-f(y)|\geq C|x-y|,\qquad\text{for all }\,x,y\in U.$
Remark IV.14.
When $m>n$, Corollary IV.13 is sharp in the sense that there exists $f\in W_{\textup{imm}}^{1,\infty}(B^{n},\mathbb{R}^{m})$ with $\nabla^{2}f\in L^{n,\infty}(B^{n})$ but $f$ is not injective in any neighborhood of the origin. We construct as follows.
Without loss of generality assume $m=n+1$. Define $\tilde{f}_{0}\in C^{\infty}(\mathbb{S}^{n-1},\mathbb{R}^{n})$ by
$\tilde{f}_{0}(x_{1},\dots,x_{n})\coloneqq\Big(\big(2+\cos(\pi x_{n})\big)x_{1},\dots,\big(2+\cos(\pi x_{n})\big)x_{n-1},\mkern 1.0mu\sin(\pi x_{n})\Big).$
Then $\tilde{f}_{0}$ is a non-injective smooth immersion from $\mathbb{S}^{n-1}$ to $\mathbb{R}^{n}$. Hence, there exists a non-injective smooth immersion $\tilde{f}\colon\mathbb{S}^{n-1}\to\mathbb{S}^{n}$. Now for $x\in B^{n}$, we define
$f(x)=\begin{dcases}|x|\mkern 2.0mu\tilde{f}\Big(\frac{x}{|x|}\Big),&\text{if }x\neq 0,\\ 0,&\text{otherwise}.\end{dcases}$
Then $f\in W_{\textup{imm}}^{1,\infty}(B^{n},\mathbb{R}^{n+1})$, and there exists a constant $C>0$ such that $|\nabla^{2}f(x)|\leq C|x|^{-1}$ for all $x\in B^{n}\setminus\{0\}$. This implies that $\nabla^{2}f\in L^{n,\infty}(B^{n})$, but $f$ is not injective in any neighborhood of 0.
In the special case $m=n$, a sharp theorem states that any $W^{2,2}$ weak immersion (also called “BLD maps” in this case) is a local homeomorphism everywhere, see [78, Theorem 1.1]. However, when $n\geq 3$, it remains open whether one can replace $W^{2,2}$ by $W^{2,(2,q)}$ for any $2<q<\infty$.
In the proof of [168, Corollary 3.43], it is mentioned that in the case $n=2$, if we assume in addition that $\vec{\Phi}\in W^{2,2}(\Sigma)$, then for any $x\in\mathbb{R}^{m}$, $\theta_{x}$ exists and is in $\mathbb{N}_{0}$. Here we prove a stronger result.
Theorem IV.15.
Let $\Sigma$ be an $n$-dimensional closed smooth manifold, $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$ with $d\vec{\Phi}\in VMO(\Sigma)$, then we have
$\theta_{y}=\text{Card}(\vec{\Phi}^{-1}(y)),\qquad\text{for all }\,y\in\mathbb{R}^{m}.$ (IV.30)
Proof.
By Theorem IV.12, for any point $y\in{\mathbb{R}}^{m}$, $\vec{\Phi}^{-1}(\{y\})$ is at most finite. Thus it suffices to show that for a bi-Lipschitz map $f\in W^{1,\infty}_{\textup{imm}}(B^{n},\mathbb{R}^{m})$ with $\nabla f\in VMO(B^{n})$, $f(0)=0$, there holds
$\lim_{r\rightarrow 0^{+}}\frac{\int_{f^{-1}(B_{r}(0))}1\,d\textup{vol}_{g_{f}}}{r^{n}|B^{n}|}=1,$ (IV.31)
where $g_{f}\coloneqq f^{*}g_{\text{std}}$. For $r\in(0,1)$, we define $f_{r}\colon B^{n}\rightarrow\mathbb{R}^{m}$ by
$f_{r}(w)\coloneqq\frac{f(r\omega)}{r}.$ (IV.32)
Then $\{f_{r}\}$ is uniformly bounded in $W^{1,\infty}(B^{n})$. Moreover, we have the following limit
$\fint_{B^{n}}\left|\nabla f_{r}-\fint_{B^{n}}\nabla f_{r}\right|=\fint_{B_{r}(0)}\left|\nabla f-\fint_{B_{r}(0)}\nabla f\right|\xrightarrow[r\to 0^{+}]{}0.$ (IV.33)
Since $f$ is bi-Lipschitz on $B^{n}$, there exists a positive constant $C_{0}$ such that
$|x|\leq C_{0}|f(x)|,\qquad\text{ for all }x\in B^{n}.$ (IV.34)
Let $\{r_{k}\}_{k=1}^{\infty}\subset\mathbb{R}_{+}$ be an arbitrary sequence with $r_{k}\rightarrow 0^{+}$, then there exists a subsequence $\{r_{k_{i}}\}_{i=1}^{\infty}$ and an $m\times n$ matrix $M$ such that
$\fint_{B^{n}}\nabla f_{C_{0}r_{k_{i}}}\xrightarrow[i\to\infty]{}M.$ (IV.35)
Now it suffices to prove
$\lim_{i\rightarrow\infty}\frac{\int_{f^{-1}\left(B_{r_{k_{i}}}(0)\right)}1\,d\textup{vol}_{g_{f}}}{r_{k_{i}}^{n}|B^{n}|}=1.$
Denote $\tilde{f}_{i}\coloneqq f_{C_{0}r_{k_{i}}}$. By (IV.33), (IV.35) we have
$\lim_{i\rightarrow\infty}\int_{B^{n}}\left|\nabla\tilde{f}_{i}-M\right|=0.$ (IV.36)
Let $\tilde{f}\colon B^{n}\rightarrow\mathbb{R}^{m}$ defined by $\tilde{f}(\omega)=M\omega$. Since $\{\tilde{f}_{i}\}$ is bounded in $W^{1,\infty}(B^{n})$ and $\tilde{f}_{i}(0)=0$ for any $i$, by (IV.36) we obtain that
$\tilde{f}_{i}\xrightarrow[i\rightarrow\infty]{}\tilde{f}\quad\text{in }W^{1,p}(B^{n}),\qquad\text{for all }1\leq p<\infty.$
By (IV.34) and definition of $f_{r}$, for any $\omega\in B^{n}$, $i\in\mathbb{N}$, we have
$\left|\omega\right|\leq C_{0}\big|\tilde{f}_{i}(\omega)\big|.$
Hence for any $\omega\in B^{n}$, we have
$|\omega|\leq C_{0}\big|\tilde{f}(\omega)\big|=C_{0}|M\,\omega|.$ (IV.37)
In particular, this implies $M$ has full rank. By (IV.34) we have $f^{-1}\big(B_{r_{k_{i}}}(0)\big)\subset B_{C_{0}r_{k_{i}}}(0)$. We can then apply change of variables formula to get
$(C_{0}r_{k_{i}})^{-n}\int_{f^{-1}\left(B_{r_{k_{i}}}(0)\right)}1\,d\textup{vol}_{g_{f}}=\int_{\tilde{f}_{i}^{-1}\big(B_{C_{0}^{-1}}(0)\big)}\sqrt{\det\big(\nabla\tilde{f}_{i}^{T}\nabla\tilde{f}_{i}\big)}.$ (IV.38)
By the definition of weak immersions, there exists a constant $C>0$ independent of $i$ such that for all $i\in\mathbb{N}$,
$C^{-1}\leq\det\big(\nabla\tilde{f}_{i}^{T}\nabla\tilde{f}_{i}\big)\leq C\qquad\text{a.e. on }B^{n}.$
And since $\nabla\tilde{f}_{i}\rightarrow M$ in $L^{p}(B^{n})$ for any $p\in[1,\infty)$, we have
$\left\|\sqrt{\det\big(\nabla\tilde{f}_{i}^{T}\nabla\tilde{f}_{i}\big)}-\sqrt{\det\big(M^{T}M\big)}\right\|_{L^{1}(B^{n})}\xrightarrow[i\rightarrow\infty]{}0.$ (IV.39)
By Sobolev embedding we have $\tilde{f}_{i}\rightarrow\tilde{f}$ in $C^{0}(\overline{B^{n}})$  [^6] In general, $f_{r}$ defined in (IV.32) does not has a $C^{0}$ limit as $r\to 0$ under the same assumptions. For instance, we may take $\bar{f}(x_{1},x_{2})\coloneqq\begin{cases}0&\text{if }(x_{1},x_{2})=0,\\[5.69054pt] \Big(x_{1},x_{2},100^{-1}\,x_{1}\cos\big(\log(1-\log(x_{1}^{2}+x_{2}^{2}))\big)\Big)&\text{otherwise}.\end{cases}$ Then $\bar{f}\in W^{2,2}_{\textup{imm}}(D^{2},\mathbb{R}^{3})$ hence $\nabla\bar{f}\in VMO(D^{2})$, but for any $\omega\in\{(x_{1},x_{2})\in D^{2}:x_{1}\neq 0\}$, $\bar{f}_{r}(\omega)\coloneqq r^{-1}\bar{f}(r\omega)$ diverges as $r\rightarrow 0$. Moreover, we compute the contingent cone [8, Definition 4.1.1] and Clarke tangent cone [8, Definition 4.1.5] of $\bar{f}(D^{2})$ at 0: $T_{\bar{f}(D^{2})}(0)=\big\{(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}:|x_{3}|\leq|x_{1}|\big\},\qquad C_{\bar{f}(D^{2})}(0)=\{0\}\times\mathbb{R}\times\{0\}.$ Neither of the two tangent cones above is a $2$-plane in $\mathbb{R}^{3}$. However, if we assume $f\colon D^{2}\rightarrow\mathbb{R}^{3}$ is a $C^{1}$ embedding, then it holds that $T_{f(D^{2})}(0)=C_{f(D^{2})}(0)=\text{Im}(\nabla f(0)).$ , and since $\mathcal{L}^{n}\left(\tilde{f}^{-1}\big(\partial B_{C_{0}^{-1}}(0)\big)\right)=0$, by [55, Theorem 3.2.11] we obtain
$\mathcal{L}^{n}\left(\tilde{f}_{i}^{-1}\big(B_{C_{0}^{-1}}(0)\big)\right)\rightarrow\mathcal{L}^{n}\left(\tilde{f}^{-1}\big(B_{C_{0}^{-1}}(0)\big)\right).$ (IV.40)
Since $\tilde{f}(\omega)=M\omega$, we have
$\begin{aligned} \mathcal{L}^{n}\left(\tilde{f}^{-1}\big(B_{C_{0}^{-1}}(0)\big)\right)&=C_{0}^{-n}\mathcal{L}^{n}\left\{\omega\in\mathbb{R}^{n}:\langle M^{T}M\omega,\omega\rangle\leq 1\right\}.\end{aligned}$ (IV.41)
Since $M^{T}M$ is positive definite, there exist $Q\in SO(n)$ and $D=\text{diag}(\lambda_{1},\dots,\lambda_{n})$ with $\lambda_{1},\dots,\lambda_{n}>0$ such that $M^{T}M=QDQ^{-1}$. It follows that
$\begin{aligned} \mathcal{L}^{n}\left\{\omega\in\mathbb{R}^{n}:\langle M^{T}M\omega,\omega\rangle\leq 1\right\}&=\mathcal{L}^{n}\left\{\omega\in\mathbb{R}^{n}:\langle DQ^{-1}\omega,Q^{-1}\omega\rangle\leq 1\right\}\\[7.11317pt] &=\mathcal{L}^{n}\left\{v\in\mathbb{R}^{n}:\langle Dv,v\rangle\leq 1\right\}\\ &=\mathcal{L}^{n}\bigg\{(v_{1},\dots,v_{n})\in\mathbb{R}^{n}:\sum_{j=1}^{n}\lambda_{j}v_{j}^{2}\leq 1\bigg\}.\end{aligned}$ (IV.42)
By a linear change of variable, we have
$\begin{aligned} \mathcal{L}^{n}\bigg\{(v_{1},\dots,v_{n})\in\mathbb{R}^{n}:\sum_{j=1}^{n}\lambda_{j}v_{j}^{2}\leq 1\bigg\}&=\prod_{j=1}^{n}\lambda_{j}^{-\frac{1}{2}}|B^{n}|=\det(M^{T}M)^{-\frac{1}{2}}|B^{n}|.\end{aligned}$ (IV.43)
Combining (IV.38)–(IV.43), we finally conclude
$\lim_{i\rightarrow\infty}\frac{\int_{f^{-1}\big(B_{r_{k_{i}}}(0)\big)}1\,d\textup{vol}_{g_{f}}}{r_{k_{i}}^{n}|B^{n}|}=\frac{C_{0}^{n}\mathcal{L}^{n}\left(\tilde{f}^{-1}\big(B_{C_{0}^{-1}}(0)\big)\right)\sqrt{\det(M^{T}M)}}{|B^{n}|}=1.$
Thus, the limit (IV.31) is proved. ∎
IV.4 Approximation by smooth immersions
In this section, we prove that every weak immersion $\vec{\Phi}$ can always be approximated by smooth immersions $\vec{\Phi}_{k}$ and that the conformal structures induced by $\vec{\Phi}_{k}$ converge to the one of $\vec{\Phi}$. To do so, we first prove in Section IV.4.1 that the conformal charts obtained by Coulomb frames converges. Combining this with an argument from Uhlenbeck, we prove in Section IV.4.2 that the conformal structures converges as well. We summarize the approximations results that we obtain in Section IV.4.3.
IV.4.1 Convergence of isothermal coordinates
The goal of this section is to prove Theorem IV.17, namely that if a sequence of weak immersions $(\vec{\Phi}_{k})_{k\in\mathbb{N}}$ converges to some $\vec{\Phi}_{\infty}$ in the strong $W^{2,2}_{\textup{loc}}$-topology, assuming a control on the metrics $g_{\vec{\Phi}_{k}}$, then the conformal charts induced by $\vec{\Phi}_{k}$ converge to conformal charts for $\vec{\Phi}$.
In order to prove the convergence of the Coulomb frames, we first need to prove some convergence result for equations having varying coefficients.
Lemma IV.16.
Let $n\geq 2$, $U$ be a bounded open subset of $\mathbb{R}^{n}$, $\{a^{ij}_{k}\}\subset L^{\infty}(U)$, $\{f_{k}\}\subset W^{-1,2}(U)$. Suppose that there exists a constant $\Lambda>0$ such that for all $\xi\in\mathbb{R}^{n}$, $k\in\mathbb{N}$, and a.e. $x\in U$, it holds
$\Lambda^{-1}|\xi|^{2}\leq\sum_{i,j}^{n}a^{ij}_{k}(x)\,\xi_{i}\xi_{j}\leq\Lambda|\xi|^{2}.$
Assume in addition that $a^{ij}_{k}\rightarrow a^{ij}$ a.e. for some $a^{ij}\in L^{\infty}(U)$ and $f_{k}\rightarrow f$ in $W^{-1,2}(U)$ for some $f\in W^{-1,2}(U)$. Let $u_{k}\in W^{1,2}(U)$ be a solution to the equation
$\partial_{i}\big(a^{ij}_{k}\,\partial_{j}u_{k}\big)=f_{k}\quad\text{in }U.$ (IV.44)
Suppose that $u_{k}\rightharpoonup u$ weakly in $W^{1,2}(U)$. Then we have $u_{k}\rightarrow u$ in $W_{\textup{loc}}^{1,2}(U)$.
Proof.
Let $V\Subset U$ be a bounded open subset with smooth boundary. Let $\xi\in C^{\infty}_{c}(U)$ be a cut-off function such that $\xi\geq 0$ on $U$ and $\xi=1$ on $V$. Denote by $\left\langle,\right\rangle$ the duality pairing between $W^{-1,2}(U)$ and $W^{1,2}_{0}(U)$. Using $(u_{k}-u)\,\xi$ as a test function in the equation (IV.44), we obtain
$\begin{aligned} &\Lambda^{-1}\int_{V}|\nabla(u_{k}-u)|^{2}\\ &\leq\int_{U}a^{ij}_{k}\,\xi\,\partial_{j}(u_{k}-u)\,\partial_{i}(u_{k}-u)\\ &=-\left\langle f_{k},\xi(u_{k}-u)\right\rangle-\int_{U}\big(a^{ij}_{k}\,\partial_{j}u_{k}\,\partial_{i}\xi\,(u_{k}-u)+a^{ij}_{k}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\big).\end{aligned}$ (IV.45)
Since $\xi(u_{k}-u)\rightharpoonup 0$ weakly in $W_{0}^{1,2}(U)$ and $\|f_{k}-f\|_{W^{-1,2}(U)}\rightarrow 0$ as $k\rightarrow\infty$, we have
$\begin{aligned} \left|\left\langle f_{k},\xi(u_{k}-u)\right\rangle\right|&\leq\left|\left\langle f_{k}-f,\xi(u_{k}-u)\right\rangle\right|+\left|\left\langle f,\xi(u_{k}-u)\right\rangle\right|\\[8.53581pt] &\leq\left\|f_{k}-f\right\|_{W^{-1,2}(U)}\left\|\xi(u_{k}-u)\right\|_{W^{1,2}(U)}+\left|\left\langle f,\xi(u_{k}-u)\right\rangle\right|\\[5.69054pt] &\xrightarrow[k\to\infty]{}0.\end{aligned}$ (IV.46)
By Rellich–Kondrachov compactness theorem, we also get
$\begin{aligned} \left|\int_{U}a^{ij}_{k}\,\partial_{j}u_{k}\,\partial_{i}\xi\,(u_{k}-u)\right|&\leq\left\|a^{ij}_{k}\right\|_{L^{\infty}(U)}\left\|\partial_{j}u_{k}\right\|_{L^{2}(U)}\left\|\partial_{i}\xi\,(u_{k}-u)\right\|_{L^{2}(U)}\xrightarrow[k\to\infty]{}0.\end{aligned}$ (IV.47)
Concerning the last term, by weak convergence $u_{k}\rightharpoonup u$ in $W^{1,2}(U)$ and dominated convergence theorem, we obtain
$\begin{aligned} &\left|\int_{U}a^{ij}_{k}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|&\\ &\leq\left|\int_{U}a^{ij}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|+\int_{U}\left|(a^{ij}_{k}-a^{ij})\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|\\ &\leq\left|\int_{U}a^{ij}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|+\left\|(a^{ij}_{k}-a^{ij})\,\xi\,\partial_{j}u\right\|_{L^{2}(U)}\left\|\partial_{i}(u_{k}-u)\right\|_{L^{2}(U)}\\ &\xrightarrow[k\to\infty]{}0.\end{aligned}$ (IV.48)
Combining (LABEL:est-diff-W12)–(LABEL:3rd-est-Duk-diff) we finally conclude $u_{k}\rightarrow u$ in $W^{1,2}_{\textup{loc}}(U)$. ∎
Using the proof of Theorem IV.4, we are now ready to prove strong convergence of conformal charts for a convergent sequence of weak immersions with bounded induced metric.
Theorem IV.17.
Let $\vec{\Phi}_{k}\in W_{\textup{imm}}^{2,2}(D^{2},\mathbb{R}^{m})$ for $k\in\mathbb{N}\cup\{\infty\}$. Assume that $\vec{\Phi}_{k}\rightarrow\vec{\Phi}_{\infty}$ in $W^{2,2}(D^{2})$ as $k\rightarrow\infty$, and that there exists a constant $\Lambda>0$ such that for any $k\in\mathbb{N}\cup\{\infty\}$, any $X\in\mathbb{R}^{2}$ and a.e. $p\in D^{2}$, there holds
$\Lambda^{-1}|X|_{\mathbb{R}^{2}}\leq|d(\vec{\Phi}_{k})_{p}(X)|_{\mathbb{R}^{m}}\leq\Lambda|X|_{\mathbb{R}^{2}}.$ (IV.49)
Then there exists a neighborhood $U\subset D^{2}$ of 0 and $W^{1,\infty}\cap W^{2,2}$ diffeomorphisms $\varphi_{k}\colon U\rightarrow\varphi_{k}(U)\subset\mathbb{R}^{2}$ such that $\vec{\Phi}_{k}\circ\varphi_{k}^{-1}$ is (weakly) conformal on $\varphi_{k}(U)$ for any $k\in\mathbb{N}\cup\{\infty\}$, $\varphi_{k}\rightarrow\varphi_{\infty}$ in $W^{2,2}(U)$ and $\|\varphi_{k}\|_{W^{1,\infty}(U)}\leq C$, $\|\varphi_{k}^{-1}\|_{W^{1,\infty}(\varphi_{k}(U))}\leq C$ for a constant $C>0$.
Proof.
We follow the proof of Theorem IV.4. For $k\in\mathbb{N}\cup\{\infty\}$, denote
$g_{k}=g_{ij,k}\ dx^{i}\otimes dx^{j},\qquad g_{ij,k}=\langle\partial_{i}\vec{\Phi}_{k},\partial_{j}\vec{\Phi}_{k}\rangle,$
$(g^{ij}_{k})=(g_{ij,k})^{-1},\qquad\det g_{k}=\det(g_{ij,k}).$
Then we have $g_{k}\rightarrow g_{\infty}$ in $W^{1,2}(D^{2})$ as $k\rightarrow\infty$, and $\{g_{k}\}$, $\{g^{ij}_{k}\}$ are bounded in $L^{\infty}(D^{2})$.
For $k\in\mathbb{N}\cup\{\infty\}$, we apply Gram–Schmidt process to $(\partial_{1}\vec{\Phi}_{k},\partial_{2}\vec{\Phi}_{k})$ and get
$\vec{f}_{1,k}\coloneqq\frac{\partial_{1}\vec{\Phi}_{k}}{|\partial_{1}\vec{\Phi}_{k}|},\qquad\vec{f}_{2,k}\coloneqq\frac{\partial_{2}\vec{\Phi}_{k}-(\partial_{2}\vec{\Phi}_{k}\cdot\vec{f}_{1,k})\vec{f}_{1,k}}{|\partial_{2}\vec{\Phi}_{k}-(\partial_{2}\vec{\Phi}_{k}\cdot\vec{f}_{1,k})\vec{f}_{1,k}|}.$ (IV.50)
By (IV.49) we have $|\partial_{1}\vec{\Phi}_{k}|,|\partial_{2}\vec{\Phi}_{k}|\in[\Lambda^{-1},\Lambda]$, and
$|\partial_{2}\vec{\Phi}_{k}-(\partial_{2}\vec{\Phi}_{k}\cdot\vec{f}_{1,k})\vec{f}_{1,k}|\in[\Lambda^{-1},\Lambda(1+\Lambda^{2})].$
Hence it holds that $(\vec{f}_{1,k},\vec{f}_{2,k})\in L^{\infty}\cap W^{1,2}(D^{2},V_{2}({\mathbb{R}}^{m}))$, where $V_{2}({\mathbb{R}}^{m})$ denotes the space of orthonormal 2-frames in ${\mathbb{R}}^{m}$. Since $(\vec{\Phi}_{k})_{k\in\mathbb{N}}$ converges in $W^{2,2}$, we deduce that $\vec{f}_{i,k}\rightarrow\vec{f}_{i,\infty}$ in $W^{1,2}(D^{2})$ as $k\rightarrow\infty$ for $i\in\{1,2\}$. For $\theta\in W^{1,2}(D^{2},{\mathbb{R}})$, we denote $(\vec{f}^{\,\theta}_{1,k},\vec{f}^{\,\theta}_{2,k})$ the rotation by the angle $\theta$ of the original frame $(\vec{f}_{1,k},\vec{f}_{2,k})$:
$\vec{f}^{\,\theta}_{1,k}+i\vec{f}^{\,\theta}_{2,k}=e^{i\theta}\left(\vec{f}_{1,k}+i\vec{f}_{2,k}\right).$
Then we still have $\left(\vec{f}^{\,\theta}_{1,k},\vec{f}^{\,\theta}_{2,k}\right)\in L^{\infty}\cap W^{1,2}(D^{2},V_{2}(\mathbb{R}^{m}))$.
We look now for a rotation of this orthonormal frame realizing the following absolute minimum
$\begin{array}[]{l}\displaystyle\inf_{\theta\in W_{0}^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}|\langle\vec{f}^{\,\theta}_{1,k},d\vec{f}^{\,\theta}_{2,k}\rangle|^{2}_{g_{k}}\ d\textup{vol}_{g_{k}}\\[14.22636pt] \displaystyle=\inf_{\theta\in W_{0}^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}\sum_{i,j=1}^{2}g^{ij}_{k}\,\langle\vec{f}^{\,\theta}_{1,k},\partial_{i}\vec{f}^{\,\theta}_{2,k}\rangle\,\langle\vec{f}^{\,\theta}_{1,k},\partial_{j}\vec{f}^{\,\theta}_{2,k}\rangle\,\sqrt{\det g_{k}}\ dx^{1}\wedge dx^{2}.\end{array}$ (IV.51)
For $i\in\{1,2\}$, we have
$\langle\vec{f}^{\,\theta}_{1,k},\partial_{i}\vec{f}^{\,\theta}_{2,k}\rangle=\partial_{i}\theta+\langle\vec{f}_{1,k},\partial_{i}\vec{f}_{2,k}\rangle.$ (IV.52)
Hence the following energy is strictly convex in $W_{0}^{1,2}(D^{2},{\mathbb{R}})$:
$E(\theta)\coloneqq\int_{D^{2}}|\langle\vec{f}^{\,\theta}_{1,k},d\vec{f}^{\,\theta}_{2,k}\rangle|^{2}_{g_{k}}\,d\textup{vol}_{g_{k}}=\int_{D^{2}}|d\theta+\langle\vec{f}_{1,k},d\vec{f}_{2,k}\rangle|^{2}_{g_{k}}\,d\textup{vol}_{g_{k}}.$
By a standard application of Mazur’s lemma, we have a unique $\theta_{k}\in W_{0}^{1,2}(D^{2},\mathbb{R})$ that achieves the minimum of $E$. It is the unique solution in $W_{0}^{1,2}(D^{2},\mathbb{R})$ for the Euler–Lagrange equation:
$\forall\phi\in W_{0}^{1,2}(D^{2},{\mathbb{R}}),\qquad\int_{D^{2}}d\phi\wedge*_{g_{k}}\left(d\theta_{k}+\langle\vec{f}_{1,k},d\vec{f}_{2,k}\rangle\right)=0.$ (IV.53)
In particular, taking $\phi=\theta_{k}$ in (IV.53), we get
$\Lambda^{-4}\int_{D^{2}}|\nabla\theta_{k}|^{2}\,dx^{2}\leq\int_{D^{2}}|d\theta_{k}|^{2}_{g_{k}}\,d\textup{vol}_{g_{k}}\leq\int_{D^{2}}|d\vec{f}_{2,k}|_{g_{k}}^{2}\,d\textup{vol}_{g_{k}}\leq\Lambda^{4}\int_{D^{2}}|\nabla\vec{f}_{2,k}|^{2}.$
Thus the sequence $\{\theta_{k}\}$ is bounded in $W_{0}^{1,2}(D^{2})$ and each subsequence contains a further subsequence $\{\theta_{k_{l}}\}$ weakly converging to some $\theta^{\prime}\in W^{1,2}_{0}(D^{2})$. Then using (IV.53), Hölder’s inequality and dominated convergence theorem, we have
$\forall\phi\in W_{0}^{1,2}(D^{2},{\mathbb{R}}),\qquad\int_{D^{2}}d\phi\wedge*_{g_{\infty}}\left(d\theta^{\prime}+\langle\vec{f}_{1,\infty},d\vec{f}_{2,\infty}\rangle\right)=0.$
Hence it holds $\theta_{\infty}=\theta^{\prime}$ and we have $\theta_{k}\rightharpoonup\theta_{\infty}$ weakly in $W^{1,2}(D^{2})$. Also by dominated convergence, we have
$d\left(*_{g_{k}}\langle\vec{f}_{1,k},d\vec{f}_{2,k}\rangle\right)\rightarrow d\left(*_{g_{\infty}}\langle\vec{f}_{1,\infty},d\vec{f}_{2,\infty}\rangle\right)\qquad\text{in $W^{-1,2}(D^{2})$. }$
Then by replacing the cutoff function $\varphi$ by $1$ on $D^{2}$, we use the same argument as in Lemma IV.16 to obtain $\theta_{k}\rightarrow\theta_{\infty}$ in $W^{1,2}(D^{2})$. Denoting $(\vec{e}_{1,k},\vec{e}_{2,k})\coloneqq(\vec{f}^{\,\theta_{k}}_{1,k},\vec{f}^{\,\theta_{k}}_{2,k})$, we have $\vec{e}_{i,k}\rightarrow\vec{e}_{i,\infty}$ as $k\rightarrow\infty$ for $i\in\{1,2\}$. We deduce from (IV.53) that the following equation holds in the sense of distribution:
$d\Big(*_{g_{k}}\langle\vec{e}_{1,k},d\vec{e}_{2,k}\rangle\Big)=0.$
Therefore by the weak Poincaré lemma, there exists $\lambda_{k}\in W^{1,2}(D^{2})$ with $\int_{D^{2}}\lambda_{k}=0$ such that
$d\lambda_{k}=*_{g_{k}}\langle\vec{e}_{1,k},d\vec{e}_{2,k}\rangle.$
Hence we have the convergence $\nabla\lambda_{k}\rightarrow\nabla\lambda_{\infty}$ in $L^{2}(D^{2})$, from which we deduce that $\lambda_{k}\rightarrow\lambda_{\infty}$ in $W^{1,2}(D^{2})$ by Poincaré inequality. In order to deduce boundedness in $C^{0}$, we use the following equation:
$\Delta_{g_{k}}\lambda_{k}=*_{g_{k}}d*_{g_{k}}d\lambda_{k}=-*_{g_{k}}d\langle\vec{e}_{1,k},d\vec{e}_{2,k}\rangle.$
Let $\lambda_{k}^{\prime}\in\bigcap_{1<p<2}W^{1,p}_{0}(D^{2})$ be the solution to the following equation:
$\begin{cases}\displaystyle\,\partial_{i}\left(\sqrt{\det g_{k}}\ g_{k}^{ij}\,\partial_{j}\lambda^{\prime}_{k}\right)=\partial_{2}\vec{e}_{1,k}\cdot\partial_{1}\vec{e}_{2,k}-\partial_{1}\vec{e}_{1,k}\cdot\partial_{2}\vec{e}_{2,k}\quad&\mbox{ in }D^{2},\\[8.53581pt] \displaystyle\lambda_{k}^{\prime}=0\quad&\mbox{ on }\partial D^{2}.\end{cases}$
By Proposition IV.8, we obtain that $\{\lambda_{k}^{\prime}\}$ is bounded in $C^{0}(\overline{D^{2}})\cap W^{1,2}_{0}(D^{2},{\mathbb{R}})$. Concerning the difference $\lambda_{k}^{\prime}-\lambda_{k}$, we have the following equation:
$\partial_{i}\left(\sqrt{\det g_{k}}\ g_{k}^{ij}\,\partial_{j}(\lambda^{\prime}_{k}-\lambda_{k})\right)=0.$
Thanks to [134], there exists $p=p(\Lambda)\in(2,\infty)$ and a positive constant $C$ independent of $k$, such that
$\|\lambda^{\prime}_{k}-\lambda_{k}\|_{C^{0}(\overline{D_{1/2}(0)})}\leq C\|\lambda^{\prime}_{k}-\lambda_{k}\|_{W^{1,p}(D_{1/2}(0))}\leq C(\Lambda)\,\|\lambda^{\prime}_{k}-\lambda_{k}\|_{W^{1,2}(D^{2})}\leq C.$
It follows that $\{\lambda_{k}\}_{k\in\mathbb{N}}$ is bounded in $C^{0}(\overline{D_{1/2}(0)})$. For $i\in\{1,2\}$, we define pullback of $\vec{e}_{i,k}$ and the dual $1$-forms
$\begin{dcases}e_{i,k}\coloneqq g^{jl}_{k}\,\langle\vec{e}_{i,k},\partial_{j}\vec{\Phi}_{k}\rangle\,\frac{\partial}{\partial x^{l}}\in L^{\infty}\cap W^{1,2}(D_{1/2}(0),TD_{1/2}(0)),\\[5.69054pt] e_{i,k}^{\ast}\coloneqq\langle\vec{e}_{i,k},\partial_{j}\vec{\Phi_{k}}\rangle\,dx^{j}=\langle\vec{e}_{i,k},d\vec{\Phi}_{k}\rangle\in L^{\infty}\cap W^{1,2}(D_{1/2}(0),T^{*}D_{1/2}(0)).\end{dcases}$ (IV.54)
Then by the same proof of Theorem IV.4, we have
$d(e^{-\lambda_{k}}\,e_{i,k}^{*})=0.$
Using the weak Poincaré lemma, we obtain the existence of $\varphi^{i}_{k}\in W^{1,\infty}\cap W^{2,2}(D_{1/2}(0),{\mathbb{R}})$ with $\int_{D_{1/2}(0)}\varphi^{i}_{k}=0$ for $i\in\{1,2\}$ such that
$d\varphi^{i}_{k}=e^{-\lambda_{k}}\,e_{i,k}^{*}.$ (IV.55)
Let $\varphi_{k}\coloneqq(\varphi^{1}_{k},\varphi^{2}_{k})\in W^{1,\infty}\cap W^{2,2}(D_{1/2}(0),{\mathbb{R}^{2}})$. From (IV.54) and (IV.55) we deduce that for a.e. $p\in D_{1/2}(0)$ and any $v\in T_{p}(D_{1/2}(0))$,
$e^{-\|\lambda_{k}\|_{C(\overline{D_{1/2}(0)})}}\,\Lambda^{-1}|v|\leq|d(\varphi_{k})_{p}(v)|=e^{-\lambda_{k}}|d\vec{\Phi}_{k}(v)|\leq e^{\|\lambda_{k}\|_{C(\overline{D_{1/2}(0)})}}\,\Lambda|v|.$ (IV.56)
Since $\lambda_{k}\rightarrow\lambda_{\infty}$, $e_{i,k}^{*}\rightarrow e^{*}_{i,\infty}$ in $W^{1,2}(D^{2})$ as $k\rightarrow\infty$, and $\{\lambda_{k}\}$ is bounded in $C(\overline{D_{1/2}(0)})$, by dominated convergence theorem and Poincaré inequality we have $\varphi_{k}\rightarrow\varphi_{\infty}$ in $W^{2,2}(D_{1/2}(0))$. Combining (IV.56) and Lemma IV.12, it follows that there exists an open neighborhood $U\subset D_{1/2}(0)$ of 0 independent of $k$ such that $\varphi_{k}$ is open and injective on $U$, and its inverse $\varphi_{k}^{-1}\in W^{1,\infty}\cap W^{2,2}(\varphi_{k}(U),\mathbb{R}^{2})$. We obtain that $\varphi_{k}$ is an isothermal chart of $\vec{\Phi}_{k}$ by the same proof of Theorem IV.4. ∎
IV.4.2 Convergence of conformal structure
As a consequence of Theorem IV.17, we deduce the convergence of the underlying conformal structure for sequences of metrics induced by weak immersions which converge in some Sobolev space. To begin with, we consider the case where $g_{k}\rightarrow g_{\infty}$ smoothly.
Lemma IV.18.
Let $\Sigma$ be an oriented connected $2$-dimensional closed smooth manifold. Assume that $(g_{k})_{k\in\mathbb{N}\cup\{\infty\}}$ is a sequence of Riemannian metrics on $\Sigma$, and
$g_{k}\xrightarrow[k\rightarrow\infty]{}g_{\infty}\quad\text{in }\,C^{\infty}(\Sigma,T^{*}\Sigma\otimes T^{*}\Sigma).$
Then there exists a sequence $(h_{k})_{k\in\mathbb{N}\cup\{\infty\}}$ of metrics of constant Gaussian curvature 1,11,-1 or 0 such that $h_{k}$ is conformal to $g_{k}$ and $h_{k}\rightarrow h_{\infty}$ in $C^{\infty}(\Sigma,T^{*}\Sigma\otimes T^{*}\Sigma)$ as $k\rightarrow\infty$.
Proof.
Let $\chi(\Sigma)$ denote the Euler characteristic of $\Sigma$, $K_{h_{k}}$ be the Gaussian curvature of $h_{k}$.
(Case I) $\chi(\Sigma)\leq-2\Rightarrow K_{h_{k}}=-1$.
By a theorem of Poincaré and its corollary (see [185, Theorem 1.6.2]), in this case, for any $s\in\mathbb{N}$, $s\geq 3$ and any metric $g\in W^{s,2}(\Sigma,T^{*}\Sigma\otimes T^{*}\Sigma)$, there exists a unique metric $h\in W^{s,2}(\Sigma,T^{*}\Sigma\otimes T^{*}\Sigma)$ such that $h$ is weakly conformal to $g$ and $h$ has constant Gaussian curvature 1-1. Moreover, the map taking $g$ to $h$ is continuous in $W^{s,2}$. It follows that $h_{k}$ is uniquely determined by $g_{k}$, and converges to $h_{\infty}$ smoothly since $g_{k}\rightarrow g_{\infty}$ in $C^{\infty}(\Sigma,T^{*}\Sigma\otimes T^{*}\Sigma)$.
(Case II) $\chi(\Sigma)=0\Rightarrow K_{h_{k}}=0$.
Let $h_{k}=e^{2\lambda_{k}}g_{k}$ for some $\lambda_{k}\in C^{\infty}(\Sigma)$. We denote the Gaussian curvature and negative Laplace–Beltrami operator associated to $g_{k}$ by $K_{g_{k}}$ and $\Delta_{g_{k}}$ respectively, the condition that $h_{k}$ has constant Gaussian curvature 0 is then equivalent to the equation (see e.g. [168, Theorem 2.6])
$\Delta_{g_{k}}\lambda_{k}=K_{g_{k}}.$ (IV.57)
By Gauss–Bonnet theorem, we have
$\int_{\Sigma}K_{g_{k}}\,d\textup{vol}_{g_{k}}=2\pi\chi(\Sigma)=0.$
Hence for each $k\in\mathbb{N}\cup\{\infty\}$, there exists a smooth solution $\lambda_{k}$ to (IV.57) and it is unique up to additive constants (see [9, Theorem 4.7]). By adding suitable constants, we can without loss of generality assume that
$\int\lambda_{k}\,d\textup{vol}_{g_{\infty}}=0\quad\text{for any }k\in\mathbb{N}\cup\{\infty\}.$ (IV.58)
Since $g_{k}\rightarrow g_{\infty}$ in $C^{\infty}$, we have $K_{g_{k}}\rightarrow K_{g_{\infty}}$ in $C^{\infty}(\Sigma)$ as $k\rightarrow\infty$, and in particular, $\{K_{g_{k}}\}_{k\in\mathbb{N}\cup\{\infty\}}$ is bounded in $C^{0}(\Sigma)$. By (IV.57) and the convergence $g_{k}\rightarrow g_{\infty}$, there exists a constant $C_{1}>0$ independent of $k$ such that
$C_{1}\int_{\Sigma}|d\lambda_{k}|_{g_{\infty}}^{2}\,d\textup{vol}_{g_{\infty}}$ $\leq\int_{\Sigma}|d\lambda_{k}|_{g_{k}}^{2}\,d\textup{vol}_{g_{k}}$
$=-\int_{\Sigma}\lambda_{k}\,\Delta_{g_{k}}\lambda_{k}\,d\textup{vol}_{g_{k}}$
$=\int_{\Sigma}-\lambda_{k}K_{g_{k}}\,d\textup{vol}_{g_{k}}.$
Furthermore, by (IV.58), Cauchy–Schwarz inequality, and Poincaré’s inequality, there exist constants $C_{2},C_{3}>0$ independent of $k$ such that
$\int_{\Sigma}-\lambda_{k}K_{g_{k}}\,d\textup{vol}_{g_{k}}$ $\leq\left(\int_{\Sigma}\lambda_{k}^{2}\,d\textup{vol}_{g_{k}}\right)^{\frac{1}{2}}\left(\int_{\Sigma}K_{g_{k}}^{2}\,d\textup{vol}_{g_{k}}\right)^{\frac{1}{2}}$
$\leq C_{2}\left(\int_{\Sigma}\lambda_{k}^{2}\,d\textup{vol}_{g_{\infty}}\right)^{\frac{1}{2}}$
$\leq C_{3}\left(\int_{\Sigma}|d\lambda_{k}|_{g_{\infty}}^{2}\,d\textup{vol}_{g_{\infty}}\right)^{\frac{1}{2}}.$
Consequently, $\{\lambda_{k}\}_{k\in\mathbb{N}\cup\{\infty\}}$ is bounded in $W^{1,2}(\Sigma)$. Now we apply the interior estimates as in [58, Section 6.3.1] to the equation (IV.57) and obtain that $\{\lambda_{k}\}_{k\in\mathbb{N}\cup\{\infty\}}$ is bounded in $W^{s,2}(\Sigma)$ for any $s\in\mathbb{N}$. A standard compactness argument then implies $\lambda_{k}\rightarrow\lambda_{\infty}$ in $C^{\infty}(\Sigma)$ as $k\rightarrow\infty$. Finally, by setting $h_{k}=e^{2\lambda_{k}}g_{k}$, we obtain the required convergent sequence of metrics of Gaussian curvature 0.
(Case III) $\chi(\Sigma)=2\Rightarrow K_{h_{k}}=1$.
We identify $\Sigma=\mathbb{C}\cup\{\infty\}$ with the canonical smooth structure and orientation. Define $\psi(z)\coloneqq 1/z$ on $\Sigma$, and $\phi\colon\mathbb{C}\rightarrow S^{2}\setminus\{(0,0,1)\}$ be the inverse of the stereographic projection, i.e.
$\phi(x,y)\coloneqq\left(\frac{2x}{1+x^{2}+y^{2}},\frac{2y}{1+x^{2}+y^{2}},\frac{-1+x^{2}+y^{2}}{1+x^{2}+y^{2}}\right).$
Let $h\coloneqq\phi^{*}g_{S^{2}}$, where $g_{S^{2}}$ is the standard round metric on $S^{2}$. Then $h$ can be extended to a Riemannian metric on $\Sigma$ with constant Gaussian curvature $1$, and is conformal to $g_{\text{std}}=dx^{2}+dy^{2}$ on $\mathbb{C}$. Given $k\in\mathbb{N}\cup\{\infty\}$, there exist two functions $\sigma_{k}\colon\mathbb{C}\rightarrow(0,\infty)$ and $\mu_{k}\colon\mathbb{C}\rightarrow D_{1}(0)\subset\mathbb{C}$ such that
$g_{k}=(g_{k})_{11}\,dx^{2}+2(g_{k})_{12}\,dxdy+(g_{k})_{22}\,dy^{2}=\sigma_{k}|dz+\mu_{k}\,d\bar{z}|^{2},$
where $dz=dx+idy$ and $d\bar{z}=dx-idy$. Moreover, such $\mu_{k}$ is uniquely determined by $g_{k}$ and can be computed as (see e.g. [152, Section 1.2.1])
$\mu_{k}=\frac{(g_{k})_{11}-(g_{k})_{22}+2i\,(g_{k})_{12}}{(g_{k})_{11}+(g_{k})_{22}+2\sqrt{(g_{k})_{11}(g_{k})_{22}-(g_{k})_{12}^{2}}}.$
In particular, we have $\mu_{k}\rightarrow\mu_{\infty}$ in $C^{\infty}_{\textup{loc}}(\mathbb{C})$. On the other hand, we obtain that for all $\omega\in\mathbb{C}$,
$(\psi^{-1})^{*}(g_{k})_{\omega}=\frac{\sigma_{k}\circ\psi^{-1}(\omega)}{|\omega|^{4}}\big|d\omega+\mu_{k}\circ\psi^{-1}(\omega)\,\omega^{2}{\bar{\omega}}^{-2}d\bar{\omega}\big|^{2}.$
Since $(\psi^{-1})^{*}g_{k}$ is smooth on $\mathbb{C}$, it follows that $\mu_{k}\circ\psi^{-1}(\omega)\,\omega^{2}{\bar{\omega}}^{-2}$ can be extended to a smooth function on $\mathbb{C}$ and converges to $\mu_{\infty}\circ\psi^{-1}(\omega)\,\omega^{2}{\bar{\omega}}^{-2}$ in $C^{\infty}_{\textup{loc}}(\mathbb{C})$. In particular, it holds that $\|\mu_{k}\|_{L^{\infty}(\mathbb{C})}<1$ and $\|\mu_{k}-\mu_{\infty}\|_{L^{\infty}(\mathbb{C})}\rightarrow 0$ as $k\rightarrow\infty$.
For $k\in\mathbb{N}\cup\{\infty\}$, let $f_{k}$ be the unique orientation-preserving homeomorphism of $\Sigma=\mathbb{C}\cup\{\infty\}$ onto itself satisfying the relations
$\begin{cases}\partial_{\bar{z}}f_{k}=\mu_{k}\,\partial_{z}f_{k},\\[5.69054pt] f_{k}(0)=0,\ f_{k}(1)=1,\ f_{k}(\infty)=\infty.\end{cases}$
In particular, $f_{k}$ is quasiconformal with complex dilatation $\mu_{k}$ (see Definition III.6). By considering the quasiconformal maps $f_{k}\circ f_{\infty}^{-1}$ and applying [90, Proposition 4.36], we have $f_{k}\rightarrow f_{\infty}$ locally uniformly on $\mathbb{C}$. Define $\sigma_{k}\in L^{\infty}(\mathbb{C},\mathbb{R}^{2\times 2})$ be the unique matrix-valued function on $\mathbb{C}$ such that $\sigma_{k}$ is positive definite a.e., $\text{det}(\sigma_{k})=1$, and
$\mu_{k}=\frac{(\sigma_{k})_{22}-(\sigma_{k})_{11}-2i\,(\sigma_{k})_{12}}{(\sigma_{k})_{11}+(\sigma_{k})_{22}+2}.$
Such $\sigma_{k}$ depends uniquely and smoothly on $\mu_{k}$, hence $\sigma_{k}\rightarrow\sigma_{\infty}$ in $C^{\infty}_{\textup{loc}}(\mathbb{C},\mathbb{R}^{2\times 2})$. By [163, Theorem 5.1], $f_{k}$ satisfies the elliptic equation
[…]
Hence standard elliptic estimates as in Case II imply that $f_{k}\rightarrow f_{\infty}$ in $C^{\infty}_{\textup{loc}}(\mathbb{C})$. On the other hand, note $\psi\circ f_{k}\circ\psi^{-1}$ is quasiconformal with complex dilatation $\mu_{k}\circ\psi^{-1}(\omega)\,\omega^{2}{\bar{\omega}}^{-2}$ (see [90, Proposition 4.13]), and
$\mu_{k}\circ\psi^{-1}(\omega)\,\omega^{2}{\bar{\omega}}^{-2}\xrightarrow[k\rightarrow\infty]{}\mu_{\infty}\circ\psi^{-1}(\omega)\,\omega^{2}{\bar{\omega}}^{-2}\qquad\text{ in }C_{\textup{loc}}^{\infty}(\mathbb{C}).$
By the same argument as before, we have
$\psi\circ f_{k}\circ\psi^{-1}\xrightarrow[k\rightarrow\infty]{}\psi\circ f_{\infty}\circ\psi^{-1}\qquad\text{ in }C_{\textup{loc}}^{\infty}(\mathbb{C}).$
Hence $f_{k}\rightarrow f_{\infty}$ in $C^{\infty}(\Sigma)$ as $k\rightarrow\infty$. Besides, by the proof of [4, Lemma 5.B.3], both $f_{k}$ and $\psi\circ f_{k}\circ\psi^{-1}$ have no critical points on $\mathbb{C}$, hence each $f_{k}$ is a diffeomorphism on $\Sigma$. Finally, for all $k\in\mathbb{N}\cup\{\infty\}$, let $h_{k}\coloneqq f_{k}^{*}h$, then $h_{k}$ is a Riemannian metric with constant Gaussian curvature $1$ on $\Sigma$, and $h_{k}\rightarrow h_{\infty}$ smoothly on $\Sigma$. Since $h$ is conformal to $g_{\text{std}}$ and $f_{k}^{*}g_{\text{std}}$ is conformal to $g_{k}$ on $\mathbb{C}$ (see for instance [90, Section 1.5.1]), we have $h_{k}=f_{k}^{*}h$ is conformal to $g_{k}$ on $\mathbb{C}\subset\Sigma$. The conformality holds on $\Sigma$ since $h_{k}$, $g_{k}$ are both Riemannian metrics on $\Sigma$, and $\mathbb{C}$ is dense in $\Sigma$. ∎
Then we consider the general case in which $\{g_{k}\}$ are metrics induced by weak immersions. The strategy of proof is an adaptation of [186, Proposition 3.2]. For a sequence of functions $f_{k}$ and a compact set $K\subset\mathbb{R}^{n}$, we write $f_{k}\rightarrow f$ in $C^{\infty}(K)$ if each $f_{k}$ is defined on a neighborhood of $K$ (possibly depending on $k$) and for any $l\in\mathbb{N}_{0}$, $\|f_{k}-f\|_{C^{l}(K)}\rightarrow 0$ as $k\rightarrow\infty$.
Theorem IV.19.
Let $\Sigma$ be an oriented connected $2$-dimensional closed smooth manifold with a reference Riemannian metric $\tilde{g}$. Assume $\vec{\Phi}_{k}\rightarrow\vec{\Phi}_{\infty}$ in $W^{2,2}(\Sigma,\mathbb{R}^{m})$ as $k\rightarrow\infty$, and there exists a constant $\Lambda>0$ such that for any $k\in\mathbb{N}\cup\{\infty\}$, there holds
$\Lambda^{-1}|X|_{\tilde{g}}\leq|d(\vec{\Phi}_{k})_{p}(X)|_{\mathbb{R}^{m}}\leq\Lambda|X|_{\tilde{g}},\qquad\text{for a.e. }p\in\Sigma\text{ and all }\mkern 2.0muX\in T_{p}\Sigma.$ (IV.59)
Let $g_{k}\coloneqq\vec{\Phi}_{k}^{*}\,g_{\text{std}}$ denote the metric induced by $\vec{\Phi}_{k}$. Assume $\tilde{g}$ and $g_{\infty}$ are weakly conformal (i.e. $\Sigma$ is equipped with the complex structure induced by $\vec{\Phi}_{\infty}$ as in Corollary IV.5), then there exists a sequence $(h_{k})_{k\in\mathbb{N}\cup\{\infty\}}$ of metrics smooth in the complex structure induced by $\vec{\Phi}_{k}$, and a sequence $(\Psi_{k})_{k\in\mathbb{N}}$ of bi-Lipschitz homeomorphisms on $\Sigma$ such that
[…]
Proof.
By Theorem IV.17, there exists a covering of $\Sigma$ by open subsets $\{U_{\alpha}\}_{\alpha=1}^{l}$ with orientation-preserving bi-Lipschitz diffeomorphisms $\varphi_{k,\alpha}\colon U_{\alpha}\rightarrow\varphi_{k,\alpha}(U_{\alpha})\subset\mathbb{R}^{2}$ (where $k\in\mathbb{N}\cup\{\infty\}$) such that $\vec{\Phi}_{k}\circ\varphi_{k,\alpha}^{-1}$ is weakly conformal, $\varphi_{k,\alpha}\rightarrow\varphi_{\infty,\alpha}$ in $W^{2,2}(U_{\alpha})$ as $k\rightarrow\infty$ and $\|\varphi_{k,\alpha}\|_{W^{1,\infty}(U_{\alpha})}+\|\varphi_{k,\alpha}^{-1}\|_{W^{1,\infty}(\varphi_{k,\alpha}(U_{\alpha}))}$ is uniformly bounded. Consequently, for any compact subset $K$ of $\varphi_{\infty,\alpha}(U_{\alpha})$, $K\subset\varphi_{k,\alpha}(U_{\alpha})$ for large enough $k$ and $\varphi_{k,\alpha}^{-1}\rightarrow\varphi_{\infty,\alpha}^{-1}$ in $W^{2,2}(K)$. Then for any $1\leq\alpha,\beta\leq l$ and compact subset $K$ of $\varphi_{\infty,\beta}(U_{\alpha}\cap U_{\beta})$, the transition maps converge as well:
$\varphi_{k,\alpha,\beta}\coloneqq\varphi_{k,\alpha}\circ\varphi_{k,\beta}^{-1}\xrightarrow[k\rightarrow\infty]{}\varphi_{\infty,\alpha}\circ\varphi_{\infty,\beta}^{-1}=:\varphi_{\infty,\alpha,\beta}\quad\text{in $W^{2,2}(K)$.}$ (IV.60)
Furthermore, since each $\varphi_{k,\alpha,\beta}$ is holomorphic, the convergence holds in $C^{\infty}(K)$.
We want to prove that for $k$ large enough, there exists an open cover $\{V_{\alpha}\}_{\alpha=1}^{l}$ of $\Sigma$ with $V_{\alpha}\Subset U_{\alpha}$ and $C^{\infty}$ injective immersions $\rho_{k,\alpha}\colon\varphi_{\infty,\alpha}(\overline{V_{\alpha}})\subset\mathbb{R}^{2}\rightarrow\varphi_{k,\alpha}(U_{\alpha})\subset\mathbb{R}^{2}$ satisfying
$\begin{cases}(i)\ \rho_{k,\alpha}\xrightarrow[k\rightarrow\infty]{}id&\text{ in }C^{\infty}(\varphi_{\infty,\alpha}(\overline{V_{\alpha}}),\mathbb{R}^{2}),\\[8.53581pt] (ii)\ \varphi_{k,\beta,\alpha}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha,\beta}=\rho_{k,\beta}&\text{ on }\varphi_{\infty,\beta}(\overline{V_{\alpha}}\cap\overline{V_{\beta}}).\end{cases}$ (IV.61)
Once these maps $\rho_{k,\alpha}$ are defined, we consider the map $\Psi_{k}\coloneqq\varphi_{k,\alpha}^{-1}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha}$ on $\overline{V_{\alpha}}$. This definition is independent of $\alpha$ by (IV.61). By construction, it holds $\Psi_{k}\rightarrow id$ in $C^{0}(\Sigma,\Sigma)$. For $k$ large enough, $\Psi_{k}$ is an immersion, hence surjective (its image is both open and closed). As for global injectivity, assume there exist sequences $\{k_{i}\}_{i=1}^{\infty}\subset\mathbb{N}$, $\{p_{i,1}\}_{i=1}^{\infty}$, $\{p_{i,2}\}_{i=1}^{\infty}\subset\Sigma$ such that $\lim_{i\rightarrow\infty}k_{i}=\infty$, $\Psi_{k_{i}}(p_{i,1})=\Psi_{k_{i}}(p_{i,2})$. We can without loss of generality assume that $p_{i,1}\rightarrow p_{1}$ and $p_{i,2}\rightarrow p_{2}$ as $i\rightarrow\infty$, but we have $\Psi_{k}\rightarrow id$ in $C^{0}(\Sigma,\Sigma)$ as $k\rightarrow\infty$, hence $p_{1}=p_{2}\in V_{\alpha}$ for some $\alpha$. For $i$ large enough we get $p_{i,1}=p_{i,2}$ since $\Psi_{k}$ is injective on $V_{\alpha}$.
We now proceed to the construction of $\{V_{\alpha}\}$ and $\{\rho_{k,\alpha}\}$. Let $\{U_{\alpha}^{\prime}\}_{\alpha=1}^{l}$ be an open cover of $\Sigma$ such that $U_{\alpha}^{\prime}$ has smooth boundary and $U_{\alpha}^{\prime}\Subset U_{\alpha}$ (for existence of such sets, see the proof of [52, Proposition 8.2.1]). We claim that there exist open sets $\{V_{\alpha,j}\}_{1\leq\alpha\leq j\leq l}$ (write $V_{\alpha,\alpha-1}=V_{\alpha,\alpha}$), $k_{0}\in\mathbb{N}$, and $C^{\infty}$ injective maps $\rho_{k,\alpha}\colon\varphi_{\infty,\alpha}(\overline{V_{\alpha,\alpha}})\rightarrow\varphi_{k,\alpha}(U_{\alpha})$ for $k\geq k_{0}$ such that for any $1\leq j\leq l$ and $1\leq\alpha,\beta\leq j$, it holds
$(i)\ V_{\alpha,j}\subset V_{\alpha,j-1}\subset U_{\alpha}^{\prime},$ (IV.62a)
$(ii)\ \Sigma=\Big(\bigcup\limits_{\gamma\leq j}V_{\gamma,j}\Big)\cup\Big(\bigcup\limits_{\gamma>j}U^{\prime}_{\gamma}\Big),$ (IV.62b)
$(iii)\ \rho_{k,\alpha}\xrightarrow[k\rightarrow\infty]{}id$ in $C^{\infty}(\varphi_{\infty,\alpha}(\overline{V_{\alpha,\alpha}}),\mathbb{R}^{2})$, (IV.62c)
$(iv)\ \varphi_{k,\beta,\alpha}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha,\beta}=\rho_{k,\beta}$ on $\varphi_{\infty,\beta}(\overline{V_{\alpha,j}}\cap\overline{V_{\beta,j}})$. (IV.62d)
The proof is by induction on $j$. For $j=1$, we define $V_{1,1}=U_{1}^{\prime}$ and $\rho_{k,1}\colon\varphi_{\infty,1}(\overline{U_{1}^{\prime}})\rightarrow\mathbb{R}^{2}$ as the identity map. For $k$ large enough, the image of $\rho_{k,1}$ is contained in $\varphi_{k,1}(U_{1})$ since $\varphi_{k,1}\rightarrow\varphi_{\infty,1}$ in $C(\overline{U_{1}^{\prime}},\mathbb{R}^{2})$ and $U_{1}^{\prime}\Subset U_{1}$. Let $1\leq j_{0}\leq l$, suppose that we have constructed $\{V_{\alpha,j}\}_{1\leq\alpha\leq j\leq j_{0}}$ and $\{\rho_{k,\alpha}\}_{1\leq\alpha\leq j_{0}}$ such that (IV.62a)–(IV.62d) hold for any $1\leq j\leq j_{0}$ and $1\leq\alpha,\beta\leq j$. We now construct $\rho_{k,j_{0}+1}$ and $\{V_{\alpha,j_{0}+1}\}_{1\leq\alpha\leq j_{0}+1}$ using the induction hypothesis.
For $1\leq\alpha\leq j_{0}$ we define
$\omega_{k,j_{0}}\coloneqq\varphi_{k,j_{0}+1,\alpha}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha,j_{0}+1}\quad\text{on $\varphi_{\infty,j_{0}+1}\left(\overline{V_{\alpha,j_{0}}}\cap U_{j_{0}+1}\right)$.}$
By our induction hypothesis (IV.62d) for $j\leq j_{0}$, this map is well-defined on $\varphi_{\infty,j_{0}+1}\Big(\big(\bigcup\limits_{\alpha\leq j_{0}}\overline{V_{\alpha,j_{0}}}\big)\cap U_{j_{0}+1}\Big)$ and independent of $\alpha$. By the assumption $\Sigma=\big(\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}}\big)\cup\big(\bigcup\limits_{\alpha>j_{0}}U^{\prime}_{\alpha}\big)$, we have
$\Big(\Sigma\setminus\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}}\Big)\cap\Big(\Sigma\setminus\bigcup\limits_{\alpha>j_{0}}U^{\prime}_{\alpha}\Big)=\emptyset.$
Hence we can choose a smooth cut-off $\xi_{j_{0}}\in C^{\infty}(\Sigma)$ such that $\xi_{j_{0}}=0$ on a neighborhood of $\Sigma\setminus\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}}$ and $\xi_{j_{0}}=1$ on a neighborhood of $\Sigma\setminus\bigcup\limits_{\alpha>j_{0}}U^{\prime}_{\alpha}$. Now we define $V_{j_{0}+1,j_{0}+1}\coloneqq U_{j_{0}+1}^{\prime}\Subset U_{j_{0}+1}$, and for $\alpha\leq j_{0}$
$V_{\alpha,j_{0}+1}\coloneqq V_{\alpha,j_{0}}\cap\,\text{int}\Big\{x\in\Sigma:\xi_{j_{0}}(x)=1\Big\}.$
The condition (IV.62a) then holds for $j=j_{0}+1$. By the definition of $\{V_{\alpha,j_{0}+1}\}_{\alpha\leq j_{0}+1}$ we have
$\Big(\bigcup\limits_{\alpha\leq j_{0}+1}V_{\alpha,j_{0}+1}\Big)\cup\Big(\bigcup\limits_{\alpha>j_{0}+1}U^{\prime}_{\alpha}\Big)$ $=\Big(\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}+1}\Big)\cup\Big(\bigcup\limits_{\alpha>j_{0}}U^{\prime}_{\alpha}\Big)$
$=\Big(\Big(\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}}\Big)\cap\,\text{int}\Big\{x\in\Sigma:\xi_{j_{0}}(x)=1\Big\}\Big)\cup\Big(\bigcup\limits_{\alpha>j_{0}}U^{\prime}_{\alpha}\Big).$
The definition of $\xi_{j_{0}}$ also implies
$\Sigma\setminus\bigcup\limits_{\alpha>j_{0}}U^{\prime}_{\alpha}\subset\text{int}\Big\{x\in\Sigma:\xi_{j_{0}}(x)=1\Big\}.$
We then obtain
$\Big(\bigcup\limits_{\alpha\leq j_{0}+1}V_{\alpha,j_{0}+1}\Big)\cup\Big(\bigcup\limits_{\alpha>j_{0}+1}U^{\prime}_{\alpha}\Big)$ $\supset\Big(\Big(\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}}\Big)\setminus\bigcup\limits_{\alpha>j_{0}}U^{\prime}_{\alpha}\Big)\cup\Big(\bigcup\limits_{\alpha>j_{0}}U^{\prime}_{\alpha}\Big)$
$=\Big(\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}}\Big)\cup\Big(\bigcup\limits_{\alpha>j_{0}}U^{\prime}_{\alpha}\Big)$
$=\Sigma.$
Hence the condition (IV.62b) holds for $j=j_{0}+1$. Since $\xi_{j_{0}}$ vanishes on a neighborhood of $\Sigma\setminus\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}}$, and $\omega_{k,j_{0}}\circ\varphi_{\infty,j_{0}+1}\in C^{\infty}\big(\big(\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}}\big)\cap U_{j_{0}+1}\big)$ we have $(\xi_{j_{0}}\circ\varphi^{-1}_{\infty,j_{0}+1})\,\omega_{k,j_{0}}\in C^{\infty}(\varphi_{\infty,j_{0}+1}(U_{j_{0}+1}))$. Then we define
$\rho_{k,j_{0}+1}\coloneqq(\xi_{j_{0}}\circ\varphi^{-1}_{\infty,j_{0}+1})\,\omega_{k,j_{0}}+(1-\xi_{j_{0}}\circ\varphi^{-1}_{\infty,j_{0}+1})\,id\in C^{\infty}(\varphi_{\infty,j_{0}+1}(U_{j_{0}+1}),\mathbb{R}^{2}).$
Let $\alpha\leq j_{0}$. Since $\xi_{j_{0}}=1$ on $\overline{V_{\alpha,j_{0}+1}}$, the following holds on $\varphi_{\infty,j_{0}+1}(U_{j_{0}+1}\cap\overline{V_{\alpha,j_{0}+1}})$:
$\rho_{k,j_{0}+1}=\omega_{k,j_{0}}=\varphi_{k,j_{0}+1,\alpha}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha,j_{0}+1}.$
Since we have $\overline{V_{j_{0}+1,j_{0}+1}}=\overline{U^{\prime}_{j_{0}+1}}\subset U_{j_{0}+1}$, the condition (IV.62d) is thus satisfied for $\beta=j=j_{0}+1$, $\alpha\leq j_{0}$. By using (IV.62a) and the identity $\varphi_{k,\alpha,\beta}=\varphi_{k,\beta,\alpha}^{-1}$, we then obtain (IV.62d) for $j=j_{0}+1$ and any $\alpha,\beta\leq j_{0}+1$.
Next we prove $\rho_{k,j_{0}+1}\rightarrow id$ in $C^{\infty}(\varphi_{\infty,j_{0}+1}(\overline{U^{\prime}_{j_{0}+1}}),\mathbb{R}^{2})$ as $k\rightarrow\infty$. Fixing $\alpha\leq j_{0}$, let $W$ be an open set such that $\varphi_{\infty,\alpha}(\overline{V_{\alpha,j_{0}}}\cap\overline{U_{j_{0}+1}^{\prime}})\subset W\Subset\varphi_{\infty,\alpha}(U_{\alpha}\cap U_{j_{0}+1})$. The convergence (IV.60) implies that
$\varphi_{k,j_{0}+1,\alpha}\xrightarrow[k\rightarrow\infty]{}\varphi_{\infty,j_{0}+1,\alpha}\quad\text{in $C^{\infty}(\overline{W})$. }$
By our induction hypothesis (IV.62c) for $\alpha\leq j_{0}$, we also have
$\rho_{k,\alpha}\xrightarrow[k\rightarrow\infty]{}id\quad\text{in $C^{\infty}(\varphi_{\infty,\alpha}(\overline{V_{\alpha,j_{0}}}))$.}$
Then the definition of $\omega_{k,j_{0}}$ implies
$\omega_{k,j_{0}}\xrightarrow[k\rightarrow\infty]{}id\quad\text{in $C^{\infty}(\varphi_{\infty,j_{0}+1}(\overline{V_{\alpha,j_{0}}}\cap\overline{U^{\prime}_{j_{0}+1}})$.}$
Hence we have
$\rho_{k,j_{0}+1}\xrightarrow[k\rightarrow\infty]{}id\quad\text{in $C^{\infty}\Big(\varphi_{\infty,j_{0}+1}\Big(\Big(\bigcup_{\alpha\leq j_{0}}\overline{V_{\alpha,j_{0}}}\Big)\cap\overline{U^{\prime}_{j_{0}+1}}\Big)\Big)$.}$
On the other hand, it holds that $\rho_{k,j_{0}+1}=id$ on $\varphi_{\infty,j_{0}+1}\big(U_{j_{0}+1}\setminus\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}}\big)$. Therefore, we have
$\rho_{k,j_{0}+1}\xrightarrow[k\rightarrow\infty]{}id\quad\text{in $C^{\infty}\big(\varphi_{\infty,j_{0}+1}(\overline{U^{\prime}_{j_{0}+1}})\big)$.}$
Hence the condition (IV.62c) holds for $j=j_{0}+1$. In particular, $\rho_{k,j_{0}+1}$ is an immersion on $\varphi_{\infty,j_{0}+1}(\overline{V_{j_{0}+1,j_{0}+1}})=\varphi_{\infty,j_{0}+1}(\overline{U^{\prime}_{j_{0}+1}})$ for $k$ large enough. Also, since $\partial U^{\prime}_{j_{0}+1}$ is smooth, we have the following embedding (see [58, Theorem 5.8.4]):
$W^{1,\infty}(\varphi_{\infty,j_{0}+1}(U^{\prime}_{j_{0}+1}))\subset C^{0,1}(\varphi_{\infty,j_{0}+1}(\overline{U^{\prime}_{j_{0}+1}})).$
Consequently, there exists $k_{0}\in\mathbb{N}$ such that for any $k\geq k_{0}$ and $x,y\in\varphi_{\infty,j_{0}+1}(\overline{U^{\prime}_{j_{0}+1}})$ it holds
$|(\rho_{k,j_{0}+1}-id)(x)-(\rho_{k,j_{0}+1}-id)(y)|\leq\frac{1}{2}\,|x-y|.$
Hence we have
$|\rho_{k,j_{0}+1}(x)-\rho_{k,j_{0}+1}(y)|\geq\frac{1}{2}\,|x-y|.$
In particular, this implies $\rho_{k,j_{0}+1}$ is injective on $\varphi_{\infty,j_{0}+1}(\overline{U^{\prime}_{j_{0}+1}})$ for $k\geq k_{0}$. Moreover, since $V_{j_{0}+1,j_{0}+1}=U^{\prime}_{j_{0}+1}\Subset U_{j_{0}+1}$ and $\varphi_{k,j_{0}+1}\rightarrow\varphi_{\infty,j_{0}+1}$ in $C^{0}(\overline{U^{\prime}_{j_{0}+1}})$ as $k\rightarrow\infty$, for $k$ large enough we have
$\rho_{k,j_{0}+1}\big(\varphi_{\infty,j_{0}+1}(\overline{V_{j_{0}+1,j_{0}+1}})\big)\subset\varphi_{k,j_{0}+1}(U_{j_{0}+1}).$
This completes the induction. Finally, for $1\leq\alpha\leq l$, we take $V_{\alpha}\coloneqq V_{\alpha,l}$, the constructions of $\{\rho_{k,\alpha}\}$ and $\{V_{\alpha}\}$ are thus accomplished.
Now since $\Sigma=\bigcup\limits_{1\leq\alpha\leq l}V_{\alpha}$, we can choose smooth non-negative cut-off functions $\phi_{\alpha}\in C^{\infty}_{c}(\varphi_{\infty,\alpha}(V_{\alpha}))$ such that $\sum\limits_{\alpha=1}^{l}\phi_{\alpha}\circ\varphi_{\infty,\alpha}>0$ on $\Sigma$. Since $\varphi_{k,\alpha}\rightarrow\varphi_{\infty,\alpha}$ in $C^{0}(\overline{V_{\alpha}})$ as $k\rightarrow\infty$, we have $\phi_{\alpha}\in C_{c}^{\infty}(\varphi_{k,\alpha}(V_{\alpha}))$ for $k\in\mathbb{N}\cup\{\infty\}$ large enough. For such $k$ we define
$g^{\prime}_{k}\coloneqq\sum_{\alpha=1}^{l}\varphi_{k,\alpha}^{*}(\phi_{\alpha}\,g_{\text{std}}).$
Then $g_{k}^{\prime}$ is a smooth metric under the complex structure induced by $\vec{\Phi}_{k}$ as defined in Corollary IV.5. Since $\varphi^{*}_{k,\alpha}\,g_{\text{std}}$ is weakly conformal to $g_{k}$ on $V_{\alpha}$, we have $g_{k}^{\prime}$ is weakly conformal to $g_{k}$. Since we assumed $\tilde{g}$ and $g_{\infty}$ are weakly conformal, we obtain that $\varphi_{\infty,\alpha}$ is weakly conformal from $(V_{\alpha},g)$ to $(\varphi_{\infty,\alpha}(V_{\alpha}),g_{\text{std}})$. Hence $\varphi_{\infty,\alpha}$ is smooth by Corollary IV.5 (i). Let $1\leq\alpha\leq l$, $x\in V_{\alpha}$, then by (IV.60) and (IV.61) we have
$(\Psi_{k}^{*}g^{\prime}_{k})_{x}$ $=\Big((\varphi_{k,\alpha}^{-1}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha})^{*}\sum_{x\in V_{\beta}}\varphi_{k,\beta}^{*}(\phi_{\beta}\,g_{\text{std}})\Big)_{x}$
$=\Big(\sum_{x\in V_{\beta}}(\varphi_{k,\beta,\alpha}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha})^{*}(\phi_{\beta}\,g_{\text{std}})\Big)_{x}$
$\xrightarrow[k\to\infty]{}\Big(\sum_{x\in V_{\beta}}(\varphi_{\infty,\beta})^{*}(\phi_{\beta}\,g_{\text{std}})\Big)_{x}=(g^{\prime}_{\infty})_{x}\;\text{smoothly in $x$}.$
Consequently, we can apply Lemma IV.18 to the sequence $\{\Psi_{k}^{*}g_{k}^{\prime}\}$ (denote $\Psi_{\infty}\coloneqq id$) to obtain a sequence of Riemannian metrics $h_{k}^{\prime}$} of constant Gaussian curvature 1,11,-1 or 0 such that $h_{k}^{\prime}$ is conformal to $\Psi_{k}^{*}g_{k}^{\prime}$ and $h_{k}^{\prime}$ converges to $h_{\infty}^{\prime}$ smoothly. Finally, we take $h_{k}\coloneqq(\Psi_{k}^{-1})^{*}h_{k}^{\prime}$, then $\Psi_{k}^{*}h_{k}$ converges to $h_{\infty}$ smoothly, $h_{k}$ is weakly conformal to $g_{k}^{\prime}$ and hence $g_{k}$. By the definition of $\Psi_{k}$, we also have $h_{k}$ is smooth in the complex structure induced by $\vec{\Phi}_{k}$ (i.e. $(\varphi_{k,\alpha}^{-1})^{*}h_{k}$ is smooth), and $h_{k}$ has the same constant Gaussian curvature as $h_{k}^{\prime}$. ∎
Remark IV.20.
Let $K>0$. Assume $\{g_{k}\}$ is replaced by a sequence of metrics $\{\tilde{g_{k}}\}$ such that for any $k$ and any coordinate chart $\varphi\colon U\colon\Sigma\rightarrow\mathbb{R}^{2}$, it holds
$\det((\tilde{g_{k}})_{ij})>K|((\tilde{g_{k}})_{ij})|^{2}\quad\text{a.e.,}$
where we denote
$(\tilde{g_{k}})_{ij}\coloneqq\tilde{g_{k}}\Big((d\varphi)^{-1}\Big(\frac{\partial}{\partial x^{i}}\Big),(d\varphi)^{-1}\Big(\frac{\partial}{\partial x^{j}}\Big)\Big).$
Assume $\tilde{g_{k}}\rightarrow\tilde{g}_{\infty}$ a.e. Then we can construct the conformal charts $\{\varphi_{k,\alpha}\}$ as in Remark IV.6, and by using the same argument as in Theorem IV.19, we find $\{h_{k}\}$, $\{\Psi_{k}\}$ satisfying Theorem IV.19 (i), (ii), (iv), but $\Psi_{k}$ may not be Lipschitz. Indeed, by [90, Proposition 4.36] and [91], we have $\varphi_{k,\alpha}\rightarrow\varphi_{\infty,\alpha}$, $\varphi^{-1}_{k,\alpha}\rightarrow\varphi^{-1}_{\infty,\alpha}$ in $W^{1,p}_{\textup{loc}}$ for some $p>2$, hence $\Psi_{k}\rightarrow id$, $\Psi^{-1}_{k}\rightarrow id$ in $W^{1,p}(\Sigma,\Sigma)$, in particular, in $C^{0}(\Sigma,\Sigma)$.
IV.4.3 Approximation results
The first goal of this section is to prove Theorem IV.23, namely that a $W^{1,\infty}$-immersion can be approximated by smooth immersions as soon as its first derivative belongs to $VMO$. Then we prove Theorem I.3 at the end. We define a weak $W^{1,\infty}$-immersion with value into manifolds.
Definition IV.21.
Let $N$ be an $n$-dimensional smooth manifold without boundary, $U\subset N$ be open and precompact, $M$ be an $\ell$-dimensional manifold smoothly embedded in $\mathbb{R}^{m}$. We say $f\in W^{1,\infty}_{\textup{imm}}(U,M)$ if $f\in W^{1,\infty}_{\textup{imm}}(U,\mathbb{R}^{m})$ and $f(x)\in M$ for any $x\in U$.
We estimate the mean oscillation of a composition and the distance with a smooth approximation.
Lemma IV.22.
Let $U\subset\mathbb{R}^{n}$ be open and bounded. Let $f\in W^{1,\infty}(U)$ with $\nabla f\in\text{VMO}(U)$. Let $\varphi\colon\overline{U}\rightarrow\overline{W}\subset\mathbb{R}^{n}$ be a $C^{1}$ diffeomorphism, where $W=\varphi(U)$.
(i) For any $r>0$, we have $\nabla(f\circ\varphi^{-1})\in VMO(W)$, and there exists a positive constant $C=C(n,\|\nabla\varphi\|_{L^{\infty}(U)},\|\nabla\varphi^{-1}\|_{L^{\infty}(W)})$ such that
$\beta_{r}(\nabla(f\circ\varphi^{-1}))\leq C\left(\beta_{r}(\nabla f)+\beta_{r}(\nabla\varphi)\|\nabla f\|_{L^{\infty}(U)}\right).$ (IV.63)
(ii) Let $\eta$ be the standard mollifier. For $\varepsilon>0$ we let $\eta_{\varepsilon}(x)=\varepsilon^{-n}\eta(\varepsilon^{-1}x)$. For $x\in U$ and $\varepsilon<\operatorname{dist}(x,U^{c})$, we define $f_{\varepsilon}(x)\coloneqq\int_{U}f(y)\,\eta_{\varepsilon}(x-y)\,dy$.
Let $K>0$ and assume in addition that $\partial U$ is Lipschitz. Then there exists a positive constant $C=C(\|\nabla\varphi\|_{L^{\infty}(U)},\|\nabla\varphi^{-1}\|_{L^{\infty}(W)},U,K)>0$ such that for any $x_{1},x_{0}\in U$ and $\varepsilon>0$ satisfying $|x_{1}-x_{0}|\leq K\varepsilon$ and $\varepsilon<\min\big(\operatorname{dist}(\varphi(x_{1}),\partial W),\operatorname{dist}(x_{0},\partial U)\big)$, there holds
$\begin{aligned} &\left|\nabla(f_{\varepsilon}\circ\varphi^{-1})(\varphi(x_{0}))-\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla(f\circ\varphi^{-1})\right|\\[2.84526pt] &\leq C\Big(\beta_{\varepsilon}(\nabla f)+\big(\omega_{\varepsilon}(\nabla\varphi)+\varepsilon\big)\|\nabla f\|_{L^{\infty}(U)}\Big),\end{aligned}$ (IV.64)
where $\omega_{\varepsilon}(\nabla\varphi)\coloneqq\sup\{|\nabla\varphi(x)-\nabla\varphi(y)|:x,y\in U,|x-y|\leq\varepsilon\}$.
Proof.
(i) Since $\varphi$ is quasiconformal, by Lemmas III.13 and III.14, we have $\nabla f\circ\varphi^{-1}$ is in $VMO(W)$. Moreover, there exists a constant $C=C\big(n,\|\nabla\varphi\|_{L^{\infty}(U)},\|\nabla\varphi^{-1}\|_{L^{\infty}(W)}\big)$ such that the following estimate holds
$\beta_{r}\big(\nabla f\circ\varphi^{-1}\big)\leq C\,\beta_{r}(\nabla f).$
Hence for every ball $B_{r}\subset W$, we have
$\fint_{B_{r}}\left|\nabla(f\circ\varphi^{-1})-\fint_{B_{r}}\nabla f\circ\varphi^{-1}\fint_{B_{r}}\nabla\varphi^{-1}\right|$
$\leq\ \fint_{B_{r}}\left|\left(\nabla f\circ\varphi^{-1}-\fint_{B_{r}}\nabla f\circ\varphi^{-1}\right)\nabla\varphi^{-1}\right|+\fint_{B_{r}}\left|\nabla f\circ\varphi^{-1}\left(\nabla\varphi^{-1}-\fint_{B_{r}}\nabla\varphi^{-1}\right)\right|$
$\leq\|\nabla\varphi^{-1}\|_{L^{\infty}(W)}\ \beta_{r}\big(\nabla f\circ\varphi^{-1})+\beta_{r}(\nabla\varphi^{-1})\|\nabla f\|_{L^{\infty}(U)}$
$\leq C\big(\beta_{r}(\nabla f)+\beta_{r}(\nabla\varphi)\|\nabla f\|_{L^{\infty}(U)}\big),$
The estimate (IV.63) is thus proved.
(ii) Let $y\in B_{\varepsilon}(\varphi(x_{1}))$, we have
$|\varphi^{-1}(y)-x_{1}|\leq\|\nabla\varphi^{-1}\|_{L^{\infty}(W)}\,|y-\varphi(x_{1})|\leq\varepsilon\,\|\nabla\varphi^{-1}\|_{L^{\infty}(W)}.$
Hence it holds that $|x_{0}-\varphi^{-1}(y)|\leq\varepsilon\,(K+\|\nabla\varphi^{-1}\|_{L^{\infty}(W)})$. Note that $(\nabla\varphi)^{-1}\in C^{0}(\overline{U})$ and for all $r>0$, we have
$\omega_{r}\big((\nabla\varphi)^{-1}\big)\leq C\,\omega_{r}(\nabla\varphi).$
Since $\partial U$ is Lipschitz, we can extend $(\nabla\varphi)^{-1}$ to $g\in C_{c}(\mathbb{R}^{n},\mathbb{R}^{n\times n})$ such that for any $r>0$,
$\omega_{r,\mathbb{R}^{n}}(g)\leq C\Big(\omega_{r}\big((\nabla\varphi)^{-1}\big)+r\,\|(\nabla\varphi)^{-1}\|_{L^{\infty}(U)}\Big)\leq C\Big(\omega_{r}(\nabla\varphi)+r\Big).$
In particular, this implies that there exists a constant $C=C(\|\nabla\varphi\|_{L^{\infty}(U)},\|\nabla\varphi^{-1}\|_{L^{\infty}(W)},U,K)>0$ such that for any $y\in B_{\varepsilon}(\varphi(x_{1}))$, it holds that
$\big\|\nabla\varphi^{-1}(y)-\nabla\varphi^{-1}(\varphi(x_{0}))\big\|$ $=\big\|\big(\nabla\varphi(\varphi^{-1}(y))\big)^{-1}-\big(\nabla\varphi(x_{0})\big)^{-1}\big\|$
$\leq\big(K+\|\nabla\varphi^{-1}\|_{L^{\infty}(W)}+1\big)\,\omega_{\varepsilon,\mathbb{R}^{n}}(g)$
$\leq C\big(\omega_{\varepsilon}(\nabla\varphi)+\varepsilon\big).$
Hence we have
$\begin{aligned} &\left|\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla(f\circ\varphi^{-1})-\nabla\varphi^{-1}(\varphi(x_{0}))\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}\right|\\[5.69054pt] &\leq C\big(\omega_{\varepsilon}(\nabla\varphi)+\varepsilon\big)\,\|\nabla f\|_{L^{\infty}(U)}.\end{aligned}$ (IV.65)
Since $\operatorname{dist}(\varphi(x_{1}),\partial W)>\varepsilon$, we have $\operatorname{dist}(x_{1},\partial U)>\varepsilon\,\|\nabla\varphi\|^{-1}_{L^{\infty}(U)}$. Now we can apply the proof of Riemann’s Theorem [162, Theorem 2] and Lemma III.13 to get
$\bigg|\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}-\fint_{B_{\varepsilon\,\|\nabla\varphi\|^{-1}_{L^{\infty}}}(x_{1})}\nabla f\bigg|\leq C\,\beta_{\varepsilon}(\nabla f).$ (IV.66)
Since $\partial U$ is Lipschitz, by Corollary III.15 we can extend $\nabla f$ to a function $\tilde{f}\in VMO(\mathbb{R}^{n},\mathbb{R}^{n})$ with
$\forall r>0,\qquad\beta_{r}(\tilde{f})\leq C\big(\beta_{r}(\nabla f)+r\,\|\nabla f\|_{L^{\infty}(U)}\big).$
Hence it holds that
$\begin{aligned} \bigg|\fint_{B_{\varepsilon\,\|\nabla\varphi\|^{-1}_{L^{\infty}}}(x_{1})}\nabla f-\fint_{B_{\varepsilon}(x_{0})}\nabla f\bigg|\leq C\,\beta_{\varepsilon}(\tilde{f})\leq C\big(\beta_{\varepsilon}(\nabla f)+\varepsilon\,\|\nabla f\|_{L^{\infty}(U)}\big).\end{aligned}$ (IV.67)
We also have that
$\begin{aligned} \left|\nabla f_{\varepsilon}(x_{0})-\fint_{B_{\varepsilon}(x_{0})}\nabla f\right|&=\left|\int_{B_{\varepsilon}(x_{0})}\eta_{\varepsilon}(x_{0}-x)\,\left(\nabla f(x)-\fint_{B_{\varepsilon}(x_{0})}\nabla f\,dx\right)\right|\\[2.84526pt] &\leq C\,\fint_{B_{\varepsilon}(x_{0})}\left|\nabla f-\fint_{B_{\varepsilon}(x_{0})}\nabla f\right|\\[2.84526pt] &\leq C\,\beta_{\varepsilon}(\nabla f).\end{aligned}$ (IV.68)
Combining (IV.66), (IV.67), (IV.68) we obtain by the triangle inequality
$\begin{aligned} &\left|\nabla\varphi^{-1}(\varphi(x_{0}))\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}-\nabla(f_{\varepsilon}\circ\varphi^{-1})(\varphi(x_{0}))\right|\\[2.84526pt] &\leq\|\nabla\varphi^{-1}\|_{L^{\infty}}\left|\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}-\nabla f_{\varepsilon}(x_{0})\right|\\[2.84526pt] &\leq C\big(\beta_{\varepsilon}(\nabla f)+\varepsilon\,\|\nabla f\|_{L^{\infty}(U)}\big).\end{aligned}$ (IV.69)
Finally, (LABEL:vmocov) follows from (LABEL:dvpdiff) and (LABEL:modiff4). ∎
We are now ready to establish an approximation result required for the proof of Theorem I.3. The underlying idea goes back to [178], see also [100, Remark 2.1]. For completeness, we present a detailed proof.
Theorem IV.23.
[…]
$\Lambda^{-1}|v|_{\tilde{g}}\leq|d(f_{k})(v)|_{\mathbb{R}^{m}}\leq\Lambda|v|_{\tilde{g}},\qquad\text{for all }\,v\in TU\,\text{ and all }\,k\in\mathbb{N}.$ (IV.70)
Moreover, if $f\in W^{2,p}(U)$ for any $1\leq p<\infty$, $f_{k}$ can be chosen to strongly converge to $f$ in $W^{2,p}(U)$.
Proof.
Since $\partial U$ is Lipschitz and $\in W^{1,\infty}(U), ff$ has a unique Lipschitz continuous representative defined on $\overline{U}$. In particular, $f(\overline{U})$ is compact. Since $M$ is smooth, there exists a bounded tubular neighborhood $W$ of $M$ with a retraction map $\pi\in C^{\infty}(W,M)$ such that $\pi(y)=y$ for any $y\in M$. Since $f(U)$ is compact, there exists a precompact neighborhood $W^{\prime}$ of $f(U)$ strictly contained in $W$. Since $\overline{U}$ is compact, there exist coordinate systems $\{(U_{j},\phi_{j})\}_{j=1}^{s}$ of $N$ such that $\overline{U}\subset\bigcup_{j=1}^{s}U_{j}$. Since $\partial U$ is Lipschitz, we can further assume that for each $U_{j}\nsubset U$, we have $\phi_{j}(U_{j})=B^{n}\subset\mathbb{R}^{n}$, $0\in\phi_{j}(\partial U)$ and, for some Lipschitz map $\gamma_{j}\colon\mathbb{R}^{n-1}\rightarrow\mathbb{R}$,
$\phi_{j}(U\cap U_{j})=\{(x_{1},\dots,x_{n})\in B^{n}:x_{n}>\gamma_{j}(x_{1},\dots,x_{n-1})\}.$
Let $\{\xi_{j}\}_{j=1}^{s}$ be a $C^{\infty}$ partition of unity such that $\operatorname{supp}(\xi_{j})\subset V_{j}\Subset U_{j}$ and $\sum_{j=1}^{s}\xi_{j}=1$ on $U$, where each $V_{j}$ is a precompact open subset of $U_{j}$. Let $\tilde{f}_{j}\coloneqq f\xi_{j}$, then $\tilde{f}_{j}\in W^{1,\infty}(U,\mathbb{R}^{k})$ and $d\tilde{f}_{j}\in VMO(U)$. Let
$0<\varepsilon<\min_{1\leq j\leq s}\left[\frac{1}{K_{j}+1}\min\Big(\text{dist}\big(\phi_{j}(\operatorname{supp}(\xi_{j})),\phi_{j}(\partial V_{j})\big),\text{dist}\big(\phi_{j}(V_{j}),\phi_{j}(U_{j})^{c}\big)\Big)\right].$ (IV.71)
Define $\eta_{\varepsilon}$ as in Lemma IV.22 (ii). For $1\leq j\leq s$, if $U_{j}\subset U$, for $x\in V_{j}$ we define
$F_{j,\varepsilon}(x)\coloneqq\int_{B_{\varepsilon}(\phi_{j}(x))}\tilde{f}_{j}(\phi_{j}^{-1}(z))\,\eta_{\varepsilon}(\phi_{j}(x)-z)\,dz.$
Then it holds that $F_{j,\varepsilon}\in C_{c}^{\infty}(V_{j},\mathbb{R}^{k})\subset C_{c}^{\infty}(U,\mathbb{R}^{k})$. If $U_{j}\cap\partial U\neq\emptyset$, let $K_{j}\coloneqq\text{Lip}(\gamma_{j})+2$. We denote $e_{n}\coloneqq(0,\dots,0,1)\in\mathbb{R}^{n}$. Then for any $z\in\phi_{j}(\overline{V}_{j}\cap\overline{U})$, $\varepsilon>0$, there holds
$B_{\varepsilon}(z+K_{j}\,\varepsilon\,e_{n})\cap\phi_{j}(U_{j}\cap\partial U)=\emptyset.$
By choice of $\varepsilon$ in (IV.71), it holds that
$B_{\varepsilon}(z+K_{j}\,\varepsilon\,e_{n})\subset\phi_{j}(U_{j}\cap U).$
We define for $x\in\phi_{j}(\overline{V}_{j}\cap\overline{U})$:
$F_{j,\varepsilon}(x)\coloneqq\int_{B_{\varepsilon}(0)}\tilde{f}_{j}\big(\phi_{j}^{-1}(\phi_{j}(x)+K_{j}\,\varepsilon\,e_{n}-z)\big)\,\eta_{\varepsilon}(z)\,dz.$
Thus, we have $F_{j,\varepsilon}=0$ in a neighborhood of $\partial V_{j}\cap U$. We extend the map $F_{j,\varepsilon}$ to a function in $C^{\infty}(\overline{U})$ by setting $F_{j,\varepsilon}=0$ on $\overline{U}\setminus V_{j}$.
We now define $F_{\varepsilon}\coloneqq\sum_{j=1}^{s}F_{j,\varepsilon}$. Then it holds that $F_{\varepsilon}\rightarrow f$ in $C(\overline{U})$ as $\varepsilon\rightarrow 0$, and $F_{\varepsilon}(\overline{U})\subset W^{\prime}$ for $\varepsilon$ small enough. Let $f_{\varepsilon}\coloneqq\pi\circ F_{\varepsilon}$. The family $\{f_{\varepsilon}\}_{\varepsilon>0}$ is bounded in $W^{1,\infty}(U,M)$ for $\varepsilon$ sufficiently small. Since $f=\pi\circ f$, the Sobolev convergence of $f_{\varepsilon}$ to $f$ follows from the same convergence of $F_{\varepsilon}$ to $f$. Now it suffices to prove that there exists a constant $\Lambda>0$ such that for all $\varepsilon$ small enough, we have
$\forall v\in T\Sigma,\qquad\Lambda^{-1}|v|_{\tilde{g}}\leq|d(f_{\varepsilon})(v)|_{\mathbb{R}^{m}}\leq\Lambda|v|_{\tilde{g}}$. (IV.72)
Let $1\leq j,j^{\prime}\leq s$, since $\overline{V}_{j}\cap\overline{V}_{j^{\prime}}\subset U_{j}\cap U_{j^{\prime}}$, we can find a pre-compact region $U_{j,j^{\prime}}$ with smooth boundary (see [52, Proposition 8.2.1]) such that
$\overline{V_{j}}\cap\overline{V_{j^{\prime}}}\subset U_{j,j^{\prime}}\Subset U_{j}\cap U_{j^{\prime}}.$
If $U_{j}\cap U_{j^{\prime}}=\emptyset$, we set $U_{j,j^{\prime}}=\emptyset$. Now we can apply (LABEL:vmocov) to the function $\tilde{f}_{j}\circ\phi_{j^{\prime}}^{-1}$ and transition map $\varphi=\phi_{j}\circ\phi_{j^{\prime}}^{-1}$ defined on $\phi_{j^{\prime}}(U_{j,j^{\prime}})$. In the case where $U_{j^{\prime}}\cap\partial U\neq\emptyset$, we extend the following map to a VMO function defined on an open subset of $\phi_{j^{\prime}}(U_{j^{\prime}})$ strictly containing $\phi_{j^{\prime}}(U_{j,j^{\prime}})$, as in the proof of Lemma IV.22:
$\nabla(\tilde{f}_{j}\circ\phi_{j^{\prime}}^{-1})\colon\phi_{j^{\prime}}(U_{j^{\prime}}\cap U)\rightarrow\mathbb{R}^{k\times n}.$
For $z\in\phi_{j}(V_{j}\cap\overline{U})$, we set $z^{\varepsilon}\coloneqq z$ if $V_{j}\subset U$, and $z^{\varepsilon}\coloneqq z+\varepsilon\,K_{j}\,e_{n}$ otherwise. Then we obtain the following estimate where the right-hand side is independent of $z$:
$\left|\nabla(F_{j^{\prime},\varepsilon}\circ\phi_{j}^{-1})(z)-\fint_{B_{\varepsilon}(z^{\varepsilon})}\nabla(\tilde{f}_{j^{\prime}}\circ\phi_{j}^{-1})\right|=\underset{\varepsilon\to 0}{o}(1).$
Consequently, it holds that
$\begin{aligned} &\left|\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-\fint_{B_{\varepsilon}(z^{\varepsilon})}\nabla(f\circ\phi_{j}^{-1})\right|\\ &\leq\sum_{j^{\prime}=1}^{s}\left|\nabla(F_{j^{\prime},\varepsilon}\circ\phi_{j}^{-1})(z)-\fint_{B_{\varepsilon}(z^{\varepsilon})}\nabla(\tilde{f}_{j^{\prime}}\circ\phi_{j}^{-1})\right|\\ &=\underset{\varepsilon\to 0}{o}(1).\end{aligned}$ (IV.73)
Since $df\in\text{VMO}(U)$, we also have
$\fint_{B_{\varepsilon}(z^{\varepsilon})}\left|\nabla(f\circ\phi_{j}^{-1})-\fint_{B_{\varepsilon}(z^{\varepsilon})}\nabla(f\circ\phi_{j}^{-1})\right|=\underset{\varepsilon\to 0}{o}(1).$
Together with (LABEL:osdiff), then we get the following estimate where the right-hand side is independent of $z\in\phi_{j}(V_{j}\cap\overline{U})$:
$\fint_{B_{\varepsilon}(z^{\varepsilon})}\left|\nabla(f\circ\phi_{j}^{-1})-\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)\right|=\underset{\varepsilon\to 0}{o}(1).$ (IV.74)
Since $f\in W_{\textup{imm}}^{1,\infty}(U,M)$, there exists a constant $C_{j}>0$ such that for a.e. $z\in\phi_{j}(V_{j})$ and for any $v\in\mathbb{S}^{n-1}$ it holds
$\big|d(f\circ\phi_{j}^{-1})_{z}(v)\big|\geq C_{j}.$ (IV.75)
Since $\overline{W^{\prime}}$ is compact and $\pi$ is smooth in an open set containing $\overline{W^{\prime}}$, the map $\nabla\pi\colon\overline{W^{\prime}}\rightarrow\mathbb{R}^{k\times k}$ is Lipschitz. We define
$\alpha\coloneqq\|\nabla\pi\|_{C^{0}(\overline{W^{\prime}})}+1+\text{Lip}(\nabla\pi|_{\overline{W^{\prime}}})\,\|\nabla(f\circ\phi_{j}^{-1})\|_{L^{\infty}(\phi_{j}(V_{j}\cap\overline{U}))}.$ (IV.76)
By (IV.74), there exists $\varepsilon_{0}>0$ such that for any $\varepsilon\in(0,\varepsilon_{0})$, we have on the one hand that $F_{\varepsilon}\in C^{\infty}(\overline{U},M)$ is well-defined with $F_{\varepsilon}(\overline{U})\subset W^{\prime}$. On the other hand for any $z\in\phi_{j}(\operatorname{supp}(\xi_{j})\cap\overline{U})$, there holds $B_{\varepsilon}(z^{\varepsilon})\subset\phi_{j}(V_{j}\cap\overline{U})$, and there exists $z^{\prime}\in B_{\varepsilon}(z^{\varepsilon})$ satisfying $f\circ\phi_{j}^{-1}$ is differentiable at $z^{\prime}$ with the following inequalities:
  • oscillation of $\nabla(f\circ\phi_{j}^{-1})$:
    $\big|\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-\nabla(f\circ\phi_{j}^{-1})(z^{\prime})\big|\leq(2\alpha)^{-1}C_{j}.$ (IV.77)
  • $C^{0}$ distance:
    $\begin{aligned} &\big|(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z^{\prime})\big|\\[5.69054pt] &\leq\big|(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z)\big|+\big|(f\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z^{\prime})\big|\\[5.69054pt] &\leq(2\alpha)^{-1}C_{j}.\end{aligned}$ (IV.78)
The last inequality follows from the facts that $f\circ\phi_{j}^{-1}$ is Lipschitz continuous on $\phi_{j}(V_{j}\cap U)$ and $\|F_{\varepsilon}-f\|_{C(\overline{U})}\rightarrow 0$ as $\varepsilon\rightarrow 0$.
Given a point $z^{\prime}$ where $f\circ\phi_{j}^{-1}$ is differentiable, it holds $d(f\circ\phi_{j}^{-1})_{z^{\prime}}(v)\in T_{f\circ\phi_{j}^{-1}(z^{\prime})}M$ for any $v\in\mathbb{R}^{n}$ since $f(U)\subset M$. In particular, we have
$\big(d\pi\circ d(f\circ\phi_{j}^{-1})\big)_{z^{\prime}}(v)=d(f\circ\phi_{j}^{-1})_{z^{\prime}}(v).$
Hence, it holds for any $z\in\phi_{j}(\operatorname{supp}(\xi_{j})\cap\overline{U})$, $0<\varepsilon<\varepsilon_{0}$, $v\in\mathbb{S}^{n-1}$ that
$\begin{aligned} &\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)\big|\\[5.69054pt] &\geq\ \big|d(f\circ\phi_{j}^{-1})_{z^{\prime}}(v)\big|-\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)-\big(d\pi\circ d(f\circ\phi_{j}^{-1})\big)_{z^{\prime}}(v)\big|.\end{aligned}$ (IV.79)
By the triangle inequality, we have for any $v\in\mathbb{S}^{n-1}$
$\begin{aligned} &\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)-\big(d\pi\circ d(f\circ\phi_{j}^{-1})\big)_{z^{\prime}}(v)\big|\\[8.53581pt] &\leq\big|\nabla\pi_{(F_{\varepsilon}\circ\phi_{j}^{-1})(z)}\big(\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})_{z}-\nabla(f\circ\phi_{j}^{-1})_{z^{\prime}}\big)(v)\big|\\[7.96677pt] &\quad\quad+\big|\big(\nabla\pi_{(F_{\varepsilon}\circ\phi_{j}^{-1})(z)}-\nabla\pi_{(f\circ\phi_{j}^{-1})(z^{\prime})}\big)\nabla(f\circ\phi_{j}^{-1})_{z^{\prime}}(v)\big|\\[8.53581pt] &\leq\big|\nabla\pi\big((F_{\varepsilon}\circ\phi_{j}^{-1})(z)\big)\big|\ \big|\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-\nabla(f\circ\phi_{j}^{-1})(z^{\prime})\big|\\[9.95845pt] &\quad\quad+\text{Lip}(\nabla\pi|_{\overline{W^{\prime}}})\ \big|(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z^{\prime})\big|\ \big|\nabla(f\circ\phi_{j}^{-1})(z^{\prime})\big|.\end{aligned}$ (IV.80)
Combining (IV.76)–(LABEL:c0dis) and (LABEL:dpitri), we obtain
$\begin{aligned} &\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)-\big(d\pi\circ d(f\circ\phi_{j}^{-1})\big)_{z^{\prime}}(v)\big|\\[7.11317pt] &\leq(2\alpha)^{-1}C_{j}\Big(\|\nabla\pi\|_{C(\overline{W^{\prime}})}+\text{Lip}(\nabla\pi|_{\overline{W^{\prime}}})\,\|\nabla(f\circ\phi_{j}^{-1})\|_{L^{\infty}(\phi_{j}(V_{j}\cap\overline{U}))}\Big)\\[7.11317pt] &\leq 2^{-1}C_{j}.\end{aligned}$ (IV.81)
By (IV.75), (LABEL:dpi>), (LABEL:<c_j) we finally conclude that for any $z\in\phi_{j}(\operatorname{supp}(\xi_{j})\cap\overline{U})$, $0<\varepsilon<\varepsilon_{0}$, $v\in\mathbb{S}^{n-1}$, it holds
$\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)\big|\geq 2^{-1}C_{j}.$
The estimate (IV.72) is obtained since $\overline{U}\subset\bigcup_{j=1}^{s}\operatorname{supp}(\xi_{j})$. ∎
We are now ready to prove Theorem I.3.
Proof of Theorem I.3.
Let $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$ be weakly conformal. Let $\tilde{g}$ be a reference Riemannian metric on $\Sigma$. By Theorem IV.23, there exists a sequence of $C^{\infty}$ immersions $\vec{\Phi}_{k}$ satisfying $\vec{\Phi}_{k}\rightarrow\vec{\Phi}$ in $W^{2,2}(\Sigma,\mathbb{R}^{m})$ and there exists a constant $\Lambda>0$ such that
$\Lambda^{-1}|X|_{\tilde{g}}\leq|d(\vec{\Phi}_{k})(X)|\leq\Lambda|X|_{\tilde{g}},\qquad\text{for all }X\in T\Sigma\,\text{ and all }\,k\in\mathbb{N}.$ (IV.82)
Write $\vec{\Phi}_{\infty}\coloneqq\vec{\Phi}$, and we construct $\{\Psi_{k}\}$, $\{h_{k}\}$ as in Theorem IV.19. Since each $\vec{\Phi}_{k}$ is smooth, from the construction we see that each $\Psi_{k}$ is also smooth, and we define $\vec{\Phi}_{k}^{\prime}\coloneqq\vec{\Phi}_{k}\circ\Psi_{k}$, $h_{k}^{\prime}\coloneqq\Psi_{k}^{*}h_{k}$, $g_{k}^{\prime}\coloneqq\vec{\Phi}_{k}^{\prime*}g_{\text{std}}$. Then we have $g_{k}^{\prime}=\Psi_{k}^{*}(\vec{\Phi}_{k}^{*}g_{\text{std}})$ is conformal to $h_{k}^{\prime}$ since $h_{k}$ is conformal to $\vec{\Phi}_{k}^{*}g_{\text{std}}$, and $h_{k}^{\prime}\rightarrow h$ in $C^{\infty}(\Sigma,T^{*}\Sigma\otimes T^{*}\Sigma)$. Moreover, by Theorem IV.19 (iii) we have $\vec{\Phi}_{k}^{\prime}\rightarrow\vec{\Phi}$ in $W^{2,2}(\Sigma,\mathbb{R}^{m})$. Write $g^{\prime}_{k}=e^{2\alpha_{k}}h_{k}^{\prime}$.
Now it remains to show that $\alpha_{k}\rightarrow\alpha$ in $C^{0}(\Sigma)$. Let $p\in\Sigma$. Since $h_{k}^{\prime}\rightarrow h$ in $C^{\infty}(\Sigma)$, by a similar argument as in Theorem IV.17 (replacing $\vec{e}_{i,k}$ by orthonormal frames on $T\Sigma$, and $d\vec{e}_{i,k}$ by covariant derivatives of these vector fields with respect to $h^{\prime}_{k}$), there exists a neighborhood $U$ of $p$ and $C^{\infty}$ diffeomorphisms $\varphi_{k}$ from $U$ to $\varphi_{k}(U)\subset\mathbb{R}^{2}$ such that for any $k\in\mathbb{N}\cup\{\infty\}$, it holds that $\varphi_{k}(p)=0$, $D^{2}\Subset\varphi_{k}(U)$, and $\varphi_{k}^{*}g_{\text{std}}$ is conformal to $h_{k}^{\prime}$ (write $h_{\infty}^{\prime}\coloneqq h$). Moreover, we have $\varphi_{k}\rightarrow\varphi_{\infty}$ in $C^{\infty}_{\textup{loc}}(U)$. Since $h_{k}^{\prime}$ is conformal to $g_{k}^{\prime}$, we also obtain that $\vec{\Phi}_{k}^{\prime}\circ\varphi_{k}^{-1}$ is conformal. Write
$(\varphi_{k}^{-1})^{*}h_{k}^{\prime}=e^{2\tilde{\lambda}_{k}}(dx_{1}^{2}+dx_{2}^{2}),\quad(\varphi_{k}^{-1})^{*}g_{k}^{\prime}=(\vec{\Phi}_{k}^{\prime}\circ\varphi^{-1}_{k})^{*}g_{\text{std}}=e^{2\lambda_{k}}(dx_{1}^{2}+dx_{2}^{2}).$
Then $\tilde{\lambda}_{k}\rightarrow\tilde{\lambda}_{\infty}$ in $C^{\infty}(\overline{D^{2}})$ since $\varphi_{k}^{-1}\rightarrow\varphi_{\infty}^{-1}$ in $C^{\infty}(\overline{D^{2}})$ and $h_{k}^{\prime}\rightarrow h_{\infty}^{\prime}=h$ in $C^{\infty}(\Sigma)$. Since $\{\lambda_{k}\}$ is bounded in $L^{\infty}(D^{2})$ by (IV.82) and Theorem IV.19 (v), and $\vec{\Phi}_{k}^{\prime}\circ\varphi_{k}^{-1}\rightarrow\vec{\Phi}\circ\varphi_{\infty}^{-1}$ in $W^{2,2}(D^{2},\mathbb{R}^{m})$, we have $\lambda_{k}\rightarrow\lambda_{\infty}$ in $W^{1,2}(D^{2})$. Now by applying the argument in Corollary III.11 to $\{\vec{\Phi}_{k}^{\prime}\circ\varphi_{k}^{-1}\}$, we have $\lambda_{k}\rightarrow\lambda_{\infty}$ in $W^{2,1}_{\textup{loc}}(D^{2})$. Since $\alpha_{k}=(\lambda_{k}-\tilde{\lambda}_{k})\circ\varphi_{k}$ on $U$, and $p$ is arbitrarily chosen on $\Sigma$, we conclude that $\alpha_{k}\rightarrow\alpha$ in $W^{2,1}(\Sigma)$, hence also in $C^{0}(\Sigma)$ (see [80, Theorem 3.3.4 & 3.3.10]). ∎
V Willmore surfaces
Motivated by the generalization of Willmore surfaces in higher dimensions, we provide in this section, a new proof of the regularity of Willmore surfaces which does not involve the choice of conformal coordinates. In Section V.1, we write the Euler–Langrange equation of Willmore surfaces in divergence form. In Section V.2, we prove the regularity of Willmore surfaces.
V.1 The Euler–Lagrange equation and conservation laws
In this section, we show that the Euler–Lagrange system of the Willmore functional can be written as a div-curl system. To do so, we need Noether theorem to rewrite it in divergence form.
Theorem V.1 (Noether’s theorem [154]).
Let $l\colon\mathbb{R}^{m}\times\mathbb{R}^{m\times n}\rightarrow\mathbb{R}$ for $l(z,p)$ being $C^{1}$ in $z$ and $C^{2}$ in $p$. Let $X$ be a tangent vector field on $\mathbb{R}^{m}$, $F(t,z)$ denote the flow of $X$ at time $t$ with $F(0,z)=z$. We say $X$ is an infinitesimal symmetry of $l$ if
[…]
Let $u$ be a critical point of $L(u)\coloneq\int_{B^{n}}l(u,\nabla u)\,d\mathcal{L}^{n}$, i.e. for any $\omega\in C^{\infty}_{c}(B^{n})$, it holds
$\left.\frac{d}{dt}\,L(u+t\omega)\right|_{t=0}=0.$
Then for any infinitesimal symmetry $X$ of $l$, we have
$\mbox{div}\,\left(\left(\frac{\partial l}{\partial p}\cdot X\right)\circ u\right)=0.$ (V.1)
In the same paper, Noether in fact considered higher-order Lagrangians of the form
$L=\int_{B^{n}}l(u,\nabla u,\dots,\nabla^{k}u)\,d\mathcal{L}^{n}\quad(k\geq 2),$
where $l$ is a smooth function. While the classical conservation law (V.1) may not hold for $k\geq 2$, symmetries of the Lagrangian density $l$ still generate conservation laws in the form $\text{div }J=0$, where $J$ is a Noether current expressible in terms of $u$ and its derivatives.
Let $\Sigma$ be a $2$-dimensional closed smooth orientable manifold. The notions of Gauss map, second fundamental form and mean curvature associated to $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(\Sigma,\mathbb{R}^{3})$ have been defined in Notation xvii.
Definition V.2.
Let $\vec{\Phi}\in W^{2,2}_{\text{imm}}(\Sigma,\mathbb{R}^{3})$. We define the Willmore energy:
$W(\vec{\Phi})\coloneqq\int_{\Sigma}\big|\vec{H}\big|^{2}\,d\textup{vol}_{g}.$
The map $\vec{\Phi}$ is said to be a critical point for $W$ if for any $\vec{w}\in C^{\infty}(\Sigma,\mathbb{R}^{3})$, there holds
$\left.\frac{d}{dt}\,W(\vec{\Phi}+t\,\vec{w})\right|_{t=0}=0.$
Such a critical point is called a weak Willmore immersion. Similarly, we can define weak Willmore immersions in $W^{2,2}_{\text{imm}}(D^{2},\mathbb{R}^{3})$.
Throughout Sections V and VI, we use the Einstein summation convention, and we leave out the symbols $\otimes$ for sections of $\bigwedge\mathbb{R}^{m}\otimes\bigwedge T^{*}\Sigma$ ($m\in\mathbb{N}^{+}$). For instance, we write $\partial_{i}\vec{\Phi}\,dx^{j}=\partial_{i}\vec{\Phi}\otimes dx^{j}$.
The Noether current associated to translations.
Now we compute the Euler-Lagrange equation satisfied by weak Willmore immersions. In fact, the divergence-form equation can be seen as a consequence of the pointwise invariance of $H^{2}\,d\textup{vol}_{g}$ by translations in the ambient space, as pointed out in [14] (see also [120]).
Theorem V.3 ([166, Theorem 1.5]).
A weak immersion $\vec{\Phi}\in W^{2,2}_{\text{imm}}(\Sigma,\mathbb{R}^{3})$ is a weak Willmore immersion if and only if the following equation holds in $\mathcal{D}^{\prime}\big(\Sigma,\mathbb{R}^{3}\otimes\bigwedge^{2}T^{*}\Sigma\big)$:
$d*_{g}\left(-2H\,d\vec{n}+d\vec{H}-H^{2}\,d\vec{\Phi}\right)=0.$ (V.2)
Proof.
Let $(x^{1},x^{2})$ be local coordinates for $\Sigma$ that maps onto $D^{2}$ and $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(D^{2},\mathbb{R}^{3})$. Let $\vec{w}\in C^{\infty}(\overline{D^{2}},\mathbb{R}^{3})$. We consider the variation
$\vec{\Phi}_{t}=\vec{\Phi}+t\vec{w}.$
Denote $g_{t}=g_{\vec{\Phi}_{t}}$, $\vec{n}_{t}=\vec{n}_{\vec{\Phi}_{t}}$, etc. Since $g^{ij}_{t}\,g_{jk,t}=\delta_{ik}$, taking derivative with respect to $t$, we have
$\begin{dcases}\left.\frac{d}{dt}\,g_{ij,t}\right|_{t=0}=\left.\frac{d}{dt}\,\partial_{i}\vec{\Phi}_{t}\right|_{t=0}\cdot\partial_{j}\vec{\Phi}+\left.\frac{d}{dt}\,\partial_{j}\vec{\Phi}_{t}\right|_{t=0}\cdot\partial_{i}\vec{\Phi}=\partial_{i}\vec{w}\cdot\partial_{j}\vec{\Phi}+\partial_{j}\vec{w}\cdot\partial_{i}\vec{\Phi},\\[5.69054pt] \left.\frac{d}{dt}\,g_{t}^{ij}\right|_{t=0}=-g^{ik}g^{\ell j}\left.\frac{d}{dt}\,g_{k\ell,t}\right|_{t=0}=-g^{ik}g^{\ell j}(\partial_{k}\vec{w}\cdot\partial_{\ell}\vec{\Phi}+\partial_{\ell}\vec{w}\cdot\partial_{k}\vec{\Phi}).\end{dcases}$ (V.3)
By Jacobi’s formula and writing $g=(g_{ij})$ for simplicity, it follows that
$\begin{aligned} \left.\frac{d}{dt}(\det g_{t})^{\frac{1}{2}}\right|_{t=0}&=\frac{1}{2}(\det g)^{\frac{1}{2}}\,\text{Tr}\left(g^{-1}\left.\frac{d}{dt}\,g_{t}\right|_{t=0}\right)\\ &=\frac{1}{2}(\det g)^{\frac{1}{2}}\,g^{ij}(\partial_{i}\vec{w}\cdot\partial_{j}\vec{\Phi}+\partial_{j}\vec{w}\cdot\partial_{i}\vec{\Phi})\\[5.69054pt] &=(\det g)^{\frac{1}{2}}\,g^{ij}\,\partial_{j}\vec{w}\cdot\partial_{i}\vec{\Phi}.\end{aligned}$ (V.4)
Since $|\vec{n}|=1$, it holds
$\vec{n}_{t}\cdot\frac{d}{dt}\,\vec{n}_{t}=\frac{1}{2}\,\frac{d}{dt}\,|\vec{n}_{t}|^{2}=0.$
Thus the variation of the Gauss map is given by
$\left.\frac{d}{dt}\,\vec{n}_{t}\right|_{t=0}=g^{ij}\left(\partial_{i}\vec{\Phi}\cdot\left.\frac{d}{dt}\,\vec{n}_{t}\right|_{t=0}\right)\partial_{j}\vec{\Phi}=-g^{ij}\left(\vec{n}\cdot\left.\frac{d}{dt}\,\partial_{i}\vec{\Phi}_{t}\right|_{t=0}\right)\,\partial_{j}\vec{\Phi}=-g^{ij}\,(\vec{n}\cdot\partial_{i}\vec{w})\,\partial_{j}\vec{\Phi}.$
We now compute the variation of the mean curvature:
$\begin{aligned} 2\left.\frac{d}{dt}\,H_{t}\right|_{t=0}&=g^{ij}\left.\frac{d}{dt}\,{\mathrm{I\!I}}_{ij,t}\right|_{t=0}+{\mathrm{I\!I}}_{ij}\left.\frac{d}{dt}\,g_{t}^{ij}\right|_{t=0}\\[2.84526pt] &=-g^{ij}\,\partial_{j}\vec{\Phi}\cdot\left.\frac{d}{dt}\,\partial_{i}\vec{n}_{t}\right|_{t=0}-g^{ij}\,\partial_{i}\vec{n}\cdot\partial_{j}\vec{w}+\mathrm{I\!I}_{ij}\left.\frac{d}{dt}\,g_{t}^{ij}\right|_{t=0}.\end{aligned}$ (V.5)
Thanks to the fact $\Delta_{g}\vec{\Phi}\perp\partial_{l}\vec{\Phi}$, we have
$\begin{aligned} (\det g)^{\frac{1}{2}}\,g^{ij}\,\partial_{j}\vec{\Phi}\cdot\left.\frac{d}{dt}\,\partial_{i}\vec{n}_{t}\right|_{t=0}&=-(\det g)^{\frac{1}{2}}\,g^{ij}\,\partial_{j}\vec{\Phi}\cdot\partial_{i}\big(g^{k\ell}(\vec{n}\cdot\partial_{k}\vec{w})\,\partial_{\ell}\vec{\Phi}\big)\\ &=-\partial_{i}\left(\big((\det g)^{\frac{1}{2}}\,g^{ij}\,\partial_{j}\vec{\Phi}\big)\cdot\big(g^{k\ell}(\vec{n}\cdot\partial_{k}\vec{w})\,\partial_{\ell}\vec{\Phi}\big)\right)\\[2.84526pt] &=-\partial_{i}\left((\det g)^{\frac{1}{2}}\,g^{ij}\,g_{j\ell}\,g^{k\ell}(\vec{n}\cdot\partial_{k}\vec{w})\right)\\[2.84526pt] &=-\partial_{i}\left((\det g)^{\frac{1}{2}}\,g^{ik}(\vec{n}\cdot\partial_{k}\vec{w})\right)\\[2.84526pt] &=-\partial_{i}\left((\det g)^{\frac{1}{2}}\,g^{ij}(\vec{n}\cdot\partial_{j}\vec{w})\right).\end{aligned}$ (V.6)
Since the second fundamental form is symmetric and $\partial_{i}\vec{n}\cdot\vec{n}=0$, we also have
$\begin{aligned} \mathrm{I\!I}_{ij}\left.\frac{d}{dt}\,g^{ij}_{t}\right|_{t=0}&=-2\,\mathrm{I\!I}_{ij}\,g^{ik}g^{\ell j}\,\partial_{k}\vec{w}\cdot\partial_{\ell}\vec{\Phi}\\ &=2\,\partial_{k}\vec{w}\cdot\left(g^{ik}(\partial_{i}\vec{n}\cdot\partial_{j}\vec{\Phi})\,g^{\ell j}\,\partial_{\ell}\vec{\Phi}\right)\\[5.69054pt] &=2\,\partial_{k}\vec{w}\cdot\big(g^{ik}\,\partial_{i}\vec{n}\big)\\[7.11317pt] &=2\,g^{ik}\,\partial_{k}\vec{w}\cdot\partial_{i}\vec{n}\\[7.11317pt] &=2\,g^{ij}\,\partial_{j}\vec{w}\cdot\partial_{i}\vec{n}.\end{aligned}$ (V.7)
Combing (V.4)–(V.7), we obtain the pointwise a.e. variation
$\begin{aligned} &\left.\frac{d}{dt}\big(H_{t}^{2}\mkern 1.0mu(\det g_{t})^{\frac{1}{2}}\big)\right|_{t=0}\\ &=2H\left.\frac{d}{dt}\,H_{t}\right|_{t=0}(\det g)^{\frac{1}{2}}+H^{2}\left.\frac{d}{dt}(\det g_{t})^{\frac{1}{2}}\right|_{t=0}\\ &=H\mkern 2.0mu\partial_{i}\left((\det g)^{\frac{1}{2}}\mkern 1.0mug^{ij}(\vec{n}\cdot\partial_{j}\vec{w})\right)+H\mkern 1.0mu(\det g)^{\frac{1}{2}}\,g^{ij}\mkern 1.0mu\partial_{i}\vec{n}\cdot\partial_{j}\vec{w}+H^{2}\mkern 1.0mu(\det g)^{\frac{1}{2}}\mkern 1.0mug^{ij}\mkern 1.0mu\partial_{j}\vec{w}\cdot\partial_{i}\vec{\Phi}.\end{aligned}$ (V.8)
For any $1\leq j\leq 2$, $a\in L^{\infty}\cap W^{1,2}(D^{2})$ and $f\in L^{2}(D^{2})$, we have
$\|a\mkern 2.0mu\partial_{j}f\|_{L^{1}+W^{-1,2}(D^{2})}$ $\leq\|\partial_{j}(af)\|_{W^{-1,2}(D^{2})}+\|f\mkern 2.0mu\partial_{j}a\|_{L^{1}(D^{2})}$
$\leq\|af\|_{L^{2}(D^{2})}+\|\partial_{j}a\|_{L^{2}(D^{2})}\|f\|_{L^{2}(D^{2})}$
$\leq\|a\|_{L^{\infty}\cap W^{1,2}(D^{2})}\mkern 1.0mu\|f\|_{L^{2}(D^{2})}.$
Hence, there exists a universal constant $C>0$ such that for all $a\in L^{\infty}\cap W^{1,2}(D^{2})$ and $T\in L^{1}+W^{-1,2}(D^{2})$, it holds that
$\|aT\|_{L^{1}+W^{-1,2}(D^{2})}\leq C\mkern 1.0mu\|a\|_{L^{\infty}\cap W^{1,2}(D^{2})}\mkern 1.0mu\|T\|_{L^{1}+W^{-1,2}(D^{2})}.$ (V.9)
By (V.8) and (V.9), we obtain in $W^{-1,2}+L^{1}(D^{2})$ that
$\begin{aligned} &\left.\frac{d}{dt}\big(H_{t}^{2}\mkern 1.0mu(\det g_{t})^{\frac{1}{2}}\big)\right|_{t=0}\\ &=-\partial_{i}H\,(\det g)^{\frac{1}{2}}\,g^{ij}(\vec{n}\cdot\partial_{j}\vec{w})+\partial_{i}\left(H\mkern 1.0mu(\det g)^{\frac{1}{2}}\mkern 1.0mug^{ij}(\vec{n}\cdot\partial_{j}\vec{w})\right)\\ &\quad+\left\langle H\,d\vec{n}+H^{2}\,d\vec{\Phi},d\vec{w}\right\rangle_{g}(\det g)^{\frac{1}{2}}\\ &=\left\langle-\vec{n}\,dH+H\,d\vec{n}+H^{2}\,d\vec{\Phi},d\vec{w}\right\rangle_{g}(\det g)^{\frac{1}{2}}+\partial_{i}\left(H\mkern 1.0mu(\det g)^{\frac{1}{2}}\mkern 1.0mug^{ij}(\vec{n}\cdot\partial_{j}\vec{w})\right).\end{aligned}$ (V.10)
We define
$\vec{V}\coloneqq\vec{n}\,dH-H\mkern 1.0mud\vec{n}-H^{2}\mkern 1.0mud\vec{\Phi}=-2H\mkern 1.0mud\vec{n}+d\vec{H}-H^{2}\mkern 1.0mud\vec{\Phi}.$ (V.11)
From (V.10) it follows that in $W^{-2,2}+W^{-1,1}(D^{2},\bigwedge^{2}T^{*}D^{2})$,
$\begin{aligned} \left.\frac{d}{dt}(H_{t}^{2}\,d\textup{vol}_{g_{t}})\right|_{t=0}&=(*_{g}\,\vec{V})\mathbin{\dot{\wedge}}d\vec{w}+\partial_{i}\left(H\mkern 1.0mu(\det g)^{\frac{1}{2}}\mkern 1.0mug^{ij}(\vec{n}\cdot\partial_{j}\vec{w})\right)dx^{1}\wedge dx^{2}\\ &=\vec{w}\cdot d*_{g}\vec{V}-d*_{g}\big(\vec{V}\cdot\vec{w}-\vec{H}\cdot d\vec{w}\big).\end{aligned}$ (V.12)
Using integration by parts and denoting by $\big\langle\cdot,\cdot\big\rangle$ the canonical pairing between $\mathcal{D}^{\prime}(D^{2},\mathbb{R}^{3}\otimes\bigwedge^{2}T^{*}D^{2})$ and $C_{c}^{\infty}(D^{2},\mathbb{R}^{3})$, for $\vec{w}\in C_{c}^{\infty}(D^{2},\mathbb{R}^{3})$ we then obtain [^7] We first prove (V.13) for immersions $\vec{\Phi}\in C^{\infty}(\overline{D^{2}},\mathbb{R}^{3})$. The general case then follows from Theorem IV.23 together with (V.8).
$\left.\frac{d}{dt}\,W(\vec{\Phi}_{t})\right|_{t=0}=\big\langle d*_{g}\vec{V},\vec{w}\big\rangle.$ (V.13)
Therefore, by a partition of unity argument, we have $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(\Sigma,\mathbb{R}^{3})$ is Willmore if and only if $d*_{g}\vec{V}=0$, i.e., the equation (V.2) holds in $\mathcal{D}^{\prime}\big(\Sigma,\mathbb{R}^{3}\otimes\bigwedge^{2}T^{*}\Sigma\big)$. ∎
To prove the regularity of weak Willmore immersions, by using local coordinates $(x^{1},x^{2})$, we can still assume $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(D^{2},\mathbb{R}^{3})$. By Definition IV.1, there exists $\Lambda\geq 1$ such that (see Notation xvii)
$\Lambda^{-1}|v|_{\mathbb{R}^{2}}^{2}\leq|d\vec{\Phi}_{x}(v)|_{\mathbb{R}^{3}}^{2}\leq\Lambda|v|_{\mathbb{R}^{2}}^{2},\qquad\text{for a.e. }x\in D^{2}\text{ and all }v\in T_{x}D^{2}.$ (V.14)
Defining $\vec{V}$ as in (V.11), we have
$*_{g}\,\vec{V}\in L^{1}+W^{-1,2}(D^{2},\mathbb{R}^{3}\otimes T^{*}D^{2}).$
By (V.2) and weak Poincaré lemma (see [53, Chapter I, Theorem 2.24]), there exists $\vec{L}\in\mathcal{D}^{\prime}(D^{2},\mathbb{R}^{3})$ such that
$d\vec{L}=*_{g}\,\vec{V}.$ (V.15)
In particular, it holds that $d\vec{L}\in W^{-1,2}+L^{1}(D^{2})$, and by Corollary III.5, we have $\vec{L}\in L^{2}_{\text{loc}}(D^{2})$. The quantity $\vec{L}$ alone is not sufficient to deduce some additional regularity since $\vec{L}$ does not possess additional derivative a priori. Using the pointwise invariance of $H^{2}\,d\textup{vol}_{g}$ by dilations and rotations, we deduce some conservation laws related to $\vec{L}$ and $\vec{\Phi}$.
The Noether currents associated to dilations and rotations.
Theorem V.4 ([168, Theorem 5.59]).
Let $\vec{\Phi}\in W^{2,2}_{\text{imm}}(D^{2},\mathbb{R}^{3})$ be a weak Willmore immersion. Then there exists $\vec{L}\in L^{2}_{\text{loc}}(D^{2},\mathbb{R}^{3})$ satisfying (V.15) [^8] Up to additive constants, this $\vec{L}$ is 1/2-1/2 of the one defined in [168, Theorem 5.59]. There is a typo in [168, Equation (5.214b)]: the ++ sign should be changed to -. and moreover, we have
$\begin{gathered}d\left(\vec{L}\cdot d\vec{\Phi}\right)=0,\quad d\left(\vec{L}\times d\vec{\Phi}+H\,d\vec{\Phi}\right)=0.\end{gathered}$ (V.17)
Proof.
We have already proved $\vec{L}\in L^{2}_{\text{loc}}(D^{2},\mathbb{R}^{3})$, hence it remains to prove the relations (V.17). We first consider the variation $\vec{\Phi}_{t}=(1+t)\vec{\Phi}$ with
$\vec{w}=\left.\frac{d}{dt}\mkern 1.0mu\vec{\Phi}_{t}\right|_{t=0}=\vec{\Phi}.$
Denote $H_{t}$ and $g_{t}$ as in the proof of Theorem V.3. Since $H_{t}^{2}\,dvol_{g_{t}}=H^{2}\,d\textup{vol}_{g}$, combining (V.12) and (V.15), we obtain that
$0$ $=\frac{d}{dt}(H_{t}^{2}\,d\textup{vol}_{g_{t}})\Big|_{t=0}$
$=-\mkern 1.0mud*_{g}\big(\vec{V}\cdot\vec{w}-\vec{H}\cdot d\vec{w}\big)$
$=-\mkern 1.0mud(d\vec{L}\cdot\vec{w})$
$=d\vec{L}\mathbin{\dot{\wedge}}d\vec{\Phi}$
$=d(\vec{L}\cdot d\vec{\Phi}).$
To prove the second conservation law in (V.17), for $\vec{a}\in\mathbb{R}^{3}$ we consider the variation $\vec{\Phi}_{t}$ satisfying
$\begin{dcases}\frac{d}{dt}\mkern 1.0mu\vec{\Phi}_{t}=\vec{a}\times\vec{\Phi}_{t},\qquad t\in\mathbb{R},\\[1.29167pt] \vec{\Phi}_{0}=\vec{\Phi}.\end{dcases}$
For this variation, the equations (V.8), (V.10), and (V.12) remain valid. Moreover, for each $t\in\mathbb{R}$, there exists $Q_{t}\in\text{SO}(3)$ depending only on $\vec{a}$ and $t$ such that $\vec{\Phi}_{t}=Q_{t}\circ\vec{\Phi}$ on $D^{2}$. Consequently, we have $H_{t}^{2}\,dvol_{g_{t}}=H^{2}\,d\textup{vol}_{g}$ for all $t\in\mathbb{R}$. It follows that
$0$ $=\frac{d}{dt}(H_{t}^{2}\,d\textup{vol}_{g_{t}})\Big|_{t=0}$
$=-\mkern 2.0mud*_{g}\big(\vec{V}\cdot\vec{w}-\vec{H}\cdot d\vec{w}\big)$
$=-\mkern 2.0mud\big(d\vec{L}\cdot(\vec{a}\times\vec{\Phi})-\vec{H}\cdot(\vec{a}\times*_{g}\,d\vec{\Phi})\mkern 1.0mu\big)$
$=\vec{a}\cdot d\big(d\vec{L}\times\vec{\Phi}-\vec{H}\times*_{g}\,d\vec{\Phi}\big)$
$=-\mkern 2.0mu\vec{a}\cdot d\big(\vec{L}\times d\vec{\Phi}+\vec{H}\times*_{g}\,d\vec{\Phi}\big).$
Since $\vec{n}\times*_{g}\,d\vec{\Phi}=d\vec{\Phi}$, and $\vec{a}\in\mathbb{R}^{3}$ is arbitrary, we then obtain
$d\big(\vec{L}\times d\vec{\Phi}+H\mkern 1.0mud\vec{\Phi}\big)=0.$
This completes the proof. ∎
By (V.17) and weak Poincaré lemma, there exist $S\in W^{1,2}_{\textup{loc}}(D^{2})$ and $\vec{R}\in W^{1,2}_{\textup{loc}}(D^{2},\mathbb{R}^{3})$ such that
$dS=\vec{L}\cdot d\vec{\Phi},\qquad\text{ and }\qquad d\vec{R}=\vec{L}\times d\vec{\Phi}+H\mkern 1.0mud\vec{\Phi}.$
From these relations, we derive a system on $S,\vec{R}$ which does not involve $\vec{L}$ anymore.
Theorem V.5 ([168, Theorem 5.60]).
Let $\vec{\Phi}\in W^{2,2}_{\text{imm}}(D^{2},\mathbb{R}^{3})$ be a weak Willmore immersion, and define $\vec{L}$ as in Theorem V.4. Then there exist $S\in W^{1,2}_{\text{loc}}(D^{2})$ and $\vec{R}\in W^{1,2}_{\text{loc}}(D^{2},\mathbb{R}^{3})$ satisfying
$dS=\vec{L}\cdot d\vec{\Phi},\qquad\text{ and }\qquad d\vec{R}=\vec{L}\times d\vec{\Phi}+H\,d\vec{\Phi}.$ (V.18)
Moreover, we have
$dS=-*_{g}d\vec{R}\cdot\vec{n},\qquad\text{ and }\qquad d\vec{R}=*_{g}\big(\vec{n}\times d\vec{R}+dS\ \vec{n}\big).$ (V.19)
Proof.
First, we prove the second equation in (V.19). Since $d\vec{\Phi}\times\vec{n}=*_{g}\,d\vec{\Phi}$, taking the cross product of $\vec{L}\times d\vec{\Phi}$ with $\vec{n}$, we obtain
$\vec{n}\times(\vec{L}\times d\vec{\Phi})$ $=-\vec{L}\times(d\vec{\Phi}\times\vec{n})-d\vec{\Phi}\times(\vec{n}\times\vec{L})$
$=-\vec{L}\times*_{g}\,d\vec{\Phi}+(\vec{n}\cdot d\vec{\Phi})\,\vec{L}-(\vec{L}\cdot d\vec{\Phi})\,\vec{n}$
$=-*_{g}(\vec{L}\times d\vec{\Phi})-dS\ \vec{n}.$
It follows that
$\vec{n}\times d\vec{R}+dS\ \vec{n}$ $=\vec{n}\times(\vec{L}\times d\vec{\Phi}+H\,d\vec{\Phi})+dS\ \vec{n}$
$=-*_{g}(\vec{L}\times d\vec{\Phi})+H\,\vec{n}\times d\vec{\Phi}$
$=-*_{g}(\vec{L}\times d\vec{\Phi}+H\,d\vec{\Phi})$
$=-*_{g}d\vec{R}.$
We then obtain the second equation in (V.19):
$d\vec{R}=*_{g}\big(\vec{n}\times d\vec{R}+dS\ \vec{n}).$
The first equation in (V.19) follows immediately:
$-*_{g}d\vec{R}\cdot\vec{n}=(\vec{n}\times d\vec{R}+dS\ \vec{n})\cdot\vec{n}=dS.$
By differentiating the relations (V.19), we obtain a $\mathrm{div}-\operatorname{curl}$ system in $\vec{R}$ and $S$.
Theorem V.6 ([168, Corollary 5.61]).
Let $\vec{\Phi}\in W^{2,2}_{\text{imm}}(D^{2},\mathbb{R}^{3})$ be a weak Willmore immersion, and define $\vec{L},S,\vec{R}$ as before. Then we have (see Notation xiii)
$\Delta_{g}S=*_{g}\,\big(d\,\vec{n}\mathbin{\dot{\wedge}}d\vec{R}\big),$ (V.20a)
$\Delta_{g}\vec{R}=*_{g}\,\big(d\vec{R}\times d\vec{n}+dS\wedge d\vec{n}\big),$ (V.20b)
$\Delta_{g}\vec{\Phi}=*_{g}\,\big(dS\wedge d\vec{\Phi}+d\vec{R}\overset{\times}{\wedge}d\vec{\Phi}\big).$ (V.20c)
Proof.
Equations (V.20a) and (V.20b) follow from the relations (V.19), by applying the operator $d^{*_{g}}$:
$\left\{\begin{aligned} &\Delta_{g}S=*_{g}\,d*_{g}dS=*_{g}\,d(d\vec{R}\cdot\vec{n})=*_{g}\,(d\vec{n}\mathbin{\dot{\wedge}}d\vec{R}),\\[8.53581pt] &\Delta_{g}\vec{R}=*_{g}\,d*_{g}d\vec{R}=-*_{g}d(\vec{n}\times d\vec{R}+dS\,\vec{n})=*_{g}\,(d\vec{R}\times d\vec{n}+dS\wedge d\vec{n}).\end{aligned}\right$.
Thanks to (V.18), we have
$d\vec{R}\overset{\times}{\wedge}d\vec{\Phi}$ $=(\vec{L}\times d\vec{\Phi}+H\,d\vec{\Phi})\overset{\times}{\wedge}d\vec{\Phi}$
$=d\vec{\Phi}\wedge(\vec{L}\cdot d\vec{\Phi})-(d\vec{\Phi}\mathbin{\dot{\wedge}}d\vec{\Phi})\,\vec{L}+H\,d\vec{\Phi}\overset{\times}{\wedge}d\vec{\Phi}$
$=d\vec{\Phi}\wedge dS+2H(\partial_{1}\vec{\Phi}\times\partial_{2}\vec{\Phi})\,dx^{1}\wedge dx^{2}$
$=d\vec{\Phi}\wedge dS+2\vec{H}*_{g}1$
$=-dS\wedge d\vec{\Phi}+(\Delta_{g}\vec{\Phi})*_{g}1.$
This is (V.20c). ∎
The Noether current associated to inversions.
Although $H^{2}\,d\textup{vol}_{g}$ is not pointwise invariant by inversions, the Lagrangian $(H^{2}-K)\,d\textup{vol}_{g}$ is; see for instance [34]. In this section, we derive the associated conservation law (V.28), which is equivalent to (V.20c).
Let $\vec{\Phi}\colon D^{2}\to\mathbb{R}^{3}$ be a Willmore immersion. As in [120], we consider the variation $\vec{\Phi}_{t}\coloneqq\varphi_{t}\circ\vec{\Phi}$ where, for a given $\vec{a}\in\mathbb{R}^{3}$, the map $\varphi_{t}$ is given by
$\varphi_{t}(x)\coloneqq\frac{\frac{x}{|x|^{2}}+t\vec{a}}{\big|\frac{x}{|x|^{2}}+t\vec{a}\big|^{2}}.$
Then we have
$\vec{w}\coloneqq\left.\frac{d}{dt}\,\vec{\Phi}_{t}\right|_{t=0}=|\vec{\Phi}|^{2}\vec{a}-2\mkern 1.0mu(\vec{\Phi}\cdot\vec{a})\mkern 1.0mu\vec{\Phi}.$
Since $\vec{\Phi}$ is Willmore, by (V.12) and (V.15), we obtain that
$\left.\frac{d}{dt}(H_{t}^{2}\,d\textup{vol}_{g_{t}})\right|_{t=0}=\mkern 2.0mud*_{g}\big(\vec{H}\cdot d\vec{w}-\vec{V}\cdot\vec{w}\big).$ (V.21)
Denote by $\text{adj}(\mathrm{I\!I}_{t})$ the adjugate matrix of $(\mathrm{I\!I}_{ij,t})$, i.e. the matrix satisfying the pointwise identity
$\text{adj}(\mathrm{I\!I}_{t})^{ij}\mkern 1.0mu\mathrm{I\!I}_{jk,t}=\det(\mathrm{I\!I}_{ij,t})\mkern 1.0mu\delta^{i}_{k}.$
We have
$\begin{aligned} \left.\frac{d}{dt}\det(\mathrm{I\!I}_{ij,t})\right|_{t=0}&=\text{adj}(\mathrm{I\!I})^{ij}\left.\frac{d}{dt}\mathrm{I\!I}_{ij,t}\right|_{t=0}\\ &=\text{adj}(\mathrm{I\!I})^{ij}\Big(\partial_{i}\vec{\Phi}\cdot\partial_{j}\big(g^{k\ell}\,(\vec{n}\cdot\partial_{k}\vec{w})\,\partial_{\ell}\vec{\Phi}\big)-\partial_{i}\vec{n}\cdot\partial_{j}\vec{w}\Big).\end{aligned}$ (V.22)
Hence by (V.4),
$\left.\frac{d}{dt}\big(K_{t}\det(g_{t})^{\frac{1}{2}}\big)\right|_{t=0}$ $=\left.\frac{d}{dt}\Big(\det(\mathrm{I\!I}_{ij,t})\,(\det g_{t})^{-\frac{1}{2}}\Big)\right|_{t=0}$
$=(\det g)^{-\frac{1}{2}}\,\text{adj}(\mathrm{I\!I})^{ij}\Big(\partial_{i}\vec{\Phi}\cdot\partial_{j}\big(g^{k\ell}\,(\vec{n}\cdot\partial_{k}\vec{w})\,\partial_{\ell}\vec{\Phi}\big)-\partial_{i}\vec{n}\cdot\partial_{j}\vec{w}\Big)$
$\quad-\det(\mathrm{I\!I}_{ij})\,(\det g)^{-\frac{1}{2}}g^{ij}\mkern 1.0mu\partial_{j}\vec{w}\cdot\partial_{i}\vec{\Phi}.$
We compute
$(\det g)^{-\frac{1}{2}}\,\text{adj}(\mathrm{I\!I})^{ij}\partial_{i}\vec{n}\cdot\partial_{j}\vec{w}$ $=-(\det g)^{-\frac{1}{2}}\,\text{adj}(\mathrm{I\!I})^{ij}(g^{k\ell}\mathrm{I\!I}_{ik}\partial_{\ell}\vec{\Phi})\cdot\partial_{j}\vec{w}$
$=-(\det g)^{-\frac{1}{2}}\det(\mathrm{I\!I}_{ij})g^{ij}\,\partial_{i}\vec{\Phi}\cdot\partial_{j}\vec{w}.$
Then (V.22) implies
$\left.\frac{d}{dt}\big(K_{t}\det(g_{t})^{\frac{1}{2}}\big)\right|_{t=0}=(\det g)^{-\frac{1}{2}}\,\text{adj}(\mathrm{I\!I})^{ij}\partial_{i}\vec{\Phi}\cdot\partial_{j}\big(g^{k\ell}\,(\vec{n}\cdot\partial_{k}\vec{w})\,\partial_{\ell}\vec{\Phi}\big).$ (V.23)
We will now show that
$\partial_{j}\Big((\det g)^{-\frac{1}{2}}\,\text{adj}(\mathrm{I\!I})^{ij}\partial_{i}\vec{\Phi}\Big)\cdot\partial_{\ell}\vec{\Phi}=0.$ (V.24)
Define $A_{g}\colon T_{x}^{*}(D^{2})\to T_{x}^{*}(D^{2})$ by
$A_{g}(dx^{i})\coloneqq(\det g)^{-1}\text{adj}(\mathrm{I\!I})^{ik}g_{kj}\,dx^{j}.$
Denoting by $g$, $g^{-1}$ the matrix $(g_{ij})$ and $(g^{ij})$ respectively, we have
$(\det g)^{-1}\text{adj}(\mathrm{I\!I})^{ik}g_{kj}$ $=(\det g)^{-1}\big(\text{adj}(\mathrm{I\!I})g\big)^{i}_{j}$
$=\big(\text{adj}(g^{-1}\mathrm{I\!I})\big)^{i}_{j}$
$=\text{Tr}(g^{-1}\mathrm{I\!I})\delta^{i}_{j}-\big(g^{-1}\mathrm{I\!I}\big)^{i}_{j}$
$=2H\delta_{j}^{i}-g^{ik}\mathrm{I\!I}_{kj}.$
Then Lemma III.1 implies
$A_{g}(d\vec{\Phi})$ $=(\det g)^{-1}\text{adj}(\mathrm{I\!I})^{ik}g_{kj}\mkern 1.0mu\partial_{i}\vec{\Phi}\,dx^{j}$
$=2H\partial_{j}\vec{\Phi}\,dx^{j}-g^{ik}\mathrm{I\!I}_{kj}\mkern 1.0mu\partial_{i}\vec{\Phi}\,dx^{j}$
$=2Hd\vec{\Phi}+d\vec{n}$
$=*_{g}\,d\vec{n}\times\vec{n}.$
Hence we have
$\partial_{j}\Big((\det g)^{-\frac{1}{2}}\,\text{adj}(\mathrm{I\!I})^{ij}\partial_{i}\vec{\Phi}\Big)dx^{1}\wedge dx^{2}$ $=d*_{g}A_{g}(d\vec{\Phi})$
$=d(\vec{n}\times d\vec{n})$
$=d\vec{n}\times d\vec{n}$
$=K\vec{n}\,d\textup{vol}_{g}.$
Thus, the identity (V.24) is proved. Since $\varphi_{t}$ is conformal, we have by pointwise conformal invariance that $(H_{t}^{2}-K_{t})\,d\textup{vol}_{g_{t}}=(H^{2}-K)\,d\textup{vol}_{g}$. Then by combining (V.21)–(V.24), we obtain that
$\begin{aligned} 0&=\left.\frac{d}{dt}\Big((H_{t}^{2}-K_{t})\,d\textup{vol}_{g_{t}}\Big)\right|_{t=0}\\ &=d*_{g}\Big(\vec{H}\cdot d\vec{w}-\vec{w}\cdot\vec{V}\Big)-\partial_{j}\Big((\det g)^{-\frac{1}{2}}\,\text{adj}(\mathrm{I\!I})^{ij}\vec{n}\cdot\partial_{i}\vec{w}\Big)dx^{1}\wedge dx^{2}\\ &=d*_{g}\Big(\vec{H}\cdot d\vec{w}+*_{g}\,d\vec{L}\cdot\vec{w}-A_{g}(\vec{n}\cdot d\vec{w})\Big).\end{aligned}$ (V.25)
By the expression $\vec{w}=|\vec{\Phi}|^{2}\vec{a}-2\mkern 1.0mu(\vec{\Phi}\cdot\vec{a})\mkern 1.0mu\vec{\Phi}$, we have
$\vec{n}\cdot d\vec{w}$ $=2\mkern 1.0mu\vec{n}\cdot\Big((d\vec{\Phi}\cdot\vec{\Phi})\mkern 1.0mu\vec{a}-(d\vec{\Phi}\cdot\vec{a})\mkern 1.0mu\vec{\Phi}\Big)-2\mkern 1.0mu(\vec{\Phi}\cdot\vec{a})\mkern 1.0mu\vec{n}\cdot d\vec{\Phi}$
$=2\mkern 1.0mu\vec{a}\cdot\Big((d\vec{\Phi}\cdot\vec{\Phi})\vec{n}-(\vec{n}\cdot\vec{\Phi})\mkern 1.0mud\vec{\Phi})\Big)$
$=2\mkern 1.0mu\vec{a}\cdot\Big((d\vec{\Phi}\times\vec{n})\times\vec{\Phi}\Big)$
$=2\mkern 1.0mu\vec{a}\cdot(*_{g}\,d\vec{\Phi}\times\vec{\Phi}).$
We also have
$A_{g}(\vec{n}\cdot d\vec{w})$ $=2\mkern 1.0mu\vec{a}\cdot\Big(\big(A_{g}(d\vec{\Phi})\times\vec{n}\big)\times\vec{\Phi}\Big)$
$=2\mkern 1.0mu\vec{a}\cdot\Big(\big(*_{g}d\vec{n}\times\vec{n})\times\vec{n}\big)\times\vec{\Phi}\Big)$
$=-2\mkern 1.0mu\vec{a}\cdot(*_{g}\,d\vec{n}\times\vec{\Phi}).$
The identity (V.25) then implies
$\vec{a}\cdot d\Big(2\mkern 1.0muH\vec{\Phi}\times d\vec{\Phi}-2\mkern 2.0mud\vec{n}\times\vec{\Phi}-d\vec{L}|\vec{\Phi}|^{2}+2(d\vec{L}\cdot\vec{\Phi})\vec{\Phi}\Big)=0.$
Since $\vec{a}$ is arbitrarily chosen, we get
$d\Big(2H\vec{\Phi}\times d\vec{\Phi}-2\mkern 2.0mud\vec{n}\times\vec{\Phi}-d\vec{L}|\vec{\Phi}|^{2}+2(d\vec{L}\cdot\vec{\Phi})\vec{\Phi}\Big)=0$ (V.26)
Moreover, we have
$\begin{aligned} &-d\vec{L}|\vec{\Phi}|^{2}+2(d\vec{L}\cdot\vec{\Phi})\vec{\Phi}\\[4.30554pt] &=d\Big(2(\vec{L}\cdot\vec{\Phi})\vec{\Phi}-\vec{L}|\vec{\Phi}|^{2}\Big)+2\vec{L}\,(d\vec{\Phi}\cdot\vec{\Phi})-2(\vec{L}\cdot\vec{\Phi})d\vec{\Phi}-2(\vec{L}\cdot d\vec{\Phi})\vec{\Phi}\\ &=d\Big(2(\vec{L}\cdot\vec{\Phi})\vec{\Phi}-\vec{L}|\vec{\Phi}|^{2}\Big)+2\vec{\Phi}\times(\vec{L}\times d\vec{\Phi})-2dS\,\vec{\Phi}\\ &=d\Big(2(\vec{L}\cdot\vec{\Phi})\vec{\Phi}-\vec{L}|\vec{\Phi}|^{2}\Big)+2\vec{\Phi}\times d\vec{R}-2H\vec{\Phi}\times d\vec{\Phi}-2dS\,\vec{\Phi}.\end{aligned}$ (V.27)
Combining (V.26)–(V.27), we then obtain
$2\,d\Big(-d\vec{n}\times\vec{\Phi}+\vec{\Phi}\times d\vec{R}-dS\,\vec{\Phi}\Big)=0.$
It follows that
$d\Big(*_{g}d\vec{\Phi}-\vec{R}\times d\vec{\Phi}-S\,d\vec{\Phi}\Big)=0.$ (V.28)
This is the conservation law corresponding to inversions.
V.2 Regularity of Willmore surfaces without conformal coordinates
The goal of this section is to provide a new proof of the regularity of Willmore surfaces which does not involve the choice of conformal coordinates. The difficulty lies into proving the continuity of the Gauss map, see Theorem V.9. The higher regularity then follows from standard bootstrap argument from the fact that $\vec{\Phi}$ can be represented locally as a graph, see Corollary V.11.
Definition V.7 (Morrey spaces).
Let $1\leq p<\infty$, $0\leq\lambda\leq n$, for a measurable function $f\colon U\rightarrow\mathbb{R}$, we say $f\in M^{p,\lambda}(U)$ if
$\sup_{r>0,x_{0}\in U}r^{-\lambda}\int_{B_{r}(x_{0})\cap U}|f(y)|^{p}\,dy<\infty,$
and we define the Morrey norm
$\|f\|_{M^{p,\lambda}(U)}\coloneqq\left(\sup_{r>0,x_{0}\in U}r^{-\lambda}\int_{B_{r}(x_{0})\cap U}|f(y)|^{p}\,dy\right)^{\frac{1}{p}}.$
We start by proving some Morrey decay on the conserved quantities, which provides a Morrey decay on the mean curvature.
Theorem V.8.
Let $\vec{\Phi}\in W^{2,2}_{\text{imm}}(D^{2},\mathbb{R}^{3})$ be a weak Willmore immersion, and define $\vec{L},S,\vec{R}$ as before. Let $\Lambda$ be the constant in (V.14) associated to $\vec{\Phi}$. Then there exists $\gamma=\gamma(\Lambda)>0$ such that $\nabla S,\nabla\vec{R},\vec{H}\in M^{2,\gamma}_{\textup{loc}}(D^{2})$.
Proof.
By dilation and translation, it suffices to prove $\nabla S,\nabla\vec{R},\vec{H}\in M^{2,\gamma}(D_{1/2}(0))$. By definition of a weak immersion, the coefficients $a^{ij}\coloneqq g^{ij}(\det g)^{1/2}$ satisfies the uniform ellipticity condition (III.13). Fix a positive small number $\varepsilon_{0}$ which will be determined later. Since $\vec{\Phi}\in W^{2,2}(D^{2})$, we have $\nabla\vec{n}\in L^{2}(D^{2})$, hence there exists $r_{0}\in(0,\frac{1}{4})$ depending on $\epsilon_{0}$ and $\vec{\Phi}$ such that
$\sup_{p\in D_{1/2}}\int_{D_{r_{0}}(p)}|\nabla\vec{n}|^{2}\,dx^{1}\wedge dx^{2}<\varepsilon_{0}.$ (V.29)
Let $p\in D_{1/2}(0)$, $r\in(0,r_{0})$. By Lemma III.3, we can define $\Psi_{S}\in W^{1,2}_{0}(D_{r}(p))$ to be the unique solutions of (see Section III.1)
$\begin{cases}\partial_{i}(a^{ij}\,\partial_{j}\Psi_{S})=\nabla^{\perp}\vec{n}\cdot\nabla\vec{R}&\quad\mbox{in }D_{r}(p),\\[8.53581pt] \Psi_{S}=0&\quad\mbox{on }\partial D_{r}(p).\end{cases}$ (V.30)
We can also define $\vec{\Psi}_{\vec{R}}\in W^{1,2}_{0}(D_{r}(p))$ to be the unique solutions of
$\begin{cases}\partial_{i}(a^{ij}\,\partial_{j}\vec{\Psi}_{\vec{R}})=\nabla^{\perp}\vec{R}\times\nabla\vec{n}+\nabla\vec{n}\ \nabla^{\perp}S&\quad\mbox{in }D_{r}(p),\\[8.53581pt] \vec{\Psi}_{\vec{R}}=0&\quad\mbox{on }\partial D_{r}(p).\end{cases}$ (V.31)
Let $v_{S}\coloneqq S-\Psi_{S}$ and $\vec{v}_{\vec{R}}\coloneqq\vec{R}-\vec{\Psi}_{\vec{R}}$. We combine the relations (V.20a) and (V.30) together with
$(\det g)^{\frac{1}{2}}*_{g}\big(d\vec{n}\mathbin{\dot{\wedge}}d\vec{R}\big)=\nabla^{\perp}\vec{n}\cdot\nabla\vec{R}.$
We obtain
$\partial_{i}(a^{ij}\,\partial_{j}v_{S})=0\quad\mbox{ in }D_{r}(p).$ (V.32)
As a result, it holds
$\begin{aligned} \int_{D_{r}(p)}|dS|_{g}^{2}\,d\textup{vol}_{g}&=\int_{D_{r}(p)}|dv_{S}|_{g}^{2}\,d\textup{vol}_{g}+\int_{D_{r}(p)}|d\Psi_{S}|_{g}^{2}\,d\textup{vol}_{g}+2\int_{D_{r}(p)}\langle dv_{S},d\Psi_{S}\rangle_{g}\,d\textup{vol}_{g}\\ &\geq\int_{D_{r}(p)}|dv_{S}|_{g}^{2}\,d\textup{vol}_{g}+2\int_{D_{r}(p)}a^{ij}\,\partial_{j}v_{S}\,\partial_{i}\Psi_{S}\,dx^{1}\wedge dx^{2}\\ &=\int_{D_{r}(p)}|dv_{S}|_{g}^{2}\,d\textup{vol}_{g}-2\int_{D_{r}(p)}\partial_{i}(a^{ij}\,\partial_{j}v_{S})\Psi_{S}\,dx^{1}\wedge dx^{2}\\ &=\int_{D_{r}(p)}|dv_{S}|_{g}^{2}\,d\textup{vol}_{g}.\end{aligned}$ (V.33)
Similarly, we have the system
$\partial_{i}(a^{ij}\,\partial_{j}\vec{v}_{\vec{R}})=0\qquad\mbox{ in }D_{r}(p).$ (V.34)
From this, we deduce the estimate
$\int_{D_{r}(p)}|d\vec{R}|_{g}^{2}\,d\textup{vol}_{g}\geq\int_{D_{r}(p)}|d\vec{v}_{\vec{R}}|_{g}^{2}\,d\textup{vol}_{g}.$ (V.35)
Let $\rho\in(0,1)$ be determined later. It holds
$\int_{D_{\rho r}(p)}\left(|dS|_{g}^{2}+|d\vec{R}|_{g}^{2}\right)d\textup{vol}_{g}$
$\leq 2\int_{D_{\rho r}(p)}\left(|dv_{S}|_{g}^{2}+|d\vec{v}_{\vec{R}}|_{g}^{2}\right)d\textup{vol}_{g}+2\int_{D_{r}(p)}\left(|d\Psi_{S}|_{g}^{2}+|d\vec{\Psi}_{\vec{R}}|_{g}^{2}\right)d\textup{vol}_{g}.$
By (IV.1), (V.30), (V.31), (V.32), (V.34), Lemmas III.2 and III.3, there exist positive constants $C_{1},\,\alpha$ depending only on $\Lambda$ such that
$\int_{D_{\rho r}(p)}\left(|dS|_{g}^{2}+|d\vec{R}|_{g}^{2}\right)d\textup{vol}_{g}$
$\leq C_{1}\left(\rho^{2\alpha}\int_{D_{r}(p)}\left(|dv_{S}|_{g}^{2}+|d\vec{v}_{\vec{R}}|_{g}^{2}\right)d\textup{vol}_{g}+\|\nabla\vec{n}\|^{2}_{L^{2}(D_{r}(p))}\int_{D_{r}(p)}\left(|dS|_{g}^{2}+|d\vec{R}|_{g}^{2}\right)d\textup{vol}_{g}\right).$
From (V.29), (V.33) and (V.35), we deduce
$\begin{aligned} \int_{D_{\rho r}(p)}\left(|dS|_{g}^{2}+|d\vec{R}|_{g}^{2}\right)d\textup{vol}_{g}\leq C_{1}(\rho^{2\alpha}+\varepsilon_{0})\int_{D_{r}(p)}\left(|dS|_{g}^{2}+|d\vec{R}|_{g}^{2}\right)d\textup{vol}_{g}.\end{aligned}$ (V.36)
Now we choose
$\varepsilon_{0}\coloneqq\frac{1}{4C_{1}},\quad\rho\coloneqq\min\left\{\frac{1}{(4C_{1})^{1/(2\alpha)}},\frac{1}{2}\right\}.$ (V.37)
For any $p\in D_{1/2}(0)$, $s\in(0,1)$, let $k\in\mathbb{N}_{0}$ such that $\rho^{k+1}\leq s<\rho^{k}$. Thanks to (V.36) and (V.37) we have
$\int_{D_{sr_{0}}(p)}\left(|dS|_{g}^{2}+|d\vec{R}|_{g}^{2}\right)d\textup{vol}_{g}$ $\leq\int_{D_{\rho^{k}r_{0}}(p)}\left(|dS|_{g}^{2}+|d\vec{R}|_{g}^{2}\right)d\textup{vol}_{g}$
$\leq C(\Lambda)\,2^{-k-1}\int_{D_{r_{0}+\frac{1}{2}}}\left(|\nabla S|^{2}+|\nabla\vec{R}|^{2}\right)$
$\leq C(\Lambda)\,2^{-\log_{\rho}s}\int_{D_{\frac{3}{4}}}\left(|\nabla S|^{2}+|\nabla\vec{R}|^{2}\right)$
$=C(\Lambda)\,s^{-\log_{\rho}2}\int_{D_{\frac{3}{4}}}\left(|\nabla S|^{2}+|\nabla\vec{R}|^{2}\right).$
Consequently, if we let $\gamma\coloneqq-\log_{\rho}2\in(0,\infty)$ which depends on $\Lambda$ only, then we have
$\sup_{p\in D_{1/2}(0),r\leq r_{0}}r^{-\gamma}\int_{D_{r}(p)}\left(|\nabla S|^{2}+|\nabla\vec{R}|^{2}\right)<\infty.$
By (V.20c), we then have
$\sup_{p\in D_{1/2}(0),r\leq r_{0}}r^{-\gamma}\int_{D_{r}(p)}H^{2}<\infty.$
We now prove that the Gauss map is continuous.
Theorem V.9.
Let $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(D^{2},\mathbb{R}^{3})$ be a weak Willmore immersion satisfying (V.14). Then there exists $\tau=\tau(\Lambda)>0$ such that $\nabla\vec{n}\in M^{2,\tau}_{\textup{loc}}(D^{2})$. In particular, there exists $\alpha=\alpha(\Lambda)\in(0,1)$ such that $\vec{n}\in C^{0,\alpha}_{\textup{loc}}(D^{2})$.
Proof.
Let $a^{ij}\coloneqq g^{ij}(\det g)^{1/2}$. By Lemma III.1, we have
$2H\,d\vec{\Phi}=d\vec{n}+\vec{n}\times*_{g}\,d\vec{n}.$
We deduce that the following system holds in $\mathcal{D}^{\prime}(D^{2})$:
$\begin{aligned} \partial_{i}(a^{ij}\,\partial_{j}\vec{n})&=(\det g)^{\frac{1}{2}}*_{g}d*_{g}d\vec{n}\\[2.84526pt] &=(\det g)^{\frac{1}{2}}*_{g}d(\vec{n}\times d\vec{n}-2H*_{g}d\vec{\Phi})\\[2.84526pt] &=2\,\partial_{1}\vec{n}\times\partial_{2}\vec{n}-2\partial_{i}(a^{ij}\,H\,\partial_{j}\vec{\Phi}).\end{aligned}$ (V.38)
Fix a positive small number $\varepsilon_{0}$ which will be determined later. As in the proof of Theorem V.8, there exists $r_{0}\in(0,\frac{1}{4})$ depending on $\epsilon_{0}$ and $\vec{\Phi}$ such that
$\sup_{p\in D_{1/2}}\int_{D_{r_{0}}(p)}|\nabla\vec{n}|^{2}\,dx^{1}\wedge dx^{2}<\varepsilon_{0}.$
Let $p\in D_{1/2}(0)$, $r\in(0,r_{0})$. By Lemma III.3, we can define $\vec{\Psi}_{1}\in W^{1,2}_{0}(D_{r}(p))$ to be the unique solution of
$\begin{cases}\partial_{i}(a^{ij}\,\partial_{j}\vec{\Psi}_{1})=2\,\partial_{1}\vec{n}\times\partial_{2}\vec{n}\quad&\mbox{in }D_{r}(p),\\[8.53581pt] \vec{\Psi}_{1}=0\quad&\mbox{on }\partial D_{r}(p).\end{cases}$ (V.39)
We have the following a priori estimate:
$\|\nabla\vec{\Psi}_{1}\|^{2}_{L^{2}(D_{r}(p))}\leq C(\Lambda)\,\|\nabla\vec{n}\|^{4}_{L^{2}(D_{r}(p))}\leq C(\Lambda)\,\epsilon_{0}\,\|\nabla\vec{n}\|^{2}_{L^{2}(D_{r}(p))}.$
Furthermore, by Riesz representation theorem for Hilbert spaces, there exists a unique $\vec{\Psi}_{2}\in W^{1,2}_{0}(D_{r}(p))$ satisfying the estimate $\|\nabla\vec{\Psi}_{2}\|_{L^{2}(D_{r}(p))}\leq C(\Lambda)\,\|H\|_{L^{2}(D_{r}(p))}$ and the system
$\partial_{i}(a^{ij}\,\partial_{j}\vec{\Psi}_{2})=-2\,\partial_{i}(a^{ij}\,H\,\partial_{j}\vec{\Phi}).$ (V.40)
Let $\vec{\Psi}\coloneqq\vec{\Psi}_{1}+\vec{\Psi}_{2}\in W^{1,2}_{0}(D_{r}(p))$ and $\vec{v}\coloneqq\vec{n}-\vec{\Psi}$. By Theorem V.8, there exist two constants $C=C(\vec{\Phi})$ and $\gamma=\gamma(\Lambda)>0$ such that
$\|\nabla\vec{\Psi}\|^{2}_{L^{2}(D_{r}(p))}$ $\leq C(\Lambda)\big(\epsilon_{0}\,\|\nabla\vec{n}\|^{2}_{L^{2}(D_{r}(p))}+\|H\|_{L^{2}(D_{r}(p))}^{2}\big)$
$\leq C(\vec{\Phi})\left(\epsilon_{0}\,\|\nabla\vec{n}\|^{2}_{L^{2}(D_{r}(p))}+r^{\gamma}\right).$
Thanks to (V.39) and (V.40), we also have
$\partial_{i}(a^{ij}\,\partial_{j}\vec{v})=0.$
It follows that, as in the proof of Theorem V.8, it holds
$\int_{D_{r}(p)}|d\vec{n}|_{g}^{2}\,d\textup{vol}_{g}\geq\int_{D_{r}(p)}|d\vec{v}|_{g}^{2}\,d\textup{vol}_{g}.$
Let $\rho\in(0,1)$ be determined later. There exist constants $\alpha=\alpha(\Lambda)>0$ defined as in Lemma III.2, $C_{2}=C_{2}(\Lambda)>0$, $C_{3}=C_{3}(\vec{\Phi})>0$ such that
$\begin{aligned} \int_{D_{\rho r}(p)}|d\vec{n}|_{g}^{2}\,d\textup{vol}_{g}&\leq 2\int_{D_{\rho r}(p)}|d\vec{v}|_{g}^{2}\,d\textup{vol}_{g}+2\int_{D_{r}(p)}|d\vec{\Psi}|_{g}^{2}\,d\textup{vol}_{g}\\ &\leq C_{2}\left(\rho^{2\alpha}\int_{D_{r}(p)}|d\vec{v}|_{g}^{2}\,d\textup{vol}_{g}+\epsilon_{0}\,\|\nabla\vec{n}\|^{2}_{L^{2}(D_{r}(p))}+C(\vec{\Phi})\,r^{\gamma}\right)\\ &\leq C_{2}\,(\rho^{2\alpha}+\varepsilon_{0})\int_{D_{r}(p)}|d\vec{n}|_{g}^{2}\,d\textup{vol}_{g}+C_{3}\,r^{\gamma}.\end{aligned}$ (V.41)
Now we choose
$\varepsilon_{0}\coloneqq\frac{1}{4C_{2}},\quad\rho\coloneqq\min\left\{\frac{1}{(4C_{2})^{1/(2\alpha)}},\frac{1}{2^{1/\gamma}},\frac{1}{2}\right\}.$
For any $p\in D_{1/2}(0)$, $s\in(0,1)$, let $k\in\mathbb{N}_{0}$ such that $\rho^{k+1}\leq s<\rho^{k}$. Then by (V.41) it holds
$\int_{D_{sr_{0}}(p)}|d\vec{n}|_{g}^{2}\,d\textup{vol}_{g}$ $\leq\int_{D_{\rho^{k}r_{0}}(p)}|d\vec{n}|_{g}^{2}\,d\textup{vol}_{g}$
$\leq C_{3}\,r_{0}^{\gamma}\left(\rho^{\gamma(k-1)}+\frac{1}{2}\,\rho^{\gamma(k-2)}+\cdots+\frac{1}{2^{k-1}}\right)+C(\Lambda)\,2^{-k-1}\int_{D_{r_{0}+\frac{1}{2}}}|\nabla\vec{n}|^{2}.$
Since $\rho\leq\frac{1}{2}$, we obtain
$\int_{D_{sr_{0}}(p)}|d\vec{n}|_{g}^{2}\,d\textup{vol}_{g}$ $\leq C(\vec{\Phi})\bigg(2^{-\log_{\rho}s}\int_{D_{\frac{3}{4}}}|\nabla\vec{n}|^{2}+\frac{r_{0}^{\gamma}k}{2^{k-1}}\bigg)$
$\leq C(\vec{\Phi})\bigg(s^{-\log_{\rho}2}\int_{D_{\frac{3}{4}}}|\nabla\vec{n}|^{2}+r_{0}^{\gamma}\,\left(\frac{2}{3}\right)^{k-1}\bigg)$
$\leq C(\vec{\Phi})\,s^{\log_{\rho}\frac{2}{3}}.$
Therefore, if we let $\tau\coloneqq\log_{\rho}\frac{2}{3}\in(0,\infty)$ which depends on $\Lambda$ only, then we have
$\sup_{p\in D_{1/2}(0),r\leq r_{0}}r^{-\tau}\int_{D_{r}(p)}|\nabla\vec{n}|^{2}<\infty.$
The Hölder continuity of $\vec{n}$ then follows from standard knowledge of the Morrey–Campanato spaces (see for instance [150, Theorem 3.5.2] and [2]). ∎
If the Gauss map is continuous, then one can write the image of $\vec{\Phi}$ as a graph.
Lemma V.10.
Let $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(D^{2},\mathbb{R}^{m})$ ($m\geq 3$). Assume the Gauss map $\vec{n}$ is continuous, then there exist an open neighborhood $U\subset D^{2}$ of 0, a $W^{2,2}\cap W^{1,\infty}$ homeomorphism $\Psi\colon U\rightarrow\Psi(U)\subset\mathbb{R}^{2}$, and $f\in W^{2,2}(\Psi(U),\mathbb{R}^{m-2})$ such that upon rotating and relabeling the coordinate axes if necessary, we have
$\vec{\Phi}\circ\Psi^{-1}(x_{1},x_{2})=(x_{1},x_{2},f(x_{1},x_{2})),\qquad\text{for all }(x_{1},x_{2})\in\Psi(U).$
Thus $\vec{\Phi}$ is locally the graph of a $W^{2,2}$ map.
Proof.
Upon rotating and relabeling the coordinate axes, we assume $\vec{n}(0)=e_{3}\wedge\dots\wedge e_{m}$, where $(e_{1},\dots,e_{m})$ denotes the canonical oriented basis of $\mathbb{R}^{m}$. By (V.14), we have
$\Lambda^{-2}\leq\det(g_{ij})=|\partial_{1}\vec{\Phi}\wedge\partial_{2}\vec{\Phi}|^{2}\leq\Lambda^{2}.$ (V.42)
Since $\vec{n}$ is continuous, there exists a neighborhood $U_{1}$ of 0 such that $|\vec{n}-\vec{n}\mkern 1.0mu(0)|<\frac{1}{2}$ on $U_{1}$. By definition of the Gauss map, we have the following inequality on $U_{1}$:
$\left\langle\frac{\partial_{1}\vec{\Phi}\wedge\partial_{2}\vec{\Phi}}{|\partial_{1}\vec{\Phi}\wedge\partial_{2}\vec{\Phi}|},e_{1}\wedge e_{2}\right\rangle=1+\left\langle\star\,\vec{n}-\star\,\vec{n}\mkern 1.0mu(0),e_{1}\wedge e_{2}\right\rangle>\frac{1}{2}.$
Writing $\vec{\Phi}=(\Phi_{1},\dots,\Phi_{m})$, on the other hand we obtain
$\left\langle\frac{\partial_{1}\vec{\Phi}\wedge\partial_{2}\vec{\Phi}}{|\partial_{1}\vec{\Phi}\wedge\partial_{2}\vec{\Phi}|},e_{1}\wedge e_{2}\right\rangle=|\partial_{1}\vec{\Phi}\wedge\partial_{2}\vec{\Phi}|^{-1}\det(\partial_{i}\Phi_{j})_{1\leq i,j\leq 2}\leq\Lambda\det(\partial_{i}\Phi_{j})_{1\leq i,j\leq 2}.$
It follows that $\det(\partial_{i}\Phi_{j})_{1\leq i,j\leq 2}>(2\Lambda)^{-1}$ on $U_{1}$. Then by Lemma IV.7, there exists an open neighborhood $U\subset U_{1}$ of 0 such that $\Psi\coloneqq(\Phi_{1},\Phi_{2})$ is injective on $U$ with $\Psi^{-1}\in W^{1,\infty}\cap W^{2,2}(\Psi(U),\mathbb{R}^{2})$. We obtain that for any $(x_{1},x_{2})\in\Psi(U)$, there holds
$\vec{\Phi}\circ\Psi^{-1}(x_{1},x_{2})=(x_{1},x_{2},\Phi_{3}\circ\Psi^{-1}(x_{1},x_{2}),\dots,\Phi_{m}\circ\Psi^{-1}(x_{1},x_{2})).$
Consequently, a weak Willmore immersion is locally a graph, and we can prove smoothness under the graph coordinates.
Corollary V.11.
Let $\Sigma$ be a $2$-dimensional closed smooth manifold, $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(\Sigma,\mathbb{R}^{3})$. Then we have $\vec{\Phi}$ is a graph near $\vec{\Phi}(p)$ for any $p\in\Sigma$, and $\vec{\Phi}\in C^{\infty}$ under the graph coordinates.
Proof.
Let $p\in\Sigma$. Then by Theorem V.9 and Lemma V.10, upon affine transformation on $\mathbb{R}^{m}$ if necessary, we can without loss of generality assume $p=0\in\mathbb{R}^{2}$ and $\vec{\Phi}(x_{1},x_{2})=(x_{1},x_{2},f(x_{1},x_{2}))$ on $D^{2}$ for some function $f\in W^{2,2}(D^{2},\mathbb{R}^{m-2})$. Define $S,\vec{R}$ as before. Recall that by (V.38) and Theorem V.6, we have the following system of elliptic equations
$\Delta_{g}S=*_{g}\big(d\vec{n}\mathbin{\dot{\wedge}}d\vec{R}\big),$ (V.43a)
$\Delta_{g}\vec{R}=*_{g}\big(d\vec{R}\times d\vec{n}+dS\wedge d\vec{n}\big),$ (V.43b)
$2H=*_{g}\big(dS\wedge d\vec{\Phi}+d\vec{R}\overset{\times}{\wedge}d\vec{\Phi}\big),$ (V.43c)
$d*_{g}d\vec{n}=-d\big(2H*_{g}d\vec{\Phi}\big)+d\vec{n}\overset{\times}{\wedge}d\vec{n}.$ (V.43d)
Moreover, by Theorems V.8 and V.9, there exist $\tau\in(0,1)$ and $\alpha\in(0,1)$ such that
$H,\nabla\vec{R},\nabla S,\nabla\vec{n}\in M^{2,\tau}_{\textup{loc}}(D^{2}),\quad\vec{n}\in C_{\textup{loc}}^{0,\alpha}(D^{2}).$ (V.44)
Using the expression of $\vec{\Phi}$ by $f$, we have
$\vec{n}=\frac{(-\partial_{1}f,-\partial_{2}f,1)}{\sqrt{1+|\nabla f|^{2}}}.$ (V.45)
Hence by direct computation we see that $\nabla f,\nabla\vec{\Phi}\in C^{0,\alpha}_{\textup{loc}}(D^{2})$. In particular, $g=g_{\vec{\Phi}}\in C_{\textup{loc}}^{0,\alpha}(D^{2})$.
By [145, Eq. (1.5)] and standard estimates on the Riesz potential [1, Proposition 3.2(ii)], we obtain $M^{1,\tau}(D^{2})\hookrightarrow W^{-1,p}(D^{2})$ for any $p\in(2,\frac{2-\tau}{1-\tau})$. Now since $_{g}\big(d\vec{n}\mathbin{\dot{\wedge}}d\vec{R}\big)\in M^{1,\tau}_{\textup{loc}}(D^{2})$, we have $_{g}\big(d\vec{n}\mathbin{\dot{\wedge}}d\vec{R}\big)\in W^{-1,p}_{\textup{loc}}(D^{2})$ for some $p>2$. Applying [31, Theorem 3.1] to the equation (V.43a) then yields $\nabla S\in L^{p}_{\textup{loc}}(D^{2})$. Similarly, we have $\nabla\vec{R}\in L^{p}_{\textup{loc}}(D^{2})$, and hence by (V.43c), $H\in L^{p}_{\textup{loc}}(D^{2})$. Moreover, since $d(2H*_{g}d\vec{\Phi})\in W^{-1,p}_{\textup{loc}}(D^{2})$, combining (V.43d) and (V.44) with a similar argument as above implies $\nabla\vec{n}\in L^{p}_{\textup{loc}}(D^{2})$. It follows that
$d\vec{n}\mathbin{\dot{\wedge}}d\vec{R}\in\begin{cases}L^{p/2}_{\textup{loc}}(D^{2})\subset W^{-1,\frac{2p}{4-p}}_{\textup{loc}}(D^{2})\quad&\text{if }p<4,\\[8.53581pt] W^{-1,q}_{\textup{loc}}(D^{2})\text{ for any $4<q<\infty$}\quad&\text{otherwise}.\end{cases}$
Hence by (V.43a) and [31, Theorem 3.1] again, we have
$\nabla S\in\begin{cases}L_{\textup{loc}}^{(2/p-1/2)^{-1}}(D^{2})\quad&\text{if }p<4,\\[8.53581pt] L^{q}_{\textup{loc}}(D^{2})\quad&\text{otherwise.}\end{cases}$
The same also holds for $\nabla\vec{R},\,\nabla\vec{n},\,H$, and after finitely many iterations, we get $S,\vec{R},\vec{n}\in W^{1,q}_{\textup{loc}}(D^{2})$ for any $q<\infty$. By computing $\nabla\vec{n}$ in the expression (V.2), we get $f\in W^{2,q}_{\textup{loc}}(D^{2})$ hence $\vec{\Phi}\in W^{2,q}_{\textup{loc}}(D^{2})$.
Now return to the equations (V.43a)–(V.43d). The right hand side of each equation is in $L^{q}_{\textup{loc}}(D^{2})$ for any $q<\infty$, and since $S,\vec{R},\vec{n},g\in W^{1,q}_{\textup{loc}}(D^{2})$, by [38, Theorem 4.1] and the proof of [66, Theorem 8.8], we obtain $S,\vec{R},\vec{n}\in W^{2,q}_{\textup{loc}}(D^{2})$, and then $\vec{\Phi}\in W^{3,q}_{\textup{loc}}(D^{2})$… We finally get $\vec{\Phi}\in C^{\infty}$ under the graph coordinates. ∎
VI Generalized Willmore functionals for 4-dimensional submanifolds
In this part we implement the approach introduced by the third author in [164] for the Willmore energy in order to deduce conserved quantities for critical points to scaling invariant Lagrangians in 4 dimension. The motivation being that, with the help of this conservation laws, one can hope to develop a strategy to devise a proof of the smoothness of weak critical points. The interpretations by Bernard [14] of the conservation laws found in [164] as being the Noether currents associated to the generators of the invariance group has enlightened the field of the analysis of conformally invariant Lagrangians of immersions. The computations of these conservation laws in dimension 4 have been first realised by Bernard in [15]. Here, we present these computations for the simpler Lagrangien $\int|dH|_{g}^{2}\,d\textup{vol}_{g}$. The proof of the regularity of the critical points has been first given in a joint work of the three authors together with Bernard in [16].
Definition VI.1.
Let $\vec{\Phi}\in W^{3,2}_{\text{imm}}(B^{4},\mathbb{R}^{5})$. We define the energy
$E(\vec{\Phi})\coloneqq\frac{1}{2}\int_{B^{4}}|dH|_{g}^{2}\,d\textup{vol}_{g}.$
The map $\vec{\Phi}$ is said to be a weak critical point of $E$ if for any $\vec{w}\in C^{\infty}_{c}(B^{4},\mathbb{R}^{5})$, there holds
$\left.\frac{d}{dt}\,E(\vec{\Phi}+t\,\vec{w})\right|_{t=0}=0.$
The Noether current associated to translations.
We first compute the Euler–Lagrange equation satisfied by weak critical points of $E$. Similar to the 2-dimensional case, this divergence-form equation is also a consequence of the pointwise invariance of $|dH|_{g}^{2}\,d\textup{vol}_{g}$ by translations in the ambient space, see [15, Section A.2].
Lemma VI.2.
An immersion $\vec{\Phi}\in W^{3,2}_{imm}(B^{4},\mathbb{R}^{5})$ is a weak critical point of $E$ if and only if the following equation holds in $\mathcal{D}^{\prime}(B^{4},\mathbb{R}^{5}\otimes\bigwedge^{4}T^{*}B^{4})$:
$d*_{g}\Big(2\,\langle d\vec{\Phi},dH\rangle_{g}\,dH-\frac{1}{2}\,d(\vec{n}\,\Delta_{g}H)+\Delta_{g}H\,d\vec{n}-|dH|_{g}^{2}\,d\vec{\Phi}\Big)=0.$ (VI.1)
Proof.
Let $\vec{\Phi}\in W^{3,2}_{\textup{imm}}(B^{4},\mathbb{R}^{5})$ and $\vec{w}\in C^{\infty}(\overline{B^{4}},\mathbb{R}^{5})$. We consider the variation
$\vec{\Phi}_{t}=\vec{\Phi}+t\vec{w}.$
Similar to Section V.1, we denote $g_{t}\coloneqq g_{\vec{\Phi}_{t}}$, $\det g_{t}\coloneqq\det(g_{ij,\vec{\Phi}_{t}})$, etc. The same computation as in (V.3)–(V.7) implies
$\begin{dcases}\left.\frac{d}{dt}\,g_{t}^{ij}\right|_{t=0}=-g^{ik}g^{sj}(\partial_{k}\vec{w}\cdot\partial_{s}\vec{\Phi}+\partial_{s}\vec{w}\cdot\partial_{k}\vec{\Phi}),\\[2.84526pt] \left.\frac{d}{dt}(\det g_{t})^{\frac{1}{2}}\right|_{t=0}=(\det g)^{\frac{1}{2}}\,g^{ij}\,\partial_{j}\vec{w}\cdot\partial_{i}\vec{\Phi},\\[2.84526pt] 4\left.\frac{d}{dt}\,H_{t}\right|_{t=0}=(\det g)^{-\frac{1}{2}}\,\partial_{i}\left((\det g)^{\frac{1}{2}}\,g^{ij}(\vec{n}\cdot\partial_{j}\vec{w})\right)+g^{ij}\,\partial_{i}\vec{n}\cdot\partial_{j}\vec{w}.\end{dcases}$ (VI.2)
We obtain the pointwise a.e. variation
$\begin{aligned} &\frac{d}{dt}\big(|dH_{t}|_{g_{t}}^{2}\,d\textup{vol}_{g_{t}}\big)\Big|_{t=0}\\[2.15277pt] &=\left(\frac{d}{dt}\,g_{t}^{ij}\Big|_{t=0}\partial_{i}H\mkern 2.0mu\partial_{j}H+2\,\bigg\langle dH,d\Big(\frac{d}{dt}\mkern 2.0muH_{t}\Big|_{t=0}\Big)\bigg\rangle_{\!g}\right)d\textup{vol}_{g}+|dH|_{g}^{2}\mkern 2.0mu\frac{d}{dt}\,d\textup{vol}_{g_{t}}\Big|_{t=0}\\ &=-\mkern 1.0mu2\mkern 2.0mu\langle d\vec{\Phi},dH\rangle_{g}\cdot\left\langle dH,d\vec{w}\right\rangle_{g}d\textup{vol}_{g}-2\mkern 2.0mu(*_{g}\,dH)\wedge d\Big(\frac{d}{dt}\mkern 2.0muH_{t}\Big|_{t=0}\Big)\\[5.59721pt] &\quad+|dH|_{g}^{2}\,g^{ij}\mkern 1.0mu(\partial_{j}\vec{w}\cdot\partial_{i}\vec{\Phi})\,d\textup{vol}_{g}.\end{aligned}$ (VI.3)
Now for $a\in L^{\infty}\cap W^{2,2}(B^{4})$ and $f\in L^{2}(B^{4})$, we write
$a\,\partial_{i}\partial_{j}f=\partial_{i}\partial_{j}(af)-\partial_{i}(f\,\partial_{j}\mkern 1.0mua)-\partial_{j}(f\,\partial_{i}\mkern 1.0mua)+f\,\partial_{i}\partial_{j}\mkern 1.0mua.$
We have the following estimates:
$\begin{dcases}\big\|\partial_{i}\partial_{j}(af)\big\|_{W^{-2,2}(B^{4})}\leq\|af\|_{L^{2}(B^{4})}\leq\|a\|_{L^{\infty}(B^{4})}\mkern 1.0mu\|f\|_{L^{2}(B^{4})},\\[4.30554pt] \big\|\partial_{i}(f\,\partial_{j}\mkern 1.0mua)\big\|_{W^{-1,\frac{4}{3}}(B^{4})}\leq\|f\,\partial_{j}\mkern 1.0mua\|_{L^{\frac{4}{3}}(B^{4})}\leq\|\nabla a\|_{L^{4}(B^{4})}\mkern 1.0mu\|f\|_{L^{2}(B^{4})},\\[3.01385pt] \|f\,\partial_{i}\partial_{j}\mkern 1.0mua\|_{L^{1}(B^{4})}\leq\|f\|_{L^{2}(B^{4})}\|\nabla^{2}a\|_{L^{2}(B^{4})}.\end{dcases}$
Hence there exists a universal constant $C>0$ such that for all $a\in L^{\infty}\cap W^{2,2}(B^{4})$ and $T\in W^{-2,2}+L^{1}(B^{4})$, it holds that
$\|aT\|_{W^{-2,2}+L^{1}(B^{4})}\leq C\mkern 1.0mu\|a\|_{L^{\infty}\cap W^{2,2}(B^{4})}\|T\|_{W^{-2,2}+L^{1}(B^{4})}.$ (VI.4)
Combining (VI.2)–(VI.4), we obtain in $W^{-2,2}+L^{1}(B^{4},\bigwedge^{4}T^{*}B^{4})$ that
$\begin{aligned} &\frac{d}{dt}\big(|dH_{t}|_{g_{t}}^{2}\,d\textup{vol}_{g_{t}}\big)\Big|_{t=0}\\ &=-2\bigg(\langle d\vec{\Phi},dH\rangle_{g}\cdot\left\langle dH,d\vec{w}\right\rangle_{g}+\Delta_{g}H\mkern 1.0mu\frac{d}{dt}\mkern 1.0muH_{t}\Big|_{t=0}\bigg)\,d\textup{vol}_{g}+2\mkern 2.0mud\mkern 1.0mu\bigg(\frac{d}{dt}H_{t}\Big|_{t=0}*_{g}dH\bigg)\\[1.29167pt] &\quad+|dH|_{g}^{2}\,g^{ij}\mkern 1.0mu(\partial_{j}\vec{w}\cdot\partial_{i}\vec{\Phi})\,d\textup{vol}_{g}\\[1.29167pt] &=\Big(-2\mkern 2.0mu\langle d\vec{\Phi},dH\rangle_{g}\cdot\langle dH,d\vec{w}\rangle_{g}+\frac{1}{2}\mkern 1.0mu\partial_{i}(\Delta_{g}H)\mkern 1.0mug^{ij}\mkern 2.0mu\vec{n}\cdot\partial_{j}\vec{w}-\frac{1}{2}\mkern 1.0mu\Delta_{g}H\mkern 2.0mug^{ij}\mkern 2.0mu\partial_{i}\vec{n}\cdot\partial_{j}\vec{w}\Big)\,d\textup{vol}_{g}\\[1.29167pt] &\quad+d*_{g}\Big(-\frac{1}{2}\Delta_{g}H\,\vec{n}\cdot d\vec{w}+2\,\frac{d}{dt}H_{t}\Big|_{t=0}\,dH\Big)+|dH|_{g}^{2}\,g^{ij}\mkern 1.0mu(\partial_{j}\vec{w}\cdot\partial_{i}\vec{\Phi})\,d\textup{vol}_{g}\\[1.29167pt] &=\Big\langle-2\,\langle d\vec{\Phi},dH\rangle_{g}\,dH+\frac{1}{2}\mkern 1.0mu\vec{n}\,d(\Delta_{g}H)-\frac{1}{2}\mkern 1.0mu\Delta_{g}H\mkern 1.0mud\vec{n}+|dH|_{g}^{2}\,d\vec{\Phi},d\vec{w}\Big\rangle_{g}\,d\textup{vol}_{g}\\ &\quad+d*_{g}\Big(-\frac{1}{2}\Delta_{g}H\,\vec{n}\cdot d\vec{w}+2\,\frac{d}{dt}H_{t}\Big|_{t=0}\,dH\Big).\end{aligned}$ (VI.5)
We define
$\begin{aligned} \vec{V}&\coloneqq 2\,\langle d\vec{\Phi},dH\rangle_{g}\,dH-\frac{1}{2}\,d(\vec{n}\,\Delta_{g}H)+\Delta_{g}H\,d\vec{n}-|dH|_{g}^{2}\,d\vec{\Phi}\\ &=2\,\langle d\vec{\Phi},dH\rangle_{g}\,dH-\frac{1}{2}\mkern 1.0mu\vec{n}\,d(\Delta_{g}H)+\frac{1}{2}\mkern 1.0mu\Delta_{g}H\mkern 1.0mud\vec{n}-|dH|_{g}^{2}\,d\vec{\Phi}.\end{aligned}$ (VI.6)
Then we have $\vec{V}\in W^{-2,2}+L^{1}(B^{4},\mathbb{R}^{5}\otimes T^{*}B^{4})$. From (VI.5) it follows that in $W^{-3,2}+W^{-1,1}(B^{4},\bigwedge^{4}T^{*}B^{4})$,
$\begin{aligned} \frac{d}{dt}\big(|dH_{t}|_{g_{t}}^{2}\,d\textup{vol}_{g_{t}}\big)\Big|_{t=0}&=(*_{g}\,\vec{V})\mathbin{\dot{\wedge}}d\vec{w}+d*_{g}\Big(-\frac{1}{2}\Delta_{g}H\,\vec{n}\cdot d\vec{w}+2\,\frac{d}{dt}H_{t}\Big|_{t=0}\,dH\Big)\\ &=\vec{w}\cdot d*_{g}\vec{V}-d*_{g}\Big(\vec{V}\cdot\vec{w}+\frac{1}{2}\mkern 1.0mu\Delta_{g}H\,\vec{n}\cdot d\vec{w}-2\,\frac{d}{dt}H_{t}\Big|_{t=0}\,dH\Big).\end{aligned}$ (VI.7)
As in (V.13), for $\vec{w}\in C_{c}^{\infty}(B^{4},\mathbb{R}^{5})$ we obtain
$\frac{d}{dt}\,E(\vec{\Phi}_{t})\Big|_{t=0}=\frac{1}{2}\mkern 1.0mu\big\langle d*_{g}\vec{V},\vec{w}\big\rangle.$
Therefore, we conclude that $\vec{\Phi}\in W^{3,2}_{\textup{imm}}(B^{4},\mathbb{R}^{5})$ is a weak critical point of $E$ if and only if $d*_{g}\vec{V}=0$. ∎
Let $\vec{\Phi}\in W^{3,2}_{\textup{imm}}(B^{4},\mathbb{R}^{5})$ be a weak critical point of $E$, and define $\vec{V}$ as in (VI.6). By the embeddings $L^{1}(B^{4})\hookrightarrow W^{-1,(4/3,\infty)}(B^{4})\hookrightarrow W^{-2,(2,\infty)}(B^{4})$ (see Lemma III.20), we obtain $_{g}\,\vec{V}\in W^{-2,(2,\infty)}\big(B^{4},\mathbb{R}^{5}\otimes\bigwedge^{3}T^{*}B^{4}\big)$. By weak Poincaré lemma (see for instance [44, Cor. 3.4]) and Lemma VI.2, there exists $\vec{L}\in W^{-1,(2,\infty)}\big(B^{4},\mathbb{R}^{5}\otimes\bigwedge^{2}T^{*}B^{4}\big)$ such that
$d*_{g}\vec{L}=*_{g}\,\vec{V}.$ (VI.8)
The Noether currents associated to dilations and rotations.
[…]
Proposition VI.3.
Let $\vec{\Phi}\in W^{3,2}_{\text{imm}}(B^{4},\mathbb{R}^{5})$ be a weak critical point of $E$. Then there exists $\vec{L}\in W^{-1,(2,\infty)}(B^{4},\mathbb{R}^{5}\otimes\bigwedge^{2}T^{*}B^{4})$ satisfying (VI.8) and moreover, we have
[…]
Proof.
To prove (VI.9a), we first consider the variation $\vec{\Phi}_{t}=(1+t)\vec{\Phi}$ with
$\vec{w}=\frac{d}{dt}\mkern 1.0mu\vec{\Phi}_{t}\Big|_{t=0}=\vec{\Phi}.$
Denote $H_{t}$ and $g_{t}$ as in the proof of Lemma VI.2. Since $|dH_{t}|_{g_{t}}^{2}\,d\textup{vol}_{g_{t}}=|dH|_{g}^{2}\,d\textup{vol}_{g}$, combining (VI.7)–(VI.8), we obtain that
$\begin{aligned} 0&=\frac{d}{dt}\big(|dH_{t}|_{g_{t}}^{2}\,d\textup{vol}_{g_{t}}\big)\Big|_{t=0}\\ &=-\mkern 2.0mud*_{g}\Big(\vec{V}\cdot\vec{w}+\frac{1}{2}\mkern 1.0mu\Delta_{g}H\,\vec{n}\cdot d\vec{w}-2\,\frac{d}{dt}H_{t}\Big|_{t=0}\,dH\Big)\\ &=-\mkern 2.0mud\Big(\vec{\Phi}\cdot d*_{g}\vec{L}-2\,\frac{d}{dt}H_{t}\Big|_{t=0}\,*_{g}dH\Big)\\ &=d\Big((*_{g}\,\vec{L})\mathbin{\dot{\wedge}}d\vec{\Phi}+2\,\frac{d}{dt}H_{t}\Big|_{t=0}\,*_{g}dH\Big).\end{aligned}$ (VI.10)
By (VI.2), we have
$\frac{d}{dt}H_{t}\Big|_{t=0}=\frac{1}{4}\mkern 2.0mug^{ij}\mkern 2.0mu\partial_{i}\vec{n}\cdot\partial_{j}\vec{\Phi}=-H.$
Then it follows from (VI.10) together with (III.7) that
[…]
Next, we prove (VI.9b). For $\vec{a}\in\bigwedge^{2}\mathbb{R}^{5}$ we define $\vec{\Phi}_{t}$ by
[…]
By (VI.2), we compute
[…]
$=0.$
Since $\vec{a}\in\bigwedge^{2}\mathbb{R}^{5}$ is arbitrary, by (VI.11) and (III.7) we then obtain
[…]
This completes the proof.
By the embedding results in Lemma III.20, for all $1\leq j\leq 4$, $a\in L^{\infty}\cap W^{1,4}(B^{4})$, and $f\in L^{2,\infty}(B^{4})$, we have
$\|a\mkern 2.0mu\partial_{j}f\|_{W^{-1,(2,\infty)}(B^{4})}$ $\leq\|\partial_{j}(af)\|_{W^{-1,(2,\infty)}(B^{4})}+C\mkern 1.0mu\|f\mkern 2.0mu\partial_{j}a\|_{L^{\frac{4}{3},\infty}(B^{4})}$
$\leq\|af\|_{L^{2,\infty}(B^{4})}+C\mkern 1.0mu\|\partial_{j}a\|_{L^{4}(B^{4})}\|f\|_{L^{2,\infty}(B^{4})}$
$\leq C\mkern 1.0mu\|a\|_{L^{\infty}\cap W^{1,4}(B^{4})}\mkern 1.0mu\|f\|_{L^{2,\infty}(B^{4})}.$
Hence, there exists a universal constant $C>0$ such that for all $a\in L^{\infty}\cap W^{1,4}(B^{4})$ and $T\in W^{-1,(2,\infty)}(B^{4})$, it holds that
$\|aT\|_{W^{-1,(2,\infty)}(B^{4})}\leq C\mkern 1.0mu\|a\|_{L^{\infty}\cap W^{1,4}(B^{4})}\mkern 1.0mu\|T\|_{W^{-1,(2,\infty)}(B^{4})}.$ (VI.12)
Using the inequality (VI.12), we obtain that
[…]
We define the codifferential $d^{*_{g}}$ as in (III.1). Then by Proposition VI.3 and the weak Poincaré lemma [44, Cor. 3.4], there exist $S\in L^{2,\infty}(B^{4},\bigwedge^{2}T^{*}B^{4})$ and $\vec{R}\in L^{2,\infty}\big(B^{4},\bigwedge^{2}\mathbb{R}^{5}\otimes\bigwedge^{2}T^{*}B^{4}\big)$ such that
[…]
Following Remark II.23, it is natural to ask in 4 dimension this time the following question.
Open Question VI.4.
Interpret variationally the following equations: Suppose there exists $\vec{L}$ in $W^{-1,(2,\infty)}\big(B^{4},\mathbb{R}^{5}\otimes\bigwedge^{2}T^{*}B^{4}\big)$ such that
[…]
Does the system of equations (VI.14) correspond to Euler-Lagrange equations related to the variations of the Dirichlet Energy of the mean curvature under some constraint? Recall from Remark II.23 that in 2 dimension the corresponding system of equations is equivalent to the Conformal Willmore Equations obtained by taking variations of the Willmore equations with some constraint on the underlying conformal class induced by the metric.
Now we deduce the counterpart to (V.20c) for weak critical points of EE. [^9] Partly due to the fact that $|dH|_{g}^{2}\,d\textup{vol}_{g}$ is not conformally invariant, we do not obtain a divergence-form identity as in (V.28).
Lemma VI.5.
With the above notations there holds
[…]
Proof.
[…]
We also have
[…]
Therefore, by (VI.13) we obtain
[…]
$=2\mkern 1.0mu\Delta_{g}H\mkern 2.0mu\vec{n}-\langle d(H^{2}),d\vec{\Phi}\rangle_{g}.$
This completes the proof. ∎
We summarize our results of this section in the following theorem.
Theorem VI.6.
Assume $\vec{\Phi}\in W_{imm}^{3,2}(B^{4},\mathbb{R}^{5})$ is a critical point of the functional
$E(\vec{\Phi})\coloneqq\frac{1}{2}\int_{B^{4}}\,|dH|_{g}^{2}\,d\textup{vol}_{g}.$
Then there exist $\vec{L}\in W^{-1,(2,\infty)}\big(B^{4},\mathbb{R}^{5}\otimes\bigwedge^{2}T^{*}B^{4}\big)$, $S\in L^{2,\infty}(B^{4},\bigwedge^{2}T^{*}B^{4})$, and $\vec{R}\in L^{2,\infty}\big(B^{4},\bigwedge^{2}\mathbb{R}^{5}\otimes\bigwedge^{2}T^{*}B^{4}\big)$ such that the following system holds:
[…]
As mentioned at the begining of this section, this set of identities and Noether currents obtained through the application of Noether theorem is the starting point to the proof of the regularity of weak immersions $\vec{\Phi}$ critical points to $E$ (see [16]).

### v3

@@
 I Introduction
-The study of topological and geometrical properties of manifolds through the search of special metrics “equipping” these manifolds and solving some special Partial Differential Equation has a long and rich history which takes its origin maybe in the construction of special curves (like Brachistochrone curves in the XVIIth century, or Euler’s elastica in the XVIIIth century) and constant Gauss curvature metrics on closed surfaces in relation with the uniformization theorem for Riemann Surfaces in the XVIIIth century.
-The development of what could be called “intrinsic geometric analysis” has been the source of spectacular results in differential topology, differential geometry and complex geometry with the search of constant scalar curvature metric, Einstein metrics, Kähler Einstein metrics, solutions to the Ricci flow… Another branch of geometric analysis is dealing with the study of “special submanifolds” within a given Riemannian manifold and its interaction with the geometry of the manifold itself.
+The study of topological and geometrical properties of manifolds through the search of special metrics “equipping” these manifolds and solving some special Partial Differential Equations has a long and rich history which takes its origin maybe in the construction of special curves (like Brachistochrone curves in the XVIIth century, or Euler’s elastica in the XVIIIth century) and constant Gauss curvature metrics on closed surfaces in relation with the uniformization theorem for Riemann Surfaces in the XXth century.
+The development of what could be called “intrinsic geometric analysis” has been the source of spectacular results in differential topology, differential geometry, and complex geometry with the search of constant scalar curvature metrics, Einstein metrics, Kähler–Einstein metrics, solutions to the Ricci flow… Another branch of geometric analysis is dealing with the study of “special submanifolds” within a given Riemannian manifold and its interaction with the geometry of the manifold itself.
 This branch of geometric analysis is maybe rooted originally both in the calculus of variations with the variational constructions of closed geodesics, the resolution of the Lagrange-Plateau problem in Euclidean space as well in the “explicit differential geometry” of special submanifolds such as plane algebraic curves, algebraic surfaces, etc.
@@
 In the present work, we consider a branch of extrinsic variational geometric analysis that has seen substantial development in recent decades, and concerns immersions that arise as critical points of Lagrangians depending on the first and second fundamental form, and their covariant derivatives.
-The simplest and maybe most studied model for such a Lagrangian is the Willmore functional of immersed surfaces into ${\mathbb{R}}^{m}$.
-Let $\Sigma$ be closed oriented two dimensional manifold and let $\vec{\Phi}$ be a $C^{2}$ immersion of this surface into the three dimensional Euclidean space ${\mathbb{R}}^{m}$.
+The simplest and maybe most studied model for such a Lagrangian is the Willmore functional of immersed surfaces into ${\mathbb{R}}^{m}$ with $m\geq 3$.
+Let $\Sigma$ be a closed oriented two-dimensional manifold and let $\vec{\Phi}$ be a $C^{2}$ immersion of this surface into ${\mathbb{R}}^{m}$.
 Denoting by $g_{\vec{\Phi}}$ the first fundamental form of this immersion and by $\vec{\mathrm{I\!I}}_{\vec{\Phi}}$ its second fundamental form, the mean curvature of the immersion is given by
@@
 $W(\vec{\Phi})\coloneqq\int_{\Sigma}|\vec{H}_{\vec{\Phi}}|^{2}\ d\textup{vol}_{g_{\vec{\Phi}}}.$
-This energy has been initially introduced by Sophie Germain and Simeon Poisson in an attempt to generalize to two dimensional elastic membranes the famous Euler Elastica modelizing the free energy of a beam [64, 160].
+This energy has been initially introduced by Sophie Germain and Siméon Denis Poisson in an attempt to generalize to two-dimensional elastic membranes the famous Euler Elastica modeling the free energy of a beam [64, 163].
 The very first derivation of the Euler–Lagrange equation of $W$ was made by Poisson himself around 1814 in the case of a graph (see page 60 of [163] and chapter 10 of [141]) in $\mathbb{R}^{3}$.
-Using concepts which were not completely clarified at the time (such as the Gauss curvature and the Laplace–Beltrami operator), we can rewrite the Poisson’s Euler–Lagrange equation for the Lagrangian $W$ in the following form
+Using concepts which were not completely clarified at the time (such as the Gauss curvature and the Laplace–Beltrami operator), we can rewrite Poisson’s Euler–Lagrange equation for the Lagrangian $W$ in the following form
 $\Delta_{g}H_{\vec{\Phi}}+2H_{\vec{\Phi}}\,(H^{2}_{\vec{\Phi}}-K_{\vec{\Phi}})=0,$ (I.1)
 where $H_{\vec{\Phi}}=\vec{n}_{\vec{\Phi}}\cdot\vec{H}_{\vec{\Phi}}$ is the mean curvature and $\vec{n}_{\vec{\Phi}}$ the unit Gauss map to the immersion, $\Delta_{g}$ is the negative Laplace–Beltrami operator which reads in local coordinates $(x^{1},x^{2})$ $\Delta_{g}\cdot\coloneqq(\det(g))^{-1/2}\partial_{x_{i}}(g^{ij}\,(\det(g))^{1/2}\partial_{x_{j}}\cdot)$ and $K_{\vec{\Phi}}\coloneqq\mbox{det}_{g_{\vec{\Phi}}}(\vec{n}_{\vec{\Phi}}\cdot\vec{\mathrm{I\!I}}_{\vec{\Phi}})$ is the Gauss curvature.
+[^1] The Euler-Lagrange equation in higher codimension has the same cubic structure in the second fundamental form, see [190, Theorem 2.1].
 Throughout this paper, we shall often write $g=g_{\vec{\Phi}}$ when there is no ambiguity.
@@
 Comparing (I.2) with (I.1), it appears that Poisson’s Euler–Lagrange equation is a degenerate nonlinear fourth-order equation combining analytical difficulties which were way too advanced for the early XIXth century.
-Hence in a natural way, in the following 100 years, the attention has been exclusively devoted to special solutions solving a simpler second order version of (I.1):
+Hence in a natural way, in the following century, the attention has been exclusively devoted to special solutions solving a simpler second order version of (I.1):
 $H_{\vec{\Phi}}=0.$
 This is the minimal surface equation corresponding to immersions that are critical points of the area.
-Poisson equation (I.1) has been rediscovered by Wilhelm Balschke [21] and his student Gerhard Thomsen [183] a century after (around 1924) in the framework of conformal geometry.
+Poisson’s equation (I.1) was rediscovered by W.
+Schadow [^2] See the comment by Wilhelm Blaschke in [21, Exercise 7, $\S$ 83].
+and also appeared in the PhD work of Gerhard Thomsen [186] a century later (in 1923) in the framework of conformal geometry.
 Hence, the study of the Lagrangian $W$ was motivated by the merging of two theories: conformal geometry and minimal surfaces theory.
@@
 Relying for a large part on the analysis of [183], Matthias Bauer and Ernst Kuwert in [12], by proving some clever strict inequalities on connected sums of immersions of surfaces, have been able to extend Simon’s result to any genus, that is, for any genus $g$ the infimum of $W$ among all immersions of the surface of genus $g$ is achieved.
-What should be noticed regarding the analysis in [180] is that first the minimization is considered in measure theoretical sense using the theory of varifolds, then the regularity is relying a lot on comparison arguments with local competitors which are graphs satisfying the biharmonic equation (i.e.
+What should be noticed regarding the analysis in [183] is that first the minimization is considered in a measure-theoretical sense using the theory of varifolds, then the regularity relies a lot on comparison arguments with local competitors which are graphs satisfying the biharmonic equation (i.e.
 the linearised version of (I.1)).
 In this sense, the analysis in [183] is very much restricted to minimization operations.
-The third author of the present work has been considering in the early 2000 the possibility to use more classical functional analysis to perform the minimisation of $W$ with the perspective of considering more general variational arguments such as min-max operations, or the study of the associated gradient flow.
+The third author of the present work has been considering in the early 2000s the possibility to use more classical functional analysis to perform the minimization of $W$ with the perspective of considering more general variational arguments such as min-max operations, or the study of the associated gradient flow.
 One of the first difficulties was to obtain an Euler–Lagrange equation compatible with the functional space in which the variations of $W$ are studied.
@@
 $\dot{\kappa}^{2}_{\vec{\Phi}}+\frac{\kappa^{4}_{\vec{\Phi}}}{4}=c^{2}_{0}\qquad\Longleftrightarrow\qquad\frac{\dot{\kappa}_{\vec{\Phi}}}{\sqrt{{c^{2}_{0}}-\frac{\kappa^{4}_{\vec{\Phi}}}{4}}}=\pm 1.$
-But this is not the solvability of (I.3) which is addressed at this stage, but rather the compatibility between the Euler–Lagrange equation and the Lagrangian from a purely variational perspective.
+However this is not the solvability of (I.3) which is addressed at this stage, but rather the compatibility between the Euler–Lagrange equation and the Lagrangian from a purely variational perspective.
 In unit speed parameterization $\vec{\Phi}(t)$ for $s\in[0,L]$ and $L=\int dl_{\vec{\Phi}}$ we have $\kappa_{\vec{\Phi}}\,\vec{n}_{\vec{\Phi}}=\frac{d^{2}\vec{\Phi}}{ds^{2}}$ and thus it holds
@@
 This does not define a distribution for an arbitrary $\vec{\Phi}$ in $W^{2,2}([0,L],{\mathbb{R}}^{2})$.
-As paradoxal as it appears, the Euler–Lagrange equation (I.3) is not compatible with the Lagrangian itself!
+As paradoxical as it appears, the Euler–Lagrange equation (I.3) is not compatible with the Lagrangian itself!
 The same paradox is present one dimension higher for the Willmore functional. Using (I.2) we have
@@
 Naturally the function space to be considered is the space $W^{2,2}(\Sigma,{\mathbb{R}}^{m})$ and, even assuming that the metric $g_{\vec{\Phi}}$ defined on $T\Sigma$ and its inverse $g^{-1}_{\vec{\Phi}}$ would be bounded in $L^{\infty}(\Sigma)$, the non-linearity in the Euler–Lagrange (I.1) is containing terms such as
-$2\,H_{\vec{\Phi}}\,|H_{\vec{\Phi}}|^{3}\qquad\notin L^{1}_{\textup{loc}}(\Sigma).$
+$2\,H_{\vec{\Phi}}\,|H_{\vec{\Phi}}|^{2}\qquad\notin L^{1}_{\textup{loc}}(\Sigma).$
 As well, this does not define a distribution a priori.
@@
 $\Delta_{h}\alpha=e^{2\alpha}\,K_{g_{\vec{\Phi}}}-K_{h}.$ (I.6)
-The Gauss–Bonnet theorem holds, denoting $\gamma(\Sigma)$ is the genus of $\Sigma$,
+The Gauss–Bonnet theorem holds, denoting by $\gamma(\Sigma)$ the genus of $\Sigma$,
 $\int_{\Sigma}K_{g_{\vec{\Phi}}}\,d\textup{vol}_{g_{\vec{\Phi}}}=4\pi\,(1-\gamma(\Sigma)).$ (I.7)
-A proof of the first part of this result was proposed in [168] but a very last ingredient which requires $\vec{\Phi}\in W^{2,2}(\Sigma,{\mathbb{R}}^{m})$ and not only $\vec{n}_{\vec{\Phi}}\in W^{1,2}(\Sigma,G_{2}({\mathbb{R}}^{m}))$ was missing to complete the argument and to ensure that $\phi$ given by (1.72) in the middle of page 323 is a bi-Lipschitz homeomorphism.
+A proof of the first part of this result was proposed in [171] but a very last ingredient which requires $\vec{\Phi}\in W^{2,2}(\Sigma,{\mathbb{R}}^{m})$ and not only $\vec{n}_{\vec{\Phi}}\in W^{1,2}(\Sigma,G_{2}({\mathbb{R}}^{m}))$ was missing to complete the argument and to ensure that $\phi$ given by [171, Equation (1.72)] in the middle of page 323 was a bi-Lipschitz homeomorphism.
 The complete argument for proving Theorem I.1 is given in Section IV of the present paper.
@@
 moreover $\vec{\xi}_{k}$ extends as a possibly branched $W^{2,2}$ weak immersion into ${\mathbb{R}}^{m}$.
-Some weeks after the proof of Theorem I.2 has been posted on arxiv, a proof of this result has appeared on arxiv as well but assuming the sequence $\vec{\Phi}_{k}$ to be conformal with respect to some sequence of Riemann surface (see [100]).
-It is important to insist at this stage that the restriction to conformal weak immersions is missing the goal posed in [166] of developing a variational theory for Willmore.
+Some weeks after the proof of Theorem I.2 was posted on arXiv, a proof of this result appeared on arXiv as well but assuming the sequence $\vec{\Phi}_{k}$ to be conformal with respect to some sequence of Riemann surface (see [103]).
+It is important to insist at this stage that the restriction to conformal weak immersions is missing the goal posed in [169] of developing a variational theory for the Willmore energy.
 It does not fulfill the need of exploring the neighborhood of a weak immersion in order to deduce an Euler–Lagrange equation for instance and Theorem I.1 is an essential tool in that respect.
 The notion of branched $W^{2,2}$ weak immersion into ${\mathbb{R}}^{m}$ is an extension of $W^{2,2}$ weak immersion allowing for isolated branched points $\{a_{1},\dots,a_{L}\}$ at which the gradient of the weak immersion is degenerating.
-Branched $W^{2,2}$ weak immersion have an underlying conformal structure that is there exists a constant Gauss curvature metric $h$ on $\Sigma$ such that
+Branched $W^{2,2}$ weak immersions have an underlying conformal structure, that is to say, there exists a constant Gauss curvature metric $h$ on $\Sigma$ such that
 $g_{\vec{\Phi}}=e^{2\alpha}\,h.$
-But the conformal factor $\alpha$ is not bounded in $L^{\infty}$ anymore at each branched point but has a relatively well understood asymptotic expansion: in a local conformal chart $x=(x^{1},x^{2})\in D^{2}$ for $h$ containing exactly one of the branched point, say $a_{i}$, and for which $x(a_{i})=0$, there exists $d_{i}\in{\mathbb{N}}$ such that (see Lemma A.5 of [166])
-$\|\alpha(x)-d_{i}\log|x|\|_{L^{\infty}(D^{2})}<+\infty,$
+However the conformal factor $\alpha$ is not bounded in $L^{\infty}$ anymore at each branched point but has a relatively well understood asymptotic expansion: in a local conformal chart $x=(x^{1},x^{2})\in D^{2}$ for $h$ containing exactly one of the branched point, say $a_{i}$, and for which $x(a_{i})=0$, there exists $d_{i}\in{\mathbb{N}}$ such that (see Lemma A.5 of [169])
+$\|\alpha(x)-d_{i}\log|x|\|_{L^{\infty}(D^{2})}<+\infty.$
 The maximal number of branched points for a given closed two-dimensional manifold $\Sigma$ is bounded by the $L^{2}$ norm of the second fundamental form.
@@
 It is proved in [167] that the solutions to the $\vec{R}-S$ system combined with (I.14) are smooth by applying integrability by compensation to this first order elliptic system combined with the structural equation
-$2\vec{H}=\nabla S\cdot\nabla^{\perp}\vec{\Phi}+\nabla\vec{R}\,\dot{\times}\,\nabla^{\perp}\vec{\Phi}.$ (I.14)
-While the use of local conformal coordinates was centrally used in [164], Section V of the present paper is presenting a proof of the regularity without using conformal coordinates.
-The motivation being to “prepare the ground” for considering regularity questions for solution to the Euler–Lagrange equation to higher dimensional counterpart to the Willmore Lagrangian which is the subject of the second part of the paper.
+$2\vec{H}=\nabla S\cdot\nabla^{\perp}\vec{\Phi}+\nabla\vec{R}\times\nabla^{\perp}\vec{\Phi}.$ (I.14)
+While the use of local conformal coordinates was essential in [167], Section V of the present paper is presenting a proof of the regularity without using conformal coordinates.
+The motivation is to “prepare the ground” for considering regularity questions for solution to the Euler–Lagrange equation to higher dimensional counterpart to the Willmore Lagrangian, which is the subject of the second part of the paper.
 Before coming to this second part we prove the following approximation theorem (in Section IV.4.3), which is one of the main new result of the present paper.
@@
 This result is fundamental in the recent work [133], where the second and third authors study the closure of the set of weak immersions under a certain Sobolev-type bound on the second fundamental form whose motivation arises from the following context.
-In 2018, Andrea Mondino and Huy Nguyen [146] proved that the Willmore functional is the only (up to a topological constant) conformally invariant functional among all the functionals defined on hypersurfaces of 3-dimensional manifolds and depending only on the first and second fundamental form.
+In 2018, Andrea Mondino and Huy The Nguyen [149] proved that the Willmore functional is the only (up to a topological constant) conformally invariant functional among all the functionals defined on hypersurfaces of 3-dimensional manifolds and depending only on the first and second fundamental form.
 However, this is not the case in higher dimensions and codimensions.
@@
 Returning to the original idea of Germain and Poisson [64, 163], one would like to study functionals that measure in some sense the umbilicity of a hypersurface.
-Such a functional should be able to provide ”best representation” of submanifolds and in particular, this requires to have a lower bound.
+Such a functional should be able to provide “best representation” of submanifolds and in particular, this requires to have a lower bound.
 In [70], Gover–Waldron proved that conformally invariant functionals can also be stated in terms of singular Yamabe-type problem, a problem raised by Loewner–Nirenberg [116] in 1974.
-In [67, 70], the authors prove that given any separating submanifold $\Sigma^{d-1}$ of some Riemannian manifold $(M^{d},g)$, if one look for the ”best” asymptotic expansion of a defining function [^3] A defining function $s\in C^{\infty}(M)$ is defining for $\Sigma$ if $s=0$ and $|ds|_{g}=1$ on $\Sigma$.
+In [67, 70], the authors prove that given any separating submanifold $\Sigma^{d-1}$ of some Riemannian manifold $(M^{d},g)$, if one looks for the “best” asymptotic expansion of a defining function [^5] A defining function $s\in C^{\infty}(M)$ is defining for $\Sigma$ if $s=0$ and $|ds|_{g}=1$ on $\Sigma$.
 The question raised in [70] is to find the highest number $\ell\in\mathbb{N}$ and a defining function $s$ such that we have an expansion of the form $|ds|_{g}=1+O(s^{\ell})$ near $\Sigma$.
@@
 This term is also the leading-order term of the functional
-$\vec{\Phi}\in\mathrm{Imm}(\Sigma^{n};\mathbb{R}^{d})\mapsto\int_{\Sigma^{n}}\big|\nabla^{\frac{n-2}{2}}\vec{\mathrm{I\!I}}_{\vec{\Phi}}\big|^{2}_{g_{\vec{\Phi}}}\ d\textup{vol}_{g_{\vec{\Phi}}}.$
+$\vec{\Phi}\in\mathrm{Imm}(\Sigma^{n},\mathbb{R}^{d})\mapsto\int_{\Sigma^{n}}\big|\nabla^{\frac{n-2}{2}}\vec{\mathrm{I\!I}}_{\vec{\Phi}}\big|^{2}_{g_{\vec{\Phi}}}\ d\textup{vol}_{g_{\vec{\Phi}}}.$
 The study of compactness and regularity questions for such functionals has been the subject of recent works that we will describe in Section II.2.
@@
 To do so, they prove that in this case, the umbilic line should be contained in a plane intersecting the surface orthogonally.
-By an application of Cauchy–Kovalevskya theorem, they prove that the mean curvature of the parts of the surface above and below this plane, seen as submanifolds of $\mathbb{H}^{3}$, completely vanishes.
-The $\mathbb{S}^{1}$-equivariant Willmore tori have been classified by Ferus–Pedit [60] (and actually contain the construction of [158]) by analyzing the geodesic flow on these surfaces.
+By an application of Cauchy–Kovalevskaya theorem, they prove that the mean curvature of the parts of the surface above and below this plane, seen as submanifolds of $\mathbb{H}^{3}$, completely vanishes.
+The $\mathbb{S}^{1}$-equivariant Willmore tori have been classified by Ferus–Pedit [60] (and actually contains the construction of [161]) by analyzing the geodesic flow on these surfaces.
 The understanding of Willmore graphs has also been subject of numerous works.
@@
 An alternative way for such constructions is provided by variational methods, such as minimization or min-max procedures in a given class of surfaces.
-As an example of application, the Willmore conjecture for genus 1 in codimension 1 has been solved by Marques–Neves [124] in 2013 using a min-max scheme for minimal surfaces combined with Urbano theorem, but remains open for higher genus and general codimension.
-This conjecture states that up to conformal transformations and in all codimensions, the only minimizer of the Willmore functional in genus 1 is the Clifford torus, whereas in genus larger than 1, the likely candidates are the Lawson surfaces [99, Conjecture 8.4].
+As an example of application, the Willmore conjecture for genus 1 in codimension 1 has been solved by Marques–Neves [128] in 2012 using a min-max scheme for minimal surfaces combined with Urbano theorem, but remains open for higher genus and general codimension.
+This conjecture states that up to conformal transformations and in all codimensions, the only minimizer of the Willmore functional in genus 1 is the Clifford torus, whereas in genus larger than 1, the likely candidates are the Lawson surfaces [100, Conjecture 8.4].
 Even if there is some numerical evidence [87], its complete resolution remains largely open.
+The case of surfaces enjoying a priori the same symmetries as the Lawson surfaces has been recently studied in [101].
 Conjecture II.2 (Willmore conjecture in codimension 1 [100]).
@@
 A particular case of [173] is the case $m=3$, i.e.
-the codimension 1 case, which can be applied to the sphere eversion and constitutes a new step towards the $16\pi$-conjecture.
+the codimension 1 case, which can be applied to the sphere eversion and constitutes a new step towards the $16\pi$-conjecture, originally formulated by Kusner in the 1980s (see for instance [102]).
 […]
@@
 As a consequence of [173], we know that the number $\beta_{0}$ is equal to the sum of the energies of finitely many branched Willmore spheres.
+In [146], Michelat and the third author proved an energy quantization of the Gauss map in the $W^{1,(2,1)}$ topology for sequences of immersions arising in the viscosity method.
 Hence, the need for a classification of branched Willmore spheres became strongly motivated.
 As we will explain in the following Section II.1.2, Bryant [28] proved that any smooth Willmore sphere is a conformal transformation of a minimal surface in $\mathbb{R}^{3}$.
-A first generalization to the branched Willmore spheres has been obtained by Lamm–Nguyen [103] when the total number of branched points counted with multiplicites is at most 3.
+A first generalization to the branched Willmore spheres has been obtained by Lamm–Nguyen [106] when the total number of branched points counted with multiplicities is at most 3.
 Michelat and the third author [145] successfully classified all ”variational branched Willmore spheres” which include in particular weak limits of immersed Willmore spheres.
@@
 $u(z)=\frac{\mathrm{Eval}(u,p_{i})}{|z|^{2m}}+\sum_{1\geq k+l\geq 1-2m}\Re\left(\alpha_{k,l}\,z^{k}\,\bar{z}^{l}\right)+\mathrm{Log}(u,p_{i})\,\log|z|+O(1).$
-The main result of [138] can be stated as (see for instance [138, Remark 6.5])
+The main result of [142] can be stated as (see for instance [142, Remark 6.5])
 $\mathrm{Ind}_{W}(\vec{\Phi})=\dim\left\{u\in V:\mathcal{L}^{2}u=0,\ \sum_{i=1}^{n}\mathrm{Eval}(u,p_{i})\,\mathrm{Log}(u,p_{i})<0\right\}.$
 In the case where $\Sigma=\mathbb{S}^{2}$ and the ends of $\vec{\Psi}$ are planar (so that $\vec{\Phi}$ as no branch point), Hirsch, Kusner and Mäder-Baumdicker [86, 85], revisiting Michelat’s computations, proved that the Willmore index of $\vec{\Phi}$ is exactly n3n-3, and thus verified Michelat’s conjecture in the case of unbranched Willmore spheres.
-To do so, they restricted the study of the admissible variations to the following set
+To do so, they restricted themselves to the study of the following admissible variations
 […]
@@
 One of the major achievement in [28] is the classification of smooth Willmore spheres.
-Indeed, since there is no holomorphic quartic on $\mathbb{S}^{2}$, every smooth Willmore sphere must satisfy $\mathcal{Q}=0$ and thus, be a conformal transformation of a minimal surface in $\mathbb{R}^{3}$.
+Indeed, since there are no non-zero holomorphic quartic differentials on $\mathbb{S}^{2}$, every smooth Willmore sphere must satisfy $\mathcal{Q}=0$ and thus, be a conformal transformation of a minimal surface in $\mathbb{R}^{3}$.
 If $\vec{\Phi}\colon\mathbb{S}^{2}\to\mathbb{R}^{3}$ is a branched Willmore immersion (for instance obtained in a bubble tree), then the equation $\bar{\partial}\mathcal{Q}=0$ is a-priori not verified at the branch points of $\vec{\Phi}$, hence $\mathcal{Q}$ is a-priori merely meromorphic and it is not clear that Bryant’s classification extends to branched Willmore immersions.
 Lamm–Nguyen [106] computed an asymptotic expansion of $\mathcal{Q}$ across branch points and proved that $\mathcal{Q}=0$ if the number of branch points of $\vec{\Phi}$ is at most 3, when counted with multiplicities.
-This idea can also be seen as an application of Liouville theorem for holomorphic functions, as explained in [120, Theorem 2.6.3].
-Then Michelat–Rivière [143] proved that a branched Willmore sphere obtained as a weak limit of smooth Willmore surfaces or obtained in a bubbling process of a sequence of smooth Willmore surfaces (called a ”variational branched Willmore sphere”) also verify $\mathcal{Q}=0$.
-To do so, they compute a refined asymptotic expansion of $\mathcal{Q}$ near a branch point and prove that $\mathcal{Q}$ is actually bounded near branch points where one of the residue of $\vec{\Phi}$ verifies a smallness assumption.
-They proved in an earlier work [141, 142] that this condition is full-filled for variational branched Willmore spheres, see also [122] for a similar property.
-The full classification has recently been obtained in [126], by adapting the analysis of the conformal Gauss map developed in [17] in order to obtain $\varepsilon$-regularity on the conformal Gauss map and [127] in order to prove an energy quantization result for Willmore surfaces with bounded Morse index.
-The next step, which would be an interesting starting point for the Willmore conjecture, is to understand the geometry of smooth Willmore surfaces $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{3}$ when $\Sigma$ has positive genus.
-On the one hand, Palmer [155] proved that the Bryant’s quartic could be interpreted as the curvature of the conformal Gauss map.
+This idea can also be seen as an application of Liouville theorem for holomorphic functions, as explained in [124, Theorem 2.6.3].
+Then Michelat–Rivière [145] proved that a branched Willmore sphere obtained as a weak limit of smooth Willmore surfaces or obtained in a bubbling process of a sequence of smooth Willmore surfaces (called a ”variational branched Willmore sphere”) also verifies $\mathcal{Q}=0$, by showing that the limit of Bryant’s quartics must remains holomorphic.
+The full classification has recently been obtained by the second author in [130], by adapting the analysis of the conformal Gauss map developed in [17] in order to obtain $\varepsilon$-regularity on the conformal Gauss map and [131] in order to prove an energy quantization result for Willmore surfaces with bounded Morse index.
+The next step, which would be an interesting starting point for the Willmore conjecture, is to understand the geometry of smooth Willmore surfaces $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{3}$ when $\Sigma$ has a positive genus.
+On the one hand, Palmer [158] proved that Bryant’s quartic can be interpreted as the curvature of the conformal Gauss map.
 Moreover, he proved that if $\mathcal{Q}$ verifies the pointwise condition $|\mathcal{Q}|_{g_{Y}}>1$, then $\Sigma$ is a torus and $\vec{\Phi}$ is unstable.
-On the other hand, Fisher-Colbrie [61] proved that stable minimal surfaces in positively curved manifold have finite total Gauss curvature.
-Now that some regularity properties have been understood in [17, 123, 127], it would be intersting to study stability questions from the view-point of the conformal Gauss map.
+On the other hand, Fisher-Colbrie [61] proved that stable minimal surfaces in a positively curved manifold have finite total Gauss curvature.
+Now that some regularity properties have been understood in [17, 127, 131], it would be interesting to study stability questions from the viewpoint of the conformal Gauss map.
 Open Question II.8.
 Can we prove that the Gauss curvature of the conformal Gauss map of a stable Willmore surface in $\mathbb{R}^{3}$ is integrable?
-In higher codimension, the conformal Gauss map has already been studied [148, 56, 117, 115] and there are Willmore spheres which are not conformal transformation of minimal surfaces.
+In higher codimension, the conformal Gauss map has already been studied [151, 56, 121, 119] and there are Willmore spheres that are not conformal transformations of minimal surfaces.
 For instance, Montiel [151] classified all the Willmore spheres in $\mathbb{R}^{4}$ and proved that in addition to conformal transformations of minimal surfaces, Willmore spheres can also be obtained as a rational curve in $\mathbb{CP}^{3}$ via the Penrose twistor fibration $\mathbb{CP}^{3}\to\mathbb{S}^{4}$.
-This classification has also been extended in [143] to branched Willmore spheres satisfying some smallness assumption on a residue.
+This classification has also been extended in [145] to branched Willmore spheres arising as limits of smooth Willmore surfaces.
+The classification of smooth Willmore spheres in arbitrary codimension is much more involved and has recently been studied in [118].
 Open Question II.9.
 Is it true that any branched Willmore spheres in $\mathbb{R}^{4}$ belongs to one of these two classes: either a conformal transformations of a branched minimal surface of $\mathbb{R}^{4}$, or via the Penrose twistor fibration $\mathbb{CP}^{3}\to\mathbb{S}^{4}$ a branched rational curve in $\mathbb{CP}^{3}$?
-One of the difficulty to such an approach comes from the size of the umbilic set of Willmore surfaces.
-Contrarily to minimal surfaces, Babich–Bobenko [10] proved that Willmore surfaces can have curves of umbilic points.
-At these points, the geometry of the conformal Gauss map degenerate.
+One of the difficulties with such an approach comes from the size of the umbilic set of Willmore surfaces.
+Unlike minimal surfaces, Babich–Bobenko [10] proved that Willmore surfaces can have curves of umbilic points.
+At these points, the geometry of the conformal Gauss map degenerates.
 Examples of Willmore surfaces having such curves have been constructed by Babich–Bobenko [10] and by Dall’Acqua–Schätzle [51].
 However, the structure of such sets is still not fully understood.
-Schätzle [177] proved that the umbilic set of a smooth Willmore immersion is composed of isolated points and one dimensional real-analytic manifold without boundary.
+Schätzle [180] proved that the umbilic set of a smooth Willmore immersion is composed of isolated points and one dimensional real-analytic manifolds without boundary.
 In the recent work [127], Marque and the second author proved that the structure of the umbilic set provides strong information on the geometry of Willmore immersions and on the value of its energy.
-For instance if a Willmore surface has a geodesic umbilic line, then it is of Babich–Bobenko type.
+For instance, if a Willmore surface has a geodesic umbilic line, then it is of Babich–Bobenko type.
 Open Question II.10.
-What properties does the umbilic lines of closed Willmore surfaces satisfy?
+What properties do the umbilic lines of closed Willmore surfaces satisfy?
 Do they satisfy any particular equation?
@@
 A natural question would be to understand the set of Willmore surfaces having the same given Bryant’s quartic, or even the set of quartics on a given surface that can be achieved as the Bryant’s quartic of some Willmore immersion.
+For instance, we can consider two immersions $\vec{\Phi}_{1},\vec{\Phi}_{2}\colon\Sigma\to\mathbb{R}^{3}$ having the same Bryant’s quartic.
+Then one can roughly think of this restriction as a relation between the second fundamental forms of their conformal Gauss maps.
+If their induced metrics $g_{\vec{\Phi}_{1}}$ and $g_{\vec{\Phi}_{2}}$ are also conformal, then one can hope that these two conformal Gauss maps coincide, and thus $\vec{\Phi}_{1}$ and $\vec{\Phi}_{2}$ should be a conformal transformation of the other one.
 Open Question II.11.
@@
 Is it possible to relate $\vec{\Phi}_{1}$ to a conformal transformation of $\vec{\Phi}_{2}$?
-Since the Bryant’s quartic roughly describes the second fundamental form of their conformal Gauss map, the map $\vec{\Phi}_{1}$ should be a conformal transformation of $\vec{\Phi}_{2}$ in the case where the induced metrics of their conformal Gauss map coincide.
-But it seems wrong if this is not the case.
 It would be interesting to study the possible differences.
@@
 Given a complex structure on a given closed surface $\Sigma$, can we describe the set of holomorphic quartics which are the Bryant’s quartic of some Willmore immersion?
-The Poisson problem has been introduced in [47] (analog of the Plateau problem for Willmore surfaces) where the authors solve the Plateau problem for datas given for the immersion and the Gauss map, see also [161].
+The Poisson problem has been introduced in [47] (analog of the Plateau problem for Willmore surfaces), where the authors solve the Plateau problem for data given for the immersion and the Gauss map, see also [164].
 Due to the conformal invariance of the Willmore energy, it would be natural to solve the Plateau problem for conformally invariant data on the boundary.
@@
 Open Question II.14.
-What is the optimal regularity assumption on the given boundary datas in order to solve the Poisson problem? How many solutions are there?
+What is the optimal regularity assumption on the given boundary data in order to solve the Poisson problem? How many solutions are there?
 Up to now, the conformal Gauss map approach has been mainly developed for surfaces of a Euclidean space.
@@
 Geometric flows are designed to produce explicit paths leading to critical points of a given functional.
-They are useful both from a theoretic and from a practical view-point.
+They are useful both from a theoretic and from a practical viewpoint.
 For instance, Conjecture II.3 is an example of application of the Willmore flow, which roughly says that there is no smooth Willmore sphere of energy strictly between the energy of $4\pi$ and $16\pi$.
-One could also think about the Willmore flow as a first step before studying the Canham–Helfrich flow modelling the evolution of membranes of the cells, see for instance [25, 173] and the references therein.
+One could also think about the Willmore flow as a first step before studying the Canham–Helfrich flow modeling the evolution of membranes of the cells, see for instance [25, 176] and the references therein.
 The main analytical difficulty of the Willmore flow (despite the fact that it is a geometric flow, and thus has to be degenerate) is the fact that it is of order 4, meaning that there exists no comparison principle in the spirit of the mean curvature flow.
 For instance, if an initial data is included in the interior of a given sphere $S$, one can a-priori not say that the flow remains inside the interior of $S$.
-Geometric flows of order 4 have original been explored in the context of biharmonic surfaces [57].
-The Willmore flow has first been introduced in 2001 independentely by Kuwert–Schätzle [101] and Simonett [181].
+Geometric flows of order 4 have originally been introduced in the context of biharmonic surfaces [57].
+The Willmore flow was first introduced in 2001 independently by Kuwert–Schätzle [104] and Simonett [184].
 They proved the existence of a solution in short time, and proved that if the initial data has small energy, then the flow smoothly converges to the round sphere.
-In 2003, Mayer–Simonett [132] proved that the Willmore flow does not preserve embeddedness of the initial data.
+In 2003, Mayer–Simonett [136] proved that the Willmore flow does not preserve the embeddedness of the initial data.
 Blatt [22] constructed in 2009 an example of initial data of spherical type for which a singularity occurs in the Willmore flow.
-More recently, Dall’Acqua–Müller–Schätzle–Spener [50] produced examples of flows made of tori of revolutions, and also produced degenerating examples.
+More recently, Dall’Acqua–Müller–Schätzle–Spener [50] produced examples of flows made of tori of revolution, and also produced degenerating examples.
 This flow is still widely unexplored.
-The parametric approach has been introduced by Palmurella–Rivière [156, 157], where the authors show that there are only two possible kind of singularities in finite time: either a concentration of energy with a possible branch point, or a degeneration of the conformal class.
+The parametric approach has been introduced by Palmurella–Rivière [159, 160], where the authors show that there are only two possible kinds of singularity in finite time: either a concentration of energy with a possible branch point or a degeneration of the conformal class.
 Open Question II.15.
@@
 Then Laurain–Rivière [109] proved in 2018 that when the underlying complex structure degenerates, then only one residue is responsible for a possible loss of energy.
-This has recently been extended in [112], where Li–Yin–Zhou prove that this residue precisely describes the loss of energy.
+This has recently been extended in [115], where Li–Yin–Zhou proved that this residue precisely describes the loss of energy.
 In 2021, Marque [126] proved that the set of Willmore immersions with energy less than or equal to $12\pi$ is compact.
@@
 Is it possible to define a solution of the Willmore flow starting from a branched surface? From a nodal surface?
-As an application, one can think about the $16\pi$-conjecture.
-In a recent work [118], Mäder-Baumdicker–Seidel studied the possible singularities that occur for an initial data $\vec{\Phi}_{0}\colon\mathbb{S}^{2}\to\mathbb{R}^{3}$ with Willmore energy strictly less than $16\pi$.
-They proved that if $W(\vec{\Phi}_{0})\leq 12\pi$, then either the flow starting from $\vec{\Phi}_{0}$ converges to a round sphere or the flow converges to two round spheres glued by a catenoid neck.
-If $W(\vec{\Phi}_{0})<16\pi$, then other singularities might arise.
-Open Question II.18.
-Given a immersion $\vec{\Phi}_{0}\colon\mathbb{S}^{2}\to\mathbb{R}^{m}$ satisfying $W(\vec{\Phi}_{0})<16\pi$, under which conditions the Willmore flow $(\vec{\Phi}_{t})_{t>0}$ starting from $\vec{\Phi}_{0}$ does not develop any singularity as long as $12\pi<W(\vec{\Phi}_{t})<16\pi$?
+As an application, one can think about the $16\pi$-conjecture and the classification of singularities along the Willmore flow.
+Lamm–Nguyen [106] proved that if we start the flow with an initial data $\vec{\Phi}_{0}\colon\mathbb{S}^{2}\to\mathbb{R}^{3}$ verifying $W(\vec{\Phi}_{0})\leq 12\pi$, then the flow starting from $\vec{\Phi}_{0}$ can only develop catenoid-type singularities.
+In a recent work [122], Mäder-Baumdicker–Seidel identified 4 distinct homotopy classes in the set of immersed spheres with Willmore energy less than or equal to $12\pi$.
+Two of them contain round spheres and the two others lead to a notion of unavoidable singularities along the flow.
+A partial study of the set of immersions with energy less than $16\pi$ is also studied in [122].
+Open Question II.18 (Question 1.13 in [122]).
+Is it possible to classify the homotopy classes of the set of immersions with Willmore energy less than $16\pi$?
+In particular, given an immersion $\vec{\Phi}_{0}\colon\mathbb{S}^{2}\to\mathbb{R}^{3}$ satisfying $W(\vec{\Phi}_{0})<16\pi$, under which conditions the Willmore flow $(\vec{\Phi}_{t})_{t}$ starting from $\vec{\Phi}_{0}$ lies in the homotopy class of a round sphere when $W(\vec{\Phi}_{t})<12\pi$?
 II.1.4 Critical points of the Willmore energy under various constraints
-Instead of looking at critical points of the Willmore functional among the set of all the immersions, one can restrict the set of admissible variations (for example, one need to fix some geometric constraints for the study of the Canham–Helfrich functional in cell biology).
-Three main possibilities are curently studied.
-One possibility, is to study the minimizers (and more generally the critical points) of the Willmore functional under the constraints of fixing both the area of the immersed surface and the volume enclosed it encloses.
+Instead of looking at critical points of the Willmore functional among the set of all the immersions, one can restrict the set of admissible variations (for example, one needs to fix some geometric constraints for the study of the Canham–Helfrich functional in cell biology).
+Currently, three main possibilities are studied.
+One possibility is to study the minimizers (and more generally the critical points) of the Willmore functional under the constraints of fixing both the area of the immersed surface and the volume it encloses.
 Since the Willmore functional is scale-invariant, this question is equivalent to prescribing the isoperimetric ratio of the immersion $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{3}$ given by
@@
 We can therefore study the following minimization problem for any $R\geq(36\pi)^{\frac{1}{3}}$:
-$\beta^{\mathcal{I}}(R)\coloneqq\inf\left\{W(\vec{\Phi}):\vec{\Phi}\in\mathrm{Imm}(\Sigma;\mathbb{R}^{3}),\ \mathcal{I}(\vec{\Phi})=R\right\}.$
-Keller–Mondino–Rivière [96] proved in 2014 that a minimizer of $\beta^{\mathcal{I}}(R)$ exists provided that this number verifies some strict inequality.
+$\beta^{\mathcal{I}}(R)\coloneqq\inf\left\{W(\vec{\Phi}):\vec{\Phi}\in\mathrm{Imm}(\Sigma,\mathbb{R}^{3}),\ \mathcal{I}(\vec{\Phi})=R\right\}.$
+Keller–Mondino–Rivière [97] proved in 2014 that there is a minimizer of $\beta^{\mathcal{I}}(R)$ provided that this number verifies some strict inequality.
 This inequality has been proved for all $R\in\left((36\pi)^{\frac{1}{3}},+\infty\right)$ when $\Sigma=\mathbb{S}^{2}$ by Schygulla [182] in 2012, when $\Sigma=\mathbb{T}^{2}$ by Scharrer [177] in 2022 and by Ndiaye–Schätzle [156] in all the remaining cases in 2015.
@@
 Describe the solutions of $\beta^{\mathcal{I}}(R)$ for $R\in\left((36\pi)^{\frac{1}{3}},+\infty\right)$. Are they unique? Do they have some symmetries?
-Another possibility, is to study the minimization of the Willmore energy under a constraint on the total mean curvature. Given an immersion $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{3}$, we denote
-$\mathcal{T}(\vec{\Phi})\coloneqq\frac{1}{vol_{g_{\vec{\Phi}}}(\Sigma)^{\frac{1}{2}}}\int_{\Sigma}H_{\vec{\Phi}}\,d\textup{vol}_{g_{\vec{\Phi}}}.$
-By Hölder inequality, it holds
+Another possibility is to study the minimization of the Willmore energy under a constraint on the total mean curvature.
+Given an immersion $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{3}$ (where $\Sigma$ is a Riemann surface, and thus carries an orientation), we denote
+$\mathcal{T}(\vec{\Phi})\coloneqq\frac{1}{\textup{vol}_{g_{\vec{\Phi}}}(\Sigma)^{\frac{1}{2}}}\int_{\Sigma}H_{\vec{\Phi}}\,d\textup{vol}_{g_{\vec{\Phi}}}.$
+By Hölder inequality, it holds that
 $\mathcal{T}(\vec{\Phi})^{2}\leq W(\vec{\Phi}).$
 We then study the following problem for any $R\in\mathbb{R}$
-$\beta^{\mathcal{T}}(R)\coloneqq\inf\left\{W(\vec{\Phi}):\vec{\Phi}\in\mathrm{Imm}(\Sigma;\mathbb{R}^{3}),\ \mathcal{T}(\vec{\Phi})=R\right\}.$
+$\beta^{\mathcal{T}}(R)\coloneqq\inf\left\{W(\vec{\Phi})\colon\vec{\Phi}\in\mathrm{Imm}(\Sigma,\mathbb{R}^{3}),\ \mathcal{T}(\vec{\Phi})=R\right\}.$
 Scharrer–West [179] proved the existence of a minimizer for $R\in\left(0,\sqrt{2}\,\mathcal{T}(\mathbb{S}^{2})\right)\setminus\{\mathcal{T}(\mathbb{S}^{2})\}$ and studied the properties of the map $(R\in\mathbb{R}\mapsto\beta^{\mathcal{T}}(R))$.
-For the case $R=\mathcal{T}(\mathbb{S}^{2})$, they argue that the arguments developed in [96] do not apply.
+For the case $R=\mathcal{T}(\mathbb{S}^{2})$, they proved that
+$\beta^{\mathcal{T}}(\mathcal{T}(\mathbb{S}^{2}))=\inf\left\{W(\vec{\Phi}):\vec{\Phi}\in\mathrm{Imm}(\Sigma,\mathbb{R}^{3})\right\}.$
+Hence, the existence of a minimizer for $\beta^{\mathcal{T}}(\mathcal{T}(\mathbb{S}^{2}))$ should be very related to Conjecture II.2.
 Open Question II.20.
-Is $\beta^{\mathcal{T}}(\mathcal{T}(\mathbb{S}^{2}))$ reached?
 Is there a minimizer for $\beta^{\mathcal{T}}(R)$ with $R\notin\left(0,\sqrt{2}\,\mathcal{T}(\mathbb{S}^{2})\right)\setminus\{\mathcal{T}(\mathbb{S}^{2})\}$?
+In the borderline case $R=0$, we know by continuity of the map $R\mapsto\beta^{\mathcal{T}}(R)$ that $\beta^{\mathcal{T}}(0)\leq 8\pi$.
+Knowing whether there exists a minimizer for $\beta^{\mathcal{T}}(0)$ thus amounts to knowing whether $\beta^{\mathcal{T}}(0)=8\pi$ is true (see [179]).
 Open Question II.21.
@@
 The study of $\beta^{\mathcal{T}}$ is related to the study of the following functional, for a given parameter $\bar{H}\in\mathbb{R}$ called the spontaneous curvature
-$\forall\vec{\Phi}\in\mathrm{Imm}(\Sigma;\mathbb{R}^{3}),\qquad W_{\bar{H}}(\vec{\Phi})\coloneqq\int_{\Sigma}(H_{\vec{\Phi}}-\bar{H})^{2}\,d\textup{vol}_{g_{\vec{\Phi}}}.$
-The Helfrich flow has been introduced in 2006 [98], where Kohsaka–Nagasawa prove the short term existence of solution to the gradient flow of $W_{\bar{H}}$ with the enclosed volume and the area prescribed by some given constants $V_{0}$ and $A_{0}$.
+$\forall\vec{\Phi}\in\mathrm{Imm}(\Sigma,\mathbb{R}^{3}),\qquad W_{\bar{H}}(\vec{\Phi})\coloneqq\int_{\Sigma}(H_{\vec{\Phi}}-\bar{H})^{2}\,d\textup{vol}_{g_{\vec{\Phi}}}.$
+The Helfrich flow has been introduced in 2006 [99], where Kohsaka–Nagasawa proved the short-term existence of solutions to the gradient flow of $W_{\bar{H}}$ with the enclosed volume and the area prescribed by some given constants $V_{0}$ and $A_{0}$.
 The cases when the initial data has small energy or when it is radially symmetric have been studied by Rupp–Scharrer–Schlierf [176].
@@
 Let $\vec{\Phi}$ be a weak immersion of a surface $\Sigma$, as mentioned above it defines a smooth conformal class — i.e.
-a Riemann Surface — which is the one induced by the pull back metric metric $g_{\vec{\Phi}}=\vec{\Phi}^{\,\ast}g_{{\mathbb{R}}^{3}}$.
-The conformal class defined by this immersion is said to be a degenerate point of the map $\mathcal{C}$ assigning the conformal class of $(\Sigma,g_{\vec{\Phi}})$ in the Teichmüller space if and only if there exists a non zero holomorphic quadratic form [^4] An holomorphic quadratic form is an holomorphic section of the bundle given by $\wedge^{1,0}T^{\ast}\Sigma\otimes\wedge^{1,0}T^{\ast}\Sigma$ where $\Sigma$ is equipped with the conformal structure induced by $g_{\vec{\Phi}}=\vec{\Phi}^{\,\ast}g_{{\mathbb{R}}^{m}}$ $q$ such that
+a Riemann surface — which is the one induced by the pull back metric $g_{\vec{\Phi}}=\vec{\Phi}^{\,\ast}g_{{\mathbb{R}}^{3}}$.
+The conformal class defined by this immersion is said to be a degenerate point of the map $\mathcal{C}$ assigning the conformal class of $(\Sigma,g_{\vec{\Phi}})$ in the Teichmüller space if and only if there exists a non zero holomorphic quadratic form [^6] A holomorphic quadratic form is a holomorphic section of the bundle given by $\wedge^{1,0}T^{\ast}\Sigma\otimes\wedge^{1,0}T^{\ast}\Sigma$ where $\Sigma$ is equipped with the conformal structure induced by $g_{\vec{\Phi}}=\vec{\Phi}^{\,\ast}g_{{\mathbb{R}}^{m}}$ $q$ such that
 $q,\vec{h}_{0}>_{WP}=0.$ (II.1)
 In the above equation, we denoted $\vec{h}_{0}$ the Weingarten operator given by
-$\vec{h}_{0}:=-\,\partial\pi_{\vec{n}}\,{\otimes}\,\partial\vec{\Phi},$
-where $\pi_{\vec{n}}$ is the orthogonal projection linear map from ${\mathbb{R}}^{m}$ onto the normal space to the immersion $N_{x}\Sigma\coloneq(\vec{\Phi}_{\ast}T_{x}\Sigma)^{\perp}$ and $\cdot,\cdot>_{WP}$ is denoting the Weil–Peterson scalar product given in local conformal coordinates for which
+$\vec{h}_{0}\coloneq-2\,\partial\pi_{\vec{n}}\,{\otimes}\,\partial\vec{\Phi},$
+where $\pi_{\vec{n}}$ is the linear orthogonal projection map from ${\mathbb{R}}^{m}$ onto the normal space to the immersion $N_{x}\Sigma\coloneq(\vec{\Phi}_{\ast}T_{x}\Sigma)^{\perp}$ and $\cdot,\cdot>_{WP}$ denotes the Weil–Peterson scalar product given in local conformal coordinates for which
 $g_{\vec{\Phi}}=e^{2\lambda}\ [dx_{1}^{2}+dx_{2}^{2}],$
 by (see [170])
-$\left<\phi\ dz\otimes dz,\psi\ dz\otimes dz\right>:=e^{-4\lambda}\ \Re\left[\overline{\phi}\,{\psi}\right]\ .$
+$\left<\phi\ dz\otimes dz,\psi\ dz\otimes dz\right>\coloneq e^{-4\lambda}\ \Re\left[\overline{\phi}\,{\psi}\right]\ .$
 Immersions in ${\mathbb{R}}^{3}$ which are degenerate points of the map $\mathcal{C}$} have been first studied in the late XIXth century under the name of isothermic surfaces (see [168]).
-There are the surfaces, such that, away from umbilic points there exists locally conformal coordinates which give the principal directions.
-Assuming now that $(\Sigma,g_{\vec{\Phi}})$ is a non degenerate point under some constraint on the conformal class of the induced metric then there exists an holomorphic quadratic form $q$ such that
+These are the surfaces, such that, away from umbilic points there exists locally conformal coordinates which give the principal directions.
+Assuming now that $(\Sigma,g_{\vec{\Phi}})$ is a non degenerate point under some constraint on the conformal class of the induced metric then there exists a holomorphic quadratic form $q$ such that
 $d\ast_{g}\left[d\vec{H}-\,3\,\pi_{\vec{n}}(d\vec{H})+\star(\ast_{g}\ d\vec{n}\wedge\vec{H})\right]=<q,\vec{h}_{0}>_{WP}\ .$ (II.2)
@@
 The holomorphic quadratic form is a Lagrange multiplier [^7] The holomorphic quadratic form identifies to the tangent space to the moduli space (theorem 4.2.2 [94]) associated to the conformal class constraint.
-Weak immersions which are solving (II.2) are proved to be smooth ([18]).
+Weak immersions solving (II.2) are proved to be smooth ([18]).
 Remark II.23.
 In [18] it is proved that the Conformal Willmore equation (II.3) is equivalent to the existence of $\vec{L}$ satisfying the set of the two conservation laws (I.11) and (I.12)
-$\left\{\begin{array}[]{l}\displaystyle d\left(\vec{L}\cdot d\vec{\Phi}\right)=0\\[14.22636pt] \displaystyle d\left(\vec{L}\times d\vec{\Phi}+H\,d\vec{\Phi}\right)=0\end{array}\right$. (II.4)
+$\left\{\begin{array}[]{l}\displaystyle d\left(\vec{L}\cdot d\vec{\Phi}\right)=0,\\[14.22636pt] \displaystyle d\left(\vec{L}\times d\vec{\Phi}+H\,d\vec{\Phi}\right)=0,\end{array}\right$. (II.4)
 coming respectively from Noether theorem for dilations and rotations which is itself equivalent to the so-called $(\vec{R},S)$ system (I.13) combined with the structure equation (I.14).
-For the higher dimensional counterparts to Willmore functional and in particular for the Dirichlet energy of the mean curvature of a weak immersion of 4-dimensional manifolds into ${\mathbb{R}}^{5}$, we are asking in Open Question VI.4 whether the conservation laws issued from Noether for dilations and rotations do correspond to the variation of this energy under some constraint.
+For the higher dimensional counterparts to the Willmore functional and in particular for the Dirichlet energy of the mean curvature of a weak immersion of 4-dimensional manifolds into ${\mathbb{R}}^{5}$, we are asking in Open Question VI.4 whether the conservation laws issued from Noether for dilations and rotations do correspond to the variation of this energy under some constraint.
 An interesting question is to know whether or not a weak immersion which is both a critical point of the Willmore functional under some constraint on the conformal class and isothermic (i.e.
@@
 II.2 In higher dimensions
-Even though, geometric properties of generalized Willmore functionals have already been the subject to lots of works, the study of the analytical properties is relatively new.
+Even though geometric properties of generalized Willmore functionals have already been the subject to a lot of works, the study of their analytical properties is relatively new.
 We list below the few known results and propose different directions of research.
@@
 The example constructed in [129] is based on the observation that the integrand of $\mathcal{E}_{GR}$ is a negative constant for $\mathbb{S}^{2}\times\mathbb{R}^{2}$, however this is not a closed submanifold.
-As mentioned by Graham in a private communication, a simple way to obtain a degenerating example of immersions from a closed manifold is to consider an ellipsoid with two semi-axis of length 1 and two others equal to some parameter $a\in(0,+\infty)$ that we can either send to 0 or to $\infty$ (since $\mathcal{E}_{GR}$ is scale invariant, both possibilities are equivalent and we recover $\mathbb{S}^{2}\times\mathbb{R}^{2}$ by sending $a\to+\infty$).
+As mentioned by Graham in a private communication, a simple way to obtain a degenerating example of immersions from a closed manifold is to consider an ellipsoid with two semi-axes of length 1 and two others equal to some parameter $a\in(0,+\infty)$ that we can either send to 0 or to $\infty$ (since $\mathcal{E}_{GR}$ is scale invariant, both possibilities are equivalent and we recover $\mathbb{S}^{2}\times\mathbb{R}^{2}$ by sending $a\to+\infty$).
 At first glance, it seems that non-compact CMC immersions can provide an interesting source of examples for possible degeneracy of $\mathcal{E}_{GR}$ towards -$\infty$.
@@
 Open Question II.27.
-Does there exist critical points of $\mathcal{E}_{GR}$ which are not conformal transformations of minimal hypersurfaces?
+Are there critical points of $\mathcal{E}_{GR}$ which are not conformal transformations of minimal hypersurfaces?
 We now list a few natural questions regarding the two-dimensional case. First, it would be important to know if a generalization of Bryant’s quartic exists or not.
@@
 The key idea of [23, 69], which has recently been developed in [6], is the link between $Q$-curvature and generalized Willmore energies.
-They show in [69], that one can define a notion of extrinsic GJMS operators on submanifolds.
-From these operators, one deduce a notion of extrinsic $Q$-curvature.
+They show in [69] that one can define an extrinsic version of the GJMS (Graham–Jenne–Mason–Sparling) operators [73] on submanifolds.
+From these operators, one deduces a notion of extrinsic $Q$-curvature.
 In [23], the authors prove that for a 4-dimensional submanifold $\Sigma^{4}\subset M^{n}$ with $n\geq 5$, the integral of the difference between the extrinsic $Q$-curvature and the intrinsic $Q$-curvature of $\Sigma$ is a generalized Willmore energy.
 Hence, one could hope to find similarities between the study of the standard $Q$-curvature and generalized Willmore energies.
-In order to study the space of immersions of a given hypersurface $\Sigma^{4}\subset\mathbb{R}^{5}$ into $\mathbb{R}^{5}$ and to provide a notion of ”best representation”, one would need a functional which is bounded from below and, as for the 2-dimensional case, would be a sort of conformally invariant distance of hypersurfaces to totally umbilic sets.
+In order to study the space of immersions of a given hypersurface $\Sigma^{4}\subset\mathbb{R}^{5}$ into $\mathbb{R}^{5}$ and to provide a notion of “best representation”, one would need a functional which is bounded from below and, as for the 2-dimensional case, would be a sort of conformally invariant distance of hypersurfaces to totally umbilic sets.
 One possibility has been introduced in [129] motivated by the generalization of the Dirichlet energy for the conformal Gauss map in dimension 2, see also [6] where Allen–Gover propose an alternative in any codimension using the normal tractor: if $\vec{\Phi}\colon\Sigma^{4}\to\mathbb{R}^{5}$ is a smooth immersion, then
@@
 For example, one could consider conformal transformations of hypersurfaces satisfying an equation of the form $P(H)=0$ for some polynomial $P$.
-It is possible that $\mathcal{E}$ has a specific structure for which, there is no minimizers.
+It is possible that $\mathcal{E}$ has a specific structure for which there are no minimizers.
 In this case, it would be interesting to find another functional, coercive as well and which does possess minimizers.
@@
 A generalization of this inequality has recently been proved in [193].
-II.2.3 Analysis in high dimension
+II.2.3 Analysis in high dimensions
 In order to answer the previous questions, one would need to develop a refined analysis for submanifolds in the spirit of the approach presented here and most likely combined with the existing theory of varifolds.
-An important starting point already raised in [80, Remark 5.4.6], would be to generalize the study of Coulomb frames in higher dimensions but also to generalize the work of Müller–Šverák [151] on the construction of isothermal coordinates for immersed surfaces in a Euclidean space.
-The Coulomb frame are of major importance due to their relation to isothermal coordinates, see for instance Theorem IV.4 below or [80, Section 5.4].
+An important starting point already raised in [81, Remark 5.4.6], would be to generalize the study of Coulomb frames in higher dimensions but also to generalize the work of Müller–Šverák [154] on the construction of isothermal coordinates for immersed surfaces in a Euclidean space.
+The Coulomb frames are of major importance due to their relation to isothermal coordinates, see for instance Theorem IV.4 below or [81, Section 5.4].
 In higher dimensions, isothermal coordinates do not exist in general, the Cotton and Weyl tensors are the obstruction.
@@
 Given a closed orientable $n$-dimensional manifold $\Sigma$ and an immersion $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{d}$ with $\mathrm{I\!I}_{\vec{\Phi}}\in L^{n}(\Sigma^{n})$, is it possible to construct counterexamples to the existence of controlled harmonic coordinates?
-Or can we construct controlled harmonic coordinates on a ”large” subset of $\Sigma$?
-The next question is to know whether this atlas of harmonic coordinates can be ”uniformized” to obtain a global metric on $\Sigma$ conformal to $g_{\vec{\Phi}}$ together with some control following from an estimate of the form (II.7).
+Or can we construct controlled harmonic coordinates on a “large” subset of $\Sigma$?
+The next question is to know whether this atlas of harmonic coordinates can be “uniformized” to obtain a global metric on $\Sigma$ conformal to $g_{\vec{\Phi}}$ together with some control following from an estimate of the form (II.7).
 A first attempt would be to find a metric having constant scalar curvature.
-This question is at the origin of the Yamabe problem and is not solvable in full generality, the topology of $\Sigma$ plays a role.
+This question is at the origin of the Yamabe problem and is not solvable in full generality, the topology of $\Sigma$ plays a role when it has a boundary.
 An alternative has been proposed recently in [108], where the authors are able to solve locally $\Delta_{\bar{g}}\mathrm{Scal}_{\bar{g}}=0$ for some metric $\bar{g}$ conformal to $g$ in dimension 4.
@@
 $\forall u\in C^{\infty}_{c}(\Sigma),\qquad\|\nabla u\|_{L^{4}(\Sigma,g_{\vec{\Phi}})}\leq C\,\|\Delta_{g_{\vec{\Phi}}}u\|_{L^{2}(\Sigma,g_{\vec{\Phi}})}.$ (II.8)
-Sobolev inequalities are a consequence of the isoperimetric inequality, which always hold for submanifolds having small $L^{n}$-norm of its mean curvature, thanks to Michael–Simon inequality [5, 135].
+Sobolev inequalities are a consequence of the isoperimetric inequality, which always holds for submanifolds with a small $L^{n}$-norm of its mean curvature, thanks to the Michael–Simon inequality [5, 139].
 A global version of the equation $\Delta_{\bar{g}}\mathrm{Scal}_{\bar{g}}=0$ would be that the $Q$-curvature of $\bar{g}$ is pointwise controlled by $|\mathrm{Riem}^{\bar{g}}|^{2}_{\bar{g}}$.
@@
 Given an immersion $\vec{\Phi}\colon\Sigma\to\mathbb{R}^{m}$ from a closed $4$-dimensional oriented manifold $\Sigma$ verifying (II.7), can we find a metric $\bar{g}$ conformal to $g_{\vec{\Phi}}$ such that the $Q$-curvature of $\bar{g}$ is constant?
-Is pointwise controlled by $|\vec{\mathrm{I\!I}}_{\vec{\Phi}}|^{4}_{g_{\vec{\Phi}}}$?
+Is it pointwise controlled by $|\vec{\mathrm{I\!I}}_{\vec{\Phi}}|^{4}_{g_{\vec{\Phi}}}$?
 In order to know whether an equation on a manifold is regularizing or not, one needs to develop a refined PDE theory.
@@
 Such inequalities are directly related to isoperimetric problems, see for instance [33, Chapter IV].
-Gallot [62] proved that the isoperimetric inequalities hold on a complete manifold $(\Sigma^{n},g)$ as soon as $\mathrm{Ric}_{g}\in L^{p}(\Sigma,g)$ for some $p>\frac{n}{2}$.
-Moreover Micheal–Simon inequality (see Brendle [26] for the optimal constants) implies that an isoperimetric inequality always holds for immersed $n$-dimensional submanifolds of arbitrary codimension in a Euclidean space whose mean curvature has small $L^{n}$-norm.
+Gallot [62] proved that the isoperimetric inequality holds on a complete manifold $(\Sigma^{n},g)$ as soon as $\mathrm{Ric}_{g}\in L^{p}(\Sigma,g)$ for some $p>\frac{n}{2}$.
+Moreover, the Michael–Simon inequality (see Brendle [26] for the optimal constants) implies that an isoperimetric inequality always holds for immersed $n$-dimensional submanifolds of arbitrary codimension in a Euclidean space whose mean curvature has small $L^{n}$-norm.
 Open Question II.38.
-Given a complete manifold $(\Sigma^{n},g)$, does there exists a number $s>0$ such that if $\mathrm{Riem}_{g}\in W^{s,\frac{n}{s+2}}(\Sigma,g)$, then the isoperimetric inequality holds?
+Given a complete manifold $(\Sigma^{n},g)$, is there a number $s>0$ such that if $\mathrm{Riem}_{g}\in W^{s,\frac{n}{s+2}}(\Sigma,g)$, then the isoperimetric inequality holds?
 In order to develop a variational theory for submanifolds of the Euclidean space, one needs to understand the ”closure” of this space under the norm we are interested in, in the same spirit that the notions of currents or varifolds arise as limits of smooth submanifolds having bounded area.
@@
 Assume that
-$\sup_{k\in\mathbb{N}}\left\|\vec{\mathrm{I\!I}}_{\vec{\Phi}_{k}}\right\|_{W^{s,\frac{n}{s+1}}(\Sigma,g_{\vec{\Phi}})}\leq C.$
+$\sup_{k\in\mathbb{N}}\left\|\vec{\mathrm{I\!I}}_{\vec{\Phi}_{k}}\right\|_{W^{s,\frac{n}{s+1}}(\Sigma,g_{\vec{\Phi}})}<+\infty.$
 What is the limit ${\lim_{k\to+\infty}}\vec{\Phi}_{k}$? Can we develop a bubbling analysis adapted to the above energy?
@@
 Hence, one should not expect a strong limit in full generality, but rather a weak limit on $\Sigma$ minus a set having small measure.
-The case $s=\frac{n-2}{2}$ and $n$ even has been treated in [129], based on the approximation result obtained in Theorem IV.23.
+The case $s=\frac{n-2}{2}$ and $n\geq 4$ even has been treated in [133], based on the approximation result obtained in Theorem IV.23.
 Theorem II.40.
@@
 Up to rescaling, we can assume that each $\vec{\Phi}_{k}(\Sigma)$ is contained in the unit ball $\mathbb{B}^{d}(0,1)$ and has diameter 1.
-Then, there exist a closed $C^{1}$ manifold $\Sigma_{\infty}$ (possibly non-connected) and a weak immersion $\vec{\Phi}_{\infty}\colon\Sigma_{\infty}\to\mathbb{R}^{d}$ such that the following holds.
+Then, there exists a closed $C^{1}$ manifold $\Sigma_{\infty}$ (possibly non-connected) and a weak immersion $\vec{\Phi}_{\infty}\colon\Sigma_{\infty}\to\mathbb{R}^{d}$ such that the following holds.
 There exists a finite number of points $\vec{q}_{1},\ldots,\vec{q}_{K}\in\mathbb{R}^{d}$ such that for each $r>0$, we have the following convergence in the pointed Gromov–Hausdorff topology
@@
 Let $\vec{\Phi}\colon\mathbb{B}^{4}(0,1)\to\mathbb{R}^{5}$ be a weak immersion critical point of the functional
-$E(\vec{\Phi})=\int_{\mathbb{B}^{4}(0,1)}\big|\nabla H_{\vec{\Phi}}\big|^{2}_{g_{\vec{\Phi}}}\ d\textup{vol}_{g_{\vec{\Phi}}}.$
+$E(\vec{\Phi})=\int_{\mathbb{B}^{4}(0,1)}\big|dH_{\vec{\Phi}}\big|^{2}_{g_{\vec{\Phi}}}\ d\textup{vol}_{g_{\vec{\Phi}}}.$
 Then $\vec{\Phi}$ is real-analytic in harmonic coordinates.
@@
   3. (iii)
-    For a real-valued function $f$, we denote $f^{+}\coloneqq\max(f,0)$, $f^{-}\coloneqq\max(-f,0)$.
+    For a real-valued function $f$, we denote $f^{+}\coloneqq\max(f,0)$ and $f^{-}\coloneqq\max(-f,0)$.
   4. (iv)
@@
     We denote by $\mathcal{L}^{n}$ the Lebesgue measure on $\mathbb{R}^{n}$.
-For $U\subset\mathbb{R}^{n}$ with $\mathcal{L}^{n}(U)<\infty$, $f\in L^{1}(U)$, we write $\fint_{U}f\coloneqq(\mathcal{L}^{n}(U))^{-1}\int_{U}f$.
+For $U\subset\mathbb{R}^{n}$ with $\mathcal{L}^{n}(U)<\infty$ and $f\in L^{1}(U)$, we write $\fint_{U}f\coloneqq(\mathcal{L}^{n}(U))^{-1}\int_{U}f$.
   7. (vii)
@@
 The Hodge star operator $_{g}$ is defined as the unique linear operator from $\bigwedge^{k}V$ to $\bigwedge^{n-k}V$ satisfying α∧*_g  β=⟨α,β⟩_g  e_1∧⋯∧e_n,   for all α,β∈⋀^kV.
-See for instance [94, Section 3.3].
+See for instance [95, Section 3.3].
 When $V=\mathbb{R}^{n}$ and $g=g_{\text{std}}$ is the standard inner product on $\mathbb{R}^{n}$, we write $\star$ instead of $_{g}$.
@@
     Let $U\subset\mathbb{R}^{n}$ be an open set and let $g$ be a Riemannian metric on $U$.
-On the space of differential $\ell$-forms on $U$, we define the codifferential (see [94, Definition 3.3.1]) by
+On the space of differential $\ell$-forms on $U$, we define the codifferential (see [95, Definition 3.3.1]) by
     $d^{*_{g}}\coloneqq(-1)^{\ell}*_{g}^{-1}d\,*_{g}=(-1)^{n(\ell+1)+1}*_{g}d\,*_{g}.$ (III.1)
@@
   14. (xiv)
-    For $\vec{u}=(\vec{u}_{1},\vec{u}_{2})$, $\vec{v}=(\vec{v}_{1},\vec{v}_{2})$, where $\vec{u}_{i},\vec{v}_{i}\in\mathbb{R}^{3}$, denote →u⋅→v≔∑_i=1^2 →u_i ⋅→v_i,  →u×→v≔∑_i=1^2 →u_i ×→v_i.
+    For $\vec{u}=(\vec{u}_{1},\vec{u}_{2})$, $\vec{v}=(\vec{v}_{1},\vec{v}_{2})$, where $\vec{u}_{i},\vec{v}_{i}\in\mathbb{R}^{3}$, we denote →u⋅→v≔∑_i=1^2 →u_i ⋅→v_i,  →u×→v≔∑_i=1^2 →u_i ×→v_i.
   15. (xv)
@@
 Set $g=g_{\vec{\Phi}}=\vec{\Phi}^{*}g_{\text{std}}$, where $g_{\textup{std}}$ is the Euclidean metric on $\mathbb{R}^{m}$.
-Denote g_ij=∂_i→Φ⋅∂_j→Φ, (g^ij)=(g_ij)^-1, detg=det(g_ij), dvol_g=(detg)^12dx^1∧⋯∧dx^n.
-The pullback metric induces a pairing on $T^{*}U$ with $\langle dx^{i},dx^{j}\rangle_{g}=g^{ij}$; we denote by $|\cdot|_{g}$ the associated pointwise norm.
+We denote g_ij=∂_i→Φ⋅∂_j→Φ, (g^ij)=(g_ij)^-1, detg=det(g_ij), dvol_g=(detg)^12dx^1∧⋯∧dx^n.
+The pullback metric induces a pairing on $T^{*}U$ with $\langle dx^{i},dx^{j}\rangle_{g}=g^{ij}$.
+We denote by $|\cdot|_{g}$ the associated pointwise norm.
 We also write $|\cdot|_{\mathbb{R}^{m}}$ for the Euclidean norm in $\mathbb{R}^{m}$ (abbreviated to $|\cdot|$ when unambiguous).
@@
     $\vec{n}\coloneqq\star\,\frac{\partial_{1}\vec{\Phi}\wedge\cdots\wedge\partial_{n}\vec{\Phi}}{|\partial_{1}\vec{\Phi}\wedge\cdots\wedge\partial_{n}\vec{\Phi}|},$
-    $\vec{\mathrm{I\!I}}(X,Y)\coloneqq\pi_{\vec{n}}X(d\vec{\Phi}(Y)),\qquad\text{for all }X,Y\in T_{p}U\text{ and all }p\in U.$
+    $\vec{\mathrm{I\!I}}(X,Y)\coloneqq\pi_{\vec{n}}X(d\vec{\Phi}(Y)),\qquad\forall X,Y\in T_{p}U,\ \forall p\in U.$
     The mean curvature vector is defined by
@@
 Lemma III.1.
-[168, Lemma 1.8] Let $\Sigma$ be a $2$-dimensional oriented smooth manifold, and $\vec{\Phi}\colon\Sigma\rightarrow\mathbb{R}^{3}$ be a smooth immersion.
+[171, Lemma 1.8] Let $\Sigma$ be a $2$-dimensional oriented smooth manifold, and $\vec{\Phi}\colon\Sigma\rightarrow\mathbb{R}^{3}$ be a smooth immersion.
 Let $_{g}$ be the Hodge star operator with respect to $g=g_{\vec{\Phi}}$.
@@
 $g_{ij}\coloneqq\partial_{i}\vec{\Phi}\cdot\partial_{j}\vec{\Phi},\quad(g^{ij})=(g_{ij})^{-1},\quad\det g=\det(g_{ij}).$
-For any $k\in\{1,2\}$, note
-$\partial_{k}(\vec{n}\cdot\vec{n})=2\,\partial_{k}\vec{n}\cdot\vec{n}=0,$
-hence
+For any $k\in\{1,2\}$, we have
+$\partial_{k}(\vec{n}\cdot\vec{n})=2\,\partial_{k}\vec{n}\cdot\vec{n}=0.$
+Hence, it holds
 $\partial_{k}\vec{n}=g^{ij}(\partial_{k}\vec{n}\cdot\partial_{j}\vec{\Phi})\,\partial_{i}\vec{\Phi}.$ (III.11)
@@
 where the third equality follows from the identity
-$(\vec{u}\times\vec{v})\times\vec{w}=(\vec{u}\cdot\vec{w})\vec{v}-(\vec{v}\cdot\vec{w})\vec{u}$
-for all $\vec{u},\vec{v},\vec{w}\in\mathbb{R}^{3}$. By (III.11), we also get
-$dx_{s}\wedge d\vec{n}$ $=\varepsilon^{sk}\,\partial_{k}n\,\alpha=-g^{ij}\epsilon^{sk}\,\mathrm{I\!I}_{kj}\,\partial_{i}\vec{\Phi}\,\alpha,$
-Therefore,
+$\forall,\vec{u},\vec{v},\vec{w}\in\mathbb{R}^{3},\qquad(\vec{u}\times\vec{v})\times\vec{w}=(\vec{u}\cdot\vec{w})\vec{v}-(\vec{v}\cdot\vec{w})\vec{u}.$
+By (III.11), we also get
+$dx_{s}\wedge d\vec{n}$ $=\varepsilon^{sk}\,\partial_{k}n\,\alpha=-g^{ij}\epsilon^{sk}\,\mathrm{I\!I}_{kj}\,\partial_{i}\vec{\Phi}\,\alpha.$
+Therefore, it holds
 $\begin{aligned} dx_{s}\wedge(d\vec{n}+\vec{n}\times*_{g}\,d\vec{n})&=-(g^{sj}\varepsilon^{ki}+g^{ij}\varepsilon^{sk})\,\mathrm{I\!I}_{kj}\,\partial_{i}\vec{\Phi}\,\alpha\\[2.84526pt] &=-\varepsilon^{si}g^{kj}\,\mathrm{I\!I}_{kj}\,\partial_{i}\vec{\Phi}\,\alpha\\[2.84526pt] &=-2H\,dx_{s}\wedge d\vec{\Phi},\end{aligned}$ (III.12)
-where the second equality follows from the identity
-$f(s)\varepsilon^{ki}+f(k)\varepsilon^{is}+f(i)\varepsilon^{sk}=0$
-which holds for any $i,k,s\in\{1,2\}$ and real-valued function $f$ defined on $\{1,2\}$. Since (III.12) holds for any $s\in\{1,2\}$, (III.10) follows immediately. ∎
+where the second equality follows from the identity, valid for any $i,k,s\in\{1,2\}$ and real-valued function $f$ defined on $\{1,2\}$
+$f(s)\varepsilon^{ki}+f(k)\varepsilon^{is}+f(i)\varepsilon^{sk}=0.$
+Since (III.12) holds for any $s\in\{1,2\}$, the relation (III.10) follows immediately. ∎
 III.3 Some estimates for elliptic equations
 We first introduce a standard Morrey estimate for elliptic equations of divergence form with no lower-order terms.
-Lemma III.2 ([77, Lemma 4.12]).
-Let $n\in\mathbb{N}^{+}$. Suppose $\{a^{ij}\}_{i,j=1}^{n}\subset L^{\infty}(B^{n})$ satisfies
-$\Lambda^{-1}|\xi|^{2}\leq a^{ij}(x)\,\xi_{i}\,\xi_{j}\leq\Lambda|\xi|^{2},\qquad\text{for a.e. }x\in B^{n}\text{ and all }\,\xi\in\mathbb{R}^{n},$ (III.13)
-where $\Lambda>0$ is a constant. Suppose $u\in W^{1,2}(B^{n})$ is a weak solution of the equation $\partial_{j}(a^{ij}\,\partial_{i}u)=0$ in $B^{n}$, i.e.
-$\int_{B^{n}}a^{ij}\,\partial_{i}u\,\partial_{j}\varphi=0,\qquad\text{for all }\,\varphi\in C^{\infty}_{c}(B^{n}).$
+Lemma III.2 ([78, Lemma 4.12]).
+Let $n\in\mathbb{N}^{+}$ and $\Lambda>0$. Suppose $\{a^{ij}\}_{i,j=1}^{n}\subset L^{\infty}(B^{n})$ satisfies
+$\Lambda^{-1}|\xi|^{2}\leq a^{ij}(x)\,\xi_{i}\,\xi_{j}\leq\Lambda|\xi|^{2},\qquad\text{for a.e. }x\in B^{n}\text{ and all }\,\xi\in\mathbb{R}^{n}.$ (III.13)
+Suppose $u\in W^{1,2}(B^{n})$ is a weak solution of the equation $\partial_{j}(a^{ij}\,\partial_{i}u)=0$ in $B^{n}$, i.e.
+$\int_{B^{n}}a^{ij}\,\partial_{i}u\,\partial_{j}\varphi=0,\qquad\forall\varphi\in C^{\infty}_{c}(B^{n}).$
 Then there exist $\alpha=\alpha(n,\Lambda)\in(0,1)$ and $C=C(n,\Lambda)>0$ such that for any $0<r<1$ there holds
@@
 For $u,v\in W^{1,2}(D^{2})$, there exists a unique solution $\phi\in W^{1,2}_{0}(D^{2})$ to the problem
-$\begin{cases}\partial_{i}(a^{ij}\,\partial_{j}\phi)=\partial_{1}u\,\partial_{2}v-\partial_{2}u\,\partial_{1}v&\mbox{in }D^{2},\\[5.69054pt] \phi=0&\mbox{on }\partial D^{2},\end{cases}$
-and we also have $\phi\in C(\overline{D^{2}})$ with the estimate
+$\begin{cases}\partial_{i}(a^{ij}\,\partial_{j}\phi)=\partial_{1}u\,\partial_{2}v-\partial_{2}u\,\partial_{1}v&\mbox{in }D^{2},\\[5.69054pt] \phi=0&\mbox{on }\partial D^{2}.\end{cases}$
+We have $\phi\in C(\overline{D^{2}})$ with the estimate
 $\|\phi\|_{L^{\infty}(D^{2})}+\|\nabla\phi\|_{L^{2}(D^{2})}\leq C(\Lambda)\,\|\nabla u\|_{L^{2}(D^{2})}\,\|\nabla v\|_{L^{2}(D^{2})}.$
-We have a stronger version of this result, see Proposition IV.8.
+We prove a stronger version of this result in Proposition IV.8 below.
 III.4 Bourgain–Brezis inequality
@@
 In the case $n=2$, F.
-Da Lio and T.
-Rivière [48] provided a proof for a stronger version of (III.14) in 2020.
+Da Lio and the third author [48] provided a proof for a stronger version of (III.14) in 2020.
 Here we denote $\dot{W}^{-1,2}(\mathbb{R}^{2})$ as the dual space of the homogeneous Sobolev space $\dot{W}^{1,2}(\mathbb{R}^{2})$.
@@
 Assume $g\in L^{1}(\mathbb{R}^{2},\mathbb{R}^{2})$, $f\in\dot{W}^{-1,2}(\mathbb{R}^{2},\mathbb{R}^{2})$, $u$ be a tempered distribution on $\mathbb{R}^{2}$ such that $\nabla u=f+g$.
-Then there exists $c\in\mathbb{R}$, and a universal constant $C$ such that
+Then there exist $c\in\mathbb{R}$ and a universal constant $C$ such that
 $\|u-c\|_{L^{2}(\mathbb{R}^{2})}\leq C(\|f\|_{\dot{W}^{-1,2}(\mathbb{R}^{2})}+\|g\|_{L^{1}(\mathbb{R}^{2})}).$
@@
 We call the function $\mu_{f}\coloneqq\partial_{\bar{z}}f/\partial_{z}f$ the complex dilatation of $f$.
-Lemma III.7 ([90, Theorem 4.30]).
-Let $\mu\in L^{\infty}(\mathbb{C},\mathbb{C})$ satisfies $\|\mu\|_{L^{\infty}(\mathbb{C})}<1$.
-Then there exists a homeomorphism $f\colon\bar{\mathbb{C}}\rightarrow\bar{\mathbb{C}}$ which is a quasiconformal map of $\mathbb{C}$ and satisfies the Beltrami equation $\partial_{\bar{z}}f=\mu\,\partial_{z}f$.
+Lemma III.7 ([91, Theorem 4.30]).
+Let $\mu\in L^{\infty}(\mathbb{C},\mathbb{C})$ be such that $\|\mu\|_{L^{\infty}(\mathbb{C})}<1$.
+Then there exists a quasiconformal homeomorphism $f\colon\bar{\mathbb{C}}\rightarrow\bar{\mathbb{C}}$ verifying the Beltrami equation $\partial_{\bar{z}}f=\mu\,\partial_{z}f$.
 Such $f$ is uniquely determined by the conditions $f(0)=0$, $f(1)=1$, $f(\infty)=f(\infty)$.
@@
 $\|\det(\nabla u)\|_{\mathcal{H}^{1}(\mathbb{R}^{n})}\leq C(n)\|u\|_{W^{1,n}(\mathbb{R}^{n})}.$
-Standard singular integral theory (see for instance [182, Theorem III.4]) implies the following results.
+Standard singular integral theory (see for instance [185, Theorem III.4]) implies the following results.
 Lemma III.10.
@@
 Similarly, for $f\in W^{1,1}_{\textup{loc}}(U)$ we say $df\in VMO(U)$ if for any coordinate system $(V,\phi)$, $\xi\in C_{c}^{\infty}(V)$ there holds $\nabla((\xi f)\circ\phi^{-1})\in VMO(\phi(V\cap U),\mathbb{R}^{n})$.
-Lemma III.13 ([92, Lemma 2.3]).
+Lemma III.13 ([93, Lemma 2.3]).
 Let $B_{R}\subset\mathbb{R}^{n}$ be a ball of radius $R$.
@@
 $\left|\fint_{B}f-\fint_{B^{\prime}}f\right|\leq C(n,K)\,\beta.$
-Then by induction we have
-$\left|\fint_{B}f-\fint_{2\tilde{B}}f\right|\leq C(n,K)\,i\beta,\qquad\text{for all }\,i\geq 0\,\text{ and all }B\in\mathcal{G}_{i}.$
+Then by induction we have for all $i\geq 0$ and all $B\in\mathcal{G}_{i}$,
+$\left|\fint_{B}f-\fint_{2\tilde{B}}f\right|\leq C(n,K)\,i\beta.$
 Finally, since $\operatorname{dist}(B,\partial B_{1}(0))=\sigma^{i}-2\sigma^{i}(1-\sigma)>2K\sigma^{i}(1-\sigma)$ for any $i\geq 0$ and $B\in\mathcal{G}_{i}$, we have
@@
 $\leq C(n,K)\sum_{i=0}^{\infty}\sigma^{i}(i+1)\beta\leq C(n,K)\beta.$
-Lemma III.14 ([162, Theorem 2], [92, Section 3]).
+Lemma III.14 ([165, Theorem 2], [93, Section 3]).
 Let $U$ be an open subset of $\mathbb{R}^{n}$, $f\in VMO(U)$, $\varphi\in W^{1,n}_{\textup{loc}}(U,\mathbb{R}^{n})$ be a homeomorphism onto its image.
@@
 Then we have $f\circ\varphi^{-1}\in VMO(\varphi(U))$ with
-$\beta_{r,\varphi(U)}(f\circ\varphi^{-1})\leq C(K)\,\beta_{r,U}(f)\qquad\text{for all }\,r>0.$
+$\forall r>0,\qquad\beta_{r,\varphi(U)}(f\circ\varphi^{-1})\leq C(K)\,\beta_{r,U}(f).$
 Corollary III.15 ([93]).
@@
 Then there exists $\tilde{f}\in VMO(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$ such that $\tilde{f}=f$ on $U$ satisfying the inequality
-$\beta_{r,\mathbb{R}^{n}}(\tilde{f})\leq C(U)\big(\beta_{r,U}(f)+r\|f\|_{L^{\infty}(U)}\big),\quad\text{for all }\,r>0.$
+$\forall r>0,\qquad\beta_{r,\mathbb{R}^{n}}(\tilde{f})\leq C(U)\big(\beta_{r,U}(f)+r\|f\|_{L^{\infty}(U)}\big).$
 III.8 Sobolev–Lorentz spaces
@@
 We record the following form of Hölder’s inequality for Lorentz spaces.
-Lemma III.18 ([89, Thm. 4.5]).
+Lemma III.18 ([90, Thm. 4.5]).
 Assume $f_{1}\in L^{p_{1},q_{1}}(U)$ and $f_{2}\in L^{p_{2},q_{2}}(U)$, with $p,p_{1},p_{2}\in[1,\infty)$, $q,q_{1},q_{2}\in[1,\infty]$, and $1/p=1/p_{1}+1/p_{2}$, $1/q\leq 1/q_{1}+1/q_{2}$.
@@
 $\|f\|_{W^{k,p}_{0}(U)}\coloneqq\|\nabla^{k}f\|_{L^{p}(U)}.$
-If $1<p<\infty$, $1<q\leq\infty$, and $1/p+1/p^{\prime}=1$, $1/q+1/q^{\prime}=1$, then the same argument as in [133, Section 1.1.15] implies
+If $1<p<\infty$, $1<q\leq\infty$, and $1/p+1/p^{\prime}=1$, $1/q+1/q^{\prime}=1$, then the same argument as in [137, Section 1.1.15] implies
 $W^{-k,(p,q)}(U)=\bigl(W^{k,(p^{\prime},q^{\prime})}_{0}(U)\bigr)^{\!*}.$ (III.20)
-Applying estimates for Riesz potentials, we obtain the following embedding results for Sobolev-Lorentz spaces, see for instance [145, Eqs.
+Applying estimates for Riesz potentials, we obtain the following embedding results for Sobolev–Lorentz spaces, see for instance [148, Eqs.
 (1.3)–(1.5)], [13, Ch.
@@
     If $f\in L^{1}(U)$, then $f\in W^{-1,(\frac{n}{n-1},\infty)}(U)$, with
-    $\|f\|_{W^{-1,(\frac{n}{n-1},\infty)}(U)}\leq C(n)\|f\|_{L^{1}(U)}.$
+    $\|f\|_{W^{-1,\left(\frac{n}{n-1},\infty\right)}(U)}\leq C(n)\|f\|_{L^{1}(U)}.$
   2. (ii)
     If $f\in L^{p,q}(U)$, then $f\in W^{-1,(\frac{np}{n-p},q)}(U)$, with
-    $\|f\|_{W^{-1,(\frac{np}{n-p},q)}(U)}\leq C(n,p)\|f\|_{L^{p,q}(U)}.$
+    $\|f\|_{W^{-1,\left(\frac{np}{n-p},q\right)}(U)}\leq C(n,p)\|f\|_{L^{p,q}(U)}.$
 IV Weak immersions
@@
 Definition IV.1.
-Let $\Sigma$ be an $n$-dimensional closed smooth manifold with a fixed smooth reference metric $g_{0}$, $m,k\in\mathbb{N}$, $1\leq p\leq\infty$.
+Let $\Sigma$ be an $n$-dimensional closed smooth manifold with a fixed smooth reference metric $g_{0}$.
+Fix $m,k\in\mathbb{N}$ and $1\leq p\leq\infty$.
 We say that $\vec{\Phi}$ is a $W^{k,p}$ weak immersion and write $\vec{\Phi}\in W^{k,p}_{\text{imm}}(\Sigma,\mathbb{R}^{m})$ if $\vec{\Phi}\in W^{k,p}\cap W^{1,\infty}(\Sigma,\mathbb{R}^{m})$ and there exists a constant $\Lambda>0$ such that the following holds for a.e.
@@
 Finally, we prove that a weak immersion can always be approximated by smooth immersions and that the underlying conformal structure (in the case of surfaces) converges as well in Section IV.4.
-In all this paper, we will denote $D^{2}$ the unit disk of $\mathbb{R}^{2}$.
+Throughout this paper, we denote by $D^{2}$ the unit disk of $\mathbb{R}^{2}$.
 IV.1 Existence of conformal coordinates
@@
 Definition IV.2.
-Let $\Sigma$ be a $2$-dimensional closed smooth manifold, $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$, $U\subset\Sigma$ be open.
+Let $\Sigma$ be a $2$-dimensional closed smooth manifold.
+Consider a weak immersion $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$ and $U\subset\Sigma$ an open set.
 Let $\varphi\colon U\subset\Sigma\rightarrow\varphi(U)\subset\mathbb{R}^{2}$ be a $W^{1,2}_{\textup{loc}}$ bi-Sobolev homeomorphism (i.e.
@@
 We define isothermal charts for $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(D^{2},\mathbb{R}^{m})$ similarly.
-Throughout this paper, we frequently use the notion of non-smooth metrics and their conformal class.
+Throughout this paper, we will frequently use the notion of non-smooth metrics and their conformal class.
 Definition IV.3.
 Let $\Sigma$ be an $n$-dimensional smooth manifold.
-Let $g$ be a measurable section on the bundle $T^{*}\Sigma\otimes T^{*}\Sigma$, we say $g$ is a metric if $g_{p}$ is symmetric and positive definite on $(T_{p}\Sigma)^{2}$ for a.e.
+Let $g$ be a measurable section on the bundle $T^{*}\Sigma\otimes T^{*}\Sigma$, we say that $g$ is a metric if $g_{p}$ is symmetric and positive definite on $(T_{p}\Sigma)^{2}$ for a.e.
 $p\in\Sigma$.
-Let $g_{1},g_{2}$ be two metrics, we say $g_{1}$ is (weakly) conformal to $g_{2}$ if there exists a measurable function $\lambda\colon\Sigma\rightarrow\mathbb{R}$ such that $g_{1}=e^{\lambda}g_{2}$.
+Let $g_{1},g_{2}$ be two metrics, we say that $g_{1}$ is (weakly) conformal to $g_{2}$ if there exists a measurable function $\lambda\colon\Sigma\rightarrow\mathbb{R}$ such that $g_{1}=e^{\lambda}g_{2}$.
 Our main goal in this subsection is to prove the following result:
@@
   1. (i)
-    We firstly assume $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(D^{2},\mathbb{R}^{m})$ and $\varphi_{1},\varphi_{2}\in W^{1,2}(D^{2},\mathbb{R}^{2})$ be isothermal charts as in Definition IV.2.
+    We first assume $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(D^{2},\mathbb{R}^{m})$ and $\varphi_{1},\varphi_{2}\in W^{1,2}(D^{2},\mathbb{R}^{2})$ be isothermal charts as in Definition IV.2.
 For $k\in\{1,2\}$, since $\varphi_{k}$ is a $W^{1,2}_{\textup{loc}}$ bi-Sobolev homeomorphism, by standard degree theory we have $\varphi_{k}$ is differentiable a.e.
 and either $\det(\nabla\varphi_{k})>0$ a.e.
-or $\det(\nabla\varphi_{k})<0$ a.e., see [82, Theorem 1.7 & Lemma A.28].
+or $\det(\nabla\varphi_{k})<0$ a.e., see [83, Theorem 1.7 & Lemma A.28].
 Without loss of generality we assume $\det(\nabla\varphi_{k})>0$ a.e.
-Let $g_{ij}\coloneqq\partial_{i}\vec{\Phi}\cdot\partial_{j}\vec{\Phi}$, then following the computation in [90, Section 1.5.1], since $\varphi_{k}$ is an isothermal chart, we have
+Let $g_{ij}\coloneqq\partial_{i}\vec{\Phi}\cdot\partial_{j}\vec{\Phi}$, then following the computation in [91, Section 1.5.1], since $\varphi_{k}$ is an isothermal chart, we have
     $\frac{\partial_{\bar{z}}\varphi_{k}}{\partial_{z}\varphi_{k}}=\frac{g_{11}-g_{22}+2ig_{12}}{g_{11}+g_{22}+2\sqrt{g_{11}g_{22}-g_{12}^{2}}}\qquad\text{a.e.}$ (IV.2)
-    The condition $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(D^{2},\mathbb{R}^{m})$ implies the right hand side has an $L^{\infty}$ norm strictly less than $1$.
-In particular, $\varphi_{k}$ is quasiconformal, and $\varphi_{1}\circ(\varphi_{2})^{-1}$ is holomorphic on $\varphi_{2}(D^{2})$, see Definition III.6 and [152, Corollary 1.2.8].
+    The condition $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(D^{2},\mathbb{R}^{m})$ implies the right-hand side has an $L^{\infty}$ norm strictly less than $1$.
+In particular, $\varphi_{k}$ is quasiconformal, and $\varphi_{1}\circ(\varphi_{2})^{-1}$ is holomorphic on $\varphi_{2}(D^{2})$, see Definition III.6 and [155, Corollary 1.2.8].
 Now return to the case when $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$.
@@
     When $\Sigma$ is orientable, there exists a metric $h$ of constant Gaussian curvature $1$, 1-1 or 0 and conformal under $\mathcal{V}$ by the uniformization theorem for compact Riemann surfaces (see e.g.
-[93, Section 4.4]).
+[94, Section 4.4]).
 When $\Sigma$ is non-orientable, $h$ is constructed by using a $2$-cover from a closed orientable smooth surface to $\Sigma$.
-    The fact $\alpha\in W^{2,1}(\Sigma)$ follows from Corollary III.11. As for the embedding $W^{2,1}(\Sigma)\hookrightarrow C^{0}(\Sigma)$, see for instance [80, Theorem 3.3.4 & 3.3.10].
+    The fact $\alpha\in W^{2,1}(\Sigma)$ follows from Corollary III.11. Concerning the embedding $W^{2,1}(\Sigma)\hookrightarrow C^{0}(\Sigma)$ we refer to [81, Theorem 3.3.4 & 3.3.10].
 Remark IV.6.
@@
 One major difference is about the regularity of the transition map as in Corollary IV.5 (ii): it may not be Lipschitz, but will be in $W^{1,p}$ for some $p>2$ instead.
-Using the conformal structure associated to $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$, the relation (I.6) follows from standard computations, see for instance [168, Theorem 1.6].
+Using the conformal structure associated to $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$, the relation (I.6) follows from standard computations, see for instance [171, Theorem 1.6].
 By integrating (I.6) with respect to $d\textup{vol}_{h}$ (when $\Sigma$ is orientable), we have (I.7) by the Gauss–Bonnet theorem for the smooth metric $h$.
@@
 Here we provide a different proof for the $2$-dimensional case.
-Instead of using properties of BLD maps [130, Lemma 4.6 & Theorem 4.7], we use regularity results for Beltrami equations [42].
+Instead of using properties of maps with bounded length distortion (BLD maps) [134, Lemma 4.6 & Theorem 4.7], we use regularity results for Beltrami equations [42].
 Thanks to the embedding $W^{1,2}(D^{2})\hookrightarrow VMO(D^{2})$, we obtain another proof in Corollary IV.13.
-Lemma IV.7 ([78, Theorem 1.1]).
+Lemma IV.7 ([79, Theorem 1.1]).
 Let $\phi\in W^{1,\infty}\cap W^{2,2}(D^{2},\mathbb{R}^{2})$.
@@
 Let $h\coloneqq\phi\circ f^{-1}$.
-By the same computation as in [152, Corollary 1.2.8], the map $h$ is holomorphic on $f(D_{1/2}(0))$.
+By the same computation as in [155, Corollary 1.2.8], the map $h$ is holomorphic on $f(D_{1/2}(0))$.
 Let $k\geq 1$ be the order of zero of $h$ at 0.
-Then there exists a constant $A\neq 0\in\mathbb{C}$, $0\in U_{1}\subset f(D_{1/2}(0))$, a map $g\in C^{\infty}(\overline{U_{1}},\mathbb{C})$ conformal from the bounded open set $U_{1}$ to $g(U_{1})\subset\mathbb{C}$ with $g(0)=0$ and $g^{\prime}(0)=1$ such that $h=A\,g^{k}$ on $U_{1}$.
+Then there exists a constant $A\neq 0\in\mathbb{C}$, a bounded open set $U_{1}\subset f(D_{1/2}(0))$ containing 0 and a map $g\in C^{\infty}(\overline{U_{1}},\mathbb{C})$ conformal from $U_{1}$ to $g(U_{1})\subset\mathbb{C}$ with $g(0)=0$ and $g^{\prime}(0)=1$ such that $h=A\,g^{k}$ on $U_{1}$.
 Let $f_{0}\coloneqq g\circ f$, then since $h=\phi\circ f^{-1}$, we have $\phi=h\circ f=A(g\circ f)^{k}=A\,f_{0}^{k}$ on $f^{-1}(U_{1})$.
 Since $f$ is quasiconformal and satisfies the equation $\partial_{\bar{z}}f=\mu\,\partial_{z}f$ with $\mu\in W^{1,2}(\mathbb{C})$ of compact support, by Lemma III.8 we have $f\in W^{2,q}_{\textup{loc}}(\mathbb{C})$ for any $q<2$.
-Then by Sobolev embeddings we have $f\in C^{0,\alpha}_{\textup{loc}}(\mathbb{C})$ for any $\alpha<1$.
-In particular, $f_{0}=g\circ f\in C^{0,\alpha}(f^{-1}(\overline{U_{1}}))$, and it holds that
-$|f_{0}(x)|\leq C(\alpha)|x|^{\alpha},\qquad\text{for all }x\in f^{-1}(\overline{U_{1}})\,\text{ and all }\alpha<1.$ (IV.3)
-Now we prove $k=1$. Suppose $k\geq 2$, then there exists a positive constant $C_{1}$ such that
+Fix $\alpha\in(0,1)$.
+By the Sobolev embedding we have $f\in C^{0,\alpha}_{\textup{loc}}(\mathbb{C})$.
+In particular, this implies that $f_{0}=g\circ f\in C^{0,\alpha}(f^{-1}(\overline{U_{1}}))$, and there exists a constant $C_{0}>0$ such that for all $x\in f^{-1}\big(\overline{U_{1}}\big)$, it holds that
+$|f_{0}(x)|\leq C_{0}|x|^{\alpha}.$ (IV.3)
+Now we prove that $k=1$. By contradiction, suppose that $k\geq 2$. Then there exists a positive constant $C_{1}$ such that
 $\sqrt{c}\leq\sqrt{\det(D\phi)}\leq|\partial_{z}\phi|=|kAf_{0}^{k-1}|\,|\partial_{z}f_{0}|\leq C_{1}|f_{0}|\,|\partial_{z}f_{0}|\quad\text{ a.e. on }f^{-1}(\overline{U_{1}}).$ (IV.4)
-Fix $\alpha>\frac{1}{2}$. By (IV.3) and (IV.4), there exists a positive constant $C_{2}$ such that for any $x\in f^{-1}(\overline{U_{1}})$, it holds
+By (IV.3) and (IV.4), there exists a positive constant $C_{2}$ such that for all $x\in f^{-1}(\overline{U_{1}})$, it holds that
 $|\partial_{z}f_{0}(x)|\geq C_{2}|x|^{-\alpha}.$ (IV.5)
-Since $4\alpha>2$, (IV.5) implies
-$\int_{f^{-1}(U_{1})}|\partial_{z}f_{0}|^{4}=\infty.$
-But by Sobolev embedding again, we have $f\in W^{1,p}_{\textup{loc}}(\mathbb{C})$ for any $p<\infty$, hence we obtain $f_{0}=g\circ f\in W^{1,4}(f^{-1}(U_{1}))$, a contradiction arises.
-Let $0\in U\Subset f^{-1}(U_{1})$.
-Since $k=1$, we have $\phi=A(g\circ f)$ is quasiconformal on $U$, and $\phi^{-1}\in W^{1,\infty}\cap W^{2,2}(\phi(U),\mathbb{R}^{2})$ since $D\phi^{-1}=(D\phi)^{-1}\circ\phi^{-1}$ a.e.
+Choose $p\in(1,\infty)$ such that $\alpha p\geq 2$. Then it follows from (IV.5) that
+$\int_{f^{-1}(U_{1})}|\partial_{z}f_{0}|^{p}=\infty.$ (IV.6)
+Since $f\in W^{2,q}_{\textup{loc}}(\mathbb{C})$ for any $q<2$, we obtain by Sobolev embedding that $f\in W^{1,p}_{\textup{loc}}(\mathbb{C})$.
+Hence we have $f_{0}=g\circ f\in W^{1,p}(f^{-1}(U_{1}))$, which contradicts (IV.6).
+Consider an open set $U$ such that $0\in U\Subset f^{-1}(U_{1})$.
+Since $k=1$, we have that $\phi=A(g\circ f)$.
+Hence, $\phi$ is quasiconformal on $U$, and $\phi^{-1}\in W^{1,\infty}\cap W^{2,2}(\phi(U),\mathbb{R}^{2})$ since $D\phi^{-1}=(D\phi)^{-1}\circ\phi^{-1}$ a.e.
 on $\phi(U)$.
 ∎
-We are now ready to prove Theorem IV.4
+We are now ready to prove Theorem IV.4.
 Proof of Theorem IV.4.
 We apply the Gram–Schmidt process to the frame $(\partial_{1}\vec{\Phi},\partial_{2}\vec{\Phi})$ and get
-$\vec{f}_{1}\coloneqq\frac{\partial_{1}\vec{\Phi}}{|\partial_{1}\vec{\Phi}|},\qquad\text{ and }\qquad\vec{f}_{2}\coloneqq\frac{\partial_{2}\vec{\Phi}-(\partial_{2}\vec{\Phi}\cdot\vec{f}_{1})\vec{f}_{1}}{|\partial_{2}\vec{\Phi}-(\partial_{2}\vec{\Phi}\cdot\vec{f}_{1})\vec{f}_{1}|}.$ (IV.6)
+$\vec{f}_{1}\coloneqq\frac{\partial_{1}\vec{\Phi}}{|\partial_{1}\vec{\Phi}|},\qquad\text{ and }\qquad\vec{f}_{2}\coloneqq\frac{\partial_{2}\vec{\Phi}-(\partial_{2}\vec{\Phi}\cdot\vec{f}_{1})\vec{f}_{1}}{|\partial_{2}\vec{\Phi}-(\partial_{2}\vec{\Phi}\cdot\vec{f}_{1})\vec{f}_{1}|}.$ (IV.7)
 By the definition of weak immersions, there exists a constant $\Lambda>0$ such that for a.e. $p\in D^{2}$ and any $X\in T_{p}(D^{2})$,
-$\Lambda^{-1}|X|^{2}\leq|d\vec{\Phi}_{p}(X)|^{2}\leq\Lambda|X|^{2}.$ (IV.7)
+$\Lambda^{-1}|X|^{2}\leq|d\vec{\Phi}_{p}(X)|^{2}\leq\Lambda|X|^{2}.$ (IV.8)
 In particular, we have $|\partial_{1}\vec{\Phi}|,|\partial_{2}\vec{\Phi}|\in[\Lambda^{-1},\Lambda]$ and $\partial_{2}\vec{\Phi}-(\partial_{2}\vec{\Phi}\cdot\vec{f}_{1})\vec{f}_{1}\in[\Lambda^{-1},\Lambda(1+\Lambda^{2})]$.
@@
 For $\theta\in W^{1,2}(D^{2},{\mathbb{R}})$, we denote $(\vec{f}^{\,\theta}_{1},\vec{f}^{\,\theta}_{2})$ the rotation by the angle $\theta$ of the original frame $(\vec{f}_{1},\vec{f}_{2})$:
-$\vec{f}^{\,\theta}_{1}+i\vec{f}^{\,\theta}_{2}=e^{i\theta}(\vec{f}_{1}+i\vec{f}_{2}),$
-then we still have $(\vec{f}^{\,\theta}_{1},\vec{f}^{\,\theta}_{2})\in L^{\infty}\cap W^{1,2}(D^{2},V_{2}(\mathbb{R}^{m}))$.
+$\vec{f}^{\,\theta}_{1}+i\vec{f}^{\,\theta}_{2}=e^{i\theta}(\vec{f}_{1}+i\vec{f}_{2}).$
+We have $(\vec{f}^{\,\theta}_{1},\vec{f}^{\,\theta}_{2})\in L^{\infty}\cap W^{1,2}(D^{2},V_{2}(\mathbb{R}^{m}))$.
 We denote $g=g_{\vec{\Phi}}$.
 We look now for a rotation of this orthonormal frame realizing the following absolute minimum (for notation, see xvii)
-$\begin{array}[]{l}\displaystyle\inf_{\theta\in W^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}|\langle\vec{f}^{\,\theta}_{1},d\vec{f}^{\,\theta}_{2}\rangle|^{2}_{g}\ d\textup{vol}_{g}\\[14.22636pt] \displaystyle=\inf_{\theta\in W^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}\sum_{i,j=1}^{2}g^{ij}\,\langle\vec{f}^{\,\theta}_{1},\partial_{i}\vec{f}^{\,\theta}_{2}\rangle\,\langle\vec{f}^{\,\theta}_{1},\partial_{j}\vec{f}^{\,\theta}_{2}\rangle\,\sqrt{\det g}\ dx^{1}\wedge dx^{2}.\end{array}$ (IV.8)
+$\begin{array}[]{l}\displaystyle\inf_{\theta\in W^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}|\langle\vec{f}^{\,\theta}_{1},d\vec{f}^{\,\theta}_{2}\rangle|^{2}_{g}\ d\textup{vol}_{g}\\[14.22636pt] \displaystyle=\inf_{\theta\in W^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}\sum_{i,j=1}^{2}g^{ij}\,\langle\vec{f}^{\,\theta}_{1},\partial_{i}\vec{f}^{\,\theta}_{2}\rangle\,\langle\vec{f}^{\,\theta}_{1},\partial_{j}\vec{f}^{\,\theta}_{2}\rangle\,\sqrt{\det g}\ dx^{1}\wedge dx^{2}.\end{array}$ (IV.9)
 For $k\in\{1,2\}$, we have
-$\langle\vec{f}^{\,\theta}_{1},\partial_{k}\vec{f}^{\,\theta}_{2}\rangle=\partial_{k}\theta+\langle\vec{f}_{1},\partial_{k}\vec{f}_{2}\rangle.$ (IV.9)
+$\langle\vec{f}^{\,\theta}_{1},\partial_{k}\vec{f}^{\,\theta}_{2}\rangle=\partial_{k}\theta+\langle\vec{f}_{1},\partial_{k}\vec{f}_{2}\rangle.$ (IV.10)
 Hence the following energy is strictly convex in $W^{1,2}(D^{2},{\mathbb{R}})$:
@@
 It satisfies the Euler–Lagrange equation (see Notation ix for the Hodge star operator $_{g}$):
-$\int_{D^{2}}d\phi\wedge*_{g}\left(d\theta+\langle\vec{f}_{1},d\vec{f}_{2}\rangle\right)=0,\qquad\text{for all }\,\phi\in W^{1,2}(D^{2},{\mathbb{R}}).$ (IV.10)
-Denote $(\vec{e}_{1},\vec{e}_{2})\coloneqq(\vec{f}^{\,\theta}_{1},\vec{f}^{\,\theta}_{2})$ for this value of $\theta$. From (IV.10) we obtain that
+$\forall\phi\in W^{1,2}(D^{2},{\mathbb{R}}),\qquad\int_{D^{2}}d\phi\wedge*_{g}\left(d\theta+\langle\vec{f}_{1},d\vec{f}_{2}\rangle\right)=0.$ (IV.11)
+Denote $(\vec{e}_{1},\vec{e}_{2})\coloneqq(\vec{f}^{\,\theta}_{1},\vec{f}^{\,\theta}_{2})$ for this value of $\theta$. From (IV.11) we obtain that
 $d\Big(*_{g}\langle\vec{e}_{1},d\vec{e}_{2}\rangle\Big)=0\quad\text{in }\mathcal{D}^{\prime}(D^{2}).$
 By the weak Poincaré lemma, there exists $\lambda\in W^{1,2}(D^{2})$ such that
-$d\lambda=*_{g}\langle\vec{e}_{1},d\vec{e}_{2}\rangle.$ (IV.11)
-Substituting this into (IV.10), we have
-$\int_{D^{2}}d\phi\wedge d\lambda=0,\qquad\text{for all }\,\phi\in W^{1,2}(D^{2},{\mathbb{R}}).$
+$d\lambda=*_{g}\langle\vec{e}_{1},d\vec{e}_{2}\rangle.$ (IV.12)
+Substituting this into (IV.11), we have
+$\forall\phi\in W^{1,2}(D^{2},{\mathbb{R}}),\qquad\int_{D^{2}}d\phi\wedge d\lambda=0.$
 Hence for any $\phi\in C^{\infty}(\overline{D^{2}})$, we have by Stokes theorem
 $\int_{\partial D^{2}}\lambda\,d\phi=\int_{D^{2}}d(\lambda\,d\phi)=0.$
-It follows that $\lambda\big|_{\mathbb{S}^{1}}$ is constant by fundamental lemma of the calculus of variations [95, Lemma 1.1.1].
+It follows that $\lambda\big|_{\mathbb{S}^{1}}$ is constant by the fundamental lemma of the calculus of variations [96, Lemma 1.1.1].
 Up to adding a constant we can assume $\lambda\in W^{1,2}_{0}(D^{2})$.
-By applying the operator $_{g}\,d\,*_{g}$ to the equation (IV.11), we obtain
+By applying the operator $_{g}\,d\,*_{g}$ to the equation (IV.12), we obtain
 $\Delta_{g}\lambda=*_{g}\,d*_{g}d\lambda=-*_{g}d\langle\vec{e}_{1},d\vec{e}_{2}\rangle\quad\text{in }D^{2}.$
@@
 $\begin{cases}\displaystyle\,\partial_{i}\left(\sqrt{\det g}\ g^{ij}\,\partial_{j}\lambda\right)=\partial_{2}\vec{e}_{1}\cdot\partial_{1}\vec{e}_{2}-\partial_{1}\vec{e}_{1}\cdot\partial_{2}\vec{e}_{2}\quad&\mbox{ in }D^{2},\\[8.53581pt] \displaystyle\lambda=0\quad&\mbox{ on }\partial D^{2}.\end{cases}$
-We deduce from Lemma III.3 that $\lambda\in C^{0}(\overline{D^{2}})\cap W^{1,2}_{0}(D^{2},{\mathbb{R}})$. For $i\in\{1,2\}$, we define pullback of $\vec{e}_{i}$ by $\vec{\Phi}$ and the dual $1$-forms
-$\begin{dcases}e_{i}\coloneqq g^{jk}\,\langle\vec{e}_{i},\partial_{j}\vec{\Phi}\rangle\,\dfrac{\partial}{\partial x^{k}}\in L^{\infty}\cap W^{1,2}(D^{2},TD^{2}),\\[5.69054pt] e_{i}^{\ast}\coloneqq\langle\vec{e}_{i},\partial_{j}\vec{\Phi}\rangle\,dx^{j}=\langle\vec{e}_{i},d\vec{\Phi}\rangle\in L^{\infty}\cap W^{1,2}(D^{2},T^{*}D^{2}).\end{dcases}$ (IV.12)
+We deduce from Lemma III.3 that $\lambda\in C^{0}(\overline{D^{2}})\cap W^{1,2}_{0}(D^{2},{\mathbb{R}})$.
+For $i\in\{1,2\}$, we define the pullback of $\vec{e}_{i}$ by $\vec{\Phi}$ and the dual $1$-forms
+$\begin{dcases}e_{i}\coloneqq g^{jk}\,\langle\vec{e}_{i},\partial_{j}\vec{\Phi}\rangle\,\dfrac{\partial}{\partial x^{k}}\in L^{\infty}\cap W^{1,2}(D^{2},TD^{2}),\\[5.69054pt] e_{i}^{\ast}\coloneqq\langle\vec{e}_{i},\partial_{j}\vec{\Phi}\rangle\,dx^{j}=\langle\vec{e}_{i},d\vec{\Phi}\rangle\in L^{\infty}\cap W^{1,2}(D^{2},T^{*}D^{2}).\end{dcases}$ (IV.13)
 It holds that $e_{i}^{*}(e_{j})=\delta_{ij}$ and $d\vec{\Phi}(e_{i})=\vec{e}_{i}$ for any $i,j\in\{1,2\}$.
-Since both $e_{i}^{*}$ and $e_{i}$ are in $L^{\infty}\cap W^{1,2}(D^{2})$, by a weak version of Cartan’s magic formula, we have as in the smooth case [168, Equation (1.70)]:
-$de_{i}^{*}=-e_{i}^{*}([e_{1},e_{2}])\,e_{1}^{*}\wedge e_{2}^{*}=d\lambda\wedge e_{i}^{\ast}.$ (IV.13)
+Since both $e_{i}^{*}$ and $e_{i}$ are in $L^{\infty}\cap W^{1,2}(D^{2})$, by a weak version of Cartan’s magic formula, we have as in the smooth case [171, Equation (1.70)]:
+$de_{i}^{*}=-e_{i}^{*}([e_{1},e_{2}])\,e_{1}^{*}\wedge e_{2}^{*}=d\lambda\wedge e_{i}^{\ast}.$ (IV.14)
 By direct computation we also have
@@
 Since $\lambda\in L^{\infty}(D^{2})\cap W^{1,2}(D^{2},{\mathbb{R}})$, we deduce from the chain rule that $e^{-\lambda}\in W^{1,2}(D^{2})$ and $d(e^{-\lambda})=-e^{-\lambda}\,d\lambda\in L^{2}(D^{2},T^{*}D^{2})$.
-Hence the product rule and (IV.13) lead to
+Hence the product rule and (IV.14) lead to
 $d(e^{-\lambda}\,e_{i}^{*})=0.$
 By weak Poincaré lemma, we obtain the existence of a map $\varphi^{i}\in W^{1,\infty}\cap W^{2,2}(D^{2},{\mathbb{R}})$ satisfying
-$d\varphi^{i}=e^{-\lambda}\,e_{i}^{*}.$ (IV.14)
-Let $\varphi\coloneqq(\varphi^{1},\varphi^{2})\in W^{1,\infty}\cap W^{2,2}(D^{2},{\mathbb{R}^{2}})$. By (IV.12) and (IV.14), we have
+$d\varphi^{i}=e^{-\lambda}\,e_{i}^{*}.$ (IV.15)
+Let $\varphi\coloneqq(\varphi^{1},\varphi^{2})\in W^{1,\infty}\cap W^{2,2}(D^{2},{\mathbb{R}^{2}})$. By (IV.13) and (IV.15), we have
 $\det(\nabla\varphi)=e^{-2\lambda}\det\big(\langle\vec{e}_{i},\partial_{j}\vec{\Phi}\rangle\big)=e^{-2\lambda}\sqrt{\det(\partial_{i}\vec{\Phi}\cdot\partial_{j}\vec{\Phi})}.$
-Since $(e_{1},e_{2})$ is positive and using (IV.7), we obtain that all the eigenvalues of $(\partial_{i}\vec{\Phi}\cdot\partial_{j}\vec{\Phi})_{1\leq i,j\leq 2}$ lie in $\Lambda^{-1},\Lambda]$ and thus, it holds that
+Since $(e_{1},e_{2})$ is positive and using (IV.8), we obtain that all the eigenvalues of $(\partial_{i}\vec{\Phi}\cdot\partial_{j}\vec{\Phi})_{1\leq i,j\leq 2}$ lie in $\Lambda^{-1},\Lambda]$ and thus, it holds that
 $\det(\nabla\varphi)\geq\Lambda^{-1}\,e^{-2\|\lambda\|_{L^{\infty}}}>0.$
@@
 $\partial_{i}(\vec{\Phi}\circ\varphi^{-1})\cdot\partial_{j}(\vec{\Phi}\circ\varphi^{-1})=d\vec{\Phi}\Big(\frac{\partial}{\partial\varphi^{i}}\Big)\cdot d\vec{\Phi}\Big(\frac{\partial}{\partial\varphi^{j}}\Big)=g\Big(\frac{\partial}{\partial\varphi^{i}},\frac{\partial}{\partial\varphi^{j}}\Big)=e^{2\lambda}\delta_{ij}.$
-IV.2 Constant of Wente inequality for a metric and consequence
+IV.2 Constant of Wente’s inequality for a metric and consequences
 In 1971, Wente [191] discovered that solutions to elliptic equations of the following form are more regular than expected:
@@
 $\|u\|_{L^{\infty}(D^{2})}+\|\nabla u\|_{L^{2}(D^{2})}\leq C\|\nabla a\|_{L^{2}(D^{2})}\|\nabla b\|_{L^{2}(D^{2})}.$
-Since then, the study of this structure has been intensively studied.
+Since then, equations that possess this Jacobian structure have been intensively studied.
 Bethuel–Ghidaglia [20] proved that the constant $C$ is independent of the domain.
-Ge [63], Braraket [11] and Topping [184] studied such equations on surfaces and proved that the constant are independent of the underlying surface.
+Ge [63], Braraket [11] and Topping [187] studied such equations on surfaces and proved that the constant is independent of the underlying surface.
 A weighted Wente inequality has also been studied by Da Lio, Gianocca and Rivière in [46, 65].
@@
 Proposition IV.8.
-Let $\Omega\subset\mathbb{R}^{2}$ be a smooth simply connected open set. Let $g$ be a smooth metric on $\Omega$. Let $a,b\in C^{\infty}(\Omega)$ and $u\in C^{\infty}(\Omega)$ be the solution to
-$\left\{\begin{aligned} &-\Delta_{g}u=*_{g}(da\wedge db)\quad&&\text{in }\Omega,\\ &u=0\quad&&\text{on }\partial\Omega.\end{aligned}\right$. (IV.15)
+Let $\Omega\subset\mathbb{R}^{2}$ be a smooth simply connected open set. Let $g$ be a smooth metric on $\Omega$. Let $a,b\in W^{1,2}(\Omega)$ and $u\in W^{1,1}(\Omega)$ be the solution to
+$\left\{\begin{aligned} &-\Delta_{g}u=*_{g}(da\wedge db)\quad&&\text{in }\Omega,\\ &u=0\quad&&\text{on }\partial\Omega.\end{aligned}\right$. (IV.16)
 Then, we have the estimate
@@
 Remark IV.9.
-  • The first equation of (IV.15) can also be written as
+  • The first equation of (IV.16) can also be written as
     $-\partial_{i}(g^{ij}\sqrt{\det g}\,\partial_{j}u)=\nabla a\cdot\nabla^{\perp}b.$
@@
 Hence, we can pass to the limit in Proposition IV.8.
-  • By density, we can assume $a,b\in W^{1,2}$.
 Proof.
@@
 By elliptic regularity, it holds $G_{p}\in C^{\infty}(\Omega\setminus\{p\})$.
-Thanks to Sard’s theorem (see for instance [108, Chapter 6, Theorem 6.10]), for almost every $\alpha>0$, the set $G_{p}^{-1}(\{\alpha\})$ is a regular curve in $\Omega$.
+Thanks to Sard’s theorem (see for instance [111, Chapter 6, Theorem 6.10]), for almost every $\alpha>0$, the set $G_{p}^{-1}(\{\alpha\})$ is a regular curve in $\Omega$.
 We orient this curve with the vector field
@@
 Using the coarea formula, see for instance [33, Chapter IV, Theorem 1], we obtain for almost every $0\leq\alpha<\beta$:
-$\int_{\omega_{p}(\alpha,\beta)}|dG_{p}|^{2}_{g}\,d\textup{vol}_{g}=\int_{\alpha}^{\beta}\int_{G_{p}^{-1}(\{s\})}|dG_{p}|_{g}\,d\textup{vol}_{g}\,ds=\beta-\alpha.$ (IV.16)
-Now, the solution $u$ to (IV.15) is given by
-$u(p)=\int_{\Omega}G_{p}(x)\,da\wedge db.$ (IV.17)
+$\int_{\omega_{p}(\alpha,\beta)}|dG_{p}|^{2}_{g}\,d\textup{vol}_{g}=\int_{\alpha}^{\beta}\int_{G_{p}^{-1}(\{s\})}|dG_{p}|_{g}\,d\textup{vol}_{g}\,ds=\beta-\alpha.$ (IV.17)
+Now, the solution $u$ to (IV.16) is given by
+$u(p)=\int_{\Omega}G_{p}(x)\,da\wedge db.$ (IV.18)
 For each $n\in\mathbb{N}$, there exists $\alpha_{n}\in[n,n+1]$ such that
 $\int_{G_{p}^{-1}(\{\alpha_{n}\})}|da|_{g}+|db|_{g}\,d\textup{vol}_{g}$ $\leq 2\int_{n}^{n+1}\int_{G_{p}^{-1}(s)}(|da|_{g}+|db|_{g})\,|dG_{p}|_{g}\,d\textup{vol}_{g}$
-$\leq 2\int_{\omega_{p}(n,n+1)}(|da|_{g}+|db|_{g})\,|dG_{p}|_{g}\,d\textup{vol}_{g}.$ (IV.18)
-From (IV.17), it holds
+$\leq 2\int_{\omega_{p}(n,n+1)}(|da|_{g}+|db|_{g})\,|dG_{p}|_{g}\,d\textup{vol}_{g}.$ (IV.19)
+From (IV.18), it holds
 $u(p)$ $=\sum_{n\in\mathbb{N}}\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}G_{p}(x)\,*_{g}(da\wedge db)\,d\textup{vol}_{g}$
-$=\sum_{n\in\mathbb{N}}\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}[G_{p}(x)-n]\,*_{g}(da\wedge db)\,d\textup{vol}_{g}$ (IV.19)
-$\qquad+\sum_{n\in\mathbb{N}}n\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}*_{g}(da\wedge db)\,d\textup{vol}_{g}.$ (IV.20)
-We estimate the term (IV.19) by Cauchy–Schwarz:
+$=\sum_{n\in\mathbb{N}}\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}[G_{p}(x)-n]\,*_{g}(da\wedge db)\,d\textup{vol}_{g}$ (IV.20)
+$\qquad+\sum_{n\in\mathbb{N}}n\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}*_{g}(da\wedge db)\,d\textup{vol}_{g}.$ (IV.21)
+We estimate the term (IV.20) by Cauchy–Schwarz:
 $\left|\sum_{n\in\mathbb{N}}\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}[G_{p}(x)-n]\,*_{g}(da\wedge db)\,d\textup{vol}_{g}\right|$ $\leq 2\sum_{n\in\mathbb{N}}\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}|da|_{g}\,|db|_{g}\,d\textup{vol}_{g}$
-$\leq 2\int_{\Omega}|da|_{g}\,|db|_{g}\,d\textup{vol}_{g}.$ (IV.21)
-We estimate the term (IV.20) by integration by parts:
+$\leq 2\int_{\Omega}|da|_{g}\,|db|_{g}\,d\textup{vol}_{g}.$ (IV.22)
+We estimate the term (IV.21) by integration by parts:
 $\sum_{n\in\mathbb{N}}n\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}*_{g}(da\wedge db)\,d\textup{vol}_{g}$
 $=$ $\ \sum_{n\in\mathbb{N}}n\left(\int_{G_{p}^{-1}(\alpha_{n+1})}(\partial_{\tau}b)a\,d\textup{vol}_{g}-\int_{G_{p}^{-1}(\alpha_{n})}(\partial_{\tau}b)a\,d\textup{vol}_{g}\right)$
-$=$ $-\sum_{n\in\mathbb{N}^{*}}\int_{G_{p}^{-1}(\alpha_{n})}(\partial_{\tau}b)a\,d\textup{vol}_{g}.$ (IV.22)
+$=$ $-\sum_{n\in\mathbb{N}^{*}}\int_{G_{p}^{-1}(\alpha_{n})}(\partial_{\tau}b)a\,d\textup{vol}_{g}.$ (IV.23)
 For each $n\in\mathbb{N}^{*}$, we decompose $G_{p}^{-1}(\{\alpha_{n}\})=\bigcup_{i\in I_{n}}\gamma^{i}_{n}$, where every $\gamma^{i}_{n}$ is a closed embedded curve in $\Omega$. Then, it holds
@@
 $\leq\left(\int_{G_{p}^{-1}(\{\alpha_{n}\})}|db|_{g}\,d\textup{vol}_{g}\right)\left(\int_{G_{p}^{-1}(\{\alpha_{n}\})}|da|_{g}\,d\textup{vol}_{g}\right).$
-Coming back to (IV.22), we obtain
+Coming back to (IV.23), we obtain
 $\left|\sum_{n\in\mathbb{N}}n\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}*_{g}(da\wedge db)\,d\textup{vol}_{g}\right|\leq\sum_{n\in\mathbb{N}^{*}}\left(\int_{G_{p}^{-1}(\{\alpha_{n}\})}|db|_{g}\,d\textup{vol}_{g}\right)\left(\int_{G_{p}^{-1}(\{\alpha_{n}\})}|da|_{g}\,d\textup{vol}_{g}\right).$
-Using the choice of the $\alpha_{n}$ in (IV.18), we obtain
+Using the choice of the $\alpha_{n}$ in (IV.19), we obtain
 $\left|\sum_{n\in\mathbb{N}}n\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}*_{g}(da\wedge db)\,d\textup{vol}_{g}\right|$
@@
 $\leq$ $\ 4\sum_{n\in\mathbb{N}^{*}}\|dG_{p}\|_{L^{2}(\omega_{p}(n,n+1),g)}^{2}\,\left(\|da\|_{L^{2}(\omega_{p}(n,n+1),g)}+\|db\|_{L^{2}(\omega_{p}(n,n+1),g)}\right)^{2}.$
-Thanks to (IV.16), we obtain
+Thanks to (IV.17), we obtain
 $\left|\sum_{n\in\mathbb{N}}n\int_{\omega_{p}(\alpha_{n},\alpha_{n+1})}*_{g}(da\wedge db)\,d\textup{vol}_{g}\right|$
 $\leq 4\sum_{n\in\mathbb{N}^{*}}\left(\|da\|_{L^{2}(\omega_{p}(n,n+1),g)}+\|db\|_{L^{2}(\omega_{p}(n,n+1),g)}\right)^{2}$
-$\leq 8\,\|da\|_{L^{2}(\Omega,g)}^{2}+8\|db\|_{L^{2}(\Omega,g)}^{2}=16.$ (IV.23)
-We used Cauchy–Schwarz and the equalities $\|da\|_{L^{2}(\Omega,g)}=\|db\|_{L^{2}(\Omega,g)}=1$ for the last estimate. Coming back to (IV.17), together with (IV.21) and (IV.23), we deduce that
+$\leq 8\,\|da\|_{L^{2}(\Omega,g)}^{2}+8\|db\|_{L^{2}(\Omega,g)}^{2}=16.$ (IV.24)
+We used Cauchy–Schwarz and the equalities $\|da\|_{L^{2}(\Omega,g)}=\|db\|_{L^{2}(\Omega,g)}=1$ for the last estimate. Coming back to (IV.18), together with (IV.22) and (IV.24), we deduce that
 $\|u\|_{L^{\infty}(\Omega)}\leq 18.$
-We multiply (IV.15) by $u$ and integrate by parts:
-$\int_{\Omega}|du|^{2}_{g}\,d\textup{vol}_{g}$ $=-\int_{\Omega}u\,\Delta_{g}u\,d\textup{vol}_{g}$
-$\leq\|u\|_{L^{\infty}(\Omega)}\int_{\Omega}|da|_{g}\,|db|_{g}\,d\textup{vol}_{g}\leq 18.$
+We multiply (IV.16) by $u$ and integrate by parts:
+$\int_{\Omega}|du|^{2}_{g}\,d\textup{vol}_{g}$ $=-\int_{\Omega}u\,\Delta_{g}u\,d\textup{vol}_{g}\leq\|u\|_{L^{\infty}(\Omega)}\int_{\Omega}|da|_{g}\,|db|_{g}\,d\textup{vol}_{g}\leq 18.$
 IV.2.2 Estimate of the moving frame
@@
 Proof.
-Thanks to [80, Lemma 4.1.3], there exists a frame $(\vec{e}_{1},\vec{e}_{2})\in W^{1,2}(\Omega,g)$ such that
-$\begin{cases}d^{*_{g}}\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}=0&\text{in }\Omega,\\ \left\langle\partial_{\nu}\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}=0&\text{on }\partial\Omega,\end{cases}$ (IV.24)
+Thanks to [81, Lemma 4.1.3], there exists a frame $(\vec{e}_{1},\vec{e}_{2})\in W^{1,2}(\Omega,g)$ such that
+$\begin{cases}d^{*_{g}}\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}=0&\text{in }\Omega,\\ \left\langle\partial_{\nu}\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}=0&\text{on }\partial\Omega,\end{cases}$ (IV.25)
 where $\nu$ is the unit outward-pointing normal for $g$.
-Moreover, by decomposition each $d\vec{e}_{i}$ in the sum of its tangential part along $\vec{\Phi}$ and its normal part, we obtain the following estimate
-$\int_{\Omega}|d\vec{e}_{1}|^{2}_{g,\mathbb{R}^{m}}+|d\vec{e}_{2}|^{2}_{g,\mathbb{R}^{m}}\,d\textup{vol}_{g}\leq 2\int_{\Omega}|\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}|^{2}_{g}\,d\textup{vol}_{g}+\int_{\Omega}|d\vec{n}|^{2}_{g}\,d\textup{vol}_{g}.$ (IV.25)
-By Poincaré Lemma and (IV.24), there exists $\lambda\in W^{1,2}(\Omega,g)$ such that
-$\begin{cases}d\lambda=*_{g}\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}&\text{in }\Omega,\\ \lambda=0&\text{on }\partial\Omega.\end{cases}$ (IV.26)
-Indeed, the boundary condition in (IV.24) implies that $\lambda$ is constant on $\partial\Omega$.
-Since the first relation of (IV.26) depends only on $d\lambda$ and not on $\lambda$ itself, we can add a constant in order to get the second relation of (IV.26).
-We apply $d^{*_{g}}$ to (IV.26) to obtain
-$\left\{\begin{aligned} &-\Delta_{g}\lambda=-*_{g}d\left(\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}\right)\quad\text{in }\Omega,\\ &\lambda=0\quad\text{on }\partial\Omega.\end{aligned}\right$.
-(IV.27)
+Moreover, by decomposing each $d\vec{e}_{i}$ into the sum of its tangential part along $\vec{\Phi}$ and its normal part, we obtain the following estimate
+$\int_{\Omega}|d\vec{e}_{1}|^{2}_{g,\mathbb{R}^{m}}+|d\vec{e}_{2}|^{2}_{g,\mathbb{R}^{m}}\,d\textup{vol}_{g}\leq 2\int_{\Omega}|\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}|^{2}_{g}\,d\textup{vol}_{g}+\int_{\Omega}|d\vec{n}|^{2}_{g}\,d\textup{vol}_{g}.$ (IV.26)
+By Poincaré Lemma and (IV.25), there exists $\lambda\in W^{1,2}(\Omega,g)$ such that
+$\begin{cases}d\lambda=*_{g}\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}&\text{in }\Omega,\\ \lambda=0&\text{on }\partial\Omega.\end{cases}$ (IV.27)
+Indeed, the boundary condition in (IV.25) implies that $\lambda$ is constant on $\partial\Omega$.
+Since the first relation of (IV.27) depends only on $d\lambda$ and not on $\lambda$ itself, we can add a constant in order to get the second relation of (IV.27).
+We apply $d^{*_{g}}$ to (IV.27) to obtain
+$\begin{cases}-\Delta_{g}\lambda=-*_{g}d\left(\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}\right)&\text{in }\Omega,\\ \lambda=0&\text{on }\partial\Omega.\end{cases}$ (IV.28)
 Thanks to Proposition IV.8, it holds
 $\begin{cases}\|\lambda\|_{L^{\infty}(\Omega)}\leq 18\,\|d\vec{e}_{1}\|_{L^{2}(\Omega,g)}\,\|d\vec{e}_{2}\|_{L^{2}(\Omega,g)},\\[5.69054pt] \|d\lambda\|_{L^{2}(\Omega,g)}\leq 3\sqrt{2}\,\|d\vec{e}_{1}\|_{L^{2}(\Omega,g)}\,\|d\vec{e}_{2}\|_{L^{2}(\Omega,g)}.\end{cases}$
-Denote $\varepsilon_{0}\coloneqq\|K_{g}\|_{L^{1}(\Omega,g)}$. We multiply the first relation of (IV.27) by $\lambda$ and integrate by parts:
+Denote $\varepsilon_{0}\coloneqq\|K_{g}\|_{L^{1}(\Omega,g)}$. We multiply the first relation of (IV.28) by $\lambda$ and integrate by parts:
 $\int_{\Omega}|d\lambda|^{2}_{g}\,d\textup{vol}_{g}$ $\leq\|\lambda\|_{L^{\infty}(\Omega)}\int_{\Omega}\left|d\left(\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}\right)\right|_{g}\,d\textup{vol}_{g}$
@@
 $\leq 18\,\varepsilon_{0}\,\|d\vec{e}_{1}\|_{L^{2}(\Omega,g)}\,\|d\vec{e}_{2}\|_{L^{2}(\Omega,g)}.$
-Thanks to (IV.26), we obtain
+Thanks to (IV.27), we obtain
 $\int_{\Omega}|\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}|^{2}_{g}\,d\textup{vol}_{g}$ $\leq 18\,\varepsilon_{0}\,\|d\vec{e}_{1}\|_{L^{2}(\Omega,g)}\|d\vec{e}_{2}\|_{L^{2}(\Omega,g)}\leq 9\,\varepsilon_{0}\left(\|d\vec{e}_{1}\|_{L^{2}(\Omega,g)}^{2}+\|d\vec{e}_{2}\|_{L^{2}(\Omega,g)}^{2}\right).$
-Using (IV.25), we obtain
+Using (IV.26), we obtain
 $\int_{\Omega}|\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}|^{2}_{g}\,d\textup{vol}_{g}\leq 9\,\varepsilon_{0}\left(2\int_{\Omega}|\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}|^{2}_{g}\,d\textup{vol}_{g}+\int_{\Omega}|d\vec{n}|^{2}_{g}\,d\textup{vol}_{g}\right).$
@@
 $\frac{1}{2}\int_{\Omega}|\left\langle d\vec{e}_{1},\vec{e}_{2}\right\rangle_{\mathbb{R}^{m}}|^{2}_{g}\,d\textup{vol}_{g}\leq\frac{1}{4}\int_{\Omega}|d\vec{n}|^{2}_{g}\,d\textup{vol}_{g}.$
-We conclude thanks to (IV.25). ∎
+We conclude thanks to (IV.26). ∎
 IV.3 Local injectivity and integral densities
@@
 }p\in B^{n}\text{ and all }\,v\in T_{p}(B^{n}).\end{dcases}$
-In particular, $\nabla f\in VMO(B^{n})$. Then there exist a constant $C>0$ and an open neighborhood $U$ of the origin, both depending only on $\Lambda,\,\omega,\,n,\,m$ such that
+In particular, $\nabla f\in VMO(B^{n})$. Then there exists a constant $C>0$ and an open neighborhood $U$ of the origin, both depending only on $\Lambda,\,\omega,\,n,\,m$ such that
 $|f(x)-f(y)|\geq C|x-y|,\qquad\text{for all }\,x,y\in U.$
@@
 It follows that there exists a constant $C_{0}>0$ such that for any $v\in\mathbb{R}^{n}$, $M\in K^{\prime}$, it holds
-$|Mv|\geq C_{0}|v|.$ (IV.28)
-Moreover, $C_{0},\,\varepsilon$, and $K$ can be chosen to depend on $\Lambda,\,\omega,\,n,\,m$ only.
+$|Mv|\geq C_{0}|v|.$ (IV.29)
+Moreover, the constants $C_{0},\,\varepsilon$, and $K$ can be chosen to depend on $\Lambda,\,\omega,\,n,\,m$ only.
 We choose $r_{0},\varepsilon_{0}\in(0,1/8)$ small enough (to be determined later) depending on $\Lambda,\,\omega,\,n,\,m$ only such that for every ball $B$ of radius $\leq 4r_{0}$ we have $\fint_{B}\left|\nabla f-\fint_{B}\nabla f\right|<\min(\varepsilon,\varepsilon_{0})$.
-Now fix $x_{0},y_{0}\in B_{r_{0}}(0)$, we denote $r\coloneqq|x_{0}-y_{0}|$, $B^{\prime}\coloneqq B_{2r}(x_{0})$, and
+Now fix $x_{0},y_{0}\in B_{r_{0}}(0)$.
+Let $r\coloneqq|x_{0}-y_{0}|$ and consider the ball $B^{\prime}\coloneqq B_{2r}(x_{0})$.
+We denote
 $M_{0}\coloneqq\fint_{B^{\prime}}\nabla f.$
 Since $2r<4r_{0}$ and $\nabla f\in K$ a.e., we have
-$\fint_{B^{\prime}}|\nabla f-M_{0}|<\min(\varepsilon,\varepsilon_{0}),$
-and hence $M_{0}\in K^{\prime}$.
-Now let $f_{0}(x)\coloneqq f(x)-M_{0}x$, then apply Morrey’s inequality [58, Theorem 5.6.4] in the case $p=2n$ and John–Nirenberg inequality [77, Theorem 3.5], for $\varepsilon_{0}$ small enough we have
-$\begin{aligned} |f_{0}(x_{0})-f_{0}(y_{0})|&\leq C_{1}\mkern 2.0mur^{1/2}\left(\int_{B^{\prime}}|\nabla f_{0}|^{2n}\right)^{\frac{1}{2n}}\\ &=C_{1}\mkern 2.0mur^{1/2}\left(\int_{B^{\prime}}|\nabla f-M_{0}|^{2n}\right)^{\frac{1}{2n}}\\[5.69054pt] &\leq C_{2}\mkern 2.0mur\mkern 2.0mu\varepsilon_{0}=2^{-1}C_{0}\mkern 2.0mur\\[8.53581pt] &=2^{-1}C_{0}|x_{0}-y_{0}|,\end{aligned}$ (IV.29)
-where $C_{1},\,C_{2}$ are positive constants depending only on $n$, and we choose $\varepsilon_{0}=(2C_{2})^{-1}C_{0}$. Therefore, by (IV.28) and (IV.29) we have
+$\fint_{B^{\prime}}|\nabla f-M_{0}|<\min(\varepsilon,\varepsilon_{0}).$
+Hence, it holds $M_{0}\in K^{\prime}$.
+Letting $f_{0}(x)\coloneqq f(x)-M_{0}x$, we apply Morrey’s inequality [58, Theorem 5.6.4] in the case $p=2n$ and John–Nirenberg inequality [78, Theorem 3.5], for $\varepsilon_{0}$ small enough we have
+$\begin{aligned} |f_{0}(x_{0})-f_{0}(y_{0})|&\leq C_{1}\mkern 2.0mur^{1/2}\left(\int_{B^{\prime}}|\nabla f_{0}|^{2n}\right)^{\frac{1}{2n}}\\ &=C_{1}\mkern 2.0mur^{1/2}\left(\int_{B^{\prime}}|\nabla f-M_{0}|^{2n}\right)^{\frac{1}{2n}}\\[5.69054pt] &\leq C_{2}\mkern 2.0mur\mkern 2.0mu\varepsilon_{0}=2^{-1}C_{0}\mkern 2.0mur=2^{-1}C_{0}|x_{0}-y_{0}|,\end{aligned}$ (IV.30)
+where $C_{1},\,C_{2}$ are positive constants depending only on $n$, and we choose $\varepsilon_{0}=(2C_{2})^{-1}C_{0}$. Therefore, by (IV.29) and (IV.30) we have
 $|f(x_{0})-f(y_{0})|\geq|M_{0}(x_{0}-y_{0})|-|f_{0}(x_{0})-f_{0}(y_{0})|\geq\frac{C_{0}}{2}|x_{0}-y_{0}|.$
@@
 Let $m,n\in\mathbb{N}$ with $2\leq n\leq m$, and let $1\leq q<\infty$.
-Suppose $f\in W^{1,\infty}_{\textup{imm}}(B^{n},\mathbb{R}^{m})$ and $\nabla^{2}f\in L^{n,q}(B^{n})$, then there exist a constant $C>0$ and an open neighborhood $U$ of the origin such that
+Suppose $f\in W^{1,\infty}_{\textup{imm}}(B^{n},\mathbb{R}^{m})$ and $\nabla^{2}f\in L^{n,q}(B^{n})$, then there exists a constant $C>0$ and an open neighborhood $U$ of the origin such that
 $|f(x)-f(y)|\geq C|x-y|,\qquad\text{for all }\,x,y\in U.$
@@
 When $m>n$, Corollary IV.13 is sharp in the sense that there exists $f\in W_{\textup{imm}}^{1,\infty}(B^{n},\mathbb{R}^{m})$ with $\nabla^{2}f\in L^{n,\infty}(B^{n})$ but $f$ is not injective in any neighborhood of the origin.
-We construct as follows.
-Without loss of generality assume $m=n+1$. Define $\tilde{f}_{0}\in C^{\infty}(\mathbb{S}^{n-1},\mathbb{R}^{n})$ by
+We construct it as follows.
+Without loss of generality, we can assume $m=n+1$.
+We define $\tilde{f}_{0}\in C^{\infty}(\mathbb{S}^{n-1},\mathbb{R}^{n})$ by
 $\tilde{f}_{0}(x_{1},\dots,x_{n})\coloneqq\Big(\big(2+\cos(\pi x_{n})\big)x_{1},\dots,\big(2+\cos(\pi x_{n})\big)x_{n-1},\mkern 1.0mu\sin(\pi x_{n})\Big).$
@@
 Now for $x\in B^{n}$, we define
-$f(x)=\begin{dcases}|x|\mkern 2.0mu\tilde{f}\Big(\frac{x}{|x|}\Big),&\text{if }x\neq 0,\\ 0,&\text{otherwise}.\end{dcases}$
-Then $f\in W_{\textup{imm}}^{1,\infty}(B^{n},\mathbb{R}^{n+1})$, and there exists a constant $C>0$ such that $|\nabla^{2}f(x)|\leq C|x|^{-1}$ for all $x\in B^{n}\setminus\{0\}$.
+$f(x)=\begin{dcases}|x|\mkern 2.0mu\tilde{f}\Big(\frac{x}{|x|}\Big)&\text{if }x\neq 0,\\ 0&\text{otherwise}.\end{dcases}$
+Then we have $f\in W_{\textup{imm}}^{1,\infty}(B^{n},\mathbb{R}^{n+1})$, and there exists a constant $C>0$ such that $|\nabla^{2}f(x)|\leq C|x|^{-1}$ for all $x\in B^{n}\setminus\{0\}$.
 This implies that $\nabla^{2}f\in L^{n,\infty}(B^{n})$, but $f$ is not injective in any neighborhood of 0.
-In the special case $m=n$, a sharp theorem states that any $W^{2,2}$ weak immersion (also called “BLD maps” in this case) is a local homeomorphism everywhere, see [78, Theorem 1.1].
+In the special case $m=n$, a sharp theorem states that any $W^{2,2}$ weak immersion (also called “BLD maps” in this case) is a local homeomorphism everywhere, see [79, Theorem 1.1].
 However, when $n\geq 3$, it remains open whether one can replace $W^{2,2}$ by $W^{2,(2,q)}$ for any $2<q<\infty$.
-In the proof of [168, Corollary 3.43], it is mentioned that in the case $n=2$, if we assume in addition that $\vec{\Phi}\in W^{2,2}(\Sigma)$, then for any $x\in\mathbb{R}^{m}$, $\theta_{x}$ exists and is in $\mathbb{N}_{0}$.
+In the proof of [171, Corollary 3.43], it is mentioned that in the case $n=2$, if we assume in addition that $\vec{\Phi}\in W^{2,2}(\Sigma)$, then for any $x\in\mathbb{R}^{m}$, $\theta_{x}$ exists and is in $\mathbb{N}_{0}$.
 Here we prove a stronger result.
 Theorem IV.15.
-Let $\Sigma$ be an $n$-dimensional closed smooth manifold, $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$ with $d\vec{\Phi}\in VMO(\Sigma)$, then we have
-$\theta_{y}=\text{Card}(\vec{\Phi}^{-1}(y)),\qquad\text{for all }\,y\in\mathbb{R}^{m}.$ (IV.30)
+Let $\Sigma$ be an $n$-dimensional closed smooth manifold and $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$ be such that $d\vec{\Phi}\in VMO(\Sigma)$. Then we have
+$\theta_{y}=\text{Card}(\vec{\Phi}^{-1}(y)),\qquad\text{for all }\,y\in\mathbb{R}^{m}.$ (IV.31)
 Proof.
@@
 Thus it suffices to show that for a bi-Lipschitz map $f\in W^{1,\infty}_{\textup{imm}}(B^{n},\mathbb{R}^{m})$ with $\nabla f\in VMO(B^{n})$, $f(0)=0$, there holds
-$\lim_{r\rightarrow 0^{+}}\frac{\int_{f^{-1}(B_{r}(0))}1\,d\textup{vol}_{g_{f}}}{r^{n}|B^{n}|}=1,$ (IV.31)
+$\lim_{r\rightarrow 0^{+}}\frac{\int_{f^{-1}(B_{r}(0))}1\,d\textup{vol}_{g_{f}}}{r^{n}|B^{n}|}=1,$ (IV.32)
 where $g_{f}\coloneqq f^{*}g_{\text{std}}$. For $r\in(0,1)$, we define $f_{r}\colon B^{n}\rightarrow\mathbb{R}^{m}$ by
-$f_{r}(w)\coloneqq\frac{f(r\omega)}{r}.$ (IV.32)
+$f_{r}(w)\coloneqq\frac{f(r\omega)}{r}.$ (IV.33)
 Then $\{f_{r}\}$ is uniformly bounded in $W^{1,\infty}(B^{n})$. Moreover, we have the following limit
-$\fint_{B^{n}}\left|\nabla f_{r}-\fint_{B^{n}}\nabla f_{r}\right|=\fint_{B_{r}(0)}\left|\nabla f-\fint_{B_{r}(0)}\nabla f\right|\xrightarrow[r\to 0^{+}]{}0.$ (IV.33)
+$\fint_{B^{n}}\left|\nabla f_{r}-\fint_{B^{n}}\nabla f_{r}\right|=\fint_{B_{r}(0)}\left|\nabla f-\fint_{B_{r}(0)}\nabla f\right|\xrightarrow[r\to 0^{+}]{}0.$ (IV.34)
 Since $f$ is bi-Lipschitz on $B^{n}$, there exists a positive constant $C_{0}$ such that
-$|x|\leq C_{0}|f(x)|,\qquad\text{ for all }x\in B^{n}.$ (IV.34)
+$|x|\leq C_{0}|f(x)|,\qquad\text{ for all }x\in B^{n}.$ (IV.35)
 Let $\{r_{k}\}_{k=1}^{\infty}\subset\mathbb{R}_{+}$ be an arbitrary sequence with $r_{k}\rightarrow 0^{+}$, then there exists a subsequence $\{r_{k_{i}}\}_{i=1}^{\infty}$ and an $m\times n$ matrix $M$ such that
-$\fint_{B^{n}}\nabla f_{C_{0}r_{k_{i}}}\xrightarrow[i\to\infty]{}M.$ (IV.35)
+$\fint_{B^{n}}\nabla f_{C_{0}r_{k_{i}}}\xrightarrow[i\to\infty]{}M.$ (IV.36)
 Now it suffices to prove
 $\lim_{i\rightarrow\infty}\frac{\int_{f^{-1}\left(B_{r_{k_{i}}}(0)\right)}1\,d\textup{vol}_{g_{f}}}{r_{k_{i}}^{n}|B^{n}|}=1.$
-Denote $\tilde{f}_{i}\coloneqq f_{C_{0}r_{k_{i}}}$. By (IV.33), (IV.35) we have
-$\lim_{i\rightarrow\infty}\int_{B^{n}}\left|\nabla\tilde{f}_{i}-M\right|=0.$ (IV.36)
+Denote $\tilde{f}_{i}\coloneqq f_{C_{0}r_{k_{i}}}$. By (IV.34), (IV.36) we have
+$\lim_{i\rightarrow\infty}\int_{B^{n}}\left|\nabla\tilde{f}_{i}-M\right|=0.$ (IV.37)
 Let $\tilde{f}\colon B^{n}\rightarrow\mathbb{R}^{m}$ defined by $\tilde{f}(\omega)=M\omega$.
-Since $\{\tilde{f}_{i}\}$ is bounded in $W^{1,\infty}(B^{n})$ and $\tilde{f}_{i}(0)=0$ for any $i$, by (IV.36) we obtain that
+Since $\{\tilde{f}_{i}\}$ is bounded in $W^{1,\infty}(B^{n})$ and $\tilde{f}_{i}(0)=0$ for any $i$, by (IV.37) we obtain that
 $\tilde{f}_{i}\xrightarrow[i\rightarrow\infty]{}\tilde{f}\quad\text{in }W^{1,p}(B^{n}),\qquad\text{for all }1\leq p<\infty.$
-By (IV.34) and definition of $f_{r}$, for any $\omega\in B^{n}$, $i\in\mathbb{N}$, we have
+By (IV.35) and definition of $f_{r}$, for any $\omega\in B^{n}$, $i\in\mathbb{N}$, we have
 $\left|\omega\right|\leq C_{0}\big|\tilde{f}_{i}(\omega)\big|.$
 Hence for any $\omega\in B^{n}$, we have
-$|\omega|\leq C_{0}\big|\tilde{f}(\omega)\big|=C_{0}|M\,\omega|.$ (IV.37)
-In particular, this implies $M$ has full rank. By (IV.34) we have $f^{-1}\big(B_{r_{k_{i}}}(0)\big)\subset B_{C_{0}r_{k_{i}}}(0)$. We can then apply change of variables formula to get
-$(C_{0}r_{k_{i}})^{-n}\int_{f^{-1}\left(B_{r_{k_{i}}}(0)\right)}1\,d\textup{vol}_{g_{f}}=\int_{\tilde{f}_{i}^{-1}\big(B_{C_{0}^{-1}}(0)\big)}\sqrt{\det\big(\nabla\tilde{f}_{i}^{T}\nabla\tilde{f}_{i}\big)}.$ (IV.38)
+$|\omega|\leq C_{0}\big|\tilde{f}(\omega)\big|=C_{0}|M\,\omega|.$ (IV.38)
+In particular, this implies $M$ has full rank. By (IV.35) we have $f^{-1}\big(B_{r_{k_{i}}}(0)\big)\subset B_{C_{0}r_{k_{i}}}(0)$. We can then apply the change of variables formula to get
+$(C_{0}r_{k_{i}})^{-n}\int_{f^{-1}\left(B_{r_{k_{i}}}(0)\right)}1\,d\textup{vol}_{g_{f}}=\int_{\tilde{f}_{i}^{-1}\big(B_{C_{0}^{-1}}(0)\big)}\sqrt{\det\big(\nabla\tilde{f}_{i}^{T}\nabla\tilde{f}_{i}\big)}.$ (IV.39)
 By the definition of weak immersions, there exists a constant $C>0$ independent of $i$ such that for all $i\in\mathbb{N}$,
 $C^{-1}\leq\det\big(\nabla\tilde{f}_{i}^{T}\nabla\tilde{f}_{i}\big)\leq C\qquad\text{a.e. on }B^{n}.$
-And since $\nabla\tilde{f}_{i}\rightarrow M$ in $L^{p}(B^{n})$ for any $p\in[1,\infty)$, we have
-$\left\|\sqrt{\det\big(\nabla\tilde{f}_{i}^{T}\nabla\tilde{f}_{i}\big)}-\sqrt{\det\big(M^{T}M\big)}\right\|_{L^{1}(B^{n})}\xrightarrow[i\rightarrow\infty]{}0.$ (IV.39)
-By Sobolev embedding we have $\tilde{f}_{i}\rightarrow\tilde{f}$ in $C^{0}(\overline{B^{n}})$  [^6] In general, $f_{r}$ defined in (IV.32) does not has a $C^{0}$ limit as $r\to 0$ under the same assumptions.
+Since $\nabla\tilde{f}_{i}\rightarrow M$ in $L^{p}(B^{n})$ for any $p\in[1,\infty)$, we have
+$\left\|\sqrt{\det\big(\nabla\tilde{f}_{i}^{T}\nabla\tilde{f}_{i}\big)}-\sqrt{\det\big(M^{T}M\big)}\right\|_{L^{1}(B^{n})}\xrightarrow[i\rightarrow\infty]{}0.$ (IV.40)
+By Sobolev embedding we have $\tilde{f}_{i}\rightarrow\tilde{f}$ in $C^{0}(\overline{B^{n}})$  [^8] In general, $f_{r}$ defined in (IV.33) does not have a $C^{0}$ limit as $r\to 0$ under the same assumptions.
 For instance, we may take $\bar{f}(x_{1},x_{2})\coloneqq\begin{cases}0&\text{if }(x_{1},x_{2})=0,\\[5.69054pt] \Big(x_{1},x_{2},100^{-1}\,x_{1}\cos\big(\log(1-\log(x_{1}^{2}+x_{2}^{2}))\big)\Big)&\text{otherwise}.\end{cases}$ Then $\bar{f}\in W^{2,2}_{\textup{imm}}(D^{2},\mathbb{R}^{3})$ hence $\nabla\bar{f}\in VMO(D^{2})$, but for any $\omega\in\{(x_{1},x_{2})\in D^{2}:x_{1}\neq 0\}$, $\bar{f}_{r}(\omega)\coloneqq r^{-1}\bar{f}(r\omega)$ diverges as $r\rightarrow 0$.
 Moreover, we compute the contingent cone [8, Definition 4.1.1] and Clarke tangent cone [8, Definition 4.1.5] of $\bar{f}(D^{2})$ at 0: $T_{\bar{f}(D^{2})}(0)=\big\{(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}:|x_{3}|\leq|x_{1}|\big\},\qquad C_{\bar{f}(D^{2})}(0)=\{0\}\times\mathbb{R}\times\{0\}.$ Neither of the two tangent cones above is a $2$-plane in $\mathbb{R}^{3}$.
-However, if we assume $f\colon D^{2}\rightarrow\mathbb{R}^{3}$ is a $C^{1}$ embedding, then it holds that $T_{f(D^{2})}(0)=C_{f(D^{2})}(0)=\text{Im}(\nabla f(0)).$ , and since $\mathcal{L}^{n}\left(\tilde{f}^{-1}\big(\partial B_{C_{0}^{-1}}(0)\big)\right)=0$, by [55, Theorem 3.2.11] we obtain
-$\mathcal{L}^{n}\left(\tilde{f}_{i}^{-1}\big(B_{C_{0}^{-1}}(0)\big)\right)\rightarrow\mathcal{L}^{n}\left(\tilde{f}^{-1}\big(B_{C_{0}^{-1}}(0)\big)\right).$ (IV.40)
+However, if we assume $f\colon D^{2}\rightarrow\mathbb{R}^{3}$ is a $C^{1}$ embedding, then it holds that $T_{f(D^{2})}(0)=C_{f(D^{2})}(0)=\text{Im}(\nabla f(0)).$ .
+Since $\mathcal{L}^{n}\left(\tilde{f}^{-1}\big(\partial B_{C_{0}^{-1}}(0)\big)\right)=0$, by [55, Theorem 3.2.11] we obtain
+$\mathcal{L}^{n}\left(\tilde{f}_{i}^{-1}\big(B_{C_{0}^{-1}}(0)\big)\right)\rightarrow\mathcal{L}^{n}\left(\tilde{f}^{-1}\big(B_{C_{0}^{-1}}(0)\big)\right).$ (IV.41)
 Since $\tilde{f}(\omega)=M\omega$, we have
-$\begin{aligned} \mathcal{L}^{n}\left(\tilde{f}^{-1}\big(B_{C_{0}^{-1}}(0)\big)\right)&=C_{0}^{-n}\mathcal{L}^{n}\left\{\omega\in\mathbb{R}^{n}:\langle M^{T}M\omega,\omega\rangle\leq 1\right\}.\end{aligned}$ (IV.41)
+$\begin{aligned} \mathcal{L}^{n}\left(\tilde{f}^{-1}\big(B_{C_{0}^{-1}}(0)\big)\right)&=C_{0}^{-n}\mathcal{L}^{n}\left\{\omega\in\mathbb{R}^{n}:\langle M^{T}M\omega,\omega\rangle\leq 1\right\}.\end{aligned}$ (IV.42)
 Since $M^{T}M$ is positive definite, there exist $Q\in SO(n)$ and $D=\text{diag}(\lambda_{1},\dots,\lambda_{n})$ with $\lambda_{1},\dots,\lambda_{n}>0$ such that $M^{T}M=QDQ^{-1}$. It follows that
-$\begin{aligned} \mathcal{L}^{n}\left\{\omega\in\mathbb{R}^{n}:\langle M^{T}M\omega,\omega\rangle\leq 1\right\}&=\mathcal{L}^{n}\left\{\omega\in\mathbb{R}^{n}:\langle DQ^{-1}\omega,Q^{-1}\omega\rangle\leq 1\right\}\\[7.11317pt] &=\mathcal{L}^{n}\left\{v\in\mathbb{R}^{n}:\langle Dv,v\rangle\leq 1\right\}\\ &=\mathcal{L}^{n}\bigg\{(v_{1},\dots,v_{n})\in\mathbb{R}^{n}:\sum_{j=1}^{n}\lambda_{j}v_{j}^{2}\leq 1\bigg\}.\end{aligned}$ (IV.42)
+$\begin{aligned} \mathcal{L}^{n}\left\{\omega\in\mathbb{R}^{n}:\langle M^{T}M\omega,\omega\rangle\leq 1\right\}&=\mathcal{L}^{n}\left\{\omega\in\mathbb{R}^{n}:\langle DQ^{-1}\omega,Q^{-1}\omega\rangle\leq 1\right\}\\[7.11317pt] &=\mathcal{L}^{n}\left\{v\in\mathbb{R}^{n}:\langle Dv,v\rangle\leq 1\right\}\\ &=\mathcal{L}^{n}\bigg\{(v_{1},\dots,v_{n})\in\mathbb{R}^{n}:\sum_{j=1}^{n}\lambda_{j}v_{j}^{2}\leq 1\bigg\}.\end{aligned}$ (IV.43)
 By a linear change of variable, we have
-$\begin{aligned} \mathcal{L}^{n}\bigg\{(v_{1},\dots,v_{n})\in\mathbb{R}^{n}:\sum_{j=1}^{n}\lambda_{j}v_{j}^{2}\leq 1\bigg\}&=\prod_{j=1}^{n}\lambda_{j}^{-\frac{1}{2}}|B^{n}|=\det(M^{T}M)^{-\frac{1}{2}}|B^{n}|.\end{aligned}$ (IV.43)
-Combining (IV.38)–(IV.43), we finally conclude
+$\begin{aligned} \mathcal{L}^{n}\bigg\{(v_{1},\dots,v_{n})\in\mathbb{R}^{n}:\sum_{j=1}^{n}\lambda_{j}v_{j}^{2}\leq 1\bigg\}&=\prod_{j=1}^{n}\lambda_{j}^{-\frac{1}{2}}|B^{n}|=\det(M^{T}M)^{-\frac{1}{2}}|B^{n}|.\end{aligned}$ (IV.44)
+Combining (IV.39)–(IV.44), we finally conclude
 $\lim_{i\rightarrow\infty}\frac{\int_{f^{-1}\big(B_{r_{k_{i}}}(0)\big)}1\,d\textup{vol}_{g_{f}}}{r_{k_{i}}^{n}|B^{n}|}=\frac{C_{0}^{n}\mathcal{L}^{n}\left(\tilde{f}^{-1}\big(B_{C_{0}^{-1}}(0)\big)\right)\sqrt{\det(M^{T}M)}}{|B^{n}|}=1.$
-Thus, the limit (IV.31) is proved. ∎
+Thus, the limit (IV.32) is proved. ∎
 IV.4 Approximation by smooth immersions
-In this section, we prove that every weak immersion $\vec{\Phi}$ can always be approximated by smooth immersions $\vec{\Phi}_{k}$ and that the conformal structures induced by $\vec{\Phi}_{k}$ converge to the one of $\vec{\Phi}$.
-To do so, we first prove in Section IV.4.1 that the conformal charts obtained by Coulomb frames converges.
-Combining this with an argument from Uhlenbeck, we prove in Section IV.4.2 that the conformal structures converges as well.
-We summarize the approximations results that we obtain in Section IV.4.3.
+In this section, we prove that every weak immersion $\vec{\Phi}$ can be approximated by smooth immersions $\vec{\Phi}_{k}$ and that the conformal structures induced by $\vec{\Phi}_{k}$ converge to that of $\vec{\Phi}$.
+To do so, we first prove in Section IV.4.1 that the conformal charts obtained by Coulomb frames converge.
+Combining this with an argument from Uhlenbeck, we prove in Section IV.4.2 that the conformal structures converge as well.
+We summarize the approximation results that we obtain in Section IV.4.3.
 IV.4.1 Convergence of isothermal coordinates
@@
 Let $u_{k}\in W^{1,2}(U)$ be a solution to the equation
-$\partial_{i}\big(a^{ij}_{k}\,\partial_{j}u_{k}\big)=f_{k}\quad\text{in }U.$ (IV.44)
+$\partial_{i}\big(a^{ij}_{k}\,\partial_{j}u_{k}\big)=f_{k}\quad\text{in }U.$ (IV.45)
 Suppose that $u_{k}\rightharpoonup u$ weakly in $W^{1,2}(U)$. Then we have $u_{k}\rightarrow u$ in $W_{\textup{loc}}^{1,2}(U)$.
@@
 Denote by $\left\langle,\right\rangle$ the duality pairing between $W^{-1,2}(U)$ and $W^{1,2}_{0}(U)$.
-Using $(u_{k}-u)\,\xi$ as a test function in the equation (IV.44), we obtain
-$\begin{aligned} &\Lambda^{-1}\int_{V}|\nabla(u_{k}-u)|^{2}\\ &\leq\int_{U}a^{ij}_{k}\,\xi\,\partial_{j}(u_{k}-u)\,\partial_{i}(u_{k}-u)\\ &=-\left\langle f_{k},\xi(u_{k}-u)\right\rangle-\int_{U}\big(a^{ij}_{k}\,\partial_{j}u_{k}\,\partial_{i}\xi\,(u_{k}-u)+a^{ij}_{k}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\big).\end{aligned}$ (IV.45)
+Using $(u_{k}-u)\,\xi$ as a test function in the equation (IV.45), we obtain
+$\begin{aligned} &\Lambda^{-1}\int_{V}|\nabla(u_{k}-u)|^{2}\leq\int_{U}a^{ij}_{k}\,\xi\,\partial_{j}(u_{k}-u)\,\partial_{i}(u_{k}-u)\\ &\leq-\left\langle f_{k},\xi(u_{k}-u)\right\rangle-\int_{U}\big(a^{ij}_{k}\,\partial_{j}u_{k}\,\partial_{i}\xi\,(u_{k}-u)+a^{ij}_{k}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\big).\end{aligned}$ (IV.46)
 Since $\xi(u_{k}-u)\rightharpoonup 0$ weakly in $W_{0}^{1,2}(U)$ and $\|f_{k}-f\|_{W^{-1,2}(U)}\rightarrow 0$ as $k\rightarrow\infty$, we have
-$\begin{aligned} \left|\left\langle f_{k},\xi(u_{k}-u)\right\rangle\right|&\leq\left|\left\langle f_{k}-f,\xi(u_{k}-u)\right\rangle\right|+\left|\left\langle f,\xi(u_{k}-u)\right\rangle\right|\\[8.53581pt] &\leq\left\|f_{k}-f\right\|_{W^{-1,2}(U)}\left\|\xi(u_{k}-u)\right\|_{W^{1,2}(U)}+\left|\left\langle f,\xi(u_{k}-u)\right\rangle\right|\\[5.69054pt] &\xrightarrow[k\to\infty]{}0.\end{aligned}$ (IV.46)
+$\begin{aligned} \left|\left\langle f_{k},\xi(u_{k}-u)\right\rangle\right|&\leq\left|\left\langle f_{k}-f,\xi(u_{k}-u)\right\rangle\right|+\left|\left\langle f,\xi(u_{k}-u)\right\rangle\right|\\[8.53581pt] &\leq\left\|f_{k}-f\right\|_{W^{-1,2}(U)}\left\|\xi(u_{k}-u)\right\|_{W^{1,2}(U)}+\left|\left\langle f,\xi(u_{k}-u)\right\rangle\right|\\[5.69054pt] &\xrightarrow[k\to\infty]{}0.\end{aligned}$ (IV.47)
 By Rellich–Kondrachov compactness theorem, we also get
-$\begin{aligned} \left|\int_{U}a^{ij}_{k}\,\partial_{j}u_{k}\,\partial_{i}\xi\,(u_{k}-u)\right|&\leq\left\|a^{ij}_{k}\right\|_{L^{\infty}(U)}\left\|\partial_{j}u_{k}\right\|_{L^{2}(U)}\left\|\partial_{i}\xi\,(u_{k}-u)\right\|_{L^{2}(U)}\xrightarrow[k\to\infty]{}0.\end{aligned}$ (IV.47)
+$\begin{aligned} \left|\int_{U}a^{ij}_{k}\,\partial_{j}u_{k}\,\partial_{i}\xi\,(u_{k}-u)\right|&\leq\left\|a^{ij}_{k}\right\|_{L^{\infty}(U)}\left\|\partial_{j}u_{k}\right\|_{L^{2}(U)}\left\|\partial_{i}\xi\,(u_{k}-u)\right\|_{L^{2}(U)}\xrightarrow[k\to\infty]{}0.\end{aligned}$ (IV.48)
 Concerning the last term, by weak convergence $u_{k}\rightharpoonup u$ in $W^{1,2}(U)$ and dominated convergence theorem, we obtain
-$\begin{aligned} &\left|\int_{U}a^{ij}_{k}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|&\\ &\leq\left|\int_{U}a^{ij}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|+\int_{U}\left|(a^{ij}_{k}-a^{ij})\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|\\ &\leq\left|\int_{U}a^{ij}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|+\left\|(a^{ij}_{k}-a^{ij})\,\xi\,\partial_{j}u\right\|_{L^{2}(U)}\left\|\partial_{i}(u_{k}-u)\right\|_{L^{2}(U)}\\ &\xrightarrow[k\to\infty]{}0.\end{aligned}$ (IV.48)
-Combining (LABEL:est-diff-W12)–(LABEL:3rd-est-Duk-diff) we finally conclude $u_{k}\rightarrow u$ in $W^{1,2}_{\textup{loc}}(U)$. ∎
+$\begin{aligned} \left|\int_{U}a^{ij}_{k}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|&\leq\left|\int_{U}a^{ij}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|+\int_{U}\left|(a^{ij}_{k}-a^{ij})\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|\\ &\leq\left|\int_{U}a^{ij}\,\xi\,\partial_{j}u\,\partial_{i}(u_{k}-u)\right|+\left\|(a^{ij}_{k}-a^{ij})\,\xi\,\partial_{j}u\right\|_{L^{2}(U)}\left\|\partial_{i}(u_{k}-u)\right\|_{L^{2}(U)}\\ &\xrightarrow[k\to\infty]{}0.\end{aligned}$ (IV.49)
+Combining (LABEL:est-diff-W12)–(IV.49) we finally conclude $u_{k}\rightarrow u$ in $W^{1,2}_{\textup{loc}}(U)$. ∎
 Using the proof of Theorem IV.4, we are now ready to prove strong convergence of conformal charts for a convergent sequence of weak immersions with bounded induced metric.
@@
 $p\in D^{2}$, there holds
-$\Lambda^{-1}|X|_{\mathbb{R}^{2}}\leq|d(\vec{\Phi}_{k})_{p}(X)|_{\mathbb{R}^{m}}\leq\Lambda|X|_{\mathbb{R}^{2}}.$ (IV.49)
+$\Lambda^{-1}|X|_{\mathbb{R}^{2}}\leq|d(\vec{\Phi}_{k})_{p}(X)|_{\mathbb{R}^{m}}\leq\Lambda|X|_{\mathbb{R}^{2}}.$ (IV.50)
 Then there exists a neighborhood $U\subset D^{2}$ of 0 and $W^{1,\infty}\cap W^{2,2}$ diffeomorphisms $\varphi_{k}\colon U\rightarrow\varphi_{k}(U)\subset\mathbb{R}^{2}$ such that $\vec{\Phi}_{k}\circ\varphi_{k}^{-1}$ is (weakly) conformal on $\varphi_{k}(U)$ for any $k\in\mathbb{N}\cup\{\infty\}$, $\varphi_{k}\rightarrow\varphi_{\infty}$ in $W^{2,2}(U)$ and $\|\varphi_{k}\|_{W^{1,\infty}(U)}\leq C$, $\|\varphi_{k}^{-1}\|_{W^{1,\infty}(\varphi_{k}(U))}\leq C$ for a constant $C>0$.
@@
 For $k\in\mathbb{N}\cup\{\infty\}$, we apply Gram–Schmidt process to $(\partial_{1}\vec{\Phi}_{k},\partial_{2}\vec{\Phi}_{k})$ and get
-$\vec{f}_{1,k}\coloneqq\frac{\partial_{1}\vec{\Phi}_{k}}{|\partial_{1}\vec{\Phi}_{k}|},\qquad\vec{f}_{2,k}\coloneqq\frac{\partial_{2}\vec{\Phi}_{k}-(\partial_{2}\vec{\Phi}_{k}\cdot\vec{f}_{1,k})\vec{f}_{1,k}}{|\partial_{2}\vec{\Phi}_{k}-(\partial_{2}\vec{\Phi}_{k}\cdot\vec{f}_{1,k})\vec{f}_{1,k}|}.$ (IV.50)
-By (IV.49) we have $|\partial_{1}\vec{\Phi}_{k}|,|\partial_{2}\vec{\Phi}_{k}|\in[\Lambda^{-1},\Lambda]$, and
-$|\partial_{2}\vec{\Phi}_{k}-(\partial_{2}\vec{\Phi}_{k}\cdot\vec{f}_{1,k})\vec{f}_{1,k}|\in[\Lambda^{-1},\Lambda(1+\Lambda^{2})].$
+$\vec{f}_{1,k}\coloneqq\frac{\partial_{1}\vec{\Phi}_{k}}{|\partial_{1}\vec{\Phi}_{k}|},\qquad\vec{f}_{2,k}\coloneqq\frac{\partial_{2}\vec{\Phi}_{k}-(\partial_{2}\vec{\Phi}_{k}\cdot\vec{f}_{1,k})\vec{f}_{1,k}}{|\partial_{2}\vec{\Phi}_{k}-(\partial_{2}\vec{\Phi}_{k}\cdot\vec{f}_{1,k})\vec{f}_{1,k}|}.$ (IV.51)
+By (IV.50) we have $|\partial_{1}\vec{\Phi}_{k}|,|\partial_{2}\vec{\Phi}_{k}|\in[\Lambda^{-1},\Lambda]$, and
+$\left|\partial_{2}\vec{\Phi}_{k}-(\partial_{2}\vec{\Phi}_{k}\cdot\vec{f}_{1,k})\vec{f}_{1,k}\right|\in[\Lambda^{-1},\Lambda(1+\Lambda^{2})].$
 Hence it holds that $(\vec{f}_{1,k},\vec{f}_{2,k})\in L^{\infty}\cap W^{1,2}(D^{2},V_{2}({\mathbb{R}}^{m}))$, where $V_{2}({\mathbb{R}}^{m})$ denotes the space of orthonormal 2-frames in ${\mathbb{R}}^{m}$.
@@
 We look now for a rotation of this orthonormal frame realizing the following absolute minimum
-$\begin{array}[]{l}\displaystyle\inf_{\theta\in W_{0}^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}|\langle\vec{f}^{\,\theta}_{1,k},d\vec{f}^{\,\theta}_{2,k}\rangle|^{2}_{g_{k}}\ d\textup{vol}_{g_{k}}\\[14.22636pt] \displaystyle=\inf_{\theta\in W_{0}^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}\sum_{i,j=1}^{2}g^{ij}_{k}\,\langle\vec{f}^{\,\theta}_{1,k},\partial_{i}\vec{f}^{\,\theta}_{2,k}\rangle\,\langle\vec{f}^{\,\theta}_{1,k},\partial_{j}\vec{f}^{\,\theta}_{2,k}\rangle\,\sqrt{\det g_{k}}\ dx^{1}\wedge dx^{2}.\end{array}$ (IV.51)
+$\begin{array}[]{l}\displaystyle\inf_{\theta\in W_{0}^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}|\langle\vec{f}^{\,\theta}_{1,k},d\vec{f}^{\,\theta}_{2,k}\rangle|^{2}_{g_{k}}\ d\textup{vol}_{g_{k}}\\[14.22636pt] \displaystyle=\inf_{\theta\in W_{0}^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}\sum_{i,j=1}^{2}g^{ij}_{k}\,\langle\vec{f}^{\,\theta}_{1,k},\partial_{i}\vec{f}^{\,\theta}_{2,k}\rangle\,\langle\vec{f}^{\,\theta}_{1,k},\partial_{j}\vec{f}^{\,\theta}_{2,k}\rangle\,\sqrt{\det g_{k}}\ dx^{1}\wedge dx^{2}.\end{array}$ (IV.52)
 For $i\in\{1,2\}$, we have
-$\langle\vec{f}^{\,\theta}_{1,k},\partial_{i}\vec{f}^{\,\theta}_{2,k}\rangle=\partial_{i}\theta+\langle\vec{f}_{1,k},\partial_{i}\vec{f}_{2,k}\rangle.$ (IV.52)
+$\langle\vec{f}^{\,\theta}_{1,k},\partial_{i}\vec{f}^{\,\theta}_{2,k}\rangle=\partial_{i}\theta+\langle\vec{f}_{1,k},\partial_{i}\vec{f}_{2,k}\rangle.$ (IV.53)
 Hence the following energy is strictly convex in $W_{0}^{1,2}(D^{2},{\mathbb{R}})$:
@@
 It is the unique solution in $W_{0}^{1,2}(D^{2},\mathbb{R})$ for the Euler–Lagrange equation:
-$\forall\phi\in W_{0}^{1,2}(D^{2},{\mathbb{R}}),\qquad\int_{D^{2}}d\phi\wedge*_{g_{k}}\left(d\theta_{k}+\langle\vec{f}_{1,k},d\vec{f}_{2,k}\rangle\right)=0.$ (IV.53)
-In particular, taking $\phi=\theta_{k}$ in (IV.53), we get
+$\forall\phi\in W_{0}^{1,2}(D^{2},{\mathbb{R}}),\qquad\int_{D^{2}}d\phi\wedge*_{g_{k}}\left(d\theta_{k}+\langle\vec{f}_{1,k},d\vec{f}_{2,k}\rangle\right)=0.$ (IV.54)
+In particular, taking $\phi=\theta_{k}$ in (IV.54), we get
 $\Lambda^{-4}\int_{D^{2}}|\nabla\theta_{k}|^{2}\,dx^{2}\leq\int_{D^{2}}|d\theta_{k}|^{2}_{g_{k}}\,d\textup{vol}_{g_{k}}\leq\int_{D^{2}}|d\vec{f}_{2,k}|_{g_{k}}^{2}\,d\textup{vol}_{g_{k}}\leq\Lambda^{4}\int_{D^{2}}|\nabla\vec{f}_{2,k}|^{2}.$
 Thus the sequence $\{\theta_{k}\}$ is bounded in $W_{0}^{1,2}(D^{2})$ and each subsequence contains a further subsequence $\{\theta_{k_{l}}\}$ weakly converging to some $\theta^{\prime}\in W^{1,2}_{0}(D^{2})$.
-Then using (IV.53), Hölder’s inequality and dominated convergence theorem, we have
+Then using (IV.54), Hölder’s inequality and dominated convergence theorem, we have
 $\forall\phi\in W_{0}^{1,2}(D^{2},{\mathbb{R}}),\qquad\int_{D^{2}}d\phi\wedge*_{g_{\infty}}\left(d\theta^{\prime}+\langle\vec{f}_{1,\infty},d\vec{f}_{2,\infty}\rangle\right)=0.$
@@
 Denoting $(\vec{e}_{1,k},\vec{e}_{2,k})\coloneqq(\vec{f}^{\,\theta_{k}}_{1,k},\vec{f}^{\,\theta_{k}}_{2,k})$, we have $\vec{e}_{i,k}\rightarrow\vec{e}_{i,\infty}$ as $k\rightarrow\infty$ for $i\in\{1,2\}$.
-We deduce from (IV.53) that the following equation holds in the sense of distribution:
+We deduce from (IV.54) that the following equation holds in the sense of distribution:
 $d\Big(*_{g_{k}}\langle\vec{e}_{1,k},d\vec{e}_{2,k}\rangle\Big)=0.$
@@
 It follows that $\{\lambda_{k}\}_{k\in\mathbb{N}}$ is bounded in $C^{0}(\overline{D_{1/2}(0)})$. For $i\in\{1,2\}$, we define pullback of $\vec{e}_{i,k}$ and the dual $1$-forms
-$\begin{dcases}e_{i,k}\coloneqq g^{jl}_{k}\,\langle\vec{e}_{i,k},\partial_{j}\vec{\Phi}_{k}\rangle\,\frac{\partial}{\partial x^{l}}\in L^{\infty}\cap W^{1,2}(D_{1/2}(0),TD_{1/2}(0)),\\[5.69054pt] e_{i,k}^{\ast}\coloneqq\langle\vec{e}_{i,k},\partial_{j}\vec{\Phi_{k}}\rangle\,dx^{j}=\langle\vec{e}_{i,k},d\vec{\Phi}_{k}\rangle\in L^{\infty}\cap W^{1,2}(D_{1/2}(0),T^{*}D_{1/2}(0)).\end{dcases}$ (IV.54)
+$\begin{dcases}e_{i,k}\coloneqq g^{jl}_{k}\,\langle\vec{e}_{i,k},\partial_{j}\vec{\Phi}_{k}\rangle\,\frac{\partial}{\partial x^{l}}\in L^{\infty}\cap W^{1,2}(D_{1/2}(0),TD_{1/2}(0)),\\[5.69054pt] e_{i,k}^{\ast}\coloneqq\langle\vec{e}_{i,k},\partial_{j}\vec{\Phi_{k}}\rangle\,dx^{j}=\langle\vec{e}_{i,k},d\vec{\Phi}_{k}\rangle\in L^{\infty}\cap W^{1,2}(D_{1/2}(0),T^{*}D_{1/2}(0)).\end{dcases}$ (IV.55)
 Then by the same proof of Theorem IV.4, we have
@@
 Using the weak Poincaré lemma, we obtain the existence of $\varphi^{i}_{k}\in W^{1,\infty}\cap W^{2,2}(D_{1/2}(0),{\mathbb{R}})$ with $\int_{D_{1/2}(0)}\varphi^{i}_{k}=0$ for $i\in\{1,2\}$ such that
-$d\varphi^{i}_{k}=e^{-\lambda_{k}}\,e_{i,k}^{*}.$ (IV.55)
+$d\varphi^{i}_{k}=e^{-\lambda_{k}}\,e_{i,k}^{*}.$ (IV.56)
 Let $\varphi_{k}\coloneqq(\varphi^{1}_{k},\varphi^{2}_{k})\in W^{1,\infty}\cap W^{2,2}(D_{1/2}(0),{\mathbb{R}^{2}})$.
-From (IV.54) and (IV.55) we deduce that for a.e.
+From (IV.55) and (IV.56) we deduce that for a.e.
 $p\in D_{1/2}(0)$ and any $v\in T_{p}(D_{1/2}(0))$,
-$e^{-\|\lambda_{k}\|_{C(\overline{D_{1/2}(0)})}}\,\Lambda^{-1}|v|\leq|d(\varphi_{k})_{p}(v)|=e^{-\lambda_{k}}|d\vec{\Phi}_{k}(v)|\leq e^{\|\lambda_{k}\|_{C(\overline{D_{1/2}(0)})}}\,\Lambda|v|.$ (IV.56)
+$e^{-\|\lambda_{k}\|_{C(\overline{D_{1/2}(0)})}}\,\Lambda^{-1}|v|\leq|d(\varphi_{k})_{p}(v)|=e^{-\lambda_{k}}|d\vec{\Phi}_{k}(v)|\leq e^{\|\lambda_{k}\|_{C(\overline{D_{1/2}(0)})}}\,\Lambda|v|.$ (IV.57)
 Since $\lambda_{k}\rightarrow\lambda_{\infty}$, $e_{i,k}^{*}\rightarrow e^{*}_{i,\infty}$ in $W^{1,2}(D^{2})$ as $k\rightarrow\infty$, and $\{\lambda_{k}\}$ is bounded in $C(\overline{D_{1/2}(0)})$, by dominated convergence theorem and Poincaré inequality we have $\varphi_{k}\rightarrow\varphi_{\infty}$ in $W^{2,2}(D_{1/2}(0))$.
-Combining (IV.56) and Lemma IV.12, it follows that there exists an open neighborhood $U\subset D_{1/2}(0)$ of 0 independent of $k$ such that $\varphi_{k}$ is open and injective on $U$, and its inverse $\varphi_{k}^{-1}\in W^{1,\infty}\cap W^{2,2}(\varphi_{k}(U),\mathbb{R}^{2})$.
+Combining (IV.57) and Lemma IV.12, it follows that there exists an open neighborhood $U\subset D_{1/2}(0)$ of 0 independent of $k$ such that $\varphi_{k}$ is open and injective on $U$, and its inverse $\varphi_{k}^{-1}\in W^{1,\infty}\cap W^{2,2}(\varphi_{k}(U),\mathbb{R}^{2})$.
 We obtain that $\varphi_{k}$ is an isothermal chart of $\vec{\Phi}_{k}$ by the same proof of Theorem IV.4.
 ∎
-IV.4.2 Convergence of conformal structure
+IV.4.2 Convergence of conformal structures
 As a consequence of Theorem IV.17, we deduce the convergence of the underlying conformal structure for sequences of metrics induced by weak immersions which converge in some Sobolev space.
@@
 Proof.
-Let $\chi(\Sigma)$ denote the Euler characteristic of $\Sigma$, $K_{h_{k}}$ be the Gaussian curvature of $h_{k}$.
-(Case I) $\chi(\Sigma)\leq-2\Rightarrow K_{h_{k}}=-1$.
-By a theorem of Poincaré and its corollary (see [185, Theorem 1.6.2]), in this case, for any $s\in\mathbb{N}$, $s\geq 3$ and any metric $g\in W^{s,2}(\Sigma,T^{*}\Sigma\otimes T^{*}\Sigma)$, there exists a unique metric $h\in W^{s,2}(\Sigma,T^{*}\Sigma\otimes T^{*}\Sigma)$ such that $h$ is weakly conformal to $g$ and $h$ has constant Gaussian curvature 1-1.
+Let $\chi(\Sigma)$ denote the Euler characteristic of $\Sigma$ and $K_{h_{k}}$ be the Gaussian curvature of $h_{k}$.
+(Case I) $\chi(\Sigma)\leq-2$ and $K_{h_{k}}=-1$.
+By a theorem of Poincaré and its corollary (see [188, Theorem 1.6.2]), in this case, for any $s\in\mathbb{N}$, $s\geq 3$ and any metric $g\in W^{s,2}(\Sigma,T^{*}\Sigma\otimes T^{*}\Sigma)$, there exists a unique metric $h\in W^{s,2}(\Sigma,T^{*}\Sigma\otimes T^{*}\Sigma)$ such that $h$ is weakly conformal to $g$ and $h$ has constant Gaussian curvature 1-1.
 Moreover, the map taking $g$ to $h$ is continuous in $W^{s,2}$.
 It follows that $h_{k}$ is uniquely determined by $g_{k}$, and converges to $h_{\infty}$ smoothly since $g_{k}\rightarrow g_{\infty}$ in $C^{\infty}(\Sigma,T^{*}\Sigma\otimes T^{*}\Sigma)$.
-(Case II) $\chi(\Sigma)=0\Rightarrow K_{h_{k}}=0$.
+(Case II) $\chi(\Sigma)=0$ and $K_{h_{k}}=0$.
 Let $h_{k}=e^{2\lambda_{k}}g_{k}$ for some $\lambda_{k}\in C^{\infty}(\Sigma)$.
 We denote the Gaussian curvature and negative Laplace–Beltrami operator associated to $g_{k}$ by $K_{g_{k}}$ and $\Delta_{g_{k}}$ respectively, the condition that $h_{k}$ has constant Gaussian curvature 0 is then equivalent to the equation (see e.g.
-[168, Theorem 2.6])
-$\Delta_{g_{k}}\lambda_{k}=K_{g_{k}}.$ (IV.57)
+[171, Theorem 2.6])
+$\Delta_{g_{k}}\lambda_{k}=K_{g_{k}}.$ (IV.58)
 By Gauss–Bonnet theorem, we have
 $\int_{\Sigma}K_{g_{k}}\,d\textup{vol}_{g_{k}}=2\pi\chi(\Sigma)=0.$
-Hence for each $k\in\mathbb{N}\cup\{\infty\}$, there exists a smooth solution $\lambda_{k}$ to (IV.57) and it is unique up to additive constants (see [9, Theorem 4.7]).
+Hence for each $k\in\mathbb{N}\cup\{\infty\}$, there exists a smooth solution $\lambda_{k}$ to (IV.58) and it is unique up to additive constants (see [9, Theorem 4.7]).
 By adding suitable constants, we can without loss of generality assume that
-$\int\lambda_{k}\,d\textup{vol}_{g_{\infty}}=0\quad\text{for any }k\in\mathbb{N}\cup\{\infty\}.$ (IV.58)
+$\forall k\in\mathbb{N}\cup\{\infty\},\qquad\int_{\Sigma}\lambda_{k}\,d\textup{vol}_{g_{\infty}}=0.$ (IV.59)
 Since $g_{k}\rightarrow g_{\infty}$ in $C^{\infty}$, we have $K_{g_{k}}\rightarrow K_{g_{\infty}}$ in $C^{\infty}(\Sigma)$ as $k\rightarrow\infty$, and in particular, $\{K_{g_{k}}\}_{k\in\mathbb{N}\cup\{\infty\}}$ is bounded in $C^{0}(\Sigma)$.
-By (IV.57) and the convergence $g_{k}\rightarrow g_{\infty}$, there exists a constant $C_{1}>0$ independent of $k$ such that
-$C_{1}\int_{\Sigma}|d\lambda_{k}|_{g_{\infty}}^{2}\,d\textup{vol}_{g_{\infty}}$ $\leq\int_{\Sigma}|d\lambda_{k}|_{g_{k}}^{2}\,d\textup{vol}_{g_{k}}$
-$=-\int_{\Sigma}\lambda_{k}\,\Delta_{g_{k}}\lambda_{k}\,d\textup{vol}_{g_{k}}$
-$=\int_{\Sigma}-\lambda_{k}K_{g_{k}}\,d\textup{vol}_{g_{k}}.$
-Furthermore, by (IV.58), Cauchy–Schwarz inequality, and Poincaré’s inequality, there exist constants $C_{2},C_{3}>0$ independent of $k$ such that
-$\int_{\Sigma}-\lambda_{k}K_{g_{k}}\,d\textup{vol}_{g_{k}}$ $\leq\left(\int_{\Sigma}\lambda_{k}^{2}\,d\textup{vol}_{g_{k}}\right)^{\frac{1}{2}}\left(\int_{\Sigma}K_{g_{k}}^{2}\,d\textup{vol}_{g_{k}}\right)^{\frac{1}{2}}$
-$\leq C_{2}\left(\int_{\Sigma}\lambda_{k}^{2}\,d\textup{vol}_{g_{\infty}}\right)^{\frac{1}{2}}$
+By (IV.58) and the convergence $g_{k}\rightarrow g_{\infty}$, there exists a constant $C_{1}>0$ independent of $k$ such that
+$C_{1}\int_{\Sigma}|d\lambda_{k}|_{g_{\infty}}^{2}\,d\textup{vol}_{g_{\infty}}$ $\leq\int_{\Sigma}|d\lambda_{k}|_{g_{k}}^{2}\,d\textup{vol}_{g_{k}}=-\int_{\Sigma}\lambda_{k}\,\Delta_{g_{k}}\lambda_{k}\,d\textup{vol}_{g_{k}}=\int_{\Sigma}-\lambda_{k}K_{g_{k}}\,d\textup{vol}_{g_{k}}.$
+Furthermore, by (IV.59), Cauchy–Schwarz inequality, and Poincaré inequality, there exist constants $C_{2},C_{3}>0$ independent of $k$ such that
+$\int_{\Sigma}-\lambda_{k}K_{g_{k}}\,d\textup{vol}_{g_{k}}$ $\leq\|K_{g_{k}}\|_{L^{\infty}(\Sigma)}\,\textup{vol}_{g_{k}}(\Sigma)^{\frac{1}{2}}\left(\int_{\Sigma}\lambda_{k}^{2}\,d\textup{vol}_{g_{k}}\right)^{\frac{1}{2}}$
+$\leq C_{2}\,\left(\sup_{l\in\mathbb{N}}\|K_{g_{l}}\|_{L^{\infty}(\Sigma)}\right)\,\left(\sup_{l\in\mathbb{N}}\textup{vol}_{g_{l}}(\Sigma)^{\frac{1}{2}}\right)\left(\int_{\Sigma}\lambda_{k}^{2}\,d\textup{vol}_{g_{\infty}}\right)^{\frac{1}{2}}$
 $\leq C_{3}\,\left(\int_{\Sigma}|d\lambda_{k}|_{g_{\infty}}^{2}\,d\textup{vol}_{g_{\infty}}\right)^{\frac{1}{2}}.$
 Consequently, $\{\lambda_{k}\}_{k\in\mathbb{N}\cup\{\infty\}}$ is bounded in $W^{1,2}(\Sigma)$.
-Now we apply the interior estimates as in [58, Section 6.3.1] to the equation (IV.57) and obtain that $\{\lambda_{k}\}_{k\in\mathbb{N}\cup\{\infty\}}$ is bounded in $W^{s,2}(\Sigma)$ for any $s\in\mathbb{N}$.
+Now we apply the interior estimates as in [58, Section 6.3.1] to the equation (IV.58) and obtain that $\{\lambda_{k}\}_{k\in\mathbb{N}\cup\{\infty\}}$ is bounded in $W^{s,2}(\Sigma)$ for any $s\in\mathbb{N}$.
 A standard compactness argument then implies $\lambda_{k}\rightarrow\lambda_{\infty}$ in $C^{\infty}(\Sigma)$ as $k\rightarrow\infty$.
 Finally, by setting $h_{k}=e^{2\lambda_{k}}g_{k}$, we obtain the required convergent sequence of metrics of Gaussian curvature 0.
-(Case III) $\chi(\Sigma)=2\Rightarrow K_{h_{k}}=1$.
+(Case III) $\chi(\Sigma)=2$ and $K_{h_{k}}=1$.
 We identify $\Sigma=\mathbb{C}\cup\{\infty\}$ with the canonical smooth structure and orientation.
@@
 $g_{k}=(g_{k})_{11}\,dx^{2}+2(g_{k})_{12}\,dxdy+(g_{k})_{22}\,dy^{2}=\sigma_{k}|dz+\mu_{k}\,d\bar{z}|^{2},$
-where $dz=dx+idy$ and $d\bar{z}=dx-idy$. Moreover, such $\mu_{k}$ is uniquely determined by $g_{k}$ and can be computed as (see e.g. [152, Section 1.2.1])
+where $dz=dx+idy$ and $d\bar{z}=dx-idy$. Moreover, such $\mu_{k}$ is uniquely determined by $g_{k}$ and can be computed as (see e.g. [155, Section 1.2.1])
 $\mu_{k}=\frac{(g_{k})_{11}-(g_{k})_{22}+2i\,(g_{k})_{12}}{(g_{k})_{11}+(g_{k})_{22}+2\sqrt{(g_{k})_{11}(g_{k})_{22}-(g_{k})_{12}^{2}}}.$
@@
 In particular, $f_{k}$ is quasiconformal with complex dilatation $\mu_{k}$ (see Definition III.6).
-By considering the quasiconformal maps $f_{k}\circ f_{\infty}^{-1}$ and applying [90, Proposition 4.36], we have $f_{k}\rightarrow f_{\infty}$ locally uniformly on $\mathbb{C}$.
-Define $\sigma_{k}\in L^{\infty}(\mathbb{C},\mathbb{R}^{2\times 2})$ be the unique matrix-valued function on $\mathbb{C}$ such that $\sigma_{k}$ is positive definite a.e., $\text{det}(\sigma_{k})=1$, and
+By considering the quasiconformal maps $f_{k}\circ f_{\infty}^{-1}$ and applying [91, Proposition 4.36], we have $f_{k}\rightarrow f_{\infty}$ locally uniformly on $\mathbb{C}$.
+We denote $\sigma_{k}\in L^{\infty}(\mathbb{C},\mathbb{R}^{2\times 2})$ the unique matrix-valued function on $\mathbb{C}$ such that $\sigma_{k}$ is positive definite a.e., $\text{det}(\sigma_{k})=1$, and such that
 $\mu_{k}=\frac{(\sigma_{k})_{22}-(\sigma_{k})_{11}-2i\,(\sigma_{k})_{12}}{(\sigma_{k})_{11}+(\sigma_{k})_{22}+2}.$
-Such $\sigma_{k}$ depends uniquely and smoothly on $\mu_{k}$, hence $\sigma_{k}\rightarrow\sigma_{\infty}$ in $C^{\infty}_{\textup{loc}}(\mathbb{C},\mathbb{R}^{2\times 2})$.
-By [163, Theorem 5.1], $f_{k}$ satisfies the elliptic equation
+Such $\sigma_{k}$ depends uniquely and smoothly on $\mu_{k}$, hence we have $\sigma_{k}\rightarrow\sigma_{\infty}$ in $C^{\infty}_{\textup{loc}}(\mathbb{C},\mathbb{R}^{2\times 2})$.
+By [166, Theorem 5.1], $f_{k}$ satisfies the elliptic equation
 […]
 Hence standard elliptic estimates as in Case II imply that $f_{k}\rightarrow f_{\infty}$ in $C^{\infty}_{\textup{loc}}(\mathbb{C})$.
-On the other hand, note $\psi\circ f_{k}\circ\psi^{-1}$ is quasiconformal with complex dilatation $\mu_{k}\circ\psi^{-1}(\omega)\,\omega^{2}{\bar{\omega}}^{-2}$ (see [90, Proposition 4.13]), and
+On the other hand, the map $\psi\circ f_{k}\circ\psi^{-1}$ is quasiconformal with complex dilatation $\mu_{k}\circ\psi^{-1}(\omega)\,\omega^{2}{\bar{\omega}}^{-2}$ (see [91, Proposition 4.13]).
+Moreover, it holds
 $\mu_{k}\circ\psi^{-1}(\omega)\,\omega^{2}{\bar{\omega}}^{-2}\xrightarrow[k\rightarrow\infty]{}\mu_{\infty}\circ\psi^{-1}(\omega)\,\omega^{2}{\bar{\omega}}^{-2}\qquad\text{ in }C_{\textup{loc}}^{\infty}(\mathbb{C}).$
-By the same argument as before, we have
+By the same argument as before, we obtain
 $\psi\circ f_{k}\circ\psi^{-1}\xrightarrow[k\rightarrow\infty]{}\psi\circ f_{\infty}\circ\psi^{-1}\qquad\text{ in }C_{\textup{loc}}^{\infty}(\mathbb{C}).$
-Hence $f_{k}\rightarrow f_{\infty}$ in $C^{\infty}(\Sigma)$ as $k\rightarrow\infty$.
+Hence it holds $f_{k}\rightarrow f_{\infty}$ in $C^{\infty}(\Sigma)$ as $k\rightarrow\infty$.
 Besides, by the proof of [4, Lemma 5.B.3], both $f_{k}$ and $\psi\circ f_{k}\circ\psi^{-1}$ have no critical points on $\mathbb{C}$, hence each $f_{k}$ is a diffeomorphism on $\Sigma$.
 Finally, for all $k\in\mathbb{N}\cup\{\infty\}$, let $h_{k}\coloneqq f_{k}^{*}h$, then $h_{k}$ is a Riemannian metric with constant Gaussian curvature $1$ on $\Sigma$, and $h_{k}\rightarrow h_{\infty}$ smoothly on $\Sigma$.
-Since $h$ is conformal to $g_{\text{std}}$ and $f_{k}^{*}g_{\text{std}}$ is conformal to $g_{k}$ on $\mathbb{C}$ (see for instance [90, Section 1.5.1]), we have $h_{k}=f_{k}^{*}h$ is conformal to $g_{k}$ on $\mathbb{C}\subset\Sigma$.
+Since $h$ is conformal to $g_{\text{std}}$ and $f_{k}^{*}g_{\text{std}}$ is conformal to $g_{k}$ on $\mathbb{C}$ (see for instance [91, Section 1.5.1]), we have $h_{k}=f_{k}^{*}h$ is conformal to $g_{k}$ on $\mathbb{C}\subset\Sigma$.
 The conformality holds on $\Sigma$ since $h_{k}$, $g_{k}$ are both Riemannian metrics on $\Sigma$, and $\mathbb{C}$ is dense in $\Sigma$.
@@
 Then we consider the general case in which $\{g_{k}\}$ are metrics induced by weak immersions.
-The strategy of proof is an adaptation of [186, Proposition 3.2].
+The strategy of proof is an adaptation of [189, Proposition 3.2].
 For a sequence of functions $f_{k}$ and a compact set $K\subset\mathbb{R}^{n}$, we write $f_{k}\rightarrow f$ in $C^{\infty}(K)$ if each $f_{k}$ is defined on a neighborhood of $K$ (possibly depending on $k$) and for any $l\in\mathbb{N}_{0}$, $\|f_{k}-f\|_{C^{l}(K)}\rightarrow 0$ as $k\rightarrow\infty$.
@@
 Let $\Sigma$ be an oriented connected $2$-dimensional closed smooth manifold with a reference Riemannian metric $\tilde{g}$.
-Assume $\vec{\Phi}_{k}\rightarrow\vec{\Phi}_{\infty}$ in $W^{2,2}(\Sigma,\mathbb{R}^{m})$ as $k\rightarrow\infty$, and there exists a constant $\Lambda>0$ such that for any $k\in\mathbb{N}\cup\{\infty\}$, there holds
-$\Lambda^{-1}|X|_{\tilde{g}}\leq|d(\vec{\Phi}_{k})_{p}(X)|_{\mathbb{R}^{m}}\leq\Lambda|X|_{\tilde{g}},\qquad\text{for a.e. }p\in\Sigma\text{ and all }\mkern 2.0muX\in T_{p}\Sigma.$ (IV.59)
-Let $g_{k}\coloneqq\vec{\Phi}_{k}^{*}\,g_{\text{std}}$ denote the metric induced by $\vec{\Phi}_{k}$.
-Assume $\tilde{g}$ and $g_{\infty}$ are weakly conformal (i.e.
+Assume that $\vec{\Phi}_{k}\rightarrow\vec{\Phi}_{\infty}$ in $W^{2,2}(\Sigma,\mathbb{R}^{m})$ as $k\rightarrow\infty$, and that there exists a constant $\Lambda>0$ such that for any $k\in\mathbb{N}\cup\{\infty\}$, there holds
+$\Lambda^{-1}|X|_{\tilde{g}}\leq|d(\vec{\Phi}_{k})_{p}(X)|_{\mathbb{R}^{m}}\leq\Lambda|X|_{\tilde{g}},\qquad\text{for a.e. }p\in\Sigma\text{ and all }\mkern 2.0muX\in T_{p}\Sigma.$ (IV.60)
+Let $g_{k}\coloneqq\vec{\Phi}_{k}^{*}\,g_{\text{std}}$ be the metric induced by $\vec{\Phi}_{k}$.
+Assume that $\tilde{g}$ and $g_{\infty}$ are weakly conformal (i.e.
 $\Sigma$ is equipped with the complex structure induced by $\vec{\Phi}_{\infty}$ as in Corollary IV.5), then there exists a sequence $(h_{k})_{k\in\mathbb{N}\cup\{\infty\}}$ of metrics smooth in the complex structure induced by $\vec{\Phi}_{k}$, and a sequence $(\Psi_{k})_{k\in\mathbb{N}}$ of bi-Lipschitz homeomorphisms on $\Sigma$ such that
@@
 Then for any $1\leq\alpha,\beta\leq l$ and compact subset $K$ of $\varphi_{\infty,\beta}(U_{\alpha}\cap U_{\beta})$, the transition maps converge as well:
-$\varphi_{k,\alpha,\beta}\coloneqq\varphi_{k,\alpha}\circ\varphi_{k,\beta}^{-1}\xrightarrow[k\rightarrow\infty]{}\varphi_{\infty,\alpha}\circ\varphi_{\infty,\beta}^{-1}=:\varphi_{\infty,\alpha,\beta}\quad\text{in $W^{2,2}(K)$.}$ (IV.60)
+$\varphi_{k,\alpha,\beta}\coloneqq\varphi_{k,\alpha}\circ\varphi_{k,\beta}^{-1}\xrightarrow[k\rightarrow\infty]{}\varphi_{\infty,\alpha}\circ\varphi_{\infty,\beta}^{-1}=:\varphi_{\infty,\alpha,\beta}\quad\text{in $W^{2,2}(K)$.}$ (IV.61)
 Furthermore, since each $\varphi_{k,\alpha,\beta}$ is holomorphic, the convergence holds in $C^{\infty}(K)$.
 We want to prove that for $k$ large enough, there exists an open cover $\{V_{\alpha}\}_{\alpha=1}^{l}$ of $\Sigma$ with $V_{\alpha}\Subset U_{\alpha}$ and $C^{\infty}$ injective immersions $\rho_{k,\alpha}\colon\varphi_{\infty,\alpha}(\overline{V_{\alpha}})\subset\mathbb{R}^{2}\rightarrow\varphi_{k,\alpha}(U_{\alpha})\subset\mathbb{R}^{2}$ satisfying
-$\begin{cases}(i)\ \rho_{k,\alpha}\xrightarrow[k\rightarrow\infty]{}id&\text{ in }C^{\infty}(\varphi_{\infty,\alpha}(\overline{V_{\alpha}}),\mathbb{R}^{2}),\\[8.53581pt] (ii)\ \varphi_{k,\beta,\alpha}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha,\beta}=\rho_{k,\beta}&\text{ on }\varphi_{\infty,\beta}(\overline{V_{\alpha}}\cap\overline{V_{\beta}}).\end{cases}$ (IV.61)
+$\begin{cases}(i)\ \rho_{k,\alpha}\xrightarrow[k\rightarrow\infty]{}id&\text{ in }C^{\infty}(\varphi_{\infty,\alpha}(\overline{V_{\alpha}}),\mathbb{R}^{2}),\\[8.53581pt] (ii)\ \varphi_{k,\beta,\alpha}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha,\beta}=\rho_{k,\beta}&\text{ on }\varphi_{\infty,\beta}(\overline{V_{\alpha}}\cap\overline{V_{\beta}}).\end{cases}$ (IV.62)
 Once these maps $\rho_{k,\alpha}$ are defined, we consider the map $\Psi_{k}\coloneqq\varphi_{k,\alpha}^{-1}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha}$ on $\overline{V_{\alpha}}$.
-This definition is independent of $\alpha$ by (IV.61).
+This definition is independent of $\alpha$ by (IV.62).
 By construction, it holds $\Psi_{k}\rightarrow id$ in $C^{0}(\Sigma,\Sigma)$.
 For $k$ large enough, $\Psi_{k}$ is an immersion, hence surjective (its image is both open and closed).
-As for global injectivity, assume there exist sequences $\{k_{i}\}_{i=1}^{\infty}\subset\mathbb{N}$, $\{p_{i,1}\}_{i=1}^{\infty}$, $\{p_{i,2}\}_{i=1}^{\infty}\subset\Sigma$ such that $\lim_{i\rightarrow\infty}k_{i}=\infty$, $\Psi_{k_{i}}(p_{i,1})=\Psi_{k_{i}}(p_{i,2})$.
-We can without loss of generality assume that $p_{i,1}\rightarrow p_{1}$ and $p_{i,2}\rightarrow p_{2}$ as $i\rightarrow\infty$, but we have $\Psi_{k}\rightarrow id$ in $C^{0}(\Sigma,\Sigma)$ as $k\rightarrow\infty$, hence $p_{1}=p_{2}\in V_{\alpha}$ for some $\alpha$.
+As for global injectivity, assume that there exist sequences $\{k_{i}\}_{i=1}^{\infty}\subset\mathbb{N}$, $\{p_{i,1}\}_{i=1}^{\infty}$, $\{p_{i,2}\}_{i=1}^{\infty}\subset\Sigma$ such that $\lim_{i\rightarrow\infty}k_{i}=\infty$ and $\Psi_{k_{i}}(p_{i,1})=\Psi_{k_{i}}(p_{i,2})$.
+We can without loss of generality assume that $p_{i,1}\rightarrow p_{1}$ and $p_{i,2}\rightarrow p_{2}$ as $i\rightarrow\infty$.
+Since we have $\Psi_{k}\rightarrow id$ in $C^{0}(\Sigma,\Sigma)$ as $k\rightarrow\infty$, we obtain $p_{1}=p_{2}\in V_{\alpha}$ for some $\alpha$.
 For $i$ large enough we get $p_{i,1}=p_{i,2}$ since $\Psi_{k}$ is injective on $V_{\alpha}$.
@@
 Let $\{U_{\alpha}^{\prime}\}_{\alpha=1}^{l}$ be an open cover of $\Sigma$ such that $U_{\alpha}^{\prime}$ has smooth boundary and $U_{\alpha}^{\prime}\Subset U_{\alpha}$ (for existence of such sets, see the proof of [52, Proposition 8.2.1]).
-We claim that there exist open sets $\{V_{\alpha,j}\}_{1\leq\alpha\leq j\leq l}$ (write $V_{\alpha,\alpha-1}=V_{\alpha,\alpha}$), $k_{0}\in\mathbb{N}$, and $C^{\infty}$ injective maps $\rho_{k,\alpha}\colon\varphi_{\infty,\alpha}(\overline{V_{\alpha,\alpha}})\rightarrow\varphi_{k,\alpha}(U_{\alpha})$ for $k\geq k_{0}$ such that for any $1\leq j\leq l$ and $1\leq\alpha,\beta\leq j$, it holds
-$(i)\ V_{\alpha,j}\subset V_{\alpha,j-1}\subset U_{\alpha}^{\prime},$ (IV.62a)
-$(ii)\ \Sigma=\Big(\bigcup\limits_{\gamma\leq j}V_{\gamma,j}\Big)\cup\Big(\bigcup\limits_{\gamma>j}U^{\prime}_{\gamma}\Big),$ (IV.62b)
-$(iii)\ \rho_{k,\alpha}\xrightarrow[k\rightarrow\infty]{}id$ in $C^{\infty}(\varphi_{\infty,\alpha}(\overline{V_{\alpha,\alpha}}),\mathbb{R}^{2})$, (IV.62c)
-$(iv)\ \varphi_{k,\beta,\alpha}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha,\beta}=\rho_{k,\beta}$ on $\varphi_{\infty,\beta}(\overline{V_{\alpha,j}}\cap\overline{V_{\beta,j}})$. (IV.62d)
-The proof is by induction on $j$.
+We claim that there exist open sets $\{V_{\alpha,j}\}_{1\leq\alpha\leq j\leq l}$ (write $V_{\alpha,\alpha-1}=V_{\alpha,\alpha}$), an integer $k_{0}\in\mathbb{N}$, and $C^{\infty}$ injective maps $\rho_{k,\alpha}\colon\varphi_{\infty,\alpha}(\overline{V_{\alpha,\alpha}})\rightarrow\varphi_{k,\alpha}(U_{\alpha})$ for $k\geq k_{0}$ such that for any $1\leq j\leq l$ and $1\leq\alpha,\beta\leq j$, it holds
+$(i)\ V_{\alpha,j}\subset V_{\alpha,j-1}\subset U_{\alpha}^{\prime},$ (IV.63a)
+$(ii)\ \Sigma=\Big(\bigcup\limits_{\gamma\leq j}V_{\gamma,j}\Big)\cup\Big(\bigcup\limits_{\gamma>j}U^{\prime}_{\gamma}\Big),$ (IV.63b)
+$(iii)\ \rho_{k,\alpha}\xrightarrow[k\rightarrow\infty]{}id$ in $C^{\infty}(\varphi_{\infty,\alpha}(\overline{V_{\alpha,\alpha}}),\mathbb{R}^{2})$, (IV.63c)
+$(iv)\ \varphi_{k,\beta,\alpha}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha,\beta}=\rho_{k,\beta}$ on $\varphi_{\infty,\beta}(\overline{V_{\alpha,j}}\cap\overline{V_{\beta,j}})$. (IV.63d)
+We prove this by induction on $j$.
 For $j=1$, we define $V_{1,1}=U_{1}^{\prime}$ and $\rho_{k,1}\colon\varphi_{\infty,1}(\overline{U_{1}^{\prime}})\rightarrow\mathbb{R}^{2}$ as the identity map.
 For $k$ large enough, the image of $\rho_{k,1}$ is contained in $\varphi_{k,1}(U_{1})$ since $\varphi_{k,1}\rightarrow\varphi_{\infty,1}$ in $C(\overline{U_{1}^{\prime}},\mathbb{R}^{2})$ and $U_{1}^{\prime}\Subset U_{1}$.
-Let $1\leq j_{0}\leq l$, suppose that we have constructed $\{V_{\alpha,j}\}_{1\leq\alpha\leq j\leq j_{0}}$ and $\{\rho_{k,\alpha}\}_{1\leq\alpha\leq j_{0}}$ such that (IV.62a)–(IV.62d) hold for any $1\leq j\leq j_{0}$ and $1\leq\alpha,\beta\leq j$.
+Let $1\leq j_{0}\leq l$, suppose that we have constructed $\{V_{\alpha,j}\}_{1\leq\alpha\leq j\leq j_{0}}$ and $\{\rho_{k,\alpha}\}_{1\leq\alpha\leq j_{0}}$ such that (IV.63a)–(IV.63d) hold for any $1\leq j\leq j_{0}$ and $1\leq\alpha,\beta\leq j$.
 We now construct $\rho_{k,j_{0}+1}$ and $\{V_{\alpha,j_{0}+1}\}_{1\leq\alpha\leq j_{0}+1}$ using the induction hypothesis.
@@
 $\omega_{k,j_{0}}\coloneqq\varphi_{k,j_{0}+1,\alpha}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha,j_{0}+1}\quad\text{on $\varphi_{\infty,j_{0}+1}\left(\overline{V_{\alpha,j_{0}}}\cap U_{j_{0}+1}\right)$.}$
-By our induction hypothesis (IV.62d) for $j\leq j_{0}$, this map is well-defined on $\varphi_{\infty,j_{0}+1}\Big(\big(\bigcup\limits_{\alpha\leq j_{0}}\overline{V_{\alpha,j_{0}}}\big)\cap U_{j_{0}+1}\Big)$ and independent of $\alpha$.
+By our induction hypothesis (IV.63d) for $j\leq j_{0}$, this map is well-defined on $\varphi_{\infty,j_{0}+1}\Big(\big(\bigcup\limits_{\alpha\leq j_{0}}\overline{V_{\alpha,j_{0}}}\big)\cap U_{j_{0}+1}\Big)$ and independent of $\alpha$.
 By the assumption $\Sigma=\big(\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}}\big)\cup\big(\bigcup\limits_{\alpha>j_{0}}U^{\prime}_{\alpha}\big)$, we have
@@
 $V_{\alpha,j_{0}+1}\coloneqq V_{\alpha,j_{0}}\cap\,\text{int}\Big\{x\in\Sigma:\xi_{j_{0}}(x)=1\Big\}.$
-The condition (IV.62a) then holds for $j=j_{0}+1$. By the definition of $\{V_{\alpha,j_{0}+1}\}_{\alpha\leq j_{0}+1}$ we have
+The condition (IV.63a) then holds for $j=j_{0}+1$. By the definition of $\{V_{\alpha,j_{0}+1}\}_{\alpha\leq j_{0}+1}$ we have
 $\Big(\bigcup\limits_{\alpha\leq j_{0}+1}V_{\alpha,j_{0}+1}\Big)\cup\Big(\bigcup\limits_{\alpha>j_{0}+1}U^{\prime}_{\alpha}\Big)$ $=\Big(\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}+1}\Big)\cup\Big(\bigcup\limits_{\alpha>j_{0}}U^{\prime}_{\alpha}\Big)$
@@
 $=\Sigma.$
-Hence the condition (IV.62b) holds for $j=j_{0}+1$.
+Hence the condition (IV.63b) holds for $j=j_{0}+1$.
 Since $\xi_{j_{0}}$ vanishes on a neighborhood of $\Sigma\setminus\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}}$, and $\omega_{k,j_{0}}\circ\varphi_{\infty,j_{0}+1}\in C^{\infty}\big(\big(\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}}\big)\cap U_{j_{0}+1}\big)$ we have $(\xi_{j_{0}}\circ\varphi^{-1}_{\infty,j_{0}+1})\,\omega_{k,j_{0}}\in C^{\infty}(\varphi_{\infty,j_{0}+1}(U_{j_{0}+1}))$.
@@
 $\rho_{k,j_{0}+1}=\omega_{k,j_{0}}=\varphi_{k,j_{0}+1,\alpha}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha,j_{0}+1}.$
-Since we have $\overline{V_{j_{0}+1,j_{0}+1}}=\overline{U^{\prime}_{j_{0}+1}}\subset U_{j_{0}+1}$, the condition (IV.62d) is thus satisfied for $\beta=j=j_{0}+1$, $\alpha\leq j_{0}$.
-By using (IV.62a) and the identity $\varphi_{k,\alpha,\beta}=\varphi_{k,\beta,\alpha}^{-1}$, we then obtain (IV.62d) for $j=j_{0}+1$ and any $\alpha,\beta\leq j_{0}+1$.
+Since we have $\overline{V_{j_{0}+1,j_{0}+1}}=\overline{U^{\prime}_{j_{0}+1}}\subset U_{j_{0}+1}$, the condition (IV.63d) is thus satisfied for $\beta=j=j_{0}+1$, $\alpha\leq j_{0}$.
+By using (IV.63a) and the identity $\varphi_{k,\alpha,\beta}=\varphi_{k,\beta,\alpha}^{-1}$, we then obtain (IV.63d) for $j=j_{0}+1$ and any $\alpha,\beta\leq j_{0}+1$.
 Next we prove $\rho_{k,j_{0}+1}\rightarrow id$ in $C^{\infty}(\varphi_{\infty,j_{0}+1}(\overline{U^{\prime}_{j_{0}+1}}),\mathbb{R}^{2})$ as $k\rightarrow\infty$.
 Fixing $\alpha\leq j_{0}$, let $W$ be an open set such that $\varphi_{\infty,\alpha}(\overline{V_{\alpha,j_{0}}}\cap\overline{U_{j_{0}+1}^{\prime}})\subset W\Subset\varphi_{\infty,\alpha}(U_{\alpha}\cap U_{j_{0}+1})$.
-The convergence (IV.60) implies that
+The convergence (IV.61) implies that
 $\varphi_{k,j_{0}+1,\alpha}\xrightarrow[k\rightarrow\infty]{}\varphi_{\infty,j_{0}+1,\alpha}\quad\text{in $C^{\infty}(\overline{W})$. }$
-By our induction hypothesis (IV.62c) for $\alpha\leq j_{0}$, we also have
+By our induction hypothesis (IV.63c) for $\alpha\leq j_{0}$, we also have
 $\rho_{k,\alpha}\xrightarrow[k\rightarrow\infty]{}id\quad\text{in $C^{\infty}(\varphi_{\infty,\alpha}(\overline{V_{\alpha,j_{0}}}))$.}$
 Then the definition of $\omega_{k,j_{0}}$ implies
-$\omega_{k,j_{0}}\xrightarrow[k\rightarrow\infty]{}id\quad\text{in $C^{\infty}(\varphi_{\infty,j_{0}+1}(\overline{V_{\alpha,j_{0}}}\cap\overline{U^{\prime}_{j_{0}+1}})$.}$
+$\omega_{k,j_{0}}\xrightarrow[k\rightarrow\infty]{}id\qquad\text{in $C^{\infty}\left(\varphi_{\infty,j_{0}+1}(\overline{V_{\alpha,j_{0}}}\cap\overline{U^{\prime}_{j_{0}+1}})\right)$.}$
 Hence we have
-$\rho_{k,j_{0}+1}\xrightarrow[k\rightarrow\infty]{}id\quad\text{in $C^{\infty}\Big(\varphi_{\infty,j_{0}+1}\Big(\Big(\bigcup_{\alpha\leq j_{0}}\overline{V_{\alpha,j_{0}}}\Big)\cap\overline{U^{\prime}_{j_{0}+1}}\Big)\Big)$.}$
+$\rho_{k,j_{0}+1}\xrightarrow[k\rightarrow\infty]{}id\qquad\text{in $C^{\infty}\Big(\varphi_{\infty,j_{0}+1}\Big(\Big(\bigcup_{\alpha\leq j_{0}}\overline{V_{\alpha,j_{0}}}\Big)\cap\overline{U^{\prime}_{j_{0}+1}}\Big)\Big)$.}$
 On the other hand, it holds that $\rho_{k,j_{0}+1}=id$ on $\varphi_{\infty,j_{0}+1}\big(U_{j_{0}+1}\setminus\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}}\big)$. Therefore, we have
 $\rho_{k,j_{0}+1}\xrightarrow[k\rightarrow\infty]{}id\quad\text{in $C^{\infty}\big(\varphi_{\infty,j_{0}+1}(\overline{U^{\prime}_{j_{0}+1}})\big)$.}$
-Hence the condition (IV.62c) holds for $j=j_{0}+1$.
+Hence the condition (IV.63c) holds for $j=j_{0}+1$.
 In particular, $\rho_{k,j_{0}+1}$ is an immersion on $\varphi_{\infty,j_{0}+1}(\overline{V_{j_{0}+1,j_{0}+1}})=\varphi_{\infty,j_{0}+1}(\overline{U^{\prime}_{j_{0}+1}})$ for $k$ large enough.
 Also, since $\partial U^{\prime}_{j_{0}+1}$ is smooth, we have the following embedding (see [58, Theorem 5.8.4]):
-$W^{1,\infty}(\varphi_{\infty,j_{0}+1}(U^{\prime}_{j_{0}+1}))\subset C^{0,1}(\varphi_{\infty,j_{0}+1}(\overline{U^{\prime}_{j_{0}+1}})).$
+$W^{1,\infty}\left(\varphi_{\infty,j_{0}+1}(U^{\prime}_{j_{0}+1})\right)\subset C^{0,1}\left(\varphi_{\infty,j_{0}+1}(\overline{U^{\prime}_{j_{0}+1}})\right).$
 Consequently, there exists $k_{0}\in\mathbb{N}$ such that for any $k\geq k_{0}$ and $x,y\in\varphi_{\infty,j_{0}+1}(\overline{U^{\prime}_{j_{0}+1}})$ it holds
@@
 This completes the induction. Finally, for $1\leq\alpha\leq l$, we take $V_{\alpha}\coloneqq V_{\alpha,l}$, the constructions of $\{\rho_{k,\alpha}\}$ and $\{V_{\alpha}\}$ are thus accomplished.
-Now since $\Sigma=\bigcup\limits_{1\leq\alpha\leq l}V_{\alpha}$, we can choose smooth non-negative cut-off functions $\phi_{\alpha}\in C^{\infty}_{c}(\varphi_{\infty,\alpha}(V_{\alpha}))$ such that $\sum\limits_{\alpha=1}^{l}\phi_{\alpha}\circ\varphi_{\infty,\alpha}>0$ on $\Sigma$.
+Since $\Sigma=\bigcup\limits_{1\leq\alpha\leq l}V_{\alpha}$, we can choose smooth non-negative cut-off functions $\phi_{\alpha}\in C^{\infty}_{c}(\varphi_{\infty,\alpha}(V_{\alpha}))$ such that $\sum\limits_{\alpha=1}^{l}\phi_{\alpha}\circ\varphi_{\infty,\alpha}>0$ on $\Sigma$.
 Since $\varphi_{k,\alpha}\rightarrow\varphi_{\infty,\alpha}$ in $C^{0}(\overline{V_{\alpha}})$ as $k\rightarrow\infty$, we have $\phi_{\alpha}\in C_{c}^{\infty}(\varphi_{k,\alpha}(V_{\alpha}))$ for $k\in\mathbb{N}\cup\{\infty\}$ large enough.
@@
 Hence $\varphi_{\infty,\alpha}$ is smooth by Corollary IV.5 (i).
-Let $1\leq\alpha\leq l$, $x\in V_{\alpha}$, then by (IV.60) and (IV.61) we have
+Let $1\leq\alpha\leq l$, $x\in V_{\alpha}$, then by (IV.61) and (IV.62) we have
 $(\Psi_{k}^{*}g^{\prime}_{k})_{x}$ $=\Big((\varphi_{k,\alpha}^{-1}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha})^{*}\sum_{x\in V_{\beta}}\varphi_{k,\beta}^{*}(\phi_{\beta}\,g_{\text{std}})\Big)_{x}$
@@
 Then we can construct the conformal charts $\{\varphi_{k,\alpha}\}$ as in Remark IV.6, and by using the same argument as in Theorem IV.19, we find $\{h_{k}\}$, $\{\Psi_{k}\}$ satisfying Theorem IV.19 (i), (ii), (iv), but $\Psi_{k}$ may not be Lipschitz.
-Indeed, by [90, Proposition 4.36] and [91], we have $\varphi_{k,\alpha}\rightarrow\varphi_{\infty,\alpha}$, $\varphi^{-1}_{k,\alpha}\rightarrow\varphi^{-1}_{\infty,\alpha}$ in $W^{1,p}_{\textup{loc}}$ for some $p>2$, hence $\Psi_{k}\rightarrow id$, $\Psi^{-1}_{k}\rightarrow id$ in $W^{1,p}(\Sigma,\Sigma)$, in particular, in $C^{0}(\Sigma,\Sigma)$.
+Indeed, by [91, Proposition 4.36] and [92], we have $\varphi_{k,\alpha}\rightarrow\varphi_{\infty,\alpha}$, $\varphi^{-1}_{k,\alpha}\rightarrow\varphi^{-1}_{\infty,\alpha}$ in $W^{1,p}_{\textup{loc}}$ for some $p>2$, hence $\Psi_{k}\rightarrow id$, $\Psi^{-1}_{k}\rightarrow id$ in $W^{1,p}(\Sigma,\Sigma)$, in particular, in $C^{0}(\Sigma,\Sigma)$.
 IV.4.3 Approximation results
@@
 Then we prove Theorem I.3 at the end.
-We define a weak $W^{1,\infty}$-immersion with value into manifolds.
+We define a weak $W^{1,\infty}$-immersion taking values in manifolds.
 Definition IV.21.
@@
 We say $f\in W^{1,\infty}_{\textup{imm}}(U,M)$ if $f\in W^{1,\infty}_{\textup{imm}}(U,\mathbb{R}^{m})$ and $f(x)\in M$ for any $x\in U$.
-We estimate the mean oscillation of a composition and the distance with a smooth approximation.
+We estimate the mean oscillation of a composition and the distance to a smooth approximation.
 Lemma IV.22.
@@
 (i) For any $r>0$, we have $\nabla(f\circ\varphi^{-1})\in VMO(W)$, and there exists a positive constant $C=C(n,\|\nabla\varphi\|_{L^{\infty}(U)},\|\nabla\varphi^{-1}\|_{L^{\infty}(W)})$ such that
-$\beta_{r}(\nabla(f\circ\varphi^{-1}))\leq C\left(\beta_{r}(\nabla f)+\beta_{r}(\nabla\varphi)\|\nabla f\|_{L^{\infty}(U)}\right).$ (IV.63)
+$\beta_{r}(\nabla(f\circ\varphi^{-1}))\leq C\left(\beta_{r}(\nabla f)+\beta_{r}(\nabla\varphi)\|\nabla f\|_{L^{\infty}(U)}\right).$ (IV.64)
 (ii) Let $\eta$ be the standard mollifier.
@@
 Then there exists a positive constant $C=C(\|\nabla\varphi\|_{L^{\infty}(U)},\|\nabla\varphi^{-1}\|_{L^{\infty}(W)},U,K)>0$ such that for any $x_{1},x_{0}\in U$ and $\varepsilon>0$ satisfying $|x_{1}-x_{0}|\leq K\varepsilon$ and $\varepsilon<\min\big(\operatorname{dist}(\varphi(x_{1}),\partial W),\operatorname{dist}(x_{0},\partial U)\big)$, there holds
-$\begin{aligned} &\left|\nabla(f_{\varepsilon}\circ\varphi^{-1})(\varphi(x_{0}))-\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla(f\circ\varphi^{-1})\right|\\[2.84526pt] &\leq C\Big(\beta_{\varepsilon}(\nabla f)+\big(\omega_{\varepsilon}(\nabla\varphi)+\varepsilon\big)\|\nabla f\|_{L^{\infty}(U)}\Big),\end{aligned}$ (IV.64)
+$\begin{aligned} &\left|\nabla(f_{\varepsilon}\circ\varphi^{-1})(\varphi(x_{0}))-\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla(f\circ\varphi^{-1})\right|\\[5.69054pt] &\leq C\Big(\beta_{\varepsilon}(\nabla f)+\big(\omega_{\varepsilon}(\nabla\varphi)+\varepsilon\big)\|\nabla f\|_{L^{\infty}(U)}\Big),\end{aligned}$ (IV.65)
 where $\omega_{\varepsilon}(\nabla\varphi)\coloneqq\sup\{|\nabla\varphi(x)-\nabla\varphi(y)|:x,y\in U,|x-y|\leq\varepsilon\}$.
@@
 $\leq C\big(\beta_{r}(\nabla f)+\beta_{r}(\nabla\varphi)\|\nabla f\|_{L^{\infty}(U)}\big),$
-The estimate (IV.63) is thus proved.
+The estimate (IV.64) is thus proved.
 (ii) Let $y\in B_{\varepsilon}(\varphi(x_{1}))$, we have
 $|\varphi^{-1}(y)-x_{1}|\leq\|\nabla\varphi^{-1}\|_{L^{\infty}(W)}\,|y-\varphi(x_{1})|\leq\varepsilon\,\|\nabla\varphi^{-1}\|_{L^{\infty}(W)}.$
-Hence it holds that $|x_{0}-\varphi^{-1}(y)|\leq\varepsilon\,(K+\|\nabla\varphi^{-1}\|_{L^{\infty}(W)})$. Note that $(\nabla\varphi)^{-1}\in C^{0}(\overline{U})$ and for all $r>0$, we have
+Hence it holds that $|x_{0}-\varphi^{-1}(y)|\leq\varepsilon\,(K+\|\nabla\varphi^{-1}\|_{L^{\infty}(W)})$. Moreover it holds $(\nabla\varphi)^{-1}\in C^{0}(\overline{U})$ and for all $r>0$, we have
 $\omega_{r}\big((\nabla\varphi)^{-1}\big)\leq C\,\omega_{r}(\nabla\varphi).$
@@
 Hence we have
-$\begin{aligned} &\left|\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla(f\circ\varphi^{-1})-\nabla\varphi^{-1}(\varphi(x_{0}))\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}\right|\\[5.69054pt] &\leq C\big(\omega_{\varepsilon}(\nabla\varphi)+\varepsilon\big)\,\|\nabla f\|_{L^{\infty}(U)}.\end{aligned}$ (IV.65)
+$\begin{aligned} &\left|\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla(f\circ\varphi^{-1})-\nabla\varphi^{-1}(\varphi(x_{0}))\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}\right|\\[5.69054pt] &\leq C\big(\omega_{\varepsilon}(\nabla\varphi)+\varepsilon\big)\,\|\nabla f\|_{L^{\infty}(U)}.\end{aligned}$ (IV.66)
 Since $\operatorname{dist}(\varphi(x_{1}),\partial W)>\varepsilon$, we have $\operatorname{dist}(x_{1},\partial U)>\varepsilon\,\|\nabla\varphi\|^{-1}_{L^{\infty}(U)}$.
-Now we can apply the proof of Riemann’s Theorem [162, Theorem 2] and Lemma III.13 to get
-$\bigg|\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}-\fint_{B_{\varepsilon\,\|\nabla\varphi\|^{-1}_{L^{\infty}}}(x_{1})}\nabla f\bigg|\leq C\,\beta_{\varepsilon}(\nabla f).$ (IV.66)
+Now we can apply the proof of Riemann’s Theorem [165, Theorem 2] and Lemma III.13 to get
+$\bigg|\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}-\fint_{B_{\varepsilon\,\|\nabla\varphi\|^{-1}_{L^{\infty}}}(x_{1})}\nabla f\bigg|\leq C\,\beta_{\varepsilon}(\nabla f).$ (IV.67)
 Since $\partial U$ is Lipschitz, by Corollary III.15 we can extend $\nabla f$ to a function $\tilde{f}\in VMO(\mathbb{R}^{n},\mathbb{R}^{n})$ with
@@
 Hence it holds that
-$\begin{aligned} \bigg|\fint_{B_{\varepsilon\,\|\nabla\varphi\|^{-1}_{L^{\infty}}}(x_{1})}\nabla f-\fint_{B_{\varepsilon}(x_{0})}\nabla f\bigg|\leq C\,\beta_{\varepsilon}(\tilde{f})\leq C\big(\beta_{\varepsilon}(\nabla f)+\varepsilon\,\|\nabla f\|_{L^{\infty}(U)}\big).\end{aligned}$ (IV.67)
+$\begin{aligned} \bigg|\fint_{B_{\varepsilon\,\|\nabla\varphi\|^{-1}_{L^{\infty}}}(x_{1})}\nabla f-\fint_{B_{\varepsilon}(x_{0})}\nabla f\bigg|\leq C\,\beta_{\varepsilon}(\tilde{f})\leq C\big(\beta_{\varepsilon}(\nabla f)+\varepsilon\,\|\nabla f\|_{L^{\infty}(U)}\big).\end{aligned}$ (IV.68)
 We also have that
-$\begin{aligned} \left|\nabla f_{\varepsilon}(x_{0})-\fint_{B_{\varepsilon}(x_{0})}\nabla f\right|&=\left|\int_{B_{\varepsilon}(x_{0})}\eta_{\varepsilon}(x_{0}-x)\,\left(\nabla f(x)-\fint_{B_{\varepsilon}(x_{0})}\nabla f\,dx\right)\right|\\[2.84526pt] &\leq C\,\fint_{B_{\varepsilon}(x_{0})}\left|\nabla f-\fint_{B_{\varepsilon}(x_{0})}\nabla f\right|\\[2.84526pt] &\leq C\,\beta_{\varepsilon}(\nabla f).\end{aligned}$ (IV.68)
-Combining (IV.66), (IV.67), (IV.68) we obtain by the triangle inequality
-$\begin{aligned} &\left|\nabla\varphi^{-1}(\varphi(x_{0}))\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}-\nabla(f_{\varepsilon}\circ\varphi^{-1})(\varphi(x_{0}))\right|\\[2.84526pt] &\leq\|\nabla\varphi^{-1}\|_{L^{\infty}}\left|\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}-\nabla f_{\varepsilon}(x_{0})\right|\\[2.84526pt] &\leq C\big(\beta_{\varepsilon}(\nabla f)+\varepsilon\,\|\nabla f\|_{L^{\infty}(U)}\big).\end{aligned}$ (IV.69)
+$\begin{aligned} \left|\nabla f_{\varepsilon}(x_{0})-\fint_{B_{\varepsilon}(x_{0})}\nabla f\right|&=\left|\int_{B_{\varepsilon}(x_{0})}\eta_{\varepsilon}(x_{0}-x)\,\left(\nabla f(x)-\fint_{B_{\varepsilon}(x_{0})}\nabla f\,dx\right)\right|\\[2.84526pt] &\leq C\,\fint_{B_{\varepsilon}(x_{0})}\left|\nabla f-\fint_{B_{\varepsilon}(x_{0})}\nabla f\right|\\[2.84526pt] &\leq C\,\beta_{\varepsilon}(\nabla f).\end{aligned}$ (IV.69)
+Combining (IV.67), (IV.68), (IV.69) we obtain by the triangle inequality
+$\begin{aligned} &\left|\nabla\varphi^{-1}(\varphi(x_{0}))\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}-\nabla(f_{\varepsilon}\circ\varphi^{-1})(\varphi(x_{0}))\right|\\[2.84526pt] &\leq\|\nabla\varphi^{-1}\|_{L^{\infty}}\left|\fint_{B_{\varepsilon}(\varphi(x_{1}))}\nabla f\circ\varphi^{-1}-\nabla f_{\varepsilon}(x_{0})\right|\\[2.84526pt] &\leq C\big(\beta_{\varepsilon}(\nabla f)+\varepsilon\,\|\nabla f\|_{L^{\infty}(U)}\big).\end{aligned}$ (IV.70)
 Finally, (LABEL:vmocov) follows from (LABEL:dvpdiff) and (LABEL:modiff4). ∎
 We are now ready to establish an approximation result required for the proof of Theorem I.3.
-The underlying idea goes back to [178], see also [100, Remark 2.1].
+The underlying idea goes back to [181], see also [103, Remark 2.1] for the $W^{2,2}$ case in dimension $2$.
 For completeness, we present a detailed proof.
@@
 […]
-$\Lambda^{-1}|v|_{\tilde{g}}\leq|d(f_{k})(v)|_{\mathbb{R}^{m}}\leq\Lambda|v|_{\tilde{g}},\qquad\text{for all }\,v\in TU\,\text{ and all }\,k\in\mathbb{N}.$ (IV.70)
+$\Lambda^{-1}|v|_{\tilde{g}}\leq|d(f_{k})(v)|_{\mathbb{R}^{m}}\leq\Lambda|v|_{\tilde{g}},\qquad\text{for all }\,v\in TU\,\text{ and all }\,k\in\mathbb{N}.$ (IV.71)
 Moreover, if $f\in W^{2,p}(U)$ for any $1\leq p<\infty$, $f_{k}$ can be chosen to strongly converge to $f$ in $W^{2,p}(U)$.
@@
 Since $f(U)$ is compact, there exists a precompact neighborhood $W^{\prime}$ of $f(U)$ strictly contained in $W$.
-Since $\overline{U}$ is compact, there exist coordinate systems $\{(U_{j},\phi_{j})\}_{j=1}^{s}$ of $N$ such that $\overline{U}\subset\bigcup_{j=1}^{s}U_{j}$.
+Since $\overline{U}$ is compact, there exist finitely many coordinate charts $\{(U_{j},\phi_{j})\}_{j=1}^{s}$ of $N$ such that $\overline{U}\subset\bigcup_{j=1}^{s}U_{j}$.
 Since $\partial U$ is Lipschitz, we can further assume that for each $U_{j}\nsubset U$, we have $\phi_{j}(U_{j})=B^{n}\subset\mathbb{R}^{n}$, $0\in\phi_{j}(\partial U)$ and, for some Lipschitz map $\gamma_{j}\colon\mathbb{R}^{n-1}\rightarrow\mathbb{R}$,
@@
 Let
-$0<\varepsilon<\min_{1\leq j\leq s}\left[\frac{1}{K_{j}+1}\min\Big(\text{dist}\big(\phi_{j}(\operatorname{supp}(\xi_{j})),\phi_{j}(\partial V_{j})\big),\text{dist}\big(\phi_{j}(V_{j}),\phi_{j}(U_{j})^{c}\big)\Big)\right].$ (IV.71)
+$0<\varepsilon<\min_{1\leq j\leq s}\left[\frac{1}{K_{j}+1}\min\Big(\text{dist}\big(\phi_{j}(\operatorname{supp}(\xi_{j})),\phi_{j}(\partial V_{j})\big),\text{dist}\big(\phi_{j}(V_{j}),\phi_{j}(U_{j})^{c}\big)\Big)\right].$ (IV.72)
 Define $\eta_{\varepsilon}$ as in Lemma IV.22 (ii). For $1\leq j\leq s$, if $U_{j}\subset U$, for $x\in V_{j}$ we define
@@
 $B_{\varepsilon}(z+K_{j}\,\varepsilon\,e_{n})\cap\phi_{j}(U_{j}\cap\partial U)=\emptyset.$
-By choice of $\varepsilon$ in (IV.71), it holds that
+By choice of $\varepsilon$ in (IV.72), it holds that
 $B_{\varepsilon}(z+K_{j}\,\varepsilon\,e_{n})\subset\phi_{j}(U_{j}\cap U).$
@@
 Now it suffices to prove that there exists a constant $\Lambda>0$ such that for all $\varepsilon$ small enough, we have
-$\forall v\in T\Sigma,\qquad\Lambda^{-1}|v|_{\tilde{g}}\leq|d(f_{\varepsilon})(v)|_{\mathbb{R}^{m}}\leq\Lambda|v|_{\tilde{g}}$. (IV.72)
+$\forall v\in T\Sigma,\qquad\Lambda^{-1}|v|_{\tilde{g}}\leq|d(f_{\varepsilon})(v)|_{\mathbb{R}^{m}}\leq\Lambda|v|_{\tilde{g}}$. (IV.73)
 Let $1\leq j,j^{\prime}\leq s$, since $\overline{V}_{j}\cap\overline{V}_{j^{\prime}}\subset U_{j}\cap U_{j^{\prime}}$, we can find a pre-compact region $U_{j,j^{\prime}}$ with smooth boundary (see [52, Proposition 8.2.1]) such that
@@
 Consequently, it holds that
-$\begin{aligned} &\left|\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-\fint_{B_{\varepsilon}(z^{\varepsilon})}\nabla(f\circ\phi_{j}^{-1})\right|\\ &\leq\sum_{j^{\prime}=1}^{s}\left|\nabla(F_{j^{\prime},\varepsilon}\circ\phi_{j}^{-1})(z)-\fint_{B_{\varepsilon}(z^{\varepsilon})}\nabla(\tilde{f}_{j^{\prime}}\circ\phi_{j}^{-1})\right|\\ &=\underset{\varepsilon\to 0}{o}(1).\end{aligned}$ (IV.73)
+$\begin{aligned} &\left|\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-\fint_{B_{\varepsilon}(z^{\varepsilon})}\nabla(f\circ\phi_{j}^{-1})\right|\\ &\leq\sum_{j^{\prime}=1}^{s}\left|\nabla(F_{j^{\prime},\varepsilon}\circ\phi_{j}^{-1})(z)-\fint_{B_{\varepsilon}(z^{\varepsilon})}\nabla(\tilde{f}_{j^{\prime}}\circ\phi_{j}^{-1})\right|\\ &=\underset{\varepsilon\to 0}{o}(1).\end{aligned}$ (IV.74)
 Since $df\in\text{VMO}(U)$, we also have
@@
 Together with (LABEL:osdiff), then we get the following estimate where the right-hand side is independent of $z\in\phi_{j}(V_{j}\cap\overline{U})$:
-$\fint_{B_{\varepsilon}(z^{\varepsilon})}\left|\nabla(f\circ\phi_{j}^{-1})-\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)\right|=\underset{\varepsilon\to 0}{o}(1).$ (IV.74)
+$\fint_{B_{\varepsilon}(z^{\varepsilon})}\left|\nabla(f\circ\phi_{j}^{-1})-\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)\right|=\underset{\varepsilon\to 0}{o}(1).$ (IV.75)
 Since $f\in W_{\textup{imm}}^{1,\infty}(U,M)$, there exists a constant $C_{j}>0$ such that for a.e. $z\in\phi_{j}(V_{j})$ and for any $v\in\mathbb{S}^{n-1}$ it holds
-$\big|d(f\circ\phi_{j}^{-1})_{z}(v)\big|\geq C_{j}.$ (IV.75)
+$\big|d(f\circ\phi_{j}^{-1})_{z}(v)\big|\geq C_{j}.$ (IV.76)
 Since $\overline{W^{\prime}}$ is compact and $\pi$ is smooth in an open set containing $\overline{W^{\prime}}$, the map $\nabla\pi\colon\overline{W^{\prime}}\rightarrow\mathbb{R}^{k\times k}$ is Lipschitz.
 We define
-$\alpha\coloneqq\|\nabla\pi\|_{C^{0}(\overline{W^{\prime}})}+1+\text{Lip}(\nabla\pi|_{\overline{W^{\prime}}})\,\|\nabla(f\circ\phi_{j}^{-1})\|_{L^{\infty}(\phi_{j}(V_{j}\cap\overline{U}))}.$ (IV.76)
-By (IV.74), there exists $\varepsilon_{0}>0$ such that for any $\varepsilon\in(0,\varepsilon_{0})$, we have on the one hand that $F_{\varepsilon}\in C^{\infty}(\overline{U},M)$ is well-defined with $F_{\varepsilon}(\overline{U})\subset W^{\prime}$.
+$\alpha\coloneqq\|\nabla\pi\|_{C^{0}(\overline{W^{\prime}})}+1+\text{Lip}(\nabla\pi|_{\overline{W^{\prime}}})\,\|\nabla(f\circ\phi_{j}^{-1})\|_{L^{\infty}(\phi_{j}(V_{j}\cap\overline{U}))}.$ (IV.77)
+By (IV.75), there exists $\varepsilon_{0}>0$ such that for any $\varepsilon\in(0,\varepsilon_{0})$, we have on the one hand that $F_{\varepsilon}\in C^{\infty}(\overline{U},M)$ is well-defined with $F_{\varepsilon}(\overline{U})\subset W^{\prime}$.
 On the other hand for any $z\in\phi_{j}(\operatorname{supp}(\xi_{j})\cap\overline{U})$, there holds $B_{\varepsilon}(z^{\varepsilon})\subset\phi_{j}(V_{j}\cap\overline{U})$, and there exists $z^{\prime}\in B_{\varepsilon}(z^{\varepsilon})$ satisfying $f\circ\phi_{j}^{-1}$ is differentiable at $z^{\prime}$ with the following inequalities:
   • oscillation of $\nabla(f\circ\phi_{j}^{-1})$:
-    $\big|\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-\nabla(f\circ\phi_{j}^{-1})(z^{\prime})\big|\leq(2\alpha)^{-1}C_{j}.$ (IV.77)
+    $\big|\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-\nabla(f\circ\phi_{j}^{-1})(z^{\prime})\big|\leq(2\alpha)^{-1}C_{j}.$ (IV.78)
   • $C^{0}$ distance:
-    $\begin{aligned} &\big|(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z^{\prime})\big|\\[5.69054pt] &\leq\big|(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z)\big|+\big|(f\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z^{\prime})\big|\\[5.69054pt] &\leq(2\alpha)^{-1}C_{j}.\end{aligned}$ (IV.78)
+    $\begin{aligned} &\big|(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z^{\prime})\big|\\[5.69054pt] &\leq\big|(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z)\big|+\big|(f\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z^{\prime})\big|\\[5.69054pt] &\leq(2\alpha)^{-1}C_{j}.\end{aligned}$ (IV.79)
 The last inequality follows from the facts that $f\circ\phi_{j}^{-1}$ is Lipschitz continuous on $\phi_{j}(V_{j}\cap U)$ and $\|F_{\varepsilon}-f\|_{C(\overline{U})}\rightarrow 0$ as $\varepsilon\rightarrow 0$.
@@
 Hence, it holds for any $z\in\phi_{j}(\operatorname{supp}(\xi_{j})\cap\overline{U})$, $0<\varepsilon<\varepsilon_{0}$, $v\in\mathbb{S}^{n-1}$ that
-$\begin{aligned} &\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)\big|\\[5.69054pt] &\geq\ \big|d(f\circ\phi_{j}^{-1})_{z^{\prime}}(v)\big|-\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)-\big(d\pi\circ d(f\circ\phi_{j}^{-1})\big)_{z^{\prime}}(v)\big|.\end{aligned}$ (IV.79)
+$\begin{aligned} &\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)\big|\\[5.69054pt] &\geq\ \big|d(f\circ\phi_{j}^{-1})_{z^{\prime}}(v)\big|-\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)-\big(d\pi\circ d(f\circ\phi_{j}^{-1})\big)_{z^{\prime}}(v)\big|.\end{aligned}$ (IV.80)
 By the triangle inequality, we have for any $v\in\mathbb{S}^{n-1}$
-$\begin{aligned} &\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)-\big(d\pi\circ d(f\circ\phi_{j}^{-1})\big)_{z^{\prime}}(v)\big|\\[8.53581pt] &\leq\big|\nabla\pi_{(F_{\varepsilon}\circ\phi_{j}^{-1})(z)}\big(\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})_{z}-\nabla(f\circ\phi_{j}^{-1})_{z^{\prime}}\big)(v)\big|\\[7.96677pt] &\quad\quad+\big|\big(\nabla\pi_{(F_{\varepsilon}\circ\phi_{j}^{-1})(z)}-\nabla\pi_{(f\circ\phi_{j}^{-1})(z^{\prime})}\big)\nabla(f\circ\phi_{j}^{-1})_{z^{\prime}}(v)\big|\\[8.53581pt] &\leq\big|\nabla\pi\big((F_{\varepsilon}\circ\phi_{j}^{-1})(z)\big)\big|\ \big|\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-\nabla(f\circ\phi_{j}^{-1})(z^{\prime})\big|\\[9.95845pt] &\quad\quad+\text{Lip}(\nabla\pi|_{\overline{W^{\prime}}})\ \big|(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z^{\prime})\big|\ \big|\nabla(f\circ\phi_{j}^{-1})(z^{\prime})\big|.\end{aligned}$ (IV.80)
-Combining (IV.76)–(LABEL:c0dis) and (LABEL:dpitri), we obtain
-$\begin{aligned} &\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)-\big(d\pi\circ d(f\circ\phi_{j}^{-1})\big)_{z^{\prime}}(v)\big|\\[7.11317pt] &\leq(2\alpha)^{-1}C_{j}\Big(\|\nabla\pi\|_{C(\overline{W^{\prime}})}+\text{Lip}(\nabla\pi|_{\overline{W^{\prime}}})\,\|\nabla(f\circ\phi_{j}^{-1})\|_{L^{\infty}(\phi_{j}(V_{j}\cap\overline{U}))}\Big)\\[7.11317pt] &\leq 2^{-1}C_{j}.\end{aligned}$ (IV.81)
-By (IV.75), (LABEL:dpi>), (LABEL:<c_j) we finally conclude that for any $z\in\phi_{j}(\operatorname{supp}(\xi_{j})\cap\overline{U})$, $0<\varepsilon<\varepsilon_{0}$, $v\in\mathbb{S}^{n-1}$, it holds
+$\begin{aligned} &\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)-\big(d\pi\circ d(f\circ\phi_{j}^{-1})\big)_{z^{\prime}}(v)\big|\\[8.53581pt] &\leq\big|\nabla\pi_{(F_{\varepsilon}\circ\phi_{j}^{-1})(z)}\big(\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})_{z}-\nabla(f\circ\phi_{j}^{-1})_{z^{\prime}}\big)(v)\big|\\[7.96677pt] &\quad\quad+\big|\big(\nabla\pi_{(F_{\varepsilon}\circ\phi_{j}^{-1})(z)}-\nabla\pi_{(f\circ\phi_{j}^{-1})(z^{\prime})}\big)\nabla(f\circ\phi_{j}^{-1})_{z^{\prime}}(v)\big|\\[8.53581pt] &\leq\big|\nabla\pi\big((F_{\varepsilon}\circ\phi_{j}^{-1})(z)\big)\big|\ \big|\nabla(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-\nabla(f\circ\phi_{j}^{-1})(z^{\prime})\big|\\[9.95845pt] &\quad\quad+\text{Lip}(\nabla\pi|_{\overline{W^{\prime}}})\ \big|(F_{\varepsilon}\circ\phi_{j}^{-1})(z)-(f\circ\phi_{j}^{-1})(z^{\prime})\big|\ \big|\nabla(f\circ\phi_{j}^{-1})(z^{\prime})\big|.\end{aligned}$ (IV.81)
+Combining (IV.77)–(LABEL:c0dis) and (LABEL:dpitri), we obtain
+$\begin{aligned} &\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)-\big(d\pi\circ d(f\circ\phi_{j}^{-1})\big)_{z^{\prime}}(v)\big|\\[7.11317pt] &\leq(2\alpha)^{-1}C_{j}\Big(\|\nabla\pi\|_{C(\overline{W^{\prime}})}+\text{Lip}(\nabla\pi|_{\overline{W^{\prime}}})\,\|\nabla(f\circ\phi_{j}^{-1})\|_{L^{\infty}(\phi_{j}(V_{j}\cap\overline{U}))}\Big)\\[7.11317pt] &\leq 2^{-1}C_{j}.\end{aligned}$ (IV.82)
+By (IV.76), (LABEL:dpi>), (LABEL:<c_j) we finally conclude that for any $z\in\phi_{j}(\operatorname{supp}(\xi_{j})\cap\overline{U})$, $0<\varepsilon<\varepsilon_{0}$, $v\in\mathbb{S}^{n-1}$, it holds
 $\big|\big(d\pi\circ d(F_{\varepsilon}\circ\phi_{j}^{-1})\big)_{z}(v)\big|\geq 2^{-1}C_{j}.$
-The estimate (IV.72) is obtained since $\overline{U}\subset\bigcup_{j=1}^{s}\operatorname{supp}(\xi_{j})$. ∎
+The estimate (IV.73) is obtained since $\overline{U}\subset\bigcup_{j=1}^{s}\operatorname{supp}(\xi_{j})$. ∎
 We are now ready to prove Theorem I.3.
@@
 By Theorem IV.23, there exists a sequence of $C^{\infty}$ immersions $\vec{\Phi}_{k}$ satisfying $\vec{\Phi}_{k}\rightarrow\vec{\Phi}$ in $W^{2,2}(\Sigma,\mathbb{R}^{m})$ and there exists a constant $\Lambda>0$ such that
-$\Lambda^{-1}|X|_{\tilde{g}}\leq|d(\vec{\Phi}_{k})(X)|\leq\Lambda|X|_{\tilde{g}},\qquad\text{for all }X\in T\Sigma\,\text{ and all }\,k\in\mathbb{N}.$ (IV.82)
+$\Lambda^{-1}|X|_{\tilde{g}}\leq|d(\vec{\Phi}_{k})(X)|\leq\Lambda|X|_{\tilde{g}},\qquad\text{for all }X\in T\Sigma\,\text{ and all }\,k\in\mathbb{N}.$ (IV.83)
 Write $\vec{\Phi}_{\infty}\coloneqq\vec{\Phi}$, and we construct $\{\Psi_{k}\}$, $\{h_{k}\}$ as in Theorem IV.19.
@@
 Since $h_{k}^{\prime}$ is conformal to $g_{k}^{\prime}$, we also obtain that $\vec{\Phi}_{k}^{\prime}\circ\varphi_{k}^{-1}$ is conformal.
-Write
+We denote
 $(\varphi_{k}^{-1})^{*}h_{k}^{\prime}=e^{2\tilde{\lambda}_{k}}(dx_{1}^{2}+dx_{2}^{2}),\quad(\varphi_{k}^{-1})^{*}g_{k}^{\prime}=(\vec{\Phi}_{k}^{\prime}\circ\varphi^{-1}_{k})^{*}g_{\text{std}}=e^{2\lambda_{k}}(dx_{1}^{2}+dx_{2}^{2}).$
-Then $\tilde{\lambda}_{k}\rightarrow\tilde{\lambda}_{\infty}$ in $C^{\infty}(\overline{D^{2}})$ since $\varphi_{k}^{-1}\rightarrow\varphi_{\infty}^{-1}$ in $C^{\infty}(\overline{D^{2}})$ and $h_{k}^{\prime}\rightarrow h_{\infty}^{\prime}=h$ in $C^{\infty}(\Sigma)$.
-Since $\{\lambda_{k}\}$ is bounded in $L^{\infty}(D^{2})$ by (IV.82) and Theorem IV.19 (v), and $\vec{\Phi}_{k}^{\prime}\circ\varphi_{k}^{-1}\rightarrow\vec{\Phi}\circ\varphi_{\infty}^{-1}$ in $W^{2,2}(D^{2},\mathbb{R}^{m})$, we have $\lambda_{k}\rightarrow\lambda_{\infty}$ in $W^{1,2}(D^{2})$.
+Then we have $\tilde{\lambda}_{k}\rightarrow\tilde{\lambda}_{\infty}$ in $C^{\infty}(\overline{D^{2}})$ since it holds $\varphi_{k}^{-1}\rightarrow\varphi_{\infty}^{-1}$ in $C^{\infty}(\overline{D^{2}})$ and $h_{k}^{\prime}\rightarrow h_{\infty}^{\prime}=h$ in $C^{\infty}(\Sigma)$.
+Since the sequence $\{\lambda_{k}\}$ is bounded in $L^{\infty}(D^{2})$ by (IV.83) and Theorem IV.19 (v), and $\vec{\Phi}_{k}^{\prime}\circ\varphi_{k}^{-1}\rightarrow\vec{\Phi}\circ\varphi_{\infty}^{-1}$ in $W^{2,2}(D^{2},\mathbb{R}^{m})$, we have $\lambda_{k}\rightarrow\lambda_{\infty}$ in $W^{1,2}(D^{2})$.
 Now by applying the argument in Corollary III.11 to $\{\vec{\Phi}_{k}^{\prime}\circ\varphi_{k}^{-1}\}$, we have $\lambda_{k}\rightarrow\lambda_{\infty}$ in $W^{2,1}_{\textup{loc}}(D^{2})$.
-Since $\alpha_{k}=(\lambda_{k}-\tilde{\lambda}_{k})\circ\varphi_{k}$ on $U$, and $p$ is arbitrarily chosen on $\Sigma$, we conclude that $\alpha_{k}\rightarrow\alpha$ in $W^{2,1}(\Sigma)$, hence also in $C^{0}(\Sigma)$ (see [80, Theorem 3.3.4 & 3.3.10]).
+Since $\alpha_{k}=(\lambda_{k}-\tilde{\lambda}_{k})\circ\varphi_{k}$ on $U$, and $p$ is arbitrarily chosen on $\Sigma$, we conclude that $\alpha_{k}\rightarrow\alpha$ in $W^{2,1}(\Sigma)$, hence also in $C^{0}(\Sigma)$ (see [81, Theorem 3.3.4 & 3.3.10]).
 ∎
 V Willmore surfaces
-Motivated by the generalization of Willmore surfaces in higher dimensions, we provide in this section, a new proof of the regularity of Willmore surfaces which does not involve the choice of conformal coordinates.
-In Section V.1, we write the Euler–Langrange equation of Willmore surfaces in divergence form.
+Motivated by the generalization of Willmore surfaces in higher dimension, we provide in this section a new proof of the regularity of Willmore surfaces which does not involve the choice of conformal coordinates.
+In Section V.1, we write the Euler–Lagrange equation of Willmore surfaces in divergence form.
 In Section V.2, we prove the regularity of Willmore surfaces.
@@
 In this section, we show that the Euler–Lagrange system of the Willmore functional can be written as a div-curl system. To do so, we need Noether theorem to rewrite it in divergence form.
-Theorem V.1 (Noether’s theorem [154]).
+Theorem V.1 (Noether theorem [157]).
 Let $l\colon\mathbb{R}^{m}\times\mathbb{R}^{m\times n}\rightarrow\mathbb{R}$ for $l(z,p)$ being $C^{1}$ in $z$ and $C^{2}$ in $p$.
@@
 In fact, the divergence-form equation can be seen as a consequence of the pointwise invariance of $H^{2}\,d\textup{vol}_{g}$ by translations in the ambient space, as pointed out in [14] (see also [124]).
-Theorem V.3 ([166, Theorem 1.5]).
+Theorem V.3 ([169, Theorem 1.5]).
 A weak immersion $\vec{\Phi}\in W^{2,2}_{\text{imm}}(\Sigma,\mathbb{R}^{3})$ is a weak Willmore immersion if and only if the following equation holds in $\mathcal{D}^{\prime}\big(\Sigma,\mathbb{R}^{3}\otimes\bigwedge^{2}T^{*}\Sigma\big)$:
@@
 $\begin{aligned} 2\left.\frac{d}{dt}\,H_{t}\right|_{t=0}&=g^{ij}\left.\frac{d}{dt}\,{\mathrm{I\!I}}_{ij,t}\right|_{t=0}+{\mathrm{I\!I}}_{ij}\left.\frac{d}{dt}\,g_{t}^{ij}\right|_{t=0}\\[2.84526pt] &=-g^{ij}\,\partial_{j}\vec{\Phi}\cdot\left.\frac{d}{dt}\,\partial_{i}\vec{n}_{t}\right|_{t=0}-g^{ij}\,\partial_{i}\vec{n}\cdot\partial_{j}\vec{w}+\mathrm{I\!I}_{ij}\left.\frac{d}{dt}\,g_{t}^{ij}\right|_{t=0}.\end{aligned}$ (V.5)
-Thanks to the fact $\Delta_{g}\vec{\Phi}\perp\partial_{l}\vec{\Phi}$, we have
+Thanks to the fact that $\Delta_{g}\vec{\Phi}\perp\partial_{l}\vec{\Phi}$, we have
 $\begin{aligned} (\det g)^{\frac{1}{2}}\,g^{ij}\,\partial_{j}\vec{\Phi}\cdot\left.\frac{d}{dt}\,\partial_{i}\vec{n}_{t}\right|_{t=0}&=-(\det g)^{\frac{1}{2}}\,g^{ij}\,\partial_{j}\vec{\Phi}\cdot\partial_{i}\big(g^{k\ell}(\vec{n}\cdot\partial_{k}\vec{w})\,\partial_{\ell}\vec{\Phi}\big)\\ &=-\partial_{i}\left(\big((\det g)^{\frac{1}{2}}\,g^{ij}\,\partial_{j}\vec{\Phi}\big)\cdot\big(g^{k\ell}(\vec{n}\cdot\partial_{k}\vec{w})\,\partial_{\ell}\vec{\Phi}\big)\right)\\[2.84526pt] &=-\partial_{i}\left((\det g)^{\frac{1}{2}}\,g^{ij}\,g_{j\ell}\,g^{k\ell}(\vec{n}\cdot\partial_{k}\vec{w})\right)\\[2.84526pt] &=-\partial_{i}\left((\det g)^{\frac{1}{2}}\,g^{ik}(\vec{n}\cdot\partial_{k}\vec{w})\right)\\[2.84526pt] &=-\partial_{i}\left((\det g)^{\frac{1}{2}}\,g^{ij}(\vec{n}\cdot\partial_{j}\vec{w})\right).\end{aligned}$ (V.6)
@@
 $\begin{aligned} \mathrm{I\!I}_{ij}\left.\frac{d}{dt}\,g^{ij}_{t}\right|_{t=0}&=-2\,\mathrm{I\!I}_{ij}\,g^{ik}g^{\ell j}\,\partial_{k}\vec{w}\cdot\partial_{\ell}\vec{\Phi}\\ &=2\,\partial_{k}\vec{w}\cdot\left(g^{ik}(\partial_{i}\vec{n}\cdot\partial_{j}\vec{\Phi})\,g^{\ell j}\,\partial_{\ell}\vec{\Phi}\right)\\[5.69054pt] &=2\,\partial_{k}\vec{w}\cdot\big(g^{ik}\,\partial_{i}\vec{n}\big)\\[7.11317pt] &=2\,g^{ik}\,\partial_{k}\vec{w}\cdot\partial_{i}\vec{n}\\[7.11317pt] &=2\,g^{ij}\,\partial_{j}\vec{w}\cdot\partial_{i}\vec{n}.\end{aligned}$ (V.7)
-Combing (V.4)–(V.7), we obtain the pointwise a.e. variation
+Combining (V.4)–(V.7), we obtain the pointwise a.e. variation
 $\begin{aligned} &\left.\frac{d}{dt}\big(H_{t}^{2}\mkern 1.0mu(\det g_{t})^{\frac{1}{2}}\big)\right|_{t=0}\\ &=2H\left.\frac{d}{dt}\,H_{t}\right|_{t=0}(\det g)^{\frac{1}{2}}+H^{2}\left.\frac{d}{dt}(\det g_{t})^{\frac{1}{2}}\right|_{t=0}\\ &=H\mkern 2.0mu\partial_{i}\left((\det g)^{\frac{1}{2}}\mkern 1.0mug^{ij}(\vec{n}\cdot\partial_{j}\vec{w})\right)+H\mkern 1.0mu(\det g)^{\frac{1}{2}}\,g^{ij}\mkern 1.0mu\partial_{i}\vec{n}\cdot\partial_{j}\vec{w}+H^{2}\mkern 1.0mu(\det g)^{\frac{1}{2}}\mkern 1.0mug^{ij}\mkern 1.0mu\partial_{j}\vec{w}\cdot\partial_{i}\vec{\Phi}.\end{aligned}$ (V.8)
@@
 In particular, it holds that $d\vec{L}\in W^{-1,2}+L^{1}(D^{2})$, and by Corollary III.5, we have $\vec{L}\in L^{2}_{\text{loc}}(D^{2})$.
-The quantity $\vec{L}$ alone is not sufficient to deduce some additional regularity since $\vec{L}$ does not possess additional derivative a priori.
+The quantity $\vec{L}$ alone is not sufficient to deduce some additional regularity since $\vec{L}$ does not possess additional derivatives a priori.
 Using the pointwise invariance of $H^{2}\,d\textup{vol}_{g}$ by dilations and rotations, we deduce some conservation laws related to $\vec{L}$ and $\vec{\Phi}$.
 The Noether currents associated to dilations and rotations.
-Theorem V.4 ([168, Theorem 5.59]).
+Theorem V.4 ([171, Theorem 5.59]).
 Let $\vec{\Phi}\in W^{2,2}_{\text{imm}}(D^{2},\mathbb{R}^{3})$ be a weak Willmore immersion.
-Then there exists $\vec{L}\in L^{2}_{\text{loc}}(D^{2},\mathbb{R}^{3})$ satisfying (V.15) [^8] Up to additive constants, this $\vec{L}$ is 1/2-1/2 of the one defined in [168, Theorem 5.59].
-There is a typo in [168, Equation (5.214b)]: the ++ sign should be changed to -.
+Then there exists $\vec{L}\in L^{2}_{\text{loc}}(D^{2},\mathbb{R}^{3})$ satisfying (V.15) [^10] Up to additive constants, this $\vec{L}$ is 1/2-1/2 of the one defined in [171, Theorem 5.59].
+There is a typo in [171, Equation (5.214b)]: the ++ sign should be changed to -.
 and moreover, we have
@@
 From these relations, we derive a system on $S,\vec{R}$ which does not involve $\vec{L}$ anymore.
-Theorem V.5 ([168, Theorem 5.60]).
+Theorem V.5 ([171, Theorem 5.60]).
 Let $\vec{\Phi}\in W^{2,2}_{\text{imm}}(D^{2},\mathbb{R}^{3})$ be a weak Willmore immersion, and define $\vec{L}$ as in Theorem V.4.
@@
 $-*_{g}d\vec{R}\cdot\vec{n}=(\vec{n}\times d\vec{R}+dS\ \vec{n})\cdot\vec{n}=dS.$
-By differentiating the relations (V.19), we obtain a $\mathrm{div}-\operatorname{curl}$ system in $\vec{R}$ and $S$.
-Theorem V.6 ([168, Corollary 5.61]).
+By differentiating the relations (V.19), we obtain a div-curl system in $\vec{R}$ and $S$.
+Theorem V.6 ([171, Corollary 5.61]).
 Let $\vec{\Phi}\in W^{2,2}_{\text{imm}}(D^{2},\mathbb{R}^{3})$ be a weak Willmore immersion, and define $\vec{L},S,\vec{R}$ as before. Then we have (see Notation xiii)
@@
 The Noether current associated to inversions.
-Although $H^{2}\,d\textup{vol}_{g}$ is not pointwise invariant by inversions, the Lagrangian $(H^{2}-K)\,d\textup{vol}_{g}$ is; see for instance [34].
+Although $H^{2}\,d\textup{vol}_{g}$ is not pointwise invariant by inversions, this is the case of the Lagrangian $(H^{2}-K)\,d\textup{vol}_{g}$, see for instance [34].
 In this section, we derive the associated conservation law (V.28), which is equivalent to (V.20c).
@@
 $\partial_{j}\Big((\det g)^{-\frac{1}{2}}\,\text{adj}(\mathrm{I\!I})^{ij}\partial_{i}\vec{\Phi}\Big)\cdot\partial_{\ell}\vec{\Phi}=0.$ (V.24)
-Define $A_{g}\colon T_{x}^{*}(D^{2})\to T_{x}^{*}(D^{2})$ by
+We define the map $A_{g}\colon T_{x}^{*}(D^{2})\to T_{x}^{*}(D^{2})$ by
 $A_{g}(dx^{i})\coloneqq(\det g)^{-1}\text{adj}(\mathrm{I\!I})^{ik}g_{kj}\,dx^{j}.$
@@
 By dilation and translation, it suffices to prove $\nabla S,\nabla\vec{R},\vec{H}\in M^{2,\gamma}(D_{1/2}(0))$.
-By definition of a weak immersion, the coefficients $a^{ij}\coloneqq g^{ij}(\det g)^{1/2}$ satisfies the uniform ellipticity condition (III.13).
-Fix a positive small number $\varepsilon_{0}$ which will be determined later.
+By definition of a weak immersion, the coefficients $a^{ij}\coloneqq g^{ij}(\det g)^{1/2}$ satisfy the uniform ellipticity condition (III.13).
+Fix a small positive number $\varepsilon_{0}$ which will be determined later.
 Since $\vec{\Phi}\in W^{2,2}(D^{2})$, we have $\nabla\vec{n}\in L^{2}(D^{2})$, hence there exists $r_{0}\in(0,\frac{1}{4})$ depending on $\epsilon_{0}$ and $\vec{\Phi}$ such that
 $\sup_{p\in D_{1/2}}\int_{D_{r_{0}}(p)}|\nabla\vec{n}|^{2}\,dx^{1}\wedge dx^{2}<\varepsilon_{0}.$ (V.29)
-Let $p\in D_{1/2}(0)$, $r\in(0,r_{0})$. By Lemma III.3, we can define $\Psi_{S}\in W^{1,2}_{0}(D_{r}(p))$ to be the unique solutions of (see Section III.1)
+Let $p\in D_{1/2}(0)$, $r\in(0,r_{0})$. By Lemma III.3, we can define $\Psi_{S}\in W^{1,2}_{0}(D_{r}(p))$ to be the unique solution of (see Section III.1)
 $\begin{cases}\partial_{i}(a^{ij}\,\partial_{j}\Psi_{S})=\nabla^{\perp}\vec{n}\cdot\nabla\vec{R}&\quad\mbox{in }D_{r}(p),\\[8.53581pt] \Psi_{S}=0&\quad\mbox{on }\partial D_{r}(p).\end{cases}$ (V.30)
-We can also define $\vec{\Psi}_{\vec{R}}\in W^{1,2}_{0}(D_{r}(p))$ to be the unique solutions of
+We also define $\vec{\Psi}_{\vec{R}}\in W^{1,2}_{0}(D_{r}(p))$ to be the unique solution of
 $\begin{cases}\partial_{i}(a^{ij}\,\partial_{j}\vec{\Psi}_{\vec{R}})=\nabla^{\perp}\vec{R}\times\nabla\vec{n}+\nabla\vec{n}\ \nabla^{\perp}S&\quad\mbox{in }D_{r}(p),\\[8.53581pt] \vec{\Psi}_{\vec{R}}=0&\quad\mbox{on }\partial D_{r}(p).\end{cases}$ (V.31)
@@
 $\begin{aligned} \partial_{i}(a^{ij}\,\partial_{j}\vec{n})&=(\det g)^{\frac{1}{2}}*_{g}d*_{g}d\vec{n}\\[2.84526pt] &=(\det g)^{\frac{1}{2}}*_{g}d(\vec{n}\times d\vec{n}-2H*_{g}d\vec{\Phi})\\[2.84526pt] &=2\,\partial_{1}\vec{n}\times\partial_{2}\vec{n}-2\partial_{i}(a^{ij}\,H\,\partial_{j}\vec{\Phi}).\end{aligned}$ (V.38)
-Fix a positive small number $\varepsilon_{0}$ which will be determined later.
+Fix a small positive number $\varepsilon_{0}$ which will be determined later.
 As in the proof of Theorem V.8, there exists $r_{0}\in(0,\frac{1}{4})$ depending on $\epsilon_{0}$ and $\vec{\Phi}$ such that
@@
 $\|\nabla\vec{\Psi}_{1}\|^{2}_{L^{2}(D_{r}(p))}\leq C(\Lambda)\,\|\nabla\vec{n}\|^{4}_{L^{2}(D_{r}(p))}\leq C(\Lambda)\,\epsilon_{0}\,\|\nabla\vec{n}\|^{2}_{L^{2}(D_{r}(p))}.$
-Furthermore, by Riesz representation theorem for Hilbert spaces, there exists a unique $\vec{\Psi}_{2}\in W^{1,2}_{0}(D_{r}(p))$ satisfying the estimate $\|\nabla\vec{\Psi}_{2}\|_{L^{2}(D_{r}(p))}\leq C(\Lambda)\,\|H\|_{L^{2}(D_{r}(p))}$ and the system
+Furthermore, by Riesz representation theorem for Hilbert spaces, there exists a unique $\vec{\Psi}_{2}\in W^{1,2}_{0}(D_{r}(p))$ satisfying the estimate $\|\nabla\vec{\Psi}_{2}\|_{L^{2}(D_{r}(p))}\leq C(\Lambda)\,\|H\|_{L^{2}(D_{r}(p))}$ and the following equation holds
 $\partial_{i}(a^{ij}\,\partial_{j}\vec{\Psi}_{2})=-2\,\partial_{i}(a^{ij}\,H\,\partial_{j}\vec{\Phi}).$ (V.40)
@@
 $\sup_{p\in D_{1/2}(0),r\leq r_{0}}r^{-\tau}\int_{D_{r}(p)}|\nabla\vec{n}|^{2}<\infty.$
-The Hölder continuity of $\vec{n}$ then follows from standard knowledge of the Morrey–Campanato spaces (see for instance [150, Theorem 3.5.2] and [2]). ∎
+The Hölder continuity of $\vec{n}$ then follows from standard knowledge of the Morrey–Campanato spaces (see for instance [153, Theorem 3.5.2] and [2]). ∎
 If the Gauss map is continuous, then one can write the image of $\vec{\Phi}$ as a graph.
@@
 Let $\Sigma$ be a $2$-dimensional closed smooth manifold, $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(\Sigma,\mathbb{R}^{3})$.
-Then we have $\vec{\Phi}$ is a graph near $\vec{\Phi}(p)$ for any $p\in\Sigma$, and $\vec{\Phi}\in C^{\infty}$ under the graph coordinates.
+Then $\vec{\Phi}$ is a graph near $\vec{\Phi}(p)$ for any $p\in\Sigma$, and $\vec{\Phi}\in C^{\infty}$ under the graph coordinates.
 Proof.
@@
 Hence by direct computation we see that $\nabla f,\nabla\vec{\Phi}\in C^{0,\alpha}_{\textup{loc}}(D^{2})$. In particular, $g=g_{\vec{\Phi}}\in C_{\textup{loc}}^{0,\alpha}(D^{2})$.
-By [145, Eq.
+By [148, Eq.
 (1.5)] and standard estimates on the Riesz potential [1, Proposition 3.2(ii)], we obtain $M^{1,\tau}(D^{2})\hookrightarrow W^{-1,p}(D^{2})$ for any $p\in(2,\frac{2-\tau}{1-\tau})$.
@@
 Now return to the equations (V.43a)–(V.43d).
-The right hand side of each equation is in $L^{q}_{\textup{loc}}(D^{2})$ for any $q<\infty$, and since $S,\vec{R},\vec{n},g\in W^{1,q}_{\textup{loc}}(D^{2})$, by [38, Theorem 4.1] and the proof of [66, Theorem 8.8], we obtain $S,\vec{R},\vec{n}\in W^{2,q}_{\textup{loc}}(D^{2})$, and then $\vec{\Phi}\in W^{3,q}_{\textup{loc}}(D^{2})$… We finally get $\vec{\Phi}\in C^{\infty}$ under the graph coordinates.
+The right-hand side of each equation is in $L^{q}_{\textup{loc}}(D^{2})$ for any $q<\infty$, and since $S,\vec{R},\vec{n},g\in W^{1,q}_{\textup{loc}}(D^{2})$, by [38, Theorem 4.1] and the proof of [66, Theorem 8.8], we obtain $S,\vec{R},\vec{n}\in W^{2,q}_{\textup{loc}}(D^{2})$, and then $\vec{\Phi}\in W^{3,q}_{\textup{loc}}(D^{2})$… We finally get $\vec{\Phi}\in C^{\infty}$ under the graph coordinates.
 ∎
@@
 This completes the proof.
-By the embedding results in Lemma III.20, for all $1\leq j\leq 4$, $a\in L^{\infty}\cap W^{1,4}(B^{4})$, and $f\in L^{2,\infty}(B^{4})$, we have
+By the embedding results in Lemma III.20, for any index $1\leq j\leq 4$ and any maps $a\in L^{\infty}\cap W^{1,4}(B^{4})$ and $f\in L^{2,\infty}(B^{4})$, we have
 $\|a\mkern 2.0mu\partial_{j}f\|_{W^{-1,(2,\infty)}(B^{4})}$ $\leq\|\partial_{j}(af)\|_{W^{-1,(2,\infty)}(B^{4})}+C\mkern 1.0mu\|f\mkern 2.0mu\partial_{j}a\|_{L^{\frac{4}{3},\infty}(B^{4})}$
@@
 […]
-Does the system of equations (VI.14) correspond to Euler-Lagrange equations related to the variations of the Dirichlet Energy of the mean curvature under some constraint?
+Does the system of equations (VI.14) correspond to Euler–Lagrange equations related to the variations of the Dirichlet Energy of the mean curvature under some constraint?
 Recall from Remark II.23 that in 2 dimension the corresponding system of equations is equivalent to the Conformal Willmore Equations obtained by taking variations of the Willmore equations with some constraint on the underlying conformal class induced by the metric.